A Bibliography of Publications in Theoretical Computer Science: 2020–2024

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

26 April 2024 Version 1.65

Title word cross-reference

(1 + 1) [439, 362, 598]. (1, λ) [2267]. (2 + ϵ)k [948]. (c, r) [44]. (H_{1}, H_{2}) [2020]. (k, l, m) [1424]. (k, n, 0, 1, 1) [883]. (k, p) [1280]. (E + Var) [844]. (μ + 1) [1021]. (n, 3) [47]. (n, k) [611, 2039, 1032]. (r, ℓ) [1725]. (s, t) [1872]. 0 · 6 [2398]. 1 [227, 419, 428, 738, 1221, 515, 884]. {1, 3} [1379]. 1/2 [570]. 2 [1392, 613, 274, 2225, 2131, 2013, 2094, 1904, 1281, 544, 1511, 2406, 2351, 1652, 1955, 146, 844, 1775, 1468, 1284, 2367, 1588, 1776, 2353, 2320, 63, 1865, 2361, 1379, 375, 704, 1731, 2327, 772, 1505, 1633]. 2^{2k} [2290]. 2^{O(k)}n [584]. 2n [1153]. 3 [700, 1336, 649, 387, 1968, 995, 1288, 1049, 1119, 1129, 1490, 884, 1589, 1988, 1243, 2061,

2299]. 4 [2013, 624, 489, 611, 1865]. 4K_{1} ; [1459]. 5 [1459, 2389]. 5k [1420, 2061]. 6 ; [489, 1865]. 7, 7, 5 [297]. 8, 8, 5 [297]. [1, j] ; [64]. ^{⋆} [2365]. ^{th} [1935]. ^{T M} [383]. _{1} [162]. _{5} ; [900]. _{R} [1297]. b [304, 2473, 1202]. β ; [2294, 1893]. C [1110]. C_{≥5} [1485]. ; C_{n}(1, 2, 3) [338]. d [287, 2072]. D_{PSNI} [1095]. ; Δ [620, 2211, 1766]. ℓ_{1} [1887]. ℓ_{p} [1921, 704]. ; ϵ [44]. F_{4} [2073, 2379, 365]. g [1531, 2149, ; 1178, 1367, 2058, 59, 1403, 1556, 2292]. ; GF (2) [996]. H ; [1686, 1348, 1309, 1768, 1811, 1269, 702, 1546]. ; i [1107]. k [888, 1793, 1322, 1735, 111, 1381, ; 559, 1645, 220, 1979, 900, 2316, 183, 1274, ; 343, 190, 1479, 2188, 544, 1941, 88, 2199, 911, ; 1463, 1832, 1007, 1099, 1205, 608, 462, 1390, ; 1969, 1183, 520, 1242, 471, 801, 303, 2253, ; 1722, 2230, 1432, 696, 2421, 284, 783, 1985,

1931, 1774, 1537, 2328, 747, 1952, 2339, 2422, 2467, 1227, 584, 749, 688, 1597, 1884, 932]. k^{2} [1220]. K_{2,t} [1477]. K_{2n} [1037]. K_{4} [1546]. L [305, 957, 265, 162, 115]. l_{2,q} [195]. l_{k} [615]. λ [250, 961, 650, 2181]. ⌊3n/2⌋ [109]. m [1109, 220, 1497, 1685, 2352]. M^{p} [513]. Z [1684]. Z_{n} × Z_{n} × Z_{n} [109]. F [1740]. L [877, 1525, 2264]. M, N [2308]. V [454]. ESR [844]. NP [383]. qcb_{0} [1933]. SuperE [844]. n [1807, 1793, 471, 1589, 1153, 1931, 1546, 1952, 986, 2352]. n = 1, 2, 3 [127]. n^{2} [1807]. n ≥ 4 [647]. O(log log N) [2325]. O(log N) [155, 44, 644]. O(ρ log n) [1461]. ω [259, 2087, 172, 466]. ωT [259]. P [77, 76, 2152, 900, 2055, 337, 2367, 80, 2151, 81, 316, 14, 2378, 1889, 2440, 103, 1599, 2494, 513, 1543]. p/q [1944]. P_{10} [2392]. P_{2} [1420]. P_{5} [2078]. π [2064]. q [2219, 1644]. r [1378, 233, 1269, 929, 1739, 332]. R^{d} [2420]. R_{g} [271, 690, 1306]. ρ [1461]. S [2472, 2135]. S_{1,2,4} [1044, 1114, 1124]. t [1751, 1948, 1202]. t/k [2352, 1289]. t/m [1931]. t/s [1152]. θ [87]. V [1057]. w [2073, 2379]. W_{2} [2394]. Z^{d} [2420]. Z_{n} [1766]. |(λ)| [961].

-abelian [462, 1776]. -acyclicity [1893]. -adhesivity [2308]. -adic [2472]. -algebras [2072]. -algorithms [2264]. -approximation [544]. -arboricity [1865]. -arc-strong [1775]. -ary [1793, 1109, 471, 1589, 1931, 1952, 2352]. -binomial [2353]. -block [783]. -bonacci [911, 1497]. -bounded [1284]. -bubble-sort [1032]. -burning [1722]. -calculus [250, 2064, 961, 650, 2181]. -caterpillars [1274]. -centdian [1599]. -center [1381, 1832, 1969, 2467]. -centers [932]. -character [2230]. -chromatic [304]. -Circle [1479]. -club [2135, 2327]. -colored [649, 1281]. -coloring [700]. -competitiveness [2225]. -complete [2211]. -completeness [383]. -component [1178, 1269, 2058, 1739]. -conditional [690, 1306]. -connected [884, 63, 704].

-connectivity [2185, 1490, 1985, 1243]. -constacyclic [2073, 2379]. -convexity [620]. -Critical [900]. -cube [471, 1589, 1952, 2352]. -cubes [1793, 1931, 986]. -cuts [1872]. -cycle [584]. -cycles [489, 1865]. -decreasing [1644]. -degenerate [343]. -description [2294]. -designs [2389]. -diagnosability [1152, 2352, 1289]. -diagnosable [1751, 1202]. -diagnosis [1152, 1931]. -dimensional [1685]. -disjoint [2351, 1432, 375]. -distance [1633]. -distinguishing [1379]. -domains [957, 265]. -dominating [190]. -domination [64, 696]. -dynamic [332]. -EA [2267, 1021, 598]. -eccentricity [1183]. -edge-connected [2094, 1904]. -edge-connectivity [1988]. -edge-weighting [1379]. -enhanced [2039]. -equilibria [844]. -equivalence [287]. -extra [1531, 2149, 1652, 1269, 1367, 1546, 1403]. -facility [1884]. -factorizations [515]. -free [1044, 1114, 1124, 900, 2078, 1768, 2392, 1686, 1348, 1459, 1740, 1309, 1811, 2020, 1485]. -game [297]. -gathering [929]. -good [2058]. -good-neighbor [59, 1556, 2292]. -hardness [1948]. -Helly [1941]. -hitting [624]. -hop [1955]. -hub [103, 1543]. -index [233]. -invariant [87]. -isometric [772]. -level [747]. -linear [1979]. -linkage [274]. -Lyndon [466, 172]. -machine [1322]. -mappability [227]. -Mark [1107]. -matching [2473, 2422]. -matchings [1202]. -matroid [2440]. -maximal [1336]. -MAXSAT [47]. -means [2421, 749, 1110]. -median [419, 428, 2494, 2339, 1597]. -minimal [2320]. -minimum [1241]. -minor [1477]. -nearest-neighbor [801]. -NN [44]. -node-connectivity [2361]. -norm [195, 615]. -operations [2450]. -order [1057]. -packing [1420]. -parities [559]. -partite [1242]. -partition [88]. -partitions [613]. -path [2061]. -planar [738]. -planarity [1280]. -plane [1221, 884]. -players [844]. -polytopes [454]. -power [183]. -prize-collecting [608]. -pseudoforest [115]. -PVCS [883]. -reductions [1525, 877]. -regular [259, 2087, 1968, 1205, 1390, 2299, 259]. -resilient [1511, 2406]. -restricted [702]. -round [1049, 1119, 1129]. -SAT [387]. -secretary [888]. -sections [1468]. -separable [1588]. -set [2188, 611]. -sided [1288]. -sink [111, 1007]. -spaces [1933]. -spanner [1948]. -spanners [1461]. -star [611]. -Steiner [544]. -subgraph [520, 688]. -submodular [1537, 2328]. -Subset [2219]. -supplier [2316, 1832]. -system [1774]. -th [1546, 513]. -threshold [1505]. -traceability [2013]. -transmitters [303]. -traveling [1735]. -tree [1220]. -tree-generated [1392]. -trees [1281]. -tromino [305]. -TSP [1463]. -Uncut [688]. -uniformity [1766]. -universal [2199]. -variable [1153]. -verifiable [1424]. -vertex [146, 948, 1420, 2061]. -well-coveredness [1725]. -wheel [1459].

0/1/2 [2136]. 0/1/2-player [2136].

1 [2390, 2310, 2469, 2359, 2434]. 1-110 [1407]. 1-162 [1634]. 1-2-3 [1226]. 10 [1617, 1764, 2000, 1936, 2100, 2247]. 11 [1664, 2121]. 12 [1438, 1515, 1385, 2033, 2402, 2322, 2479, 2368, 2443]. 13 [2167]. 14 [1465, 1729]. 14th [1529]. 15 [2074]. 16 [1975]. 16th [2186]. 17 [2213, 1895]. 18 [1625, 1777, 2008, 2105]. 19 [1526, 1474].

2 [2050, 1418, 2234, 916, 1166]. 2-Club [2050]. 2-hop [1241]. 2-player [2136]. 2-SAT [1166]. 20 [1446, 1946, 2268]. 2017 [157]. 2019 [1133]. 2021 [2161, 1998]. 2022 [1418, 1428, 1438, 1446, 1617, 1625, 1634, 1503, 1515, 1526, 1540, 1578, 1604, 1375, 1385, 1395, 1407, 1456, 1465, 1474, 1483, 1493, 1764, 1777, 1787, 1718, 1729, 1744, 1753, 1647, 1664, 1681, 1701]. 2023 [2000, 2008, 2017, 2121, 2133, 2143, 1924, 1936, 1946, 1958, 1914, 2065, 2074, 2083, 2092, 2100, 2105, 2112, 1965, 1975, 1982, 1991, 2024, 2033, 2043, 2056, 2234, 2247, 2268, 2200, 2213, 2223, 1819, 1854, 1876, 1895, 2153, 2167, 2179, 2190]. 2024 [2390, 2402, 2415, 2425, 2345, 2310, 2322, 2335, 2278, 2288, 2469, 2479, 2489, 2359, 2368, 2380, 2434, 2443, 2452, 2461]. 21 [1681, 2133, 2179, 2415, 2335, 2380, 2452]. 22 [2083, 2043]. 23 [1744]. 24 [1540, 1395, 1982, 2489]. 25 [1483, 1914, 2223]. 25th [823]. 26 [1634, 1578, 1787, 2017, 2092, 2112]. 27 [2425, 2345, 2461]. 272 [340]. 27th [1133]. 28 [1407, 1958]. 29 [1493, 1701, 2143, 2190]. 2PC [1211].

3 [1647, 2104, 301]. 3-SAT [2104, 301]. 30 [1604, 2056, 1944]. 31 [1753, 1991]. 3s [1226]. 3SAT [90].

4 [1924, 2024, 1854, 2278].

5 [1503, 1375].

6 [1718, 2153]. 60th [1247]. 65th [1628].

7 [1456, 2065]. 70th [1999, 861]. 732 [339]. 769 [129]. 770 [133]. 795 [591].

8 [1428, 2288]. 862 [1525]. 866 [2212].

9 [1965, 2200, 1819, 1876]. 943 [2094]. 960 [2379].

AADL [1351]. AAIM’19 [742]. ABC [1244]. ABE [2067]. abelian [1814, 597, 462, 1776]. aborts [1805]. above [47, 344]. ABPs [169]. absent [1643, 1562]. absolute [2414, 1151]. Absolutely [633]. Abstract [1255, 959]. abstraction [2233, 503, 1137]. abstraction-refinement [503]. abstractions [1236]. AbU [2053, 2474]. Accelerated [241]. acceptable [1492]. acceptance [549, 39]. accepted [1235, 1201]. Access [124, 883, 1213, 1709, 1859, 1141, 1260]. Accountable [175, 2466]. Accurate [1620]. achromatic [1592]. activation [2358, 1700, 517, 1972]. Acyclic [1899, 1077, 1548, 2052, 1010, 1791]. acyclicity [1893]. ad [1335, 216, 2218, 1571]. ad-hoc [1335]. adaptation [443, 1353]. adaptation-complete [1353]. Adaptive [1282, 1213, 1657, 1748, 524, 1512, 2128, 2324, 1805, 1690, 164, 1055, 1088, 709, 1774, 1918, 333]. Adaptively [1926, 2449]. adaptivity [1437]. Adding [253]. addition [1829, 2272]. additional [1981]. Additive [1461, 546, 1568, 597]. adhesivity [2308]. adic [2472]. adjacent [1865]. adjusting [2267]. adjustment [1050, 1120, 1130]. admissible [1315]. admission [1500, 1572, 1955]. admitting [2436]. adversarial [121]. adversaries [2446, 1506]. adversary [2236, 2350]. advertising [949]. Advice [2324, 867, 1500, 784, 1523]. advised [1792]. AFAs [2459]. affine [130, 514]. after [833]. again [506]. against [827, 12, 2446, 1331, 2220, 973]. age [1929]. ageing [446]. agent [121, 541, 74, 529]. agents [680, 2281, 953, 851, 73, 706, 618, 1871, 2338, 388, 192]. aggregate [762]. aggregation [857, 1672, 550, 1362, 42, 1620]. AGITCPN [1673]. agreeable [1140]. agreed [351]. Agreeing [1563]. agreement [1964, 2010, 2446, 2188, 1415, 2114, 701, 1177]. ahead [1639]. aided [292]. AIFV [916]. AIFV-2 [916]. air [1653, 347]. Aldo [461]. algebra [1760, 1621]. Algebraic [2273, 20, 286, 1984, 1066, 1157, 265, 1758, 2414]. algebras [2072]. Algorithm [545, 54, 1912, 1435, 2488, 2050, 2042, 662, 735, 1372, 2209, 792, 2395, 2482, 2317, 1043, 1113, 1123, 1328, 32, 2316, 2282, 2094, 1904, 1598, 72, 354, 623, 544, 721, 1056, 36, 373, 2261, 1271, 272, 456, 1003, 1849, 45, 586, 608, 743, 2302, 1031, 1641, 1010, 1284, 2012, 984, 1012, 1383, 1689, 2486, 1666, 1576, 115, 1152, 490, 520, 296, 2365, 44, 801, 1649, 240, 1599, 899, 553, 1880, 241, 90, 2433, 2147, 415, 424, 3, 129, 929, 29, 2396, 1179, 1186, 1869, 2494, 107, 744, 2328, 435, 1988, 1535, 644, 2015, 69, 333, 1006, 584, 37, 1369, 1691, 2429, 2348, 1029, 1864, 445]. Algorithmic [1133, 56, 2115, 538, 1038, 1495, 742, 1806, 1583, 243, 1584, 2419, 1638, 2020]. Algorithmics [1529]. Algorithms [1287, 1530, 1716, 47, 2262, 906, 2297, 950, 1155, 200, 560, 809, 2186, 473, 2327, 1195, 39, 760, 1881, 16, 2397, 1087, 1602, 227, 982, 526, 2331, 640, 1735, 836, 279, 1792, 2080, 1705, 1728, 700, 1512, 1981, 2324, 1014, 684, 1817, 1509, 103, 803, 2430, 1532, 1318, 446, 291, 753, 190, 738, 2205, 2188, 1143, 2165, 1784, 1596, 1415, 1347, 2456, 1223, 1740, 1955, 525, 1780, 1218, 1463, 275, 218, 551, 2096, 2063, 2265, 1007, 1176, 1721, 204, 2283, 2410, 2325, 1361, 1675, 1838, 2103, 34, 1110, 820, 1001, 1595, 1706, 2320, 1723, 637, 2319, 2230, 1533, 797, 2264, 1207]. algorithms [1379, 2244, 1727, 2130, 1913, 1614, 1443, 2005, 349, 1241, 681, 53, 2330, 1781, 1828, 2004, 747, 987, 758, 1496, 1843, 119, 1561, 2339, 2467, 299, 648, 1488, 362, 370, 749, 1868, 1884, 1866, 2038, 1162, 2376, 1725, 1534]. alignment [1939, 298]. all-but-one [1088]. all-true [47]. alleviation [320]. alliances [1659, 2386]. allocation [138, 1165, 1738, 571, 894, 581, 37]. allocations [2385, 2089, 2408]. allowing [2095]. Almost [2300, 1705, 2446, 1007, 571, 1697, 1209, 1696, 670, 755, 648]. almost-increasing [1696, 755]. Almost-surely [2446]. along [1315]. alphabet [1814, 911, 875, 971, 791, 1731]. alphabet-invariant [971]. alphabets [1251, 2183]. also [769]. ALT [157]. Alternating [255, 549, 243]. Alternating-time [255]. Alternation [970].

Alternative [94]. alternatives [962]. amalgamated [791]. amalgamation [673]. ambient [2441]. ambiguity [868, 840, 2193]. among [2060]. amortized [585, 548, 1012, 1580]. AMP [2321]. amplified [187]. analog [2414]. analyses [1796]. Analysis [2022, 439, 1847, 441, 391, 1268, 598, 1250, 1417, 1672, 1883, 9, 721, 1890, 1286, 1955, 502, 2321, 443, 2277, 2164, 164, 283, 930, 548, 1908, 1911, 180, 1653, 1167, 1894, 471, 1906, 560, 1534, 2058, 2246, 444, 440, 1241, 1997, 1356, 1779, 1351, 39, 362, 327, 760, 273, 1100, 188]. analytics [1637]. Analyzing [994, 616, 2340]. anarchy [130, 1384]. and-depth [361]. angular [1901]. annealing [323]. annealing-based [323]. annulus [457]. anomalous [1714]. anomaly [1707]. anomaly-based [1707]. anonymization [1000, 494]. Anonymous [311, 2449, 494, 1547, 2423, 1154, 1209, 1861]. answering [1762]. Anti [675, 1813, 14]. anti-slide [1813]. anti-spikes [14]. Anti-unification [675]. Antipodal [1109]. antiport [81]. any [1823, 351]. APCol [78]. Aperiodic [1487]. aperture [2290]. apex [2159]. Application [966, 1748, 2313, 2040, 1584, 218, 1093, 1269, 1237, 839, 1440]. applications [1191, 366, 821, 866, 276, 100, 2363, 131, 1421, 2283, 582, 2020, 1926, 104, 688, 50, 824]. applicative [257]. applied [2064]. Applying [1320]. approach [1160, 1066, 2207, 1762, 486, 1770, 1462, 36, 1737, 963, 376, 2277, 1607, 2110, 1852, 1069, 1192, 958, 761, 2421, 1485, 1074, 1856, 1237, 201]. approaches [2068, 2041]. approval [2138]. Approximability [2206, 693, 841, 1471, 931, 2212, 1726, 1879, 2448]. Approximate [1098, 234, 224, 1196, 2363, 980, 1747, 2178, 2408, 1902, 1964, 1384, 1568, 1737, 2274, 38, 154, 1434]. approximately [2270, 1218]. Approximating [146, 2333, 1700, 688, 11, 801, 1818].

Approximation [1040, 1735, 1423, 103, 822, 525, 631, 275, 204, 2410, 1666, 490, 520, 34, 1110, 1595, 1704, 1207, 1030, 696, 2396, 747, 987, 2339, 1488, 749, 1162, 1864, 2397, 1602, 1486, 363, 1728, 642, 326, 208, 1939, 2316, 2282, 2094, 1904, 190, 819, 544, 1596, 162, 2077, 1832, 608, 743, 288, 1838, 115, 480, 484, 1723, 1695, 1722, 814, 899, 2382, 1186, 1869, 744, 2328, 1574, 894, 119, 1543, 1048, 1118, 1128, 1746, 1868, 1884, 2182, 2376, 982]. approximations [1891, 1638]. April [245, 267, 880, 891, 907, 919, 1418, 1428, 1438, 1446, 2000, 2008, 2017, 2390, 2402, 2415, 2425]. APX [1722]. APX-hardness [1722]. Aquila [1859]. arbitrarily [1291]. Arbitrary [306, 685, 1878, 2091, 11, 189, 217, 968, 2140, 2362, 2244, 1224]. arbitrary-radix [2091]. Arbitrary-size [2091]. Arborescence [733]. arborescences [1905]. arboricity [381, 489, 1865]. arc [1661, 2185, 1867, 1775, 2237, 1179]. arc-connectivity [2237]. arc-partition [1661]. arc-reversals [2185]. architecture [1672, 968, 1710]. architectures [1850, 1636]. area [55, 740]. argumentation [73]. arguments [1099, 1646]. arithmetic [1750, 1552, 25, 1908]. arm [238]. Array [2367, 628, 2131, 865, 1057, 1888, 1673]. arrays [2286, 1480]. articles [823]. artificial [446]. ary [1793, 1109, 2126, 471, 1589, 1931, 1952, 2352]. ascent [57]. ascription [502]. Asian [1636]. Aspects [742, 2400, 2204, 1672, 1640, 1398, 56, 2115, 1038, 615, 280, 981]. assembly [1252, 716, 995, 1134]. assessment [71, 1174, 1234, 1897]. assessments [671, 2016, 2277]. assignment [55, 116, 47, 498, 72, 189, 1135, 40, 525, 617, 492]. assignments [108, 2255, 181, 1724]. assisted [10]. Associated [2367]. assumption [2035]. assumptions [1482, 1213, 1209]. asymmetric [1182]. Asymptotic [2126, 376, 1630, 2445]. Asynchronous [2236, 1631, 1559, 306, 685, 2446, 1479, 2344, 2264, 816, 701, 1954, 388, 192, 1148]. Asynchrony [1068]. AT-free [763, 1687]. atoms [1520]. attack [2064, 1707, 1279, 1593, 2070, 1108, 1815]. attacks [1858, 12, 2481, 568, 1434, 187]. attention [2059]. attention-driven [2059]. Attractor [1954]. attractors [2334, 2354, 708]. Attribute [913, 293, 2079, 1668, 65, 1534, 2053, 1209, 1237, 915]. Attribute-based [913, 293, 1668, 65, 2053, 1209]. attributed [1442, 2097, 674]. auction [1816, 2218]. auctions [651, 2431, 284, 578]. augmentation [910, 1361]. augmented [442, 2149, 2049, 1931, 1996]. augmenting [1705, 1337, 1195]. August [412, 421, 430, 1051, 1070, 1617, 1625, 1634, 2121, 2133, 2143]. authenticated [1861, 1334]. authentication [2493, 285]. authority [175]. authorization [856, 1674]. Automata [966, 1217, 1630, 259, 549, 596, 1312, 1987, 1967, 597, 2263, 1514, 699, 1737, 1836, 901, 318, 1175, 1502, 2366, 456, 968, 1810, 1757, 970, 1523, 321, 719, 1313, 972, 2455, 2344, 2252, 2424, 1844, 791, 465, 2468, 1609, 705, 2377, 1560, 974, 258, 1441, 1639, 2003, 2459, 2364]. automata-theoretic [1987]. automated [1849, 488, 2327, 69]. automatic [2229, 821, 2313, 1200, 1612]. automaton [197, 1244]. Automaton-ABC [1244]. Automedian [734]. automorphisms [1354]. autonomous [727, 1736, 1873]. autonomy [74]. autoreducible [384]. Availability [1912]. available [962]. AVD [1806]. average [1545, 147, 2027, 1929]. average-case [1545]. aversion [982]. avoidability [1291]. avoidance [926, 2371]. Avoiding [1558, 2210, 574]. Aware [1688, 2316, 493, 40, 371, 33]. axiomatisability [506]. axiomatization [2432, 357]. axis [701, 3, 129]. axis-parallel [3, 129].

B [564]. B-trees [564]. back [549, 118].

back-off [118]. backoff [1412].

backreferences [1825]. backup [492]. Backward [394]. bad [2240]. bags [372]. balance [1742, 758]. balance-fair [1742]. Balanced [1083, 45, 1670, 1814, 189, 580, 1802, 1554, 2445, 1953, 635, 1894, 2357, 1138, 2319, 2472, 1809]. Balanced-flow [45]. balancing [522, 2085, 814, 1265]. ball [1267]. balls [2420, 2241]. bamboo [1413]. Band [1279]. bandit [2108, 1013, 2265, 166]. bands [2002]. Bandwidth [1805]. Bandwidth-efficient [1805]. Barrier [479, 846, 2225, 679, 1147, 1864]. base [55, 1662, 831, 1666, 334]. base-stations [55]. Based [1912, 293, 1886, 1822, 1858, 1190, 2131, 2195, 764, 1065, 1850, 260, 1339, 2313, 1328, 295, 404, 1165, 1743, 291, 587, 72, 1143, 819, 503, 491, 1654, 2059, 2079, 1370, 979, 1668, 2449, 242, 890, 1331, 2456, 2358, 309, 269, 494, 502, 567, 1607, 2110, 147, 2493, 1769, 2325, 768, 1707, 1860, 177, 1919, 1637, 1174, 65, 278, 663, 1055, 180, 1383, 1653, 1367, 1708, 856, 117, 1110, 1167, 2259, 2423, 296, 913, 2320, 323, 1655, 1723, 717, 46, 1714, 560, 1534, 1785, 2058, 444, 2053, 1209, 1345, 1674, 536, 1861, 1141, 1334, 334, 472, 1356, 292, 1636, 1613, 630, 682, 1610, 744, 285]. based [476, 1824, 2119, 2399, 2349, 154, 1758, 39, 333, 1424, 1473, 2045, 2088, 1866, 2429, 2364, 175, 1180, 187, 50, 839, 1864, 1869]. bases [1762, 281, 2141]. basic [2450, 155, 1352]. basis [12, 404, 63, 1326]. batch [418, 427, 563]. batching [1972]. batching-machines [1972]. Baumslag [1249, 1487]. Bayesian [141, 573, 670, 1356]. BB84 [12]. BC [1227]. BC-subtrees [1227]. BDI [73]. be [531, 1751, 2202, 733, 2164]. Beachcombing [101]. Beating [570]. Beatty [196]. beaver [1570]. Beep [1978]. Beep-and-Sleep [1978]. beeping [208]. Beeps [124]. befriend [209]. behavior [597, 2358, 1653, 1714]. Behavioral [2474, 569]. behaviors [875]. Behavioural [1452]. behaviours [874]. Being [769, 1104]. Belief [1275, 1714, 993]. Belief-invariant [1275]. below [1812]. bend [417, 426]. best [194, 1669, 914, 1724]. best-case [1724]. Better [700, 1597]. between [858, 868, 1922, 1499, 452, 420, 429, 1944, 1591, 1201]. betweenness [1485]. Beyond [61, 259, 624, 1362, 2448, 709, 1774, 942, 2282, 1307, 2304, 544, 1694, 2457, 1893, 2355]. beyond-planar [942, 1307]. BFS [546, 2022, 1772]. bi [2261, 814, 1251]. bi-criteria [2261, 814]. Bi-immunity [1251]. biased [2465]. bichromatic [1273]. Bicluster [2004]. bicomponent [2145]. biconnected [2475]. Bicriteria [758, 1949, 29]. Bicube [714, 1289]. bicubes [1050, 1120, 1130]. bicyclic [1588]. bidders [651]. bidding [2431]. bidirectional [1732]. big [2430, 523, 1720, 1637]. bijective [1880]. bike [38]. bilinear [361, 364]. Billaud [1835]. bimonoids [1514]. bin [1789, 1890, 2301, 2257]. Binary [922, 278, 1750, 1807, 93, 698, 13, 2401, 2384, 1741, 1510]. Binary-decision-diagram-based [278]. binomial [2353]. biochemical [1233]. biological [2460, 870, 997]. biology [1022]. Biometric [2493, 1906, 285]. Biometric-based [2493, 285]. biometrics [1861]. biometrics-based [1861]. BioModels [1234]. biomolecular [321]. bipancyclicity [1951]. Bipartite [112, 2296, 1968, 515, 2473, 2196, 556, 1679, 1601, 63, 1326, 1078, 1432, 1951, 1106, 53, 693, 351]. bipartition [1224]. BipHook [1993]. birds [570]. birthday [1999, 1628, 1247, 1935, 2098, 861]. Bisection [344, 1003]. Bishop [1761]. Bisimilarity [1352, 26, 1757]. Bisimulation [392, 1298, 600, 958]. bitprobe [524, 122]. bitstring [1421]. Bivariate [2165]. black [9]. black-box [9]. Block [1257, 1206, 1943, 1883, 833, 1769, 2314, 2047, 793, 1599, 1207, 1861, 783]. block-based [1769]. block-interchanges [2047, 1207]. block-wise [833]. blockchain [1883]. Blocking [1036, 1977, 548, 560, 36]. BLUE [2148, 1048, 1118, 1128]. Board [7, 18, 422, 431, 436, 448, 460, 470, 483, 496, 508, 528, 543, 562, 576, 593, 605, 622, 639, 653, 669, 28, 49, 67, 84, 134, 156, 168, 205, 214, 225, 246, 268, 290, 314, 330, 341, 352, 359, 368, 380, 390, 396, 406, 413, 677, 695, 1018, 1025, 1034, 1042, 1052, 1061, 1071, 1081, 1090, 1097, 1112, 1122, 1132, 1146, 1159, 1171, 1185, 1198, 1215, 1230, 1245, 1263, 1278, 711, 724, 741, 751, 767, 780, 795, 811, 826, 838, 848, 859, 881, 892, 908, 920]. Board [935, 946, 956, 964, 977, 990, 999, 1009, 1293, 1303, 1579, 1605, 1618, 1626, 1635, 1648, 1665, 1682, 1702, 1719, 1730, 1745, 1754, 1765, 1317, 1324, 1333, 1341, 1350, 1359, 1366, 1376, 1386, 1396, 1408, 1419, 1778, 1788, 1429, 1439, 1447, 1457, 1466, 1475, 1484, 1494, 1504, 1516, 1527, 1541, 1800, 1820, 2101, 2106, 2113, 2122, 2134, 2144, 2154, 2168, 2180, 2191, 2201, 2214, 2224, 2235, 2248, 2269, 1840, 1855, 1877, 1896, 1915, 1925, 1937, 1947, 1959, 1966, 1976, 1983, 1992, 2001, 2009, 2018, 2025, 2034]. Board [2044, 2057, 2066, 2075, 2084, 2093, 2279, 2289, 2311, 2323, 2336, 2346, 2360, 2369, 2381, 2391, 2403, 2416, 2426, 2435, 2444, 2453, 2462, 2470, 2480, 2490]. bonacci [911, 1497]. bondage [912]. bonds [1404]. booby [374]. book [849, 924, 1001, 2220, 487]. boolean [1513, 2152, 1482, 1486, 1064, 2016, 262, 1575, 627, 491, 1511, 2406, 1619, 278, 346, 2309, 1954, 258, 1157, 1188, 2166, 1639, 1790]. Boost [807]. Boosting [91]. bootstrap [1609]. bootstrapping [2125]. border [1662]. Borel [1984]. bottleneck [1779]. Bottom [967]. Bottom-up [967]. bound [105, 1190, 952, 1337, 537, 1150, 221, 2445, 344, 1004, 2158, 1290, 1310, 294]. boundaries [1402]. Bounded [409, 325, 2067, 765, 1853, 526, 279, 184, 1885, 1686, 739, 868, 1867, 2303, 190, 210, 2137, 2108, 272, 1013, 1948, 2258, 1003, 1284, 1012, 793, 771, 490, 166, 1768, 181, 282, 1100, 682, 1488, 384, 1162]. Bounded-collusion [2067]. Bounded-degree [765, 1948, 490]. bounded-genus [2303]. bounded-treewidth [1003]. Bounding [770, 162, 1711]. Bounds [781, 131, 16, 757, 524, 516, 1789, 1995, 1943, 912, 1076, 627, 715, 1056, 162, 2108, 1513, 2090, 164, 43, 445, 1969, 713, 166, 1697, 809, 213, 169, 1818, 1412]. box [9]. boxes [1086]. Boyer [435]. BP [2440, 820, 333]. BPMC [1470]. BPP [1298]. Braided [140]. Branching [266, 26, 248, 145, 664, 263, 703, 2327]. Branching-time [266, 248, 263]. branchings [2006]. breadth [757, 1461]. breakdown [2172]. Breaker [1903]. Breaking [2225, 679, 386]. bribery [1212, 678, 2250]. bridges [1247]. Broadcast [1591, 110, 2236, 1102, 2170, 1137, 65, 2423, 2062, 97, 175]. broadcasting [326, 2463]. broken [2388]. Bron [1318]. Brownian [2266]. Bruijn [730]. BSIF [71]. bubble [1286, 2155, 1032, 1986, 1831]. bubble-sort [1286, 2155]. bucket [1268]. Budget [2150, 951, 1709, 578]. Budget-feasible [2150]. Budgeted [812, 1833, 1721, 1838]. Buffer [1028, 1284]. build [1570]. Building [1247, 717]. built [2126]. burning [1722]. burnt [2351, 1478]. Burrows [1057, 235]. busy [1570]. BWT [865, 242, 243, 806]. byzantine [1885, 2236, 2011, 1102, 2446, 2188, 2170, 2029, 1154]. byzantine-tolerant [1885, 1102].

C [198]. caching [1928, 1824]. cacti [845]. cactus [2012, 1167]. cactus-based [1167]. calculate [1588]. Calculating [105]. calculation [1453]. Calculational [959]. calculi [257]. calculus [2156, 2211, 1294, 250, 145, 2064, 2441, 1624, 1192, 961, 2053, 650, 2181]. Calibration [369]. calibrations [418, 427, 1140, 1458]. Call [1500, 1572, 257]. call-by-name [257]. Calude [1999]. can [2417, 2165, 733, 754, 2164, 1400, 2374]. Canadian [1872]. canalization [2309]. candidate [1570, 1669]. candidates [1472, 1863, 2259]. cannot [2202]. canonical [1388, 2021]. Cantor [1776]. capacitated [1723]. Capacity [2215, 1662, 781, 2077, 2467]. caps [2387]. capture [1436, 537, 957, 1261]. capturing [1582]. car [1796]. car-sharing [1796]. Carathéodory [52]. carcinoma [1671]. Card [630, 46, 536]. Card-based [630, 46, 536]. cardinality [1305, 1890, 910, 2257, 1371, 904, 1537]. cards [1994, 1153]. carrier [567, 216]. carrier-based [567]. cartesian [1546, 85, 634, 681]. cascading [404]. case [1874, 1545, 127, 647, 1852, 1724, 334, 760, 2181]. cases [169]. catalytic [1524, 1539]. categorial [1069]. Categorical [1355, 2082, 1288, 1797]. categories [2082]. caterpillar [646]. caterpillars [1274]. causal [1102, 1064]. Causality [502, 2198]. cave [1148]. Cayley [174, 661, 897, 687]. CCA [1464, 654, 1088, 1304]. CCA-secure [654, 1304]. CCA-security [1088]. CCA2-secure [1105]. celebrating [906]. Cell [316, 2414, 599]. Cell-like [316, 599]. cells [2054]. Cellular [1609, 1630, 597, 318, 1175, 2366, 2344, 1844, 2003]. centdian [1599]. center [1381, 103, 1832, 773, 910, 579, 1969, 180, 1496, 2467]. centers [1287, 932]. central [683]. Centralised [1770, 434]. centrality [404]. certain [1727]. Certificate [1790, 839]. certificate-based [839]. certificateless [1476]. Certification [1874, 1172, 1414, 2158]. certified [862]. certifying [1789, 1179, 1988]. chain [1817, 163, 1861, 362]. chain-formation [1817]. chainable [1222]. chains [1758, 1450]. challenging [1672]. change [2344, 993, 2021]. change-making [2021]. changes [512]. changing [1614]. Channel [124, 65]. channels [781, 1519, 2109]. character [833, 2230]. characterisation [2111]. characterizability [1208]. Characterization [698, 375, 2021, 1212, 2087, 629, 2414, 547, 993, 1505]. Characterizations [404, 1442, 1813, 1248]. Characterizing [2097]. characters [1938]. charged [1929]. charger [557, 798]. chargers [1557, 1655]. check [2210]. Checked [2197]. checker [959]. Checking [1734, 531, 1448, 672, 503, 2117, 1907, 356, 125, 258, 1758, 1450]. Cheeger [1683]. chemical [11]. Chernoff [1412]. cherry [790]. cherry-picking [790]. chimera [2388]. chirality [2228]. choice [1440, 1362, 163, 1653, 1906]. choices [9]. choosability [609]. choose [223]. chordal [1169, 2331, 1160, 343, 2315, 1577, 797, 829]. choreographic [21]. Chosen [1370, 187]. Chosen-ciphertext [1370, 187]. Chromatic [16, 105, 1971, 792, 1806, 1867, 1187, 304, 548]. Chronic [1860]. Ciobanu [1628]. ciphertext [1370, 65, 292, 187]. ciphertexts [175]. circle [1660, 1693, 2029, 2409, 1479]. circuit [53]. circuits [1172, 1482, 378, 398, 1519, 361, 1724, 1238, 722]. circulant [338]. circular [1814, 1867, 698, 2337, 1179]. circular-arc [1867, 1179]. CISTs [787, 1050, 1120, 1130]. class [539, 1542, 458, 256, 2284]. Classes [2476, 287, 938, 1305, 1943, 99, 918, 1099, 80, 1155, 1727, 1027, 1470]. Classical [228, 1451, 1519, 1523, 1616, 1398]. classical-quantum [1451]. Classically [1238]. classification [1910, 1708, 2151, 1620, 1636]. classifier [337, 1860, 69]. clause [786, 1756]. claw [1383, 1711]. claw-free [1711]. CLBP [71]. Clearing [666]. client [1265]. client-server [1265]. clique [354, 1012, 173]. clique-width [354]. cliques [815, 1891, 2081, 1318]. CLL [1297]. clock [76]. clock-freeness [76]. clocks [254]. clonal [1544]. Clos [277]. close [1645, 1083, 2436]. Closed [1497, 1761, 1990, 1817, 2326, 944, 2184, 584]. closed-rich [2326]. Closest [871]. closure [2023, 969, 2175, 80, 2184]. Cloud [1912, 1552, 1858, 2449, 1353, 1769, 117, 1476, 839]. cloud-based [1858]. club [2135, 2327, 2050]. Cluster [2050, 1641, 2135, 2327, 1371]. clustered [1735]. Clustering [1677, 1322, 2482, 70, 1969, 1501, 2319, 2421]. clusterings [580]. CNF [786]. CNN [1623]. co [307, 2487, 778, 2419, 1351]. co-graphs [778]. co-modeling [1351]. co-phylogenetic [307]. co-secure [2419]. co-transcriptional [2487]. coalitional [1165]. coalitions [1035, 1824]. coalitions-based [1824]. code [1798, 585, 1926, 87, 1741, 472, 2399]. code-based [472, 2399]. code-correlation [1926]. Coded [776, 2401]. codes [1552, 1644, 739, 2013, 2389, 1388, 131, 1049, 1119, 1129, 1217, 1731, 2437, 476]. codewords [1405]. Coding [463, 1690]. coeffects [2156]. cographs [1248]. Coherent [1248]. coin [1460]. coins [2021]. Collaborative [207, 953, 1824, 777, 221]. collecting [391, 608, 1666, 1048, 1118, 1128]. Collection [1879, 746, 120]. collective [12, 1615]. collector [362]. collinear [1273]. collisionless [728]. Collisions [2465]. Colloquium [823]. Collusion [979, 2362, 2067]. Collusion-resistant [979]. Color [854, 1693, 189, 1961]. Color-spanning [854]. Colored [1767, 450, 649, 1281, 1940, 565, 1389, 1673]. Colorful [733]. coloring [475, 377, 1833, 1043, 1113, 1123, 516, 700, 805, 399, 1459, 1347, 1920, 218, 1961, 518, 1577, 771, 948, 1382, 1155, 1416, 1821, 1564, 797, 1727, 2130, 1077, 1410, 332, 1633]. colorings [95, 512, 1173, 1501]. colors [1143]. colourability [2428]. Colourful [1274]. Colouring [2219, 1022, 1875, 1711, 5, 591, 1039, 2445, 1027, 2219]. colourings [1592, 1103]. column [1939]. combination [1261]. combinations [284]. combinatorial [662, 2431, 131, 51, 1584, 614, 708, 1779, 578]. Combinatorics [1643, 232, 342, 1998]. combinators [1453]. combinatory [1621]. Combining [249, 2068, 492, 1351]. comma [1731]. comma-free [1731]. commerce [2218]. commodity [1518]. common [735, 526, 244, 2372, 1696, 764, 864, 2228, 1888, 1576, 2103, 296, 755, 1434, 102]. commonality [1944]. communicating [2163, 1629, 2344]. Communication [215, 2161, 823, 2226, 606, 2342, 2451, 514, 212, 1137, 852, 643, 2032, 1837, 155, 206]. communications [1784]. Community [560, 39, 658, 760]. Community-based [560]. commutative [739, 868]. commutativity [1193]. Comp [2379]. Compact [1867, 1836, 2071, 1388, 235, 393, 1220, 265]. Compacting [65]. compaction [1443]. compactness [462]. companions [1835]. comparative [817]. Comparing [1067, 947]. comparison [628, 2418, 2082, 147]. comparison-based [147]. comparisons [1362]. compartmental [2277]. compatibility [2413]. compatible [477, 347]. Competition [2079]. Competition-based [2079]. competitive [952, 940, 1796, 949]. competitiveness [2225, 1872]. compiler [2414, 1181]. complement [1315]. complementarily [2206]. Complete [556, 1524, 827, 2356, 2211, 2055, 2274, 2064, 1353, 2196, 2388, 1242, 87, 1422, 256, 701, 357, 2374, 1546, 2119, 511, 1963, 1942]. Completeness [1471, 2197, 2484, 673, 345, 148, 896, 383, 1522, 399]. completion [2077]. complex [623]. complexes [1259]. complexities [2387, 1616]. Complexity [1999, 1191, 931, 2212, 1661, 2185, 836, 2482, 184, 1339, 671, 2016, 912, 822, 2161, 194, 579, 1361, 1138, 1379, 1727, 2130, 1871, 1879, 1622, 914, 1586, 349, 1687, 1868, 817, 475, 2048, 1853, 613, 2136, 777, 249, 1507, 1971, 1545, 1642, 1806, 1742, 19, 1023, 1330, 867, 530, 1489, 1949, 1512, 1343, 2324, 1312, 2204, 1173, 1771, 1252, 1973, 1417, 1363, 2262, 2303, 1073, 1212, 2069, 1318, 1703, 2, 339, 1749, 1143, 649, 627, 514, 161, 1725, 212, 2463, 1394, 1977, 2239, 1780, 2401, 844, 309, 2090, 1460, 2052, 2176, 2241, 147, 610, 304, 54, 1216, 283, 936]. complexity [2314, 2305, 656, 2473, 2027, 2120, 1360, 1638, 80, 346, 793, 2355, 1382, 896, 1776, 2353, 1795, 497, 1768, 2124, 2340, 2404, 2464, 152, 64, 555, 1916, 1813, 2424, 2135, 1279, 104, 1325, 317, 228, 704, 1993, 783, 1933, 1241, 1724, 626, 1100, 1082, 1177, 1188, 1469, 202, 1848, 2293, 1189, 2299, 1550, 155, 1790, 2471, 1967, 215, 823, 2382, 206]. Component [1431, 1084, 188, 1996, 1850, 478, 502, 452, 485, 401, 1178, 1269, 2058, 1739, 1831, 1289]. component-based [1850, 502]. component-composition [485]. components [1274]. componentwise [2, 339]. Composable [12, 2487]. composite [164, 2045]. composition [1569, 485, 1367, 2320, 2373, 400]. compositional [866, 2064]. comprehensive [1092]. Compressed [229, 242, 736]. compressible [1282]. compression [1950, 414, 423, 92]. compressions [1373]. Comput [591, 2212, 340, 2094, 339, 133, 1525, 129]. Computability [1999, 1354]. Computable [1222, 1638, 1733, 1267, 202]. Computation [1057, 1980, 2186, 286, 1628, 735, 1910, 2014, 1372, 179, 2303, 1255, 235, 2183, 1612, 1187, 2414, 861, 2473, 2140, 961, 125, 2062, 238, 104, 438, 1622, 1220, 1842, 1779, 1530, 1716]. computation-tree [125]. Computational [1239, 2204, 1771, 1212, 1640, 718, 2314, 2464, 704, 783, 615, 280, 2429, 1522, 1927, 817, 264, 2484, 2387, 2241, 283, 2340, 1813, 1100, 2299, 650, 1143]. computationally [1524, 733]. computations [2487, 630, 1759, 1296]. compute [662, 454, 1346, 778]. Computer [312, 1935, 2488, 905, 151]. Computing [1236, 620, 310, 1086, 1696, 865, 1348, 2454, 664, 833, 2089, 195, 43, 63, 1326, 1724, 755, 2400, 753, 2205, 623, 2040, 2449, 523, 1720, 2007, 1637, 117, 1344, 323, 637, 877, 1525, 1830, 1994, 2177, 2147, 1153, 1156, 1610, 839, 342, 878]. concatenation [320]. concave [276, 2261]. Concentration [1985]. concept [2487, 500]. concepts [76, 1250, 623, 1934]. Concurrent [1136, 264, 2198, 499, 874]. condition [286, 2210, 1652, 1164, 1403]. Conditional [670, 485, 152, 687, 271, 1367, 2365, 1880, 1084, 690, 1306, 2227]. conditional-fault [2365]. conditions [1891, 954, 1093]. Conference [312, 2186, 742]. confidentiality [1334]. configuration [1452, 836, 1166]. Conflict [793, 100, 54, 2166, 1868]. conflict-free [54, 2166]. conflicts [2219, 1028, 1093]. Confluence [1091]. confusion [770]. CONGEST [2456, 1961]. Congested [173]. Congestion [2195, 130, 1384, 1311, 2095]. congruence [26, 2199, 2175, 2442]. conjecture [1077, 772, 1226, 1633]. conjectures [660, 199, 1564]. conjugacy [783]. conjunctive [1893, 896, 384]. connect [108]. Connected [1903, 812, 1881, 1581, 1670, 1328, 486, 2094, 1904, 1749, 1922, 89, 146, 1364, 1587, 275, 204, 420, 429, 1309, 1002, 1923, 1588, 1821, 884, 490, 520, 1138, 63, 1599, 1871, 704, 2036, 1594]. connected- [520]. connectedness [1565]. connection [1005, 1880]. connections [565]. connective [2040]. Connectivity [1569, 2217, 577, 1589, 1561, 1392, 1314, 1567, 2185, 1971, 2149, 1770, 478, 697, 452, 1490, 2237, 420, 429, 54, 1174, 611, 1178, 1269, 702, 403, 1431, 176, 2155, 2361, 402, 1985, 1546, 1988, 59, 1243, 1478, 657, 986, 1996, 1831, 1897, 2292, 1054, 1289, 2227]. connectivity-preserving [1314, 1770]. conquer [1328, 1485]. consecutive [1645, 947, 228]. Consensus [124, 2004, 2384, 408]. conservation [1287, 782]. conserving [2003]. considerations [937]. considering [1863, 1319]. consistency [1780]. Conspiracy [535]. constacyclic [2073, 2379]. Constant [1580, 654, 2062, 1562, 2306, 2221, 1361, 1012, 2158, 1837, 175]. constant-randomness [2221]. Constant-round [2062]. Constant-size [654, 2158, 175]. constant-space [2221]. constants [2464]. constituent [2263]. Constrained [2397, 858, 2460, 178, 1193, 1085, 1329, 207, 680, 953, 1265, 851, 678, 1890, 1832, 910, 2107, 1137, 1870, 298, 748]. constraint [442, 1990, 2440, 2063, 2257, 2396, 1082, 1186, 1781, 1574, 1496, 616, 904, 1646]. constraints [274, 1900, 2211, 276, 41, 2069, 185, 439, 2431, 1490, 186, 2341, 814, 1614, 1774, 1537, 578, 1371, 2405, 2467, 517, 1868]. construct [984, 1164, 1880]. constructed [1108]. Constructible [1698]. Constructing [1393, 787, 1050, 1120, 1130, 1010]. construction [170, 468, 2332, 2022, 883, 1161, 663, 736, 1157, 682]. Constructions [1511, 2406, 730, 366, 979, 1732, 453, 2318]. constructive [1658, 1561]. contact [1509, 2187]. contained [2028]. containing [698]. containment [790, 1845]. content [264]. Contention [1415, 2114, 1826]. Contention-related [1415, 2114]. Context [2367, 2131, 2273, 144, 2193, 464, 947, 2046, 1586, 201]. Context-Free [2367, 2131, 2273, 2193, 947, 1586]. contexts [876]. Contextual [2367, 869]. contextuality [1613]. contig [744]. contig-based [744]. continua [1222]. Continuity [960, 1445]. Continuous [839, 504, 319, 10, 1240, 725, 958, 728, 1454]. continuous-variable [319, 10, 1240]. contraction [2012]. contractions [1036, 1977, 1808]. contracts [1065, 2270, 24]. Control [1235, 1602, 1955, 1472, 217, 2414, 875, 2140, 1213, 1859, 2427, 335, 1571]. control-flow [2427]. controllability [2152]. Controlled [1629, 20]. convergecast [1995]. Convergence [324, 328, 442, 128]. Converting [2459]. Convex [2306, 441, 662, 276, 1676, 969, 2139, 164, 241, 53]. convex/concave [276]. convexity [52, 620, 2471, 1380, 2454]. convoluted [1708]. convolutional [1623]. Cooperating [532]. cooperative [385, 1852]. cop [2465, 655, 1651]. cop-win [1651]. cope [1843]. Coppersmith [433]. cops [1142, 1651, 1016, 537, 1163, 540, 1149, 808]. Copy [1189]. core [21, 2431, 2408, 1866]. corecursive [2197]. Coreness [385]. coresets [1194]. corners [1794]. corpora [491]. Correction [339, 855]. correctness [19]. correlated [200]. correlation [1926]. corresponding [628, 458]. Corrigendum [591, 2212, 340, 2094, 2379, 133, 1525, 129, 2488]. Cost [1543, 110, 665, 418, 427, 1338, 844, 814, 29, 369, 758, 1972]. cost-parity [665]. cost-reducing [1338]. costly [159]. costs [2226, 1783, 1597]. countable [203]. Counter [269, 868, 2325, 972]. counter-based [2325]. counterexamples [158]. counterparts [259]. Counting [501, 1601, 2337, 1296, 2116, 1471, 1173, 1254, 1218, 1536, 1797, 1747]. counts [947]. couplings [22]. coupon [362]. Coupons [31]. Cournot [1338]. covalent [1404]. covariance [443, 15, 294, 917]. covariance-Hessian [15]. Cover [35, 16, 1190, 2317, 360, 1364, 1499, 631, 2457, 204, 2107, 762, 1700, 2344, 903, 1666, 520, 1595, 2408, 1264, 1951, 2249, 440, 2049, 3, 129, 29, 302, 2374, 987, 581, 2061, 1978, 1838]. cover-free [762]. coverable [375]. coverage [55, 846, 310, 1147, 484, 416, 425, 758, 1663, 1162, 1864, 899]. coveredness [1725]. Covering [1585, 552, 303, 1026, 1378, 182, 978]. covering-interdiction [978]. coverings [1887]. covers [927, 2351, 2258, 1879]. CPA [1105]. CPA/CCA2 [1105]. CPA/CCA2-secure [1105]. CPUs [355]. crash [1415, 2114]. created [1620]. creation [1401]. Crick [2451]. Cristian [1999]. criteria [2261, 814]. criterion [2308]. Critical [1594, 1093, 2078, 1458, 1802, 1554, 911, 304, 924, 2307, 900]. cross [454]. cross-polytopes [454]. crossed [1567, 853, 2058, 402, 2146]. Crossing [1307, 566, 942]. crossings [1342]. crossover [320, 1843]. crosswords [2295]. crowdsensing [2150]. crowdsourced [915]. crowdsourcing [1816, 800]. CRS [2285]. crucial [2016]. Cryptanalysis [1180]. cryptocurrency [2382]. cryptographer [535]. Cryptographic [1853, 775]. cryptography [883, 1473]. cryptosystem [1593, 2437, 1108]. CS [2059]. CS-based [2059]. CSP [1049, 1119, 1129, 2189]. CTL [503]. CTMC [1094]. cube [1372, 1923, 471, 577, 853, 1589, 402, 1952, 2352, 1897, 1054]. cube-connected [1923]. cubes [1793, 1567, 787, 1667, 2149, 478, 988, 1174, 2058, 2049, 1931, 1403, 986, 327, 1996, 2146]. cubic [2206, 1491, 1084, 1108, 1101, 1180]. CUDA [1443]. current [551]. curvatures [2205]. curve [649, 433]. curves [1794, 1791]. Cut [1259, 1305, 2303, 275, 2139, 1712, 1768, 554, 1767, 102, 363, 738]. Cut-and-project [1259]. cut/paste [102]. cuts [1872, 1683]. Cutting [1413]. cyber [1714, 1351]. cyber-physical [1714, 1351]. Cycle [1838, 35, 570, 1567, 121, 2086, 924, 2118, 1951, 2049, 828, 584, 480]. cycle-structure [1567]. Cycle/Path/Tree [35]. cycles [1667, 489, 988, 2456, 1940, 1756, 365, 1953, 1923, 1865, 2230, 1954, 644, 2349]. Cyclic [61, 927, 1511, 791]. cylinder [1600].

D [2131, 995, 2367, 1731]. D0L [2451].

DAGs [1942, 2475]. dark [1148]. Data [1355, 1797, 120, 382, 1552, 1287, 1, 340, 1672, 2430, 939, 623, 70, 985, 455, 523, 1720, 1917, 773, 796, 518, 610, 2325, 1637, 180, 1653, 746, 1425, 1433, 1220, 284, 1260, 1856, 1620, 2298, 915, 1100]. database [2493, 501, 555]. databases [1330]. datapath [1928]. dates [1508]. datestamp [856]. datestamp-based [856]. DAWGs [2183, 1732]. Days [1133]. DCell [1565]. de-anonymization [494]. deadline [1538]. deadlock [1417]. deaima [1710]. December [592, 604, 621, 638, 652, 668, 1262, 1277]. decentralized [1816, 2067]. Decidability [1390, 1319]. decidable [1829, 671, 2016]. Deciding [1810, 2374]. decipherability [60]. Decision [1883, 974, 1471, 2211, 2238, 785, 278, 1402, 1671, 357, 509]. decomposing [587]. Decomposition [2485, 2068, 803, 1553, 30, 10, 2265, 1851, 2302, 278, 2003]. decompositions [1584, 686, 2141]. decreasing [1644, 2495, 595]. decrypt [1276]. decryption [177, 568, 682]. dedicated [1999, 1868]. deduction [2040]. Deep [1910, 1857, 1708, 1714, 1803, 1710, 154, 2429, 1058]. defectives [1014]. defend [1815]. Defensive [1659]. deferred [1519]. defined [1187, 971, 876, 2377]. Definitions [366, 1415, 2366, 2318]. degeneracy [538, 1369]. degenerate [2245, 234, 343, 1369]. degrading [2384]. Degree [1000, 1329, 1867, 210, 2137, 954, 1823, 1948, 346, 490, 2028, 1227, 1162, 765]. Degree-anonymization [1000]. degree-constrained [1329]. Delaunay [2147]. Delay [1870, 1318, 882, 2096, 852, 1284, 1632]. Delay-constrained [1870]. delays [2132]. delegation [143, 1610]. deleted [556, 487]. Deletion [2367, 459, 1417, 624, 1749, 1740, 1383, 115, 490, 2124, 2135, 2249, 1535, 1828, 2327, 2050]. deletions [2315]. Delimited [1095]. delivery [207, 1728, 953, 1650, 1137, 2410, 563, 488]. Delta [1656]. demands [1709]. demonstration [2487]. dendrite [1239]. deniability [1822]. denoising [1058]. dense [260, 1339]. Densest [688]. density [295]. density-based [295]. deorientations [2185]. dependable [2215]. dependencies [1064, 2140]. dependency [1330, 1619, 353, 2429]. dependent [2192, 1783, 774]. deployment [192]. depolarizing [1238]. Depth [1400, 1086, 1974, 1949, 361]. Depth-two [1400]. Deque [465]. Derandomization [1151]. Derivation [317, 1524, 944]. derivations [866, 2072]. derivative [2027]. derivatives [1266, 971]. derived [505, 2388]. descending [659]. describes [2064]. description [253, 2294]. descriptional [317]. Descriptive [1933, 202]. Design [444, 841, 1822, 1338, 959, 835, 573, 373, 24, 1816, 2073, 2379, 38, 2361]. designable [458]. Designing [2290, 756]. designs [2389, 37]. DESN [1058]. Destructiveness [320]. desubstitutable [804]. detecting [1858, 2456, 1714]. Detection [2372, 1910, 658, 1330, 1945, 1707, 1192, 1636, 760]. deterioration [2172]. determinant [626]. determinantal [2464]. determining [1786, 1806]. Determinisability [1313]. determinism [703]. Deterministic [2329, 582, 1186, 1639, 1872, 1014, 1995, 1771, 2060, 901, 91, 1440, 456, 1233, 873, 2486, 408, 302, 2330]. deterministically [1393]. development [1544]. devices [2088]. DFR [1105]. DFS [1772]. Diagnosability [2373, 400, 1470, 1531, 1922, 1652, 271, 420, 429, 485, 2275, 403, 1152, 714, 176, 1739, 690, 1403, 687, 2352, 1306, 1556, 2051, 1831, 2292, 1289]. diagnosable [1751, 1202]. diagnosis [2418, 271, 180, 1152, 2365, 1931, 1815, 691, 1202]. diagnostic [714]. diagonal [296]. diagonal-based [296]. diagram [278]. diagrams [1047, 1117, 1127, 2147, 574]. Dial [2069]. Dial-A-Ride [2069]. Dialectica [1956]. Diameter [512, 1705, 1686, 2303, 544, 1553, 895, 288, 1361, 1711, 1195, 1371]. diameter-optimally [1705, 1195]. diameters [1050, 1120, 1130, 1195]. diamond [815, 1875, 1383]. diamond-free [815]. dichotomous [2255, 2076]. Dichotomy [1968, 304]. dictionaries [756]. dictionary [636, 752]. difference [1236]. different [1251, 1143, 924, 517]. differentiable [1829]. Differential [1266, 1766, 13, 569, 1917, 325, 188, 294, 915]. Differentially [917, 1907, 2293]. Differentiating [1030]. Differentiating-total [1030]. difficulty [1104]. diffusion [962, 113, 280]. digest [2015]. Digital [1636, 662, 612]. digits [1807, 831]. digraph [1775, 778, 2292]. Digraphs [1005, 174, 1661, 1656, 1187, 2127]. dimension [174, 338, 819, 580, 1887, 89, 88, 96, 743, 109, 583, 1204, 94]. dimensional [230, 939, 70, 2164, 1757, 1685, 150, 973, 595, 974, 917]. Direct [843, 734]. directed [274, 2080, 2496, 324, 146, 778, 1010, 2124, 2129, 1791, 2006, 1481, 1310, 1311]. directions [2203]. disagreement [2010, 409]. disc [2045]. disc-ring [2045]. discipline [650]. disclosure [1848]. Disconnected [2026]. disconnection [1590]. discount [1792, 1072]. Discovering [1789]. discovery [1330, 1638, 334]. discrepancy [1697]. Discrete [906, 335, 1699, 1161, 519]. Discrete-time [335]. discretization [2276]. discretized [2205]. discrimination [1583]. disease [1860]. Disjoint [1309, 1311, 2300, 2395, 1667, 1143, 2351, 2492, 361, 2039, 1811, 1432, 1951, 375, 2049, 1551, 1561, 1301]. Disjunctive [1930, 912, 2115, 384]. disk [127, 647, 1920, 726, 1721, 625, 643, 490, 2036, 2187]. disks [1921]. disoriented [1479]. dispersable [849]. Dispersing [1144]. dispersion [1154]. Disposability [872]. disposable [889]. dissemination [1335]. dissolution [2111]. Distance [1210, 960, 1941, 1378, 1498, 2280, 1477, 678, 451, 1938, 1955, 1502, 1085, 2127, 63, 1326, 1902, 1677, 829, 2409, 102, 1633]. distance- [1378]. distance-1 [1955]. distance-hereditary [451, 63, 1326]. distances [521, 664]. distinctness [2486]. distinguishable [160]. distinguishing [1379]. distributability [1455]. Distributed [1886, 1378, 1987, 190, 1477, 2304, 189, 1553, 1920, 33, 1577, 2244, 2433, 293, 866, 2104, 2393, 684, 1147, 1955, 218, 1137, 117, 2053, 2474, 2427, 1994, 644]. Distribution [445, 2372, 1086, 721, 2165, 1823, 319, 10, 1240, 534, 1843]. distributions [11]. distributivity [140]. diversity [1489, 2316, 232, 832, 444]. diversity-aware [2316]. diversity-based [444]. Divide [1328, 26, 787, 577, 1485, 1952, 1897, 1054]. Divide-and-conquer [1328, 1485]. divide-and-swap [787, 577, 1952, 1897, 1054]. division [2411]. divisor [1434]. DLOGTIME [2111]. DLOGTIME-uniform [2111]. DNA [269, 1624, 1258, 236, 1260]. DNF [75, 601]. DNN [1654]. DNN-enabled [1654]. DNNs’ [1402]. Do [1981, 754, 81, 1560]. doctrines [1956]. document [237]. dollar [806]. domain [660, 199]. Domains [153, 960, 260, 200, 957, 265]. Dominating [1596, 1106, 2219, 867, 2392, 486, 190, 1036, 1977, 376, 1721, 345, 1360, 1601, 2320, 1695, 2028, 1495, 897, 2182]. Domination [272, 812, 2204, 183, 1477, 1808, 2457, 56, 2115, 2297, 2030, 625, 1491, 545, 2419, 64, 1038, 1030, 696, 1435, 1687, 2036, 299, 1818, 97]. domino [1249]. Doodle [386]. double [457, 1435, 1040]. double-strip [457]. doubling [1887]. Doubly [2128, 1285]. down [1413, 1639]. downlink [1709]. DR [2282, 2261, 2037, 553]. DR-submodular [2261, 2282, 2037, 553]. dragon [1794]. dragonfly [1561]. Dragons [2438]. drawing [1221, 149]. drawings [417, 426, 2306, 740, 185, 2196]. drift [2164]. driven [2059, 1502, 1810, 972, 2455, 2053]. drone [2410]. drones [1662, 2410, 1859]. DS [803, 30]. DSA [1805]. dual [787, 355, 1050, 1120, 1130, 1726, 2079, 2257, 948, 327]. dual-attribute [2079]. dual-CISTs [787, 1050, 1120, 1130]. dual-role-facility [1726]. duality [2071, 393, 1989]. Dualization [281]. duals [1509]. Duo [656]. duplication [498]. Dyck [1285]. Dynamic [138, 211, 518, 334, 993, 1632, 110, 381, 2393, 114, 355, 1680, 222, 916, 706, 1007, 718, 2271, 614, 2189, 440, 1179, 747, 2015, 2399, 37, 332, 1889]. Dynamical [597, 1399, 1733, 1421, 2277, 2477]. dynamically [1614]. dynamicity [1822]. dynamics [2102, 962, 1518, 1891, 376, 2173, 2321, 328, 1919, 86, 335].

e-commerce [2218]. E-voting [1623]. EA [598, 439, 2267, 1021, 440, 362]. Eager [1449]. easier [531]. easily [2347]. Easy [108, 305, 1471]. EC [1805]. EC-DSA [1805]. eccentricities [941]. Eccentricity [451, 1694, 1183]. ECDSA [2347]. ECTL [266]. ECTL# [266]. Edge [2050, 812, 538, 2405, 657, 2146, 417, 426, 1000, 2306, 1160, 1173, 1667, 2094, 1904, 478, 1654, 1740, 1036, 1977, 1808, 1940, 452, 910, 2297, 743, 1700, 4, 133, 1383, 1389, 1431, 1565, 2039, 2357, 2408, 1038, 1078, 853, 1880, 1379, 2130, 1077, 1546, 1988, 987, 1824, 474, 1199, 1866, 2272]. edge-colored [1940, 1389]. edge-connectivity [1546]. edge-cover [1700]. edge-disjoint [1667, 2039]. Edge-fault-tolerant [657]. Edge-independent [2146]. edge-length [4, 133]. edge-pair [1078]. edge-pancyclicity [1389]. Edge-Vertex [812, 2306]. edge-weighted [1866]. edges [1953, 2374, 2349]. edit [1502]. editing [777, 228, 2004]. Editorial [1530, 1716, 1133, 1715, 822, 8, 68, 315, 331, 437, 594, 712, 860, 965, 991, 1019, 1062, 1231, 1246, 1397, 1517, 1606, 1627, 1841, 2161, 589, 7, 18, 422, 431, 436, 448, 460, 470, 483, 496, 508, 528, 543, 562, 576, 593, 605, 622, 639, 653, 669, 28, 49, 67, 84, 134, 156, 168, 205, 214, 225, 246, 268, 290, 314, 330, 341, 352, 359, 368, 380, 390, 396, 406, 413, 677, 695, 1018, 1025, 1034, 1042, 1052, 1061, 1071, 1081, 1090, 1097, 1112, 1122, 1132, 1146, 1159, 1171, 1185, 1198, 1215]. Editorial [1230, 1245, 1263, 1278, 711, 724, 741, 751, 767, 780, 795, 811, 826, 838, 848, 859, 881, 892, 908, 920, 935, 946, 956, 964, 977, 990, 999, 1009, 1293, 1303, 1579, 1605, 1618, 1626, 1635, 1648, 1665, 1682, 1702, 1719, 1730, 1745, 1754, 1765, 1317, 1324, 1333, 1341, 1350, 1359, 1366, 1376, 1386, 1396, 1408, 1419, 1778, 1788, 1429, 1439, 1447, 1457, 1466, 1475, 1484, 1494, 1504, 1516, 1527, 1541, 1800, 1820, 2101, 2106, 2113, 2122, 2134, 2144, 2154, 2168, 2180, 2191, 2201, 2214, 2224, 2235]. Editorial [2248, 2269, 1840, 1855, 1877, 1896, 1915, 1925, 1937, 1947, 1959, 1966, 1976, 1983, 1992, 2001, 2009, 2018, 2025, 2034, 2044, 2057, 2066, 2075, 2084, 2093, 2279, 2289, 2311, 2323, 2336, 2346, 2360, 2369, 2381, 2391, 2403, 2416, 2426, 2435, 2444, 2453, 2462, 2470, 2480, 2490, 438]. Editors’ [157]. edits [833]. EF [601]. effect [1649, 1272]. Effectful [257, 961]. effective [1851, 1673, 1714, 1856]. effectively [1047, 1117, 1127, 1847]. effectiveness [1160]. effects [1252, 1849]. Efficiency [902, 1972, 1464, 522, 262, 521, 2385, 1557, 1344, 82, 2032, 1467, 2412]. Efficient [434, 343, 557, 890, 1388, 1978, 1187, 2140, 1219, 1675, 2271, 1655, 238, 1220, 284, 536, 509, 1779, 1496, 2399, 16, 1314, 1900, 2080, 2393, 1805, 2022, 36, 756, 2089, 1536, 1961, 796, 2302, 1962, 1576, 746, 347, 637, 213, 736, 192, 2328, 1732]. Efficiently [1156, 2281]. EFX [570, 886]. Ehrenfeucht [75]. eigenvalue [2272]. Eisenstein [2290]. either [2117]. Elastic [2291, 234]. elastic-degenerate [234]. election [1472, 217, 983, 128]. elections [1212, 1011]. electrical [836]. element [2486]. elementary [1406, 1630, 251, 1790]. elements [524, 2351, 943, 278, 2141]. Eliciting [2370]. elimination [570, 1707, 670]. eliminations [1960]. ellipsoid [443]. elliptic [433]. embeddable [2254, 190]. Embedded [986, 738]. Embedding [1372, 1667, 87, 20, 1221, 365, 149, 2388, 1078, 1600]. embeddings [849, 649, 1281, 1513]. Emergence [870, 595]. empirical [2293]. Emptiness [398]. emulators [2496]. enable [446]. enabled [1858, 1654, 1859]. Encoding [1064, 1297, 1654, 995, 2166]. encryption [293, 2491, 1464, 193, 2157, 178, 1370, 1668, 2449, 890, 1331, 654, 65, 1055, 1219, 856, 1304, 913, 2362, 1345, 1674, 1334, 292, 682, 453, 476, 2318, 2125, 1101, 175, 187, 979]. endomorphism [1448]. Energy [1688, 1557, 2305, 899, 1287, 207, 680, 953, 2303, 933, 851, 729, 627, 798, 1724, 1978]. energy-constrained [207, 680, 953, 851]. enhanced [1531, 2344, 2039, 2291]. enhancement [1857, 1637]. Enhancing [2357]. enough [1083]. ensemble [1910]. ensembles [69]. entities [1738]. entity [717]. Entropic [2483]. entropy [286, 1684, 692]. enumerating [501]. Enumeration [1566, 106, 1227, 343, 2096, 1012, 1580, 2141]. environment [1816, 488, 2032, 1859, 1871, 2494]. environments [117, 1188]. envy [570, 571, 893, 1697]. envy-freeness [571, 1697]. enzyme [1261]. episturmian [2354]. EPTAS [852, 2257]. Equal [301, 387]. equality [1370, 58, 1219, 856, 1674, 1153]. equation [1236, 1683, 1180]. equations [270, 1108]. equidistributed [62]. Equilibria [386, 1275, 522, 844, 2090, 2274, 1267, 632, 152, 2095]. equilibrium [1883, 1972]. Equitable [609, 771, 1382, 797, 981]. equitably [1369]. equity [902]. equivalence [287, 1984, 764, 558, 739, 456, 776, 336, 2241]. equivalences [2474, 703]. equivalency [336]. equivalent [909]. erasing [160]. erasure [498]. Erratum [339]. Erratum/Correction [339]. Error [295, 781, 1583, 2414]. Error-tolerant [295]. errors [1473]. Essays [1973]. estimate [1244, 1014]. estimates [2229]. estimation [2165, 2325, 1857, 1279, 294, 917, 1843]. estimation-of-distribution [2165, 1843]. Eternal [1016, 1190]. Eternity [2207]. Euclidean [1817, 182, 480]. Eulerian [1656, 1580]. Evacuating [1921]. evacuation [220, 127, 647, 729, 726, 1801, 643, 2029]. evader [1148]. evaders [666]. Evaluating [23, 70]. evaluation [994, 753, 2208, 1344, 2119, 2284]. event [1763, 2053]. event-driven [2053]. Every [149]. evil [529]. evolution [441, 442, 2060, 13, 870, 443, 2246, 15, 292]. evolutional [1401]. Evolutionary [1980, 1981, 1883, 446, 876, 438, 444, 1913, 1614, 1241, 69, 362]. Evolvability [1608]. Exact [1602, 526, 2331, 1630, 1727, 1879, 1874, 2014, 530, 1737, 1176, 420, 429, 1908, 1234, 1074]. examining [1825]. exchange [52]. exchanged [853, 1589, 691]. exchanges [397]. Excluding [700]. exclusion [1508]. exclusive [2228]. Execution [1842]. Exhaustive [1046, 1116, 1126]. exist [1560]. Existence [1099, 2385, 351]. expandability [197]. expanded [2, 339]. Expansion [1139, 1002]. expansions [831]. Expansive [596]. expansivity [318, 1944]. Experimental [753]. experiments [2082]. explainability [1969]. Explaining [1258, 161]. Exploration [706, 567, 952, 2228, 539, 2216, 221, 1440, 1676, 500, 2271, 1736, 1573]. Exploring [1680, 1499, 692]. exponent [1802, 1554, 1593]. Exponential [715, 1021, 1471, 279, 2165, 1596, 2358, 1218, 1216, 2012, 1105]. exponential-time [1218, 1216, 2012]. exponents [568]. exposure [2481, 177, 568, 682]. expressibility [266]. expression [807, 1900, 298]. expressions [1552, 259, 1825, 971]. expressive [263]. Expressiveness [499, 1760]. Expressivity [2197]. Extended [568, 2087, 2030, 873, 69, 273]. Extending [266]. Extension [1364, 1363, 1283]. Extensional [2232, 1442]. Extensions [614, 2476, 1805]. external [865, 869, 2264]. external-lights [2264]. Extra [1831, 1531, 2149, 1652, 452, 420, 429, 1269, 1367, 176, 1546, 1403]. extractable [2285]. extraction [491, 1708, 1636]. extractor [1927]. extractors [285]. Extremal [1491, 1141]. extreme [1207]. Extricating [1710].

FA [1950]. face [1623]. facilities [1391, 813]. Facility [2260, 667, 116, 1726, 659, 1391, 1700, 2076, 893, 1723, 615, 747, 2409, 513, 1884]. Factor [2472, 244, 2282, 2481, 2493, 2290]. Factor-balanced [2472]. Factorisation [2477, 2485]. factorisations [583]. factorization [864, 1497]. factorizations [171, 467, 515]. factors [646, 731, 911, 889]. fails [2202, 2064]. failure [1513, 200, 1856]. failures [1415, 2114, 2095]. Fair [2411, 1742, 2316, 2089, 1460, 936, 2339, 2467]. fairest [1165, 351]. fairly [2281]. fairness [266, 1165, 1852, 1035, 2339]. fairness-based [1165]. fake [1708, 1636]. families [1660, 1569, 1163, 2111, 762, 2428, 2086, 2450, 400, 584, 2468]. family [1923, 1291]. Fan [942, 566]. Fan-crossing [942, 566]. fascinating [861]. Fast [208, 679, 231, 432, 643, 236, 681, 2347, 1759, 1369, 1043, 1113, 1123, 2417, 1945, 553, 2405, 1029]. Faster [227, 414, 423, 580, 1218, 1463, 1535, 1828, 2004, 535, 2125]. fate [997]. Fault [1392, 1498, 338, 124, 365, 714, 2155, 2058, 701, 474, 1809, 2045, 1549, 2170, 502, 762, 1491, 1174, 2029, 1367, 635, 2357, 2365, 492, 853, 2169, 1084, 1952, 657, 1815, 1168, 1986, 1427, 1202]. Fault-free [365, 1809]. Fault-Prone [124, 1549]. fault-tolerance [1174, 1952]. Fault-tolerant [1392, 1498, 338, 2155, 701, 2170, 762, 1491, 1168]. faults [661, 474]. faulty [2351, 1953, 401, 2039, 1809, 2203]. Feasibility [2277]. feasible [2150]. feature [1910, 159, 1442, 1707, 1708, 1636]. features [72, 1849]. February [83, 167, 766, 779, 794, 810, 1349, 1358, 1365, 1924, 1936, 1946, 1958, 2345, 2310, 2322, 2335]. Feedback [2251, 2209, 2331, 781, 2108, 1013, 2265, 936, 510, 166, 282, 2005, 693]. feedbacks [1506]. Few [1811, 1563, 2089, 844, 1585, 2315, 372]. Fibonacci [1644, 196, 2445, 988]. fidelity [1758]. fidelity-based [1758]. field [273]. file [1476]. Filling [2295, 1004, 2244, 744]. Filters [876, 1932]. filtration [1857]. Financial [2099]. find [108, 2165, 666]. Finding [2392, 2159, 895, 918, 1683, 1669, 1928, 2253, 2006, 634, 2349, 707, 1510, 1881, 531, 1971, 1742, 1143, 349, 644, 533]. Fine [1417, 2205]. Fine-grained [1417, 2205]. Finite [1514, 2326, 1684, 392, 868, 597, 660, 199, 211, 456, 968, 873, 970, 321, 2142, 2120, 1257, 2424, 125, 2477, 791, 1622, 2184, 1100, 2459]. Finite-image [1514]. finite-index [868]. finite-state [2142]. finite-variable [1622]. Finitely [160, 619]. firebreak [1462]. firewalls [775]. firing [1255]. First [505, 2432, 1069]. First-order [505, 2432]. fitness [72, 1843]. Fitting [939]. five [524]. Fixed [2334, 640, 276, 738, 2063, 2005, 1134, 102, 2050, 1381, 1990, 953, 2, 339, 918, 2410, 1001, 252, 228, 720, 1657]. fixed-order [1001]. Fixed-parameter [640, 276, 738, 2063, 1134, 102, 2050, 1990, 918, 228]. flat [317]. Flexible [2242, 1674]. flight [746]. flimsy [1640]. Flip [1741, 2004]. Flip-swap [1741]. flipping [1460]. flooding [1858]. floor [1882]. floor-plans [1882]. floorplan [984]. Flow [355, 2341, 1100, 1518, 1703, 1007, 45, 1085, 1928, 347, 2427]. flowshops [1538, 1574, 119, 350]. Fluid [326, 764]. FM [1057]. FM-index [1057]. Focused [1990]. focusing [69]. foes [2060]. folded [787, 2217, 2492, 365, 1152, 2058, 1084, 2146, 2045]. folding [2487]. food [1650]. forbidden [2300, 572, 540]. Forbidding [1472, 2484]. forecasting [1856]. forest [544]. forests [1566, 1809]. fork [1440, 852]. fork-join [852]. form [114, 262]. Formal [1451, 1443, 60, 2313, 1331, 261, 2450, 692]. Formalising [807]. formalisms [254, 558, 1586]. formalized [922]. formation [1748, 306, 685, 1817, 1878, 2458, 1479, 758]. forms [2002, 1151, 361]. formula [786, 2064, 75]. formulas [2476, 1619, 1291, 1189]. formulation [1845, 335]. formulations [887]. forward [1690, 1233]. foundation [1658, 2413]. Foundations [1980]. founder [2291]. four [2448, 2374]. four-vertex [2448]. Fourier [286]. FPQ [1015]. FPQ-ordering [1015]. FPT [2397, 2395, 1598, 1832, 2012, 1373]. FPT-algorithm [1598]. FPTAS [454, 1679]. Fractional [2258, 645, 2350, 719, 1242, 616]. Fragile [1512]. fragments [671, 2016, 1622]. Fraïssé [75]. framework [244, 1020, 1672, 1335, 756, 1816, 1506, 92, 472, 1848, 2088]. Fraud [333]. Free [2367, 1964, 1558, 633, 2131, 651, 1023, 2071, 942, 1044, 1114, 1124, 2392, 1686, 1348, 1862, 900, 2078, 1417, 2273, 815, 2002, 1459, 1922, 1388, 1740, 872, 763, 762, 2193, 54, 1309, 1103, 365, 510, 947, 661, 1383, 1768, 1711, 1811, 2020, 791, 1731, 601, 1485, 1141, 1586, 2166, 1687, 1809, 1199, 201, 599]. freeness [76, 1486, 571, 1697]. frequency [2325, 2349]. friends [2060, 223, 2375]. frontier [2032]. Frontiers [1999, 1529]. frugal [231]. Fu [1935]. Full [1822, 1109, 801, 909, 568]. Fully [381, 564, 2423, 2362, 293, 288, 1209, 2399, 2125]. FUN [589]. function [781, 451, 745, 2414, 2035, 1706, 553, 2246, 1153, 1204, 1781, 453, 362, 904, 1691]. function-hiding [453]. Functional [353, 1101, 1464, 716, 178, 2485, 2414, 1304, 2362, 453]. functions [441, 2011, 2482, 1555, 1829, 1943, 385, 1346, 262, 1575, 178, 831, 627, 1511, 2406, 1449, 1968, 162, 439, 1612, 2256, 2358, 2164, 1846, 2267, 725, 278, 1021, 346, 1088, 1755, 820, 1926, 1327, 241, 1913, 2309, 2396, 1157, 1505, 1790]. Fundamental [873]. Further [660, 1740]. fusion [1636]. future [1339]. fuzziness [1022]. fuzzy [1544, 70, 1860, 1708, 1110, 1671, 2494, 285, 2072, 1927]. fuzzy-based [1860]. fuzzy-clonal [1544].

Gabidulin [2437]. Gabriel

[1628, 2196, 727]. Gacs [1678, 1957, 1973]. gadgets [2136]. gain [441]. Game [117, 1906, 5, 591, 1079, 1582, 1903, 1016, 2243, 537, 522, 486, 1165, 191, 1598, 503, 1045, 1115, 1125, 2398, 30, 297, 770, 2076, 374, 2118, 1416, 1821, 983, 383, 1803, 519, 2371, 655, 154]. game-based [503, 154]. game-theoretic [1803]. games [2136, 2471, 1275, 1549, 138, 130, 1384, 2195, 1827, 1898, 665, 2280, 2260, 385, 667, 1726, 1338, 399, 532, 2192, 664, 1107, 844, 309, 2090, 2274, 1852, 768, 1053, 2160, 1035, 2473, 1203, 893, 2408, 2220, 1767, 2095, 75, 601, 1149, 2371, 774, 2370, 2409, 581, 1972, 2343]. gammoids [348]. gap [941, 1499, 636, 752]. gapped [238]. gaps [2245, 1320]. garbage [1091]. Gathering [311, 1817, 683, 493, 222, 1873, 680, 1804, 728, 727, 816, 929, 388]. Gaussian [1709]. gem [2078]. gene [1064, 1232]. General [36, 1149, 441, 1548, 2312, 2446, 587, 221, 1584, 968, 1176, 1721, 500, 302, 2371, 899]. Generalised [1039, 957, 2231]. Generalising [1592]. generalization [1940, 2095, 2019]. Generalized [171, 503, 1421, 1506, 2383, 689, 1621, 184, 1384, 1253, 2079, 573, 2484, 271, 930, 702, 1923, 1894, 1501, 1776, 2353, 1711, 853, 2005, 1869, 1243, 1616, 336, 1897, 467]. Generalizing [113]. generally [1913]. generate [996, 2230]. generated [1392, 619, 661, 1772, 687, 2378]. Generating [786, 585, 1570, 2177, 718, 1673, 1522]. Generation [2254, 791, 1046, 1116, 1126, 1328, 291, 432, 1673, 1075, 2327]. Generative [2233]. generator [862]. generators [433]. Generic [1161, 1453, 364, 549, 1050, 1120, 1130, 623, 745, 663, 1861, 283]. generically [2285]. Genericity [1464]. genetic [320, 333]. genome [1250]. genomic [744]. genus [2496, 2303, 190]. geodetic [52, 2262, 1357, 763]. Geometric [1273, 2332, 1670, 2080, 2313, 2159, 2447, 2226]. geometrical [1255, 1402]. Geometry [1398]. Gerrymandering [2174, 950]. GF [2450]. Gheorghe [861]. GHS [415, 424]. GI [843]. GI-hard [843]. girth [1378, 1711]. given [1509, 281, 1056, 195, 351, 1510]. Global [143, 124, 696, 110, 994, 951, 1311, 388]. gluing [1294]. GMMS [570]. Gödel [1956]. Golden [1203]. good [1775, 403, 983, 2058, 1913, 59, 1556, 2292]. good-neighbor [403]. Gossiping [851]. GPS [2203]. grab [2398]. graded [125]. Gradient [2293, 336]. grading [2313]. graduates [1294]. grained [1417, 2205]. Grammars [2367, 2131, 463, 868, 869, 2484, 2193, 875, 2046, 1586, 603]. Graph [2471, 607, 1717, 918, 842, 1501, 16, 817, 1581, 1329, 1833, 1436, 628, 952, 967, 764, 938, 2496, 1305, 539, 1990, 1091, 1417, 1328, 2432, 1987, 404, 2430, 939, 343, 20, 399, 1598, 1462, 941, 89, 1286, 99, 2126, 1271, 146, 2063, 2265, 951, 1221, 1721, 2428, 768, 43, 1010, 345, 727, 1155, 1416, 1242, 2252, 554, 1695, 63, 1326, 2028, 240, 2244, 465, 415, 424, 1985, 1092, 1546, 2119, 351, 584, 2045, 824]. graph-based [768, 2119]. graph-like [1092]. graph-theoretical [404]. graph-walking [2252]. graphical [1020, 1351]. Graphs [1901, 2202, 1660, 450, 1713, 95, 1964, 1378, 52, 1248, 417, 426, 1773, 1169, 2014, 2331, 363, 1971, 1206, 1498, 139, 1548, 338, 279, 658, 2306, 1592, 1786, 1806, 184, 2312, 1670, 2080, 967, 300, 2251, 1305, 566, 942, 1044, 1114, 1124, 2392, 539, 1686, 1348, 805, 646, 1173, 843, 900, 2078, 2407, 1867, 2159, 2447, 103, 1590, 815, 112, 1307, 1274, 624, 683, 1878, 2194, 489, 190, 1477, 738, 1459, 354, 1163, 2205, 1039, 2148, 1922, 2137, 2485, 451, 609, 1045, 1115, 1125, 2398, 1941, 2351, 1357, 1377, 1414, 1920, 88, 2394, 2463, 58, 1659]. graphs [2386, 272, 1940, 1875, 1948, 1536, 629, 1547, 449, 778, 218, 33, 1003, 452, 2465, 641, 2206, 194, 56, 2115, 2297, 763, 565, 420, 429, 2030, 2277, 950, 1698, 304, 625, 2315, 852, 1491, 545, 515, 1309, 1577, 1012, 2419, 1360, 4, 133, 510, 485, 2196, 556, 401, 661, 1178, 771, 1389, 1689, 1382, 1679, 1601, 884, 490, 2388, 1768, 2129, 1711, 1811, 540, 1259, 1564, 1791, 2020, 1865, 2155, 2006, 1599, 797, 1880, 1106, 1379, 1899, 282, 375, 1830, 1435, 1495, 1727, 2130, 1027, 2375, 2147, 2291, 1902, 829, 1485, 1481, 1993, 914, 828, 1077, 1603, 1179, 53, 1310, 693, 126, 1177]. graphs [1687, 644, 59, 487, 655, 1195, 897, 2036, 687, 1228, 657, 1470, 885, 106, 2299, 1224, 1543, 97, 1369, 1866, 1162, 1410, 332, 1202, 1633, 2187, 2227, 674]. gray [2230, 1798, 1644, 585, 1741]. greed [688]. greedily [2417, 2377]. greedily-defined [2377]. greedy [2350, 149, 1821, 2182, 2328]. Green [846]. Grid [311, 1736, 1559, 1314, 1372, 2306, 306, 2228, 410]. grids [1500, 2262, 2458, 1513, 376, 519, 2290, 2003, 808].

Groot [393]. ground [915]. grounded [2272]. Group [484, 1014, 295, 161, 1801, 1208, 571, 1209, 2399, 1506]. group-characterizability [1208]. Grouped [2457]. grouping [1144]. groups [174, 1249, 1558, 197, 597, 1357, 2138, 1487, 1697, 618, 1761, 791]. Grover [2433]. growth [2332, 376, 1823]. guarantee [801]. guaranteed [484]. guarantees [154, 1597]. guarded [1059]. guarding [1076]. guards [1076, 258]. Guest [2161, 157]. guided [1547]. Guwahati [906].

H [1298]. H-team [1298]. Hadamard [1511]. half [2304, 2208, 2184]. half-integral [2304]. half-space [2184]. halfplane [1076]. Hamiltonian [1392, 1886, 1953, 1830, 1481, 53, 644, 2349, 1168, 1809]. Hamming [1734, 1085]. Hamming-isometric [1734]. hand [77, 2163]. handed [148]. handling [442, 1990]. handshakes [1822]. Hanoi [2312, 1377]. happy [475, 931, 2212, 516, 2407]. Hard [305, 533, 454, 93, 108, 2159, 301, 373, 733, 2295, 2438, 1472, 2164, 2250, 618, 843]. hardcore [1679]. hardest [2046]. Hardness [479, 1939, 521, 565, 2439, 2483, 38, 797, 1543, 16, 1853, 982, 1786, 1598, 978, 272, 1948, 2297, 545, 1685, 1416, 42, 1722, 1030, 696, 1435, 2371, 894, 364, 2397, 2350]. Hardware [1276]. hash [50]. Hausdorff [1222, 1204]. having [504]. head [1641]. heaps [356]. heat [1683]. heavy [1900]. hedonic [768, 1035]. hell [531]. Helly [941, 1941]. Helly-gap [941]. help [1984]. helps [434, 1311]. her [1247]. hereditary [938, 451, 2428, 2127, 63, 1326, 1027]. heredity [114]. Hertli [90]. hesitant [1671]. Hessian [15]. heterogeneous [1287, 2156, 1728, 623, 832]. Heuristic [1688]. heuristics [715]. hexagonal [2290]. hidden [232, 2283, 1533, 1550]. hiding [2189, 1141, 453]. Hierarchical [665, 2141, 293, 2254, 1252, 45, 177, 1055, 200, 1084]. hierarchy [2116, 387, 785, 2193, 1327, 408]. high [2102, 1378, 580, 70, 1638, 1430, 1291, 917]. high-dimensional [70]. high-girth [1378]. high-level-programming [1638]. Higher [1934, 2192, 144, 1757, 1157]. higher-dimensional [1757]. Higher-order [1934, 2192, 144]. highly [1881, 689, 2246]. HIoTNs [1861]. hit [3, 129]. hitching [1675]. Hitting [279, 1026, 2397, 624, 2420, 572]. hoc [1335, 216, 1571]. Holant [1486, 2206, 2299]. hole [1459]. hole-twin [1459]. holographic [1646]. holomorphic [2482]. home [2203]. homogeneous [953, 933, 2409, 1631]. homogenization [2287]. homomorphic [2157, 1208, 1181, 2125]. Homomorphisms [2252, 2129]. homotopy [2413]. Honor [822, 1935]. Honoring [1247, 861]. honour [1973]. hop [1532, 1955, 654, 1241]. Horn [1555, 616, 1189]. Horspool [435]. Hospitals [2387]. Hospitals/Residents [2387]. Hotelling [1549]. HPMC [1368]. hub [103, 1694, 894, 1543]. Huffman [737]. hull [662, 620, 1380, 2454]. hunt [1547]. Hybrid [1253, 2263, 1168, 1211, 2342, 1280, 263, 1641, 1637, 1015, 2421, 2494, 476, 1470, 691, 2051]. Hyper [1986, 1168]. hyper-Hamiltonian [1168]. hyperbolic [1267]. hypercube [2208, 2217, 2492, 1953, 1152, 1431, 2039, 1894, 2357, 1432, 1951, 2169, 1600, 690, 474, 691, 2045]. hypercube-like [2217, 1431, 1432, 1951]. hypercubes [1531, 2420, 365, 1269, 702, 2275, 635, 1739, 1168, 1809]. hyperelliptic [300]. hypergraph [113, 291, 1703, 631, 1468, 374]. hypergraphs [1347, 954, 2258, 1584, 1683, 1103]. hypermutations [446]. hyperspaces [1909]. hyperspectral [2059]. hypotheses [2039]. hypothesis [1216, 151].

I/O [1288, 2363, 796]. I/O-efficient [796].

I/O-optimal [1288]. IBE [177]. id [219]. ideals [2072]. identical [2372, 563]. identifiable [1805]. Identification [333, 163, 1437, 1202]. identifiers [388]. identifying [2460]. identities [2362]. Identity [1331, 856, 1334, 187, 1743, 979, 1668, 2449, 890, 663, 1055, 2423, 1345, 292, 682, 175, 50]. Identity-based [1331, 856, 1334, 187, 1743, 979, 1668, 2449, 890, 663, 1055, 2423, 1345, 292, 682, 175, 50]. idle [117]. II [279, 671, 1266]. III [506, 2016]. IITG [906]. Image [71, 2059, 1514, 1636, 1058]. images [196]. Imbalance [1264]. immune [446]. immunity [1251, 1157]. Immunization [827]. Impact [2069, 322, 2431, 2160, 727]. impartial [2471]. imperative [2181]. Imperfect [996]. implementable [1137]. Implementation [966, 1355, 1101]. Implementations [1812, 411, 46]. implementing [807]. implication [500]. implicational [281, 2141]. implications [1507, 1801, 569]. implicit [249, 216]. important [1368]. Impossibilities [769]. impossibility [1415]. impossible [533, 532]. imprecise [1693, 288, 180]. improper [609]. improve [1641, 1847]. Improved [524, 792, 559, 1938, 2358, 1948, 218, 2265, 1176, 147, 812, 1383, 713, 1706, 1290, 814, 2330, 1573, 894, 119, 299, 333, 1048, 1118, 1128, 1884, 2488, 2348, 2050, 1087, 122, 952, 1817, 1150, 1056, 344, 1031, 1707, 2291, 1902, 1671, 1551, 69, 1006, 2412, 1697]. Improvement [568, 90]. Improvements [1344, 155]. Improving [1593]. in-degree [210]. In-place [1480]. inapproximability [1138]. incentive [347]. incentives [231]. Incidence [1043, 1113, 1123]. including [336]. inclusion [1087]. inclusive [1922, 1556]. incomplete [1275]. incompleteness [660, 199]. inconsistent [1762, 601]. increasing [1696, 296, 755]. incremental [1160, 2085, 241]. increments [1266]. indefinitely [804]. indefiniteness [862]. Independent [897, 642, 1044, 1114, 1124, 1784, 963, 1812, 1218, 641, 2081, 1577, 1012, 1360, 1782, 686, 1880, 282, 693, 2146]. indeterminate [761]. index [233, 868, 1057, 1291, 1830]. indexed [2202]. indexing [1645, 2183, 796, 2290]. indicating [325]. indications [879]. indices [1806, 242]. indistinguishability [654]. Individual [1405, 798, 1583, 1035]. indivisible [1738]. induced [2392, 1990, 343, 1380, 58, 1875, 123, 1811, 540, 1106, 992, 648]. Inductive [1, 340, 2366]. Inductive-data-type [1, 340]. indulgent [1804, 2384]. industrial [1426]. inefficiency [1972]. inequalities [782]. Inequity [982]. inference [394, 283, 1888]. Inferring [433, 915, 240]. infinite [2329, 306, 171, 467, 834, 1878, 835, 1934, 911, 458, 2326]. inflationary [1448]. Influence [732, 1115, 1872, 404, 209, 1045, 371, 1534, 802, 1657, 2296, 1125]. influence-based [404]. Information [215, 1335, 2161, 823, 416, 425, 742, 1848, 1454, 1275, 2411, 12, 1398, 2178, 92, 1929, 832, 1994, 1677, 957, 782, 206]. information-theoretic [92]. informational [1250]. inhibitorless [2334]. inhibitors [599]. inhibitory [1522]. initial [2276, 1093, 1325]. initializable [1480]. initialization [254]. Injective [805, 2130]. inner [2491, 1464, 2449, 1304, 453]. innovation [1338]. Input [1502, 972, 2136, 1810, 2455, 2340, 453]. Input-driven [1502, 972, 1810, 2455]. input/output [2136, 2340]. Insertion [2367, 353, 877, 1525, 603]. inspired [1518]. instability [70]. Installation [479, 1597]. instance [663]. instances [305, 2296, 1319]. instantiations [1331]. instruction [19]. Insurance [333]. Integer [2422, 398, 276, 2207, 2282, 196, 2183, 788, 553, 2377, 1781, 102]. integers [2290]. integral [2304]. integrating [2218]. integration [1636]. integrity [716, 1499]. intellectual [1552]. intelligence [1654]. intensionality [499]. Interaction [371, 1299, 2053]. Interaction-aware [371]. interactions [2321]. interactive [866, 775, 1049, 1119, 1129, 1646]. interchanges [2047, 1207]. interconnection [1269, 176, 2422, 1556, 2051]. interdiction [978]. interference [1182, 55, 1955, 1095]. interior [889]. intermediate [264]. Internal [1562, 143]. International [1529, 823, 2186, 1306, 742]. Internet [1860, 1859]. interpolation [505]. interpretation [959, 1615]. intersecting [587, 489, 1347, 1077]. intersection [1660, 2229, 454, 1423, 1670, 1336, 2447, 587, 2258, 922, 1926, 181, 2412, 650]. interval [504, 1343, 1305, 1867, 295, 191, 2087, 2454, 218, 2081, 518, 132, 1689, 1673, 1321, 1430, 104, 375, 828, 885, 106]. intervals [2397, 191, 770]. Intrinsic [2468]. Introducing [961]. Introduction [157]. intrusion [1707, 1856]. invariance [442]. invariant [1275, 392, 1656, 971, 165, 87]. invariants [161]. inverse [864, 910, 1085, 2252, 2494, 615]. inversion [57, 364, 574]. invertible [1267, 1844]. investigating [1064, 2199]. Investigation [1623]. invisibility [1608]. invisible [699]. involving [2185, 1829, 671, 618]. IoT [1672, 1688, 1641, 2474]. IoT-Wireless [1688]. irreducible [2141]. irredundant [1566]. irregularity [1786]. islands [101]. isometric [1793, 1734, 772]. isomorphism [354, 842]. isomorphisms [2232]. isoperimetry [1602]. ISP [2085]. Issue [966, 1980, 1063, 742, 1628, 824, 1530, 1716, 2400, 342, 1935, 215, 1717, 157, 1998, 2186, 1529, 438]. issues [1339]. Italian [312]. Italy [905]. item [284]. items [2281]. Iterated [2142, 371, 993]. iteration [1059].

January [6, 17, 27, 48, 66, 676, 694, 710, 723, 750, 1292, 1302, 1316, 1323, 1332, 1340, 1914, 1819, 1854, 1876, 1895, 2278, 2288, 1799, 1839]. Java [2156]. Java-like [2156]. Jérôme [1528]. job [181]. job-intersection [181]. jobs [41, 563, 1791, 1201]. join [1448, 1440, 852]. join-semilattices [1448]. Joint [1575, 610, 1824]. Jonoska [1247]. Jordan [1851]. JSON [2023]. jubilee [906]. July [389, 395, 405, 1017, 1024, 1033, 1111, 1121, 1604, 2100, 2105, 2112, 1041]. Jump [2424, 1990]. jump-and-repair [1990]. jumping [901, 2314]. June [358, 367, 379, 976, 989, 998, 1008, 1503, 1515, 1526, 1540, 1578, 2065, 2074, 2083, 2092, 2469, 2479, 2489]. just [1794]. Justifying [2138].

kangaroo [1308]. kD [1136]. Kemeny [1362]. Kempe [512]. Ker [822]. Ker-I [822]. Kerbosch [1318]. kernel [459, 344, 2110, 948, 1383, 1420, 1551, 2061]. kernelization [555]. kernels [310, 572, 126, 299, 648, 1199]. key [1555, 1370, 319, 10, 1240, 177, 1219, 1674, 1181, 1861, 568, 682, 2125, 1424]. keys [2495, 175]. KGC [1331]. kind [100]. Kindergarden [1294]. kinds [565, 1243]. kinetic [932]. Kingdomino [383]. Kleinberg [410]. Klondike [1191]. knapsack [1614, 2396, 1781, 1774, 1537, 2328]. Kneser [515]. knight [1342]. kNN [337]. kNN-P [337]. knot [1417]. knot-free [1417]. knowledge [1482, 2128, 1762, 33, 1164, 1272, 536, 388, 334, 1997]. known [1871]. Ko [822]. Kolmogorov [1853, 1023, 2098, 1848]. Kripke [2460]. Kronecker [2437]. ksmt [2211]. Kucera [1678, 1957]. Kurodoko [2171].

L [1882]. L-shaped [1882]. label

[2372, 240]. labeled [1235, 2364]. labeling [1498, 1995, 224, 850, 1591]. labelings [2126]. labellings [1206]. labels [1694, 895, 971]. laceability [474, 1168]. ladder [830]. Lagrangian [442]. lambda [23, 257]. Land [2344]. landmark [1680]. landscape [1742, 72]. language [1910, 558, 196, 1451, 491, 2007, 2124, 252, 2046]. Languages [2378, 203, 259, 160, 60, 2496, 2313, 671, 2016, 868, 1967, 2087, 2401, 1521, 1587, 871, 969, 2139, 2176, 718, 2305, 2142, 2120, 947, 973, 2450, 877, 1525, 317, 2472, 465, 1741, 974, 692, 1520, 1441, 1235, 1522]. Laplacian [113, 1223, 2272]. Lara [1760]. large [1901, 411, 1751, 557, 1637, 924, 1711, 1234, 1260]. large-scale [557, 1234, 1260]. larger [1021]. Largest [1903, 818, 401]. laser [1929]. laser-charged [1929]. latent [40]. Lattice [2071, 979, 177, 1674, 682, 1822, 1046, 1116, 1126, 2282, 1370, 500, 1213, 553, 1209, 1781, 1434, 1473, 2088]. Lattice-based [979, 177, 1674, 682, 1822, 1370, 1473, 2088]. Lattice-free [2071]. lattices [1743, 281, 1668, 1219, 913, 1345, 1454]. LAW [1133, 1166]. laws [2145, 1453]. Lawson [265]. layouts [1773]. Lazy [2162]. LCP [382, 2286, 865, 71]. LCS [1433]. LDPC [476]. Leader [494, 217, 128]. Leader-based [494]. LeadingOnes [721]. Leakage [285, 2493, 187, 839]. leakage-amplified [187]. Leakage-resilient [285, 839]. learned [1294, 985]. Learning [1984, 2270, 1910, 159, 1699, 1792, 559, 72, 162, 2108, 1282, 1013, 2256, 2283, 873, 1607, 2110, 1857, 165, 1708, 166, 1649, 1859, 1620, 1636, 69, 1847, 1006, 1473, 2429]. least [142, 150]. Lech [2478]. Left [148, 1944]. Left-handed [148]. LEGO [1294]. Leman [1271]. Lempel [1497]. length [274, 1140, 2013, 1200, 910, 4, 133, 2230, 1692, 1433, 1496, 1201]. lengths [698]. lens [1987, 1453]. letter [947]. letters [2424]. level [61, 857, 271, 1638, 2340, 747]. levels [1327]. Levin [1957, 782]. lexicographic [322, 320]. lexicographically [1423]. liability [2099]. Liar [625]. Lie [1642, 2355, 1795]. lies [191]. life [1628, 1094]. Lifelong [159]. Lift [360]. lifted [503, 2389, 670]. light [1393]. lights [685, 2264, 701, 1873]. Lightweight [242]. like [1750, 2156, 2217, 1431, 2151, 973, 1432, 1951, 316, 1092, 1108, 474, 599]. LIL [1344]. limit [2145, 1709]. Limitations [551]. limited [1662, 2430]. Limiting [1912]. limits [2277]. line [1886, 1206, 2447, 740, 813, 1801, 1585, 1028, 1221, 2383, 903, 1880, 3, 129, 929, 928]. Linear [2183, 2103, 348, 240, 3, 129, 382, 1236, 1393, 662, 1773, 442, 1562, 249, 1760, 1798, 1734, 598, 255, 679, 2211, 1979, 2207, 463, 739, 1817, 248, 2430, 498, 2389, 439, 1007, 1221, 875, 788, 1151, 251, 1689, 1595, 497, 801, 2062, 1599, 929, 703, 720, 1310, 709, 1809, 1864]. linear-broadcast [2062]. linear-memory [679]. Linear-space [2103, 382]. Linear-time [2183, 240, 248, 2430]. linearised [435]. linearity [1766, 2343]. linearizable [1136]. Linearly [661, 686]. lines [434]. link [551, 92, 2039, 1856]. link-faulty [2039]. linkage [274]. linklessly [190]. links [2382, 1809]. Lint [2019]. Lipschitz [2090, 2274]. List [377, 332, 1347, 771, 2436]. List-coloring [377, 1347]. listing [173]. lists [186, 842, 774, 414, 423]. live [2271]. lived [1232]. LNFG [1785]. Load [1265, 522, 2085, 814]. Local [79, 363, 678, 1414, 2145, 1035, 510, 717, 642, 2418, 994, 1920, 785, 2266, 1353, 1382, 2365, 2158, 973, 1209, 1837, 1731, 547, 1238, 188, 294, 915, 749, 1631, 2376, 2205]. Local-entity [717]. local-equitable [1382]. localities [1406]. localization [1436, 2243, 1619]. localize [1513]. localized [854, 2205]. locally [2071, 1751, 1667, 1049, 1119, 1129, 1174]. locating [95, 105, 792, 1462, 1491, 1564]. locating-chromatic [105, 792]. Location [2011, 2260, 667, 1726, 493, 659, 1391, 910, 1700, 2076, 893, 1723, 717, 2494, 615, 747, 894, 2409, 513, 1884]. location-aware [493]. location-based [717]. LockFree [1136]. LockFree-kD-tree [1136]. Löf [2266]. log [1161]. logarithm [1533]. logarithmic [382, 2466, 754, 361]. logarithmic-sized [2466]. Logic [203, 2116, 249, 259, 25, 392, 266, 2432, 672, 2087, 2232, 2366, 2063, 1769, 670, 251, 278, 497, 1321, 125, 357, 509, 265, 799, 898, 1970, 511, 993]. logical [1020, 1266, 1442, 1216, 1908]. logics [264, 1452, 255, 253, 263, 1622]. London [1133]. Long [1232, 1836, 219]. Longest [244, 735, 526, 1696, 864, 833, 1888, 1576, 2103, 296, 887, 2253, 238, 828, 349, 755]. longest-common-prefix [1888]. look [1639]. look-ahead [1639]. Lookahead [1908, 247, 1284]. looking [1690]. Loop [1448, 549, 1164, 1755, 2320]. Loop-checking [1448]. Loopless [2230]. loops [2487]. Loosely [1784, 128]. Loosely-stabilizing [1784, 128]. losses [1911]. lossy [1858, 92, 1088]. lotteries [830]. love [1294]. Low [131, 1971, 1105, 2088]. low-noise [1105]. low-power [2088]. Low-weight [131]. Lower [516, 1995, 2090, 445, 169, 1329, 1190, 757, 952, 1789, 537, 221, 1056, 344, 793, 713, 2158, 809, 213, 1310, 294, 1818, 1369]. lowpower [1858]. LP [1143, 1723, 1869, 35, 688]. LP-based [1869, 1143, 1723]. LP-rounding [688]. LPN [1105]. LPQ [71]. LSD [1133]. LTL [549, 1300]. Luca [461]. luminous [1559, 2228, 2216, 2244]. lumping [1234]. lunches [633]. LWE [2439, 2035, 2483, 1141, 2125, 1473, 1927]. Lyndon [467, 466, 864, 171, 85, 172]. Lyubashevsky [472].

M [2439, 2126]. M-ary [2126]. Ma [336]. Machine [1792, 937, 1322, 1346, 1056, 34, 1859, 1620, 1636, 774, 69, 1868]. machine-dependent [774]. Machines [1611, 868, 1255, 211, 2077, 1570, 873, 1400, 563, 2340, 814, 181, 372, 1989, 2177, 1181, 1194, 1972]. macro [2097]. magnitude [916]. Mahler [2478]. maintenance [381, 334, 1866]. Majority [1175, 962, 2060, 455, 1141, 1843, 322]. majority-vote [1843]. Maker [1903]. makes [806]. Makespan [181, 372, 1574]. making [1671, 2021]. malicious [1141, 2412]. malicious-majority [1141]. malignant [1671]. Mallows [1669, 2259]. Malware [1911, 1192]. Management [742, 1284, 347, 1960, 1861]. manipulation [1210]. Many [2039, 661, 2307, 1432]. Many-to-many [2039, 2307, 1432]. map [1863, 86, 364]. mappability [227]. Mapping [656]. March [825, 837, 847, 1375, 1385, 1395, 1407, 1965, 1975, 1982, 1991, 2359, 2368, 2380]. Marginal [445, 721]. Mark [1107]. marked [1253]. market [347, 1906]. marketing [30]. Markoff [468, 170]. Markov [1244, 2233, 2082, 163, 610, 776, 1758, 1450, 362]. marriage [1083]. Martin [2266]. martingale [36]. Massively [2319]. master [2436]. match [194, 914, 2284]. matched [2048, 2014]. matches [1315]. Matching [802, 928, 1409, 1545, 98, 1423, 807, 234, 2281, 2350, 1073, 1703, 2304, 1834, 2202, 1737, 58, 756, 631, 449, 2052, 2442, 1685, 2473, 1712, 636, 1367, 1863, 759, 1242, 1768, 618, 973, 1078, 1106, 1899, 634, 236, 752, 1074, 681, 53, 435, 645, 2422, 648]. matchings [2392, 2173, 2026, 186, 1962, 2307, 2374, 351, 1202]. Mathematical [1813, 336]. Mathematics [906, 1658]. matrices [286, 840, 2337]. Matrix [1237, 2254, 173, 1511, 2484, 443, 294, 917, 2272]. Matroid [2107, 1423, 2440, 2283, 186, 1186]. Matroid-constrained [2107]. Matroidal [2227]. matroids [1826, 348, 1237]. matter [434, 1770, 81]. matters [610]. MAWs [2183]. Max [2350, 1699, 1735, 1165, 738, 344, 1993, 688]. Max-Bisection [344]. Max-Cut [738]. Max-min [2350, 1165]. Maximal [2245, 731, 2184, 230, 1630, 1423, 2016, 1336, 1073, 343, 1784, 2173, 1410]. maximization [2123, 2440, 32, 732, 2282, 2079, 2261, 371, 2037, 1917, 820, 1706, 560, 416, 425, 1534, 802, 1492, 949, 1186, 2330, 1781, 709, 1774, 1918, 2328, 39, 517, 748, 904, 855, 1657, 31]. Maximize [745, 132]. Maximizing [1371, 55, 2237, 553, 2396, 1913]. Maximum [886, 1044, 1114, 1124, 641, 1002, 554, 487, 1872, 1305, 803, 521, 1143, 275, 918, 1721, 1577, 1969, 1782, 583, 2230, 1078, 53, 107, 2298, 1227, 707, 1162, 363, 2068, 733, 656]. Maximum-Duo [656]. maximum-spacing [1969]. maximum-weight [1782]. maxmin [1738]. MAXSAT [47]. May [289, 313, 329, 934, 945, 955, 1456, 1465, 1474, 1483, 1493, 2024, 2033, 2043, 2056, 2434, 2443, 2452, 2461, 531]. McEliece [2437]. McEliece-type [2437]. me [2375]. Mean [1602]. means [1110, 2421, 1731, 749]. measure [271, 853, 1546, 2038]. measurements [1519, 1611]. measures [1630, 778, 1325, 317]. Measuring [832]. mechanics [1294]. mechanism [1858, 856, 38, 1861]. Mechanisms [1726, 40, 325, 2150]. media [1708]. median [419, 428, 2494, 2339, 1048, 1118, 1128, 1597]. mediators [113]. meet [2141, 1100]. meet-irreducible [2141]. meeting [1794]. meets [25, 1464, 1945, 993]. member [1209]. Membership [1026, 671]. membrane [79, 82, 1401, 878]. memorized [1863]. memory [865, 679, 2242, 933, 2022, 1528, 494, 2340, 2427]. memory-efficient [2022]. Menger [1565, 2155, 657]. Menger-edge-connectedness [1565]. merger [2240]. merging [242]. Message [213, 2236, 1978]. Meta [1688, 72, 284]. meta-data [284]. Meta-Heuristic [1688]. meta-learning [72]. metacyclic [174]. metaheuristic [1641]. metalanguage [1059]. metalanguages [1521]. method [1244, 587, 491, 2059, 371, 2064, 1233, 1320, 1863, 1588, 1110, 2357, 1758, 1815]. methodology [1671]. methods [1900, 1890, 1908, 433, 1847]. Metric [174, 338, 2231, 248, 2432, 103, 89, 96, 743, 109, 1311, 1444, 63, 1326, 1997, 513, 1543, 1884, 94]. metrics [1641, 930, 958]. microbiome [2321]. millionaire [46]. mim [272, 2020]. mim-width [272, 2020]. min [1699, 1735, 2350, 1165]. min-max [1699, 1735]. minded [370]. minimal [1643, 1160, 2411, 1339, 2476, 463, 486, 1770, 1554, 1583, 918, 1010, 2320, 637, 316, 791, 1411, 992]. minimal/shortest [637]. minimalist [2413]. minimality [1093]. Minimisation [1763]. minimization [1270, 1028, 195, 1523, 181, 372, 2293]. minimize [1458, 1929]. Minimizing [1182, 418, 427, 933, 521, 2077, 1342]. Minimum [1693, 457, 916, 903, 1026, 35, 1329, 2303, 1073, 2094, 1904, 544, 229, 1948, 1721, 1085, 1577, 1911, 1870, 520, 1595, 477, 1695, 583, 2028, 1411, 1736, 29, 1241, 1469, 1663, 2409, 102, 479, 1879]. Minimum-width [457]. mining [1916, 1785, 799]. minmax [1496]. Minor [2388, 1477, 510, 584]. minor-closed [584]. minor-free [510]. minors [279, 572]. Minsum [111, 1007, 1864]. Mirror [2220]. MIS [2244, 679]. misinformation [1636]. mismatch [1624]. mismatches [2245]. MIST [1689]. mistake [2108, 1013, 166]. mistake-bounded [2108, 1013, 166]. mitigation [1192]. mitoticity [384]. mixability [629]. Mixed [276, 1295, 492, 760, 1713, 1773, 34, 299, 1818]. Mixed-case [760]. mixing [629]. mm [1425, 2365]. MM* [420, 429, 2275, 1152]. mobile [846, 207, 680, 1885, 851, 557, 1391, 1816, 727, 800, 2271, 1655, 1430, 2189, 2264, 816, 2338, 388, 192, 1873, 1864]. Möbius [179, 1403]. Moblot [2458]. modal [1622]. mode [901]. Model [963, 1907, 125, 1450, 1552, 1211, 122, 1322, 1426, 764, 208, 2393, 173, 2418, 672, 2458, 959, 21, 729, 503, 1922, 40, 1370, 1331, 324, 2456, 73, 1955, 1612, 2363, 1921, 1823, 2321, 217, 1225, 2466, 443, 2414, 420, 429, 1607, 2110, 485, 1638, 1679, 1152, 714, 1669, 2151, 2259, 2448, 356, 2365, 1166, 1674, 258, 1739, 188, 294, 758, 2285, 687, 1758, 1470, 1224, 2051, 1368, 1202, 979, 333]. Model-checking [125]. Modeling [2187, 1351]. Modelling [1624, 1859, 2277]. models [2042, 1244, 2460, 1442, 2108, 963, 2358, 2145, 163, 1857, 958, 2275, 2340, 74, 1791, 1234]. modes [1524]. modification [1160, 194, 914, 435, 1199]. Modified [1053, 1912, 12, 2155, 2494]. Modular [2393, 1584, 164]. Module [2483]. Module-LWE [2483]. modules [981]. modulo [2232, 1099, 346, 1889]. Molecular [2458]. Moment [1356]. Moment-based [1356]. monad [1444]. monadic [392]. monadic-second [392]. Monnot [1528]. Monochromatic [1750]. monoids [462]. Monotone [1537, 1172, 1575, 2261, 745, 2037, 2267, 1021, 553, 1913, 2396, 1724, 2330, 1781, 709, 1774]. Monotonic [1748, 1514, 581]. monotonicity [2282, 325]. Moore [435, 1989]. moral [1104]. Morphic [62, 196]. Morse [1750, 1658, 2353]. Morse-like [1750]. most [354, 1228]. motion [2136, 2266]. motivated [114, 2178]. Move [388]. Move-optimal [388]. moves [1886, 2455, 655, 1488]. MPC [679]. MR [1058]. MSVL [198]. MUL [1780]. MUL-tree [1780]. Multi [450, 2222, 480, 759, 1534, 1913, 453, 2125, 651, 857, 1518, 865, 1743, 2157, 1532, 191, 902, 212, 631, 2258, 579, 541, 972, 654, 65, 663, 903, 614, 2246, 444, 1614, 1181, 284, 29, 519, 476, 2003, 1424, 2429]. multi-agent [541]. Multi-attribute [1534]. multi-channel [65]. Multi-colored [450]. multi-commodity [1518]. multi-counter [972]. multi-cover [631, 903, 29]. multi-hop [1532, 654]. Multi-input [453]. multi-instance [663]. multi-intersection [2258]. multi-interval [191]. multi-item [284]. Multi-key [2125, 1181]. multi-level [857]. Multi-matching [759]. Multi-objective [1913, 444, 1614]. multi-party [2157, 212]. multi-player [519]. multi-recombinative [2246]. multi-secret [1424]. multi-service [579]. multi-signature [1743]. Multi-stage [2222, 614]. multi-state [2003]. multi-string [865]. multi-task [2429]. multi-unit [651]. multi-user [476]. Multicarrier [319, 10]. multichannels [100]. multiclass [166]. multicollision [586]. multicoloring [218]. multicommodity [1703]. multicovering [276]. multicriterian [1671]. multicut [184, 2239, 608]. Multidimensional [754, 1260, 994]. multigraph [2373]. multilinear [169]. multimodal [2246, 1861]. Multiplayer [532]. Multiple [1040, 570, 124, 893, 962, 863, 1939, 546, 1538, 813, 706, 1557, 195, 947, 1863, 416, 425, 1074, 681, 350, 2109, 599]. multiplex [1985]. multiplication [173, 2002, 719]. multiprocessor [1751, 2418, 2365]. Multistage [1738, 951, 1387]. multitape [456]. multivalued [2309]. multivariate [193]. multiwinner [2138]. Murru [1593]. mutation [1945]. mute [2497]. Mutual [1559, 2196, 2194, 1508, 782]. mutual-visibility [2194]. mutually [1667]. Mycielskians [1043, 1113, 1123]. Myhill [1441]. myopic [2228, 2216, 2244].

name [257]. Nash

[886, 522, 2274, 632, 2095, 1972]. Natasa [1247]. natural [1910, 2040, 1467]. navigation [699]. navigational [1660]. Near [1322, 680, 1887, 2428, 1393, 547].

Near-gathering [680]. near-linear [1393]. Near-neighbor [1887]. Near-optimal [1322]. near-triangulations [547]. nearest [1136, 195, 801]. Nearly [2199, 1969, 1347]. Negation [2023, 1893]. Negation-closure [2023]. negative [1968]. negativity [2282]. Neighbor [95, 1478, 1887, 403, 801, 1564, 2058, 1379, 59, 1556, 2292]. Neighbor-connectivity [1478]. Neighbor-locating [95]. neighbor-sum- [1379]. neighborhood [1502, 545]. neighborhoods [2375]. neighbour [1136]. Nerode [1441]. Nested [762, 2487, 1093, 759, 42, 2309]. net [1406, 1404, 1673, 2166]. nets [79, 2356, 141, 2393, 1298, 1615, 1188]. Network [1688, 841, 1248, 836, 208, 141, 684, 404, 1654, 2079, 1553, 573, 1256, 1513, 895, 45, 790, 1860, 1919, 1708, 1923, 1929, 1928, 1714, 2361, 213, 997, 704, 1623, 2226, 1356, 1620, 280, 1815, 2038, 2292, 1058, 333]. Networks [2451, 1392, 1182, 110, 2054, 2152, 2203, 1544, 982, 1910, 1858, 1273, 1064, 111, 606, 2411, 1489, 596, 1312, 11, 114, 2303, 32, 1569, 2085, 1532, 2342, 851, 100, 189, 557, 1335, 2208, 211, 161, 697, 1286, 1652, 30, 118, 850, 2037, 2217, 773, 910, 1007, 567, 2099, 600, 216, 670, 1637, 1911, 180, 611, 1269, 1367, 1072, 1588, 403, 31, 714, 1167, 1431, 1565, 2357, 746, 176, 471, 1655, 717, 876, 832, 560, 416, 425, 1534, 1432, 1951, 277, 1880, 2373, 802, 1492, 1803, 1954, 2468, 2382, 258, 1856, 1620, 1084, 1636, 473, 2166, 602, 400, 645, 1243, 1032, 1478, 1824, 2119, 154]. networks [1561, 2422, 474, 39, 517, 1986, 2352, 1427, 1306, 1556, 2051, 760, 855, 981, 1571, 2091, 1831, 2284, 2187, 1289, 1889, 1557]. Neural [333, 161, 2358, 718, 1607, 1846, 1860, 1911, 2132, 2151, 1803, 1623, 1620, 602, 154, 879, 2497, 14, 2109, 335, 1235, 1631, 1522, 2287]. neural-like [2151]. neuro [1544]. news [1708, 1636]. NextPriorityConcept [623]. NFA [2178]. nimbers [1930]. NIZKs [2285]. NN [44]. no [633, 698, 1477, 1875, 1227]. Node [1594, 2395, 219, 2361]. node-disjoint [2395]. nodes [732, 1513]. noise [2042, 1917, 1857, 1238, 1105, 1006]. noiseless [781]. noisy [443]. nominal [1759]. Non [1630, 709, 1774, 2491, 19, 858, 2211, 2123, 2282, 121, 1922, 2449, 324, 1968, 1388, 2261, 91, 1440, 2037, 796, 1905, 164, 328, 548, 1706, 1928, 1095, 553, 241, 1844, 703, 1766, 2396, 2330, 1781, 2293, 120, 1556, 1646]. non- [164]. non-adversarial [121]. non-blocking [548]. non-canonical [1388]. non-datapath [1928]. non-determinism [703]. non-deterministic [91, 1440]. non-inclusive [1922, 1556]. non-interactive [1646]. non-interference [1095]. non-linear [2211]. non-linearity [1766]. Non-maximal [1630]. Non-monotone [709, 1774, 2261, 2037, 553, 2396, 2330]. non-negative [1968]. non-negativity [2282]. non-overlapping [796]. non-spanning [1905]. non-stationary [2293]. non-strongly [241]. non-submodular [2123, 1706, 2396, 1781]. non-uniform [324, 328, 1844]. non-uniformly [120]. non-visible [858]. non-zero [2491, 2449]. non-zeroness [19]. nonadaptive [295]. noncommuting [2478]. nondeterminism [958]. Nondeterministic [2176, 248, 1346, 1810, 2124, 1296, 2459]. nonembeddability [1613]. Nonexistence [1228]. nonisomorphic [106]. Nonlinear [2151, 2109, 2059]. nonlinearity [1157]. nonsubmodular [904]. Nonuniform [1047, 1117, 1127]. norm [195, 615]. normal [585, 834, 2002, 262, 587, 582]. normality [2126]. normalized [1849, 361]. normalizing [2464]. Not-All-Equal [301, 387]. notations [1467]. Note [449, 1603, 690, 504, 489, 1013, 763, 2355, 1723, 2028, 2072]. notes [78]. notion [782].

Novel [803, 50, 1743, 2358, 271, 883, 1533, 201]. November [527, 542, 561, 575, 1197, 1214, 1229, 1764, 1777, 1787, 2234, 2247, 2268]. NP [1942, 93, 2159, 2055, 345, 1685, 1400, 2388, 1422, 2374]. NP-complete [1942, 2055, 2388, 1422, 2374]. NP-completeness [345]. NP-hard [93, 2159]. NP-hardness [1685]. NSD [1806]. NTRU [1743, 2125]. nucleolus [2473]. number [16, 286, 1713, 620, 1169, 105, 1190, 1507, 792, 1109, 1942, 1592, 1014, 1867, 1458, 2165, 1380, 2406, 988, 58, 1808, 449, 1775, 1187, 763, 304, 2410, 1012, 1005, 924, 1588, 713, 1242, 477, 150, 1722, 277, 862, 2003, 1148, 97, 1510, 1201, 273]. number-conserving [2003]. numberings [1621]. numbers [52, 1307, 1640, 2454, 43, 1313, 1467]. Numerical [2054, 1846, 70, 2378]. Nurimisaki [2171].

O [2363]. O-efficient [796]. O-optimal [1288]. obfuscation [1552, 654]. object [82]. objective [444, 1913, 1614]. objectives [164]. objects [662, 2104, 1137]. obligations [1360]. Oblivious [311, 306, 1213]. obnoxious [2076]. observation [1657]. observations [1383]. obstacle [1676, 765]. obstructive [1860]. Obtaining [2285]. occasion [1628, 1247, 861]. occur [2210]. occur-check [2210]. occurrences [1645]. October [469, 482, 495, 507, 1145, 1158, 1170, 1184, 1718, 1729, 1744, 1753, 2200, 2213, 2223]. odd [2406, 1157]. odd-variable [1157]. off [118]. Offensive [2386, 1910]. Offline [370, 1805, 1938, 2088]. offloading [1824, 2429]. once [1082, 1189]. One [25, 1142, 1651, 570, 417, 426, 266, 2243, 12, 769, 939, 726, 1523, 725, 636, 1088, 776, 87, 595, 701, 752, 519, 529, 1535, 655]. one-axis [701]. one-cop-moves [655]. one-dimensional [939, 595]. one-pass [266]. one-round [519]. one-sided [776]. One-variable [25]. One-visibility [1142, 1651, 2243]. one-way [769, 1523, 725]. one-wayness [769]. OneMax [445, 362]. ones [228]. Online [116, 1699, 619, 539, 1805, 1458, 2420, 182, 1391, 784, 2256, 1676, 2172, 132, 636, 820, 1791, 1910, 867, 1789, 32, 1738, 1039, 40, 1938, 2261, 2108, 30, 1013, 1650, 2414, 1523, 165, 1637, 1072, 1796, 166, 1649, 240, 758, 370, 2088]. Online/offline [1805, 2088]. only [12, 2111]. opacity [2364]. opaque [685, 2228]. Open [41, 937, 2220, 1932, 574]. Open-shop [41]. operation [2139, 2124]. operational [2176]. operations [2332, 1612, 2450]. Operator [672]. operators [71, 2046]. Opinion [2060, 962]. opposite [730]. oppositional [1859]. optima [2165]. Optimal [962, 1206, 1662, 1885, 2228, 539, 1014, 1338, 2013, 2188, 798, 9, 659, 737, 2486, 1154, 1430, 2264, 705, 1600, 830, 1651, 1322, 1705, 1979, 411, 2446, 2270, 995, 1780, 1547, 1003, 2428, 1826, 541, 1284, 2259, 1928, 1649, 200, 388, 1157, 2349, 280, 1148, 1288]. Optimally [2218, 2338, 1705, 3, 129, 1195]. Optimisation [836, 444]. optimising [1489]. Optimism [164]. Optimization [2272, 1544, 1387, 598, 1363, 491, 2059, 51, 1240, 1557, 1216, 164, 930, 117, 1929, 614, 1261, 488]. optimized [337]. Optimizer [1859]. Optimizing [2085, 1641, 1912, 746, 2164]. option [1031]. optional [667]. oracle [199, 1536, 829]. oracles [1660, 827, 660, 1400, 1902]. Orbit [197]. Order [1866, 505, 392, 2432, 1057, 1380, 144, 1934, 2192, 756, 1001, 104, 1741, 322]. order-preserving [756]. ORDER/MAJORITY [322]. Ordered [2427, 651, 1759]. Ordering [1967, 2, 339, 2294, 1015, 1001]. orders [916, 924]. ordinal [2260]. ordinals [2041]. ordinary [219]. organization [902, 1260]. organizational [74]. orientation [1169, 607, 1329, 1147]. orientations [1509, 2237]. oriented [1163, 1187, 808]. Orienteering [882, 1463]. oritatami [2487, 968]. Ortho [884]. Ortho-polygon [884]. orthogonal [662, 5, 591, 769, 740, 1076, 303, 666, 1434]. Ostrowski [821]. Ostrowski-automatic [821]. OT [1211]. OT-hybrid [1211]. other [886, 1507, 864, 942, 698, 1259]. outerplanar [2475, 4, 133]. outliers [1602, 819, 2467]. outperform [446]. output [2136, 219, 2340]. outsourcing [1031]. overall [1318]. Overflow [1960]. overlap [1039]. overlapping [796].

P [1524, 1239, 2131, 863, 2104, 2111, 337, 718, 1607, 2110, 1846, 1400, 2132, 1401, 2032, 879, 2497, 1632, 2109, 1235, 1631, 599, 1522, 2287]. Pachinko [534]. packet [2382]. packing [530, 1898, 1890, 2301, 552, 2257, 1420, 1551]. packings [2187]. page [1713, 1942]. page-number [1942]. Pages [245, 267, 412, 421, 430, 592, 604, 621, 638, 652, 668, 83, 167, 6, 17, 27, 48, 66, 358, 367, 379, 389, 395, 405, 289, 313, 329, 527, 542, 561, 575, 469, 482, 495, 507, 447, 880, 891, 907, 919, 1051, 1070, 1262, 1277, 766, 779, 794, 810, 676, 694, 710, 723, 750, 976, 989, 998, 1008, 1017, 1024, 1033, 1041, 1111, 1121, 825, 837, 847, 934, 945, 955, 1197, 1214, 1229, 1145, 1158, 1170, 1184, 1080, 1089, 1096, 1131, 1418, 1428, 1438, 1446, 1617, 1625, 1634, 1349]. Pages [1358, 1365, 1292, 1302, 1316, 1323, 1332, 1340, 1503, 1515, 1526, 1540, 1578, 1604, 1375, 1385, 1395, 1407, 1456, 1465, 1474, 1483, 1493, 1764, 1777, 1787, 1718, 1729, 1744, 1753, 1647, 1664, 1681, 1701, 1799, 1819, 1839, 1854, 1876, 1895, 1710]. paging [1691]. Paid [397]. pair [1078]. Paired [2351, 56, 2115, 1687]. paired-domination [56]. pairs [2300, 1775, 1093, 1505]. pairwise [1409, 1362, 1962]. palindromes [2245, 150, 2253, 238]. Palindromic [1692, 1200, 833, 1576]. Palindromization [170, 468]. pancake [2351, 1478]. pancyclicity [773, 1389, 2049, 1032]. papers [2186]. paradox [2240, 2098]. Parallel [235, 1721, 1352, 2367, 674, 203, 1773, 2077, 261, 563, 2319, 181, 372, 1880, 2338, 2147, 3, 129, 1443, 53, 1574, 1301, 1868, 1265]. parallelizable [1791]. Parallelized [904]. parallelogram [457]. Parameter [163, 2050, 640, 1990, 276, 738, 9, 918, 2063, 1714, 228, 1134, 102]. Parameterization [481, 47]. parameterizations [1833, 1752, 94, 572]. Parameterized [1881, 475, 2048, 1343, 2463, 1394, 2239, 936, 896, 1732, 752, 626, 258, 648, 735, 2286, 613, 2209, 2395, 1417, 103, 1011, 1740, 756, 2052, 2457, 656, 636, 1001, 64, 555, 348, 1264, 2135, 2249, 2005, 1188, 1828, 2004, 299, 2014]. Parameterizing [377]. parameters [2460, 890, 617, 1368, 2067]. parametric [1244, 1056, 1852]. Parametrised [104]. parametrized [1082]. Paraunitary [286]. parent [444]. Pareto [1962]. Parikh [840, 2337]. parities [559, 1006]. parity [665]. parsimony [322, 320]. part [1266, 2199]. Partial [1621, 1374, 2481, 1315, 971, 184, 2262, 1509, 1281, 525, 360, 1721, 2027, 2120, 903, 29, 388, 568, 2015, 2370, 2051]. partial-vertex-cover [360]. partially [1871]. participants [835]. partite [1242]. partition [1661, 88, 2475, 1394, 2357, 574, 102, 981]. Partitioning [1686, 2315, 345, 809, 1369, 2038]. partitions [1581, 613, 1248]. partizan [1045, 1115, 1125]. party [2157, 212, 983, 2347]. pass [266]. passenger [1653]. passing [1809]. Password [1673, 1276]. past [255, 1514]. past-finite [1514]. paste [102]. Path [1879, 35, 111, 700, 953, 2351, 455, 1875, 1810, 1797, 1007, 45, 930, 1437, 1870, 887, 125, 240, 375, 1869, 1496, 2061, 1632, 479]. Path-Collection [1879]. Paths [845, 2300, 1046, 1116, 1126, 1742, 1337, 1143, 1380, 659, 261, 1940, 1536, 2492, 1309, 2039, 1811, 1432, 1561, 1809, 640, 1752, 1311]. pathwidth [1596, 1001, 1535]. patrolling [541, 1430]. Pattern [574, 1748, 1545, 160, 234, 306, 685, 1343, 1878, 2458, 926, 269, 2442, 973, 1732, 2058, 1074, 681, 435]. Pattern-avoiding [574]. Patterns [57, 287, 414, 423, 1836, 634, 1481]. Paun [861]. paw [1199]. paw-free [1199]. Pebble [1547]. pebbles [699]. pebbling [43]. Pell [1108, 1180]. penalties [1595, 2062, 987]. penalty [1329]. pendulum [335]. people [209]. percolation [1609]. perfect [629, 1712, 547, 2374, 897]. Perfectly [2014, 2048, 1211]. perform [1611]. Performance [1955, 994, 985, 2277, 1641, 1847]. performing [360]. periods [634, 2343]. permanent [626]. Permutation [2488, 769, 106, 97, 2091, 2348]. permutations [1798, 287, 769, 734, 1290, 1207, 1075, 1766, 992, 256]. Perpetual [2216, 2228]. persistence [1068]. persistent [564, 2029, 1095]. personal [980]. perspective [938, 684, 732, 1011, 243, 2007, 1994, 1803]. Péter [1973]. Petri [79, 141, 2393, 1404, 1673, 1615, 2166]. phi [831]. phylogenetic [1489, 307, 473]. Physarum [1518, 324, 328]. Physarum-inspired [1518]. Physical [2171, 1272, 1164, 1714, 1351]. picking [790]. picture [973]. pieces [2417]. piecewise [1233]. PKE [1105]. place [1480, 1188]. placement [1662, 798, 200]. Placing [387]. plain [1357]. planar [1273, 1498, 2306, 2251, 942, 805, 1307, 1274, 740, 489, 738, 1536, 1490, 1144, 1865, 465, 1551, 1077, 126, 1687, 1410, 1633]. planarity [61, 1280, 2475, 1015]. plane [493, 1076, 1221, 480, 884, 704, 547, 1864]. planning [2136, 260, 1339, 994, 2215, 45, 74, 1871]. plans [1882]. plasticity [2497, 14]. platforms [2218]. play [2417]. player [2136, 2220, 519]. players [844]. PLC [1192]. Plug [464]. Plug-in [464]. plus [1450, 688]. PMC [1922, 485, 2275, 1152, 714, 1739, 687, 2051, 1202]. PMC/MM* [1152]. PMMS [2385]. point [2071, 2, 339, 649, 1281, 252, 2464, 720, 1993]. point-free [2071]. point-set [649, 1281]. Point-Tolerance [1993]. points [1693, 2334, 1273, 1981, 587, 2420, 1144, 288, 477, 717, 1220]. pointwise [709, 1774]. polarization [876]. polarizationless [2111]. polarizations [316]. policy [1338]. polls [386]. poly [382, 754]. poly-logarithmic [382, 754]. polygon [587, 884, 666]. polygonal [765]. polygons [1079, 1076, 303, 1573]. polyhedral [659]. Polyline [185]. polylogarithmic [585, 128]. Polymer [1891, 600]. Polynomial [2002, 354, 1223, 456, 2096, 572, 251, 1199, 300, 1044, 1114, 1124, 2392, 2016, 387, 1347, 2202, 1812, 272, 195, 247, 288, 1373, 1695, 1327, 1300, 828, 107, 1574]. Polynomial-delay [2096]. polynomial-maximal [2016]. Polynomial-time [354, 1223, 272, 247, 1695, 1574]. polynomials [996, 195, 1101]. polyomino [530]. polytope [360]. polytopes [454]. Popular [186, 2307, 2173, 2006]. Popularity [287]. Population [581, 1244, 2233, 1254, 2321, 2267, 120, 1224]. populations [870, 1021]. poset [1445]. Posets [504, 203]. position [582, 2371]. positionality [632]. Positive [1490, 969]. POSIX [807]. Post [1240, 1533, 476]. Post-processing [1240]. post-quantum [1533, 476]. postulates [993]. potential [689, 1934]. Power [969, 722, 158, 1239, 2010, 183, 2456, 253, 217, 2478, 718, 567, 263, 2455, 1437, 903, 1595, 1709, 1166, 529, 513, 2088]. power-set [253]. powers [1151, 1944, 877, 1525, 1692]. PPA [1099]. PPA- [1099]. PPAD [2274]. PPAD-complete [2274]. PPTL [1785]. Practical [2157, 46, 298, 2412, 1143]. practice [323]. Pre [318, 1708]. Pre-expansivity [318]. pre-trained [1708]. precedence [41, 672]. precise [9, 180]. preclusion [449, 1242, 645, 2422]. prediction [1860, 2376]. Preface [824, 590, 2400, 135, 342, 1935, 588, 407, 226, 1060, 1063, 308]. Prefer [730]. Prefer-opposite [730]. Prefer-same [730]. Preference [1834, 667, 186, 2436, 993]. preferences [1409, 2260, 2255, 1962, 1653, 2076]. preferences-based [1653]. preferred [501]. Prefix [1023, 735, 864, 585, 834, 1388, 1200, 1888, 2047, 1290, 1443]. Prefix-free [1023, 1388]. prefixes [1046, 1116, 1126]. Preprocessing [1646]. Presburger [25]. prescribed [1971, 1168, 1809]. presence [732, 493, 2114, 1154, 1006, 765]. presentable [1950]. presentations [619, 1771]. Presenting [393]. Preservation [656, 2309]. preserved [244, 491, 554]. preserving [1314, 1770, 1887, 756, 1476]. pressure [322, 320]. price [130, 1384, 2108, 1013, 1364, 768, 1969, 1072, 166, 397]. Priced [2174]. Pricing [37, 982, 1338, 1072]. Priestley [1909]. primality [843]. Prime [1560, 1248, 2481, 432, 1505]. primer [2073, 2379]. Primitive [923, 996, 1612, 256]. primitives [769]. Principal [188]. principle [1519]. principles [13, 1956]. Prior [651, 1058]. Prior-free [651]. priorities [2099]. Priority [2495, 127, 647, 506, 2195, 2324, 1801, 1430, 774]. priority-based [2195]. Privacy [1476, 1552, 2449, 569, 1917, 2064, 325, 188, 294, 915]. Privacy-preserving [1476]. privacy-utility [325]. private [212, 1907, 1926, 1593, 917, 2293, 155, 175, 2412]. prize [608, 1666, 1048, 1118, 1128]. prize-collecting [1666, 1048, 1118, 1128]. PRNG [1344]. proactive [1805]. Probabilistic [26, 930, 2317, 1762, 248, 963, 883, 1320, 1746]. probabilities [1344, 302]. probability [1244, 11, 2082, 670, 1613, 39, 517]. probably [1003]. probe [1725]. probing [1783]. Problem [1040, 2397, 888, 526, 931, 2212, 1249, 937, 1507, 274, 111, 184, 19, 1448, 1792, 1381, 1670, 1728, 867, 1789, 1343, 1305, 2068, 1417, 2350, 103, 2316, 183, 1752, 2094, 1904, 291, 738, 2, 339, 72, 1135, 882, 1462, 2148, 544, 1845, 902, 659, 1834, 1725, 733, 1740, 1394, 525, 1801, 1364, 1362, 1948, 456, 980, 1650, 2052, 2387, 2297, 1176, 586, 608, 204, 743, 579, 1031, 936, 788, 1712, 948, 1689, 1838, 2486, 1576, 1155, 115, 2103, 490, 31, 1110, 1595, 1706, 296, 1311, 38, 1649, 2129, 1138, 252, 887, 1723, 1722, 1599, 1533, 899, 797, 440, 90, 704, 1308, 1614]. problem [929, 705, 828, 29, 1134, 1241, 1074, 1869, 2021, 2494, 107, 693, 615, 987, 1496, 2015, 119, 1616, 2339, 2467, 513, 1550, 362, 370, 749, 760, 855, 1691, 1884, 1866, 1162, 1139, 2203, 1872, 1594, 1543]. Problems [35, 817, 475, 1378, 1471, 1486, 1507, 1699, 1661, 2185, 1735, 398, 2395, 1423, 516, 1500, 1572, 1990, 1771, 1363, 1987, 2194, 1703, 1076, 1039, 2055, 921, 1941, 2276, 373, 2261, 978, 51, 275, 871, 1490, 910, 2206, 951, 194, 1007, 2283, 1085, 1216, 1523, 1361, 930, 1685, 2178, 1700, 1969, 2341, 1782, 1796, 1870, 1373, 618, 1704, 64, 2135, 2249, 228, 1994, 1481, 914, 2005, 974, 126, 925, 1537, 1177, 1873, 1828, 2327, 1779, 1469, 2298, 747, 894, 1434, 1632, 2299, 1199, 2182, 2376, 2069]. procedural [1312]. procedure [2211, 357, 509, 2287]. process [1066, 113, 1624, 1192]. processed [1620]. processes [866, 2233, 248, 144, 1449, 145, 1440, 2464, 153, 909, 1352]. processing [1910, 244, 1098, 1056, 2077, 1240]. processor [2172, 1649]. processors [933, 876]. procurement [284]. produced [433]. product [2491, 843, 1464, 2449, 43, 1304, 2464, 2031, 2437, 1546, 453, 645, 1473]. production [1031, 1846, 82]. products [1598, 416, 425, 791]. Professor [1999, 1628, 1973]. Profit [2037, 31, 949, 385, 32, 2079, 1492]. program [391, 1252]. program-size [1252]. programmable [434, 1770]. programming [276, 2207, 21, 1521, 2007, 320, 1638, 614, 2053, 2015]. programs [142, 916, 2366, 2333, 1755, 198, 1301, 201, 799]. Progress [2246]. progressions [1750]. Progressive [1654]. Progressive-encoding-based [1654]. Project [1827, 360, 1259]. projection [357, 509, 974, 602]. projections [1853, 994]. projector [2147]. promised [2486]. promoters [599]. promoters/inhibitors [599]. Prone [124, 1549]. Proof [2222, 878, 1591, 772, 531, 1464, 224, 2487, 775, 420, 429, 1353, 2150, 1272, 536, 1301, 1970, 511, 50]. Proof-labeling [1591, 224]. proof-of-concept [2487]. Proof-of-Works [2222]. Proofs [782, 2128, 2232, 1049, 1119, 1129, 1164, 1997, 1646]. Propagation [686, 353]. proper [1169, 1305, 2407, 1590, 1005, 1179]. properly [1940]. Properties [1542, 1971, 1244, 864, 1814, 1271, 80, 1402, 1907, 1445, 2377, 1356, 201, 2222]. property [244, 1514, 462]. property-preserved [244]. proportionate [34]. proposal [1760]. propositional [2116, 505, 2476, 2232, 1622, 509]. protecting [1212, 1011]. protection [1552, 1637]. Protocol [1688, 12, 46, 1861, 1224]. Protocols [1837, 1254, 2171, 630, 120, 155]. provable [154]. providers [464, 1906]. Proving [2377]. proximal [241]. proximity [1049, 1119, 1129, 33]. proxy [654, 913, 979]. pruning [1780]. Pseudo [270, 491]. pseudo-Boolean [491]. Pseudo-solutions [270]. pseudoforest [115]. pseudorandom [178]. PSO [2494]. PSPACE [399, 1416]. PSPACE-completeness [399]. PSPACE-hardness [1416]. PTAS [510]. Public [212, 1370, 890, 1219, 2423, 1674, 1424]. public-key [1219, 1674]. pullback [1315]. pulmonary [1860]. Puncturable [1668]. punishment [2301]. Purchase [1653]. Pure [2367, 2095, 2131, 1539, 2274]. purely [1524]. pursuit [1698]. Push [1233]. Push-forward [1233]. pushdown [1502, 1810, 2455]. Pushing [434, 2455]. pushout [1092]. pushouts [673]. Puzzle [2438, 2207]. puzzles [2314, 1164, 1813]. PVCS [883].

QC [476]. QC-LDPC [476]. QCTL [1450]. QKD [12]. QSAT [1401]. Quadratic [459, 441, 193, 72, 2202, 2333]. quadrilaterals [2349]. Quality [441, 40, 994, 71, 801]. Quality-aware [40]. quantified [387, 1619]. quantifier [1619, 670, 601]. quantifier-free [601]. Quantitative [2229, 664, 1440, 1234]. Quantum [1519, 586, 2283, 2421, 1616, 1858, 1275, 1023, 769, 2157, 1294, 1398, 1451, 2040, 514, 319, 10, 1240, 1860, 1523, 2486, 323, 1533, 1238, 722, 1610, 862, 476, 1758, 1450, 1301, 155, 22]. Quasi [1347, 1205, 2231, 2321, 1444, 1930]. quasi-metric [2231, 1444]. quasi-nimbers [1930]. Quasi-polynomial [1347]. quasi-stability [2321]. Quasi-universal [1205]. quasiperiods [753]. Quaternion [602]. Quaternion-valued [602]. quaternions [602]. qubits [722]. QUBO [1845, 887]. Queries [1506, 382, 1562, 1694, 1288, 229, 1797, 394, 873, 1893, 1437, 896, 555, 1485, 1411, 1433, 1220, 284, 916]. Query [1762, 1270, 940, 2259, 2404, 854, 524, 1098, 2090, 44, 42, 1616, 155]. Query-based [2259]. Query-competitive [940]. queues [2495]. Quotients [1520, 969].

R [2125]. RAC [417, 426]. Radical [1932]. Radio [550, 110, 1206, 208, 118, 895, 713]. radius [1768]. radix [2091]. Rainbow [1389, 861, 1103, 1155, 1985, 1279]. Rainbow-Band-Separation [1279]. rainbow-free [1103]. RAM [1374]. Ramsey [556, 924, 487]. random [158, 827, 1426, 1268, 32, 731, 36, 2266, 870, 1823, 449, 2465, 930, 2035, 1679, 876, 1166, 2375, 415, 424, 1985, 302, 1260, 2494, 862, 644, 120]. randomised [250]. Randomization [784, 1196]. Randomized [2325, 480, 816, 77, 2163, 715]. randomizing [522]. Randomness [1999, 1325, 1973, 212, 2221, 1149]. Range [55, 809, 1411, 382, 229, 754, 916]. range-LCP [382]. rank [2238, 2173, 348, 1997, 1646]. rank-1 [1646]. rank-maximal [2173]. rank-metric [1997]. Ranked [237, 975]. ranked-tree-to-string [975]. ranking [1755, 1671, 630]. Ranks [992]. Rapid [1944]. Rastrigin [2246]. Rate [1912, 376, 1709, 107]. ratio [732, 2282, 745, 4, 133, 484, 893, 1704, 1371]. Rational [2478, 943, 203, 2145, 1313, 481, 1610]. rays [220, 925]. Re [1825, 1221, 654, 913, 979]. Re-embedding [1221]. re-encryption [654, 913, 979]. Re-examining [1825]. reachability [2080, 2407, 921, 2302, 554]. reachable [2255, 1010]. Reaching [2114]. reactantless [2334]. Reaction [1063, 2334, 1629, 1064, 1065, 1020, 1066, 11, 1067, 600, 464, 1068, 1069, 992, 1608]. reactions [1067]. Reactive [1321, 24]. read [494, 1137, 1082, 1189]. read-once [1082, 1189]. read/write [494, 1137]. readability [112]. Real [2221, 735, 1486, 1829, 1823, 2366, 1908, 1637, 1842, 2364]. Real-time [2221, 2366, 2364]. realistic [800]. realizability [2198, 1575]. Realization [921]. realizations [139, 1548]. Realizing [1733]. Reallocating [813]. rearrangements [277]. Reasoning [73, 1769, 2152, 144, 1353, 1216, 896, 334]. rebalanced [432]. rebalancing [38]. Rechargeable [1557, 1929, 2382]. reciprocal [194]. recognising [2031]. recognition [1271, 850, 1623, 1772, 752, 1179]. Recognizing [1942, 300, 93, 2447, 523, 2436, 2177, 1993]. Recoloring [1037]. recombination [441, 1581]. recombinative [2246]. recommendation [800]. reconciliation [1780]. reconciliations [307]. reconfigurability [1892]. Reconfiguration [1581, 108, 2251, 788, 885, 830]. Reconfiguring [123, 1905]. Reconstructing [2286]. Reconstruction [829, 2059, 239, 1468, 1151, 1485]. recovery [1637]. rectangle [2036, 414, 423]. rectangles [995]. rectangular [1509, 2417, 984]. rectilinear [552]. recurrences [1798, 2377]. recurrent [1708, 602]. Recursive [2276, 1850, 1707, 2119, 1453, 1427, 1410, 2284, 256]. recursively [1187]. red [1048, 1118, 1128, 2148]. red-blue [1048, 1118, 1128]. redescription [1916]. reduce [1808]. reduced [737]. Reducing [840, 875, 1731, 1252, 1338, 44]. reduction [819, 1887, 261, 164, 1692, 1237]. reductions [2205, 1720, 877, 1525]. reference [1074]. Refined [2387, 273, 1383]. refinement [2078, 503]. Refining [233]. reflected [1741]. refutations [1082]. regional [2387]. register [1885, 705, 1586, 1441]. registers [411]. registration [1209]. regret [2265, 1496]. regular [137, 363, 259, 658, 807, 1825, 1900, 2496, 1885, 1878, 959, 1967, 2087, 1968, 2401, 871, 452, 1205, 420, 429, 2305, 971, 1390, 2162, 1679, 761, 176, 973, 298, 1243, 1319, 657, 2299]. regular-expression [1900]. regular-like [973]. regularities [595]. regularity [868]. regulated [866]. regulation [1064, 1232]. reinforcement [492, 2429]. Reinforcing [2449]. related [76, 1507, 1305, 1703, 1415, 2114, 131, 275, 2178, 1870, 1586, 925, 2298]. relatedness [491]. Relating [2163, 2292]. relation [57, 15, 364]. Relational [874, 1760, 1330]. relations [1984, 764, 1266, 1612, 759]. Relationship [1922, 452, 942, 420, 429]. Relationships [868]. relative [662, 1216]. Relativized [1974]. relaxation [1280, 1311]. relaxations [35]. relaxed [2210]. Reliability [2208, 180, 1367, 1167, 1894, 1546, 2119, 327, 691, 1897, 2284, 1286, 471, 1603]. reliable [2236, 2170, 1228]. relocatable [1147]. Remarks [1909, 869, 1851, 2394]. remote [285]. Removing [1583]. rendezvous [2329, 2264]. rental [1792]. Rényi [191]. reordering [2338]. repair [1990, 1624]. repairs [501]. repeated [2170, 2370]. Repeatedly [2281]. repeats [269]. repetition [526, 926, 2445]. repetition-bounded [526]. repetitions [230, 1736]. replica [200]. Replicated [1355]. reports [2370]. repository [1234]. Representable [2082]. representation [2254, 1399, 1867, 33, 348, 993]. representations [2159, 1388, 1836, 884, 465]. Representing [74, 262]. Reprint [468, 467, 466]. reprogramming [997]. require [361]. requirement [798, 2409]. requirements [740]. RESCOT [1261]. rescue [2118]. Research [883]. resemblance [1911]. Residents [2387]. resilience [2446, 42, 555]. resilient [366, 1511, 2406, 285, 839]. resistance [177, 682]. resistant [979, 2362]. Resolute [1472]. Resolution [291, 1901, 2306, 1417, 100, 1826, 717, 1189]. resource [138, 571, 2095]. resources [117]. respect [1001, 1677, 1189]. Restrained [2030]. Restricted [403, 899, 1210, 2331, 310, 539, 1073, 58, 247, 842, 702, 353, 317, 974, 1188]. Restriction [1261]. restrictions [617]. result [1968]. results [888, 982, 1471, 857, 1415, 272, 1948, 641, 2297, 1756, 565, 873, 545, 345, 2419, 538, 1138, 277, 1030, 696, 1435, 1495, 228, 1586, 993, 1177, 894, 69]. revenue [1906]. reversal [1257, 1291]. reversals [2185]. reverse [775]. reversibility [1629]. Reversible [2007, 1521, 1624, 278, 252, 1844, 1520]. reversing [1615]. revisited [1957, 1805, 612, 2341, 1712, 1264, 202]. Revisiting [730, 656, 2298, 1434, 1412]. Revocable [1055, 292, 682, 890, 177]. revocation [2449, 1209]. revoke [2362]. rewritable [2393]. Rewriting [1094, 1357, 2163, 866, 20, 1315, 1092, 674]. Ricci [2205]. rich [2326, 1283]. ride [2069, 1675, 2069]. ride-hitching [1675]. ridesharing [980, 37]. Riemann [151]. right [77, 2163, 875]. right-hand [77, 2163]. Rikudo [1422]. ring [836, 1680, 1225, 2466, 2271, 1154, 364, 2045]. rings [222, 388, 192, 2290]. Ripple [1272]. Risk [152, 51, 2293]. RLS [440]. RNA [2487, 458, 2376]. rNMP [1261]. road [2129]. robber [2465, 1142, 1651, 1163]. robbers [540, 808, 1016, 537, 1149]. robot [726, 643]. Robots [311, 1559, 207, 306, 685, 2228, 1817, 683, 493, 729, 2216, 1479, 727, 2271, 1154, 1430, 2264, 816, 701, 2244, 1736, 1873]. Robust [612, 321, 2042, 841, 1748, 598, 1863]. robustly [11]. Robustness [114, 1384, 2414, 1926]. Role [1027, 1226, 1726, 683]. Roman [1435]. root [2482, 86]. rooted [458, 1002, 1838]. rotation [1511, 2406, 1421, 1157]. rotations [1000]. Rothschild [2019]. rough [1399, 720, 1237]. round [1049, 1119, 1129, 1367, 2062, 519, 1029]. round-trip [1029]. rounding [1387, 688]. Route [488]. routes [606]. Routing [2342, 1688, 1543, 858, 1672, 103, 1532]. row [1481]. rows [2338]. RSA [2481, 432, 2070, 1108, 568, 1180]. RSA-like [1108]. Rule [414, 423, 1944, 2060, 1175, 1631]. rules [77, 2163, 386, 2111, 1093, 81, 1401, 2032, 316, 2327, 2497, 1235, 1522]. Rumor [210, 855, 560, 36]. run [322, 721, 445, 705]. Running [598]. runs [1156]. Runtime [440, 715]. Russian [1994].

S [2103]. S-table [2103]. Saettone [1593]. Safe [716, 2250]. safety [2474, 1297]. sales [1653]. Salesman [1040, 1869]. salesmen [1735]. same [730, 2160]. Sample [1550, 75]. samples [601]. Sampling [819, 1300, 1938, 1961]. Sampling-based [819]. sandpiles [595]. sandwich [371]. Santa [535]. SAT [1426, 2104, 1150, 387, 1893, 1166, 601, 301]. satisfaction [132]. satisfactory [1394]. Satisfiability [1166, 1490, 1176, 2298, 1646]. satisfying [108, 1724]. Satnav [2203]. saturation [2086]. scaffold [1004, 744]. Scalability [1912]. Scalable [53]. Scale [165, 1912, 557, 1637, 1234, 1260]. Scale-invariant [165]. Scaling [2277, 355]. scapegoating [1815]. SCAR [632]. scattered [203]. scenario [2347]. scheduler [522, 120]. schedules [541]. Scheduling [1140, 1538, 563, 774, 1201, 2195, 2242, 41, 1458, 2069, 1532, 557, 1056, 902, 550, 2172, 551, 2410, 617, 852, 1031, 132, 2341, 34, 1649, 1655, 1791, 2427, 1177, 1037, 369, 119, 350, 1868, 1972]. Schelling [1053, 2160]. Schema [2023]. scheme [293, 1387, 193, 1050, 1120, 1130, 1743, 2077, 1584, 2466, 2493, 883, 1826, 288, 432, 480, 1304, 472, 1574, 2399, 749, 839]. schemes [524, 122, 1995, 224, 2342, 835, 850, 1208, 1704, 814, 444, 241, 1591, 581, 1424]. Schnyder [149]. Sci [591, 2212, 340, 2094, 339, 2379, 133, 1525, 129]. Science [2488, 905, 151, 312, 1935]. sciences [1094]. score [1939]. scoring [1045, 1115, 1125, 2398]. Scorpion [1191]. Scott [1932]. Scott-open [1932]. SDN [2085]. SEAMHR [1688]. search [1136, 715, 1801, 510, 374, 2118, 1772, 749, 2376, 2203]. searchable [2318]. searchers [220, 1148]. searches [1732]. Searching [824, 142, 121, 1148, 1836, 1938, 2405]. second [392, 1669, 104]. second-best [1669]. second-order [104]. secondary [458, 2376]. Secret [1822, 835, 1208, 1141, 529, 1424]. secretary [888, 1537]. sections [1468]. Secure [1552, 2036, 1211, 1276, 1464, 2449, 1225, 2419, 1161, 1304, 2062, 1859, 1495, 1994, 1861, 630, 453, 1105, 1688, 654]. Securely [1153]. Security [155, 1276, 12, 1805, 1672, 1331, 1637, 663, 1055, 1088, 1673, 2474, 187]. seed [1770, 517]. seeding [1110]. seeds [2439]. segment [754, 1325]. segments [1585, 1144, 1430, 3, 129, 928, 707, 1505]. Selected [823, 2186]. selection [1544, 1910, 419, 428, 72, 869, 1849, 1709, 2259, 237, 444, 2382]. selections [870]. selectors [100]. Self [2102, 2170, 995, 2267, 2384, 2320, 1571, 1874, 1748, 2011, 549, 1885, 1252, 587, 2022, 2079, 162, 443, 1010, 1960, 364]. self-adaptation [443]. Self-adjusting [2267]. Self-assembly [995, 1252]. self-bilinear [364]. self-bounding [162]. self-eliminations [1960]. self-intersecting [587]. self-intersection [587]. self-loop [549]. self-profit [2079]. self-stabilization [1874, 1748]. Self-stabilizing [2102, 2170, 2384, 2320, 1571, 2011, 1885, 2022, 1010]. Selfish [2301]. selling [1338, 370]. semantic [967, 764, 558, 491, 958]. Semantics [2189, 136, 391, 1094, 1339, 141, 1451, 497, 125, 1615, 2181]. semi [1977, 2439, 361, 720, 897, 1454, 351]. semi-Cayley [897]. semi-disjoint [361]. semi-lattices [1454]. semi-linear [720]. semi-matchings [351]. semi-uniform [2439]. semicomputable [1222, 1733]. semigroups [197]. semilattices [1448, 2231, 1100]. semiring [2477]. semirings [675, 943]. Semitotal [272, 1808]. senders [1010]. sensing [216, 2162]. sensitive [1663]. Sensitivity [514, 1630, 1943, 793, 1779]. Sensor [1688, 1557, 1182, 1544, 557, 204, 746, 1929, 1655, 1714, 1856, 1620]. sensor-based [1714]. sensors [846, 1147, 1864]. sentences [601]. separable [1588]. separating [2300, 660, 199]. separation [1769, 1954, 2148, 1279]. separations [2184]. separator [918]. separators [99, 2302]. September [447, 1080, 1089, 1096, 1131, 1647, 1664, 1681, 1701, 2153, 2167, 2179, 2190]. sequence [730, 227, 1939, 2445, 298]. sequences [789, 62, 821, 786, 19, 196, 921, 1802, 1554, 2354, 2445, 954, 1205, 1390, 57, 1776, 433, 1913, 236, 755, 202, 574, 707]. Sequential [997, 2152, 1239]. Sequentialization [1312]. serial [1746]. series [203, 1773, 2313, 2478, 2151]. series-parallel [203, 1773]. server [937, 1276, 1265, 492, 292, 1610]. server-aided [292]. servers [2225]. service [579, 1906, 1824, 1597]. Services [1912, 2215]. sessions [1295]. Set [1139, 1978, 2492, 1137, 2219, 2397, 931, 2212, 937, 2209, 2331, 1086, 867, 642, 2251, 1658, 1399, 671, 486, 624, 190, 2188, 649, 1281, 1784, 1196, 1812, 253, 631, 1267, 1585, 2096, 1756, 1747, 1721, 936, 1577, 1360, 510, 611, 1782, 2320, 540, 1695, 1430, 2028, 1926, 1354, 282, 3, 129, 1141, 2184, 2005, 29, 693, 1469, 1261, 517, 2182, 2412, 2068, 2016, 1596]. Set-constrained [1137]. Set-to-set [2492]. set-up [937]. set-valued [1747]. SETH [2202]. sets [2048, 2229, 2014, 1098, 1044, 1114, 1124, 923, 2262, 835, 734, 1036, 1977, 376, 1218, 641, 2081, 345, 1012, 1601, 1892, 1495, 804, 1237, 1842, 1759, 897, 1319, 928, 384]. setting [2449, 663, 2307, 476]. settings [1549, 310]. setwise [1362]. seven [700, 1508]. Several [1890, 2108, 2478, 126, 155, 2182]. SFTs [1487]. shallow [722]. shape [2332]. shaped [1882]. shapes [1886]. shared [933]. sharing [835, 1208, 1796, 1141, 1476, 1424]. Sharp [2108]. Shearer [1812]. shift [2073, 2379]. shifts [1684, 1771, 927, 776]. shop [937, 355, 41, 2341, 34, 1177]. Short [850, 1050, 1120, 1130, 988, 890, 895]. Shortest [640, 927, 885, 1562, 1742, 1337, 142, 1536, 930, 1870, 637, 240, 1632]. Shrinkage [479]. shuffle [93, 734, 478, 2031]. shuffle-cubes [478]. shuffled [1862]. shuffled-square-free [1862]. Siamese [1911]. sided [1288, 186, 1004, 776]. sides [77, 2163]. sieve [273]. sigma [1075]. sigma-tau [1075]. signal [1255]. signals [2370]. signature [1743, 612, 2466, 432, 663, 1533, 472, 2347, 2399]. Signatures [1482, 366, 1225, 762, 1161, 1209, 1181, 2088]. signcryption [2423, 1476, 839]. signed [1592, 1954]. silver [906]. sim [2020]. sim-width [2020]. similar [130, 618]. similarity [257]. Simon [2199, 2175, 2442, 1616]. Simple [642, 2407, 1509, 541, 578, 1387, 1173, 2308, 1328, 2238, 301, 118, 768, 1826, 1689, 252, 181, 2147, 1481, 1988, 2038]. simpletons [1412]. simplices [1733]. simplified [721, 2015]. Simulatable [2035]. simulate [1400]. simulating [968, 1238]. simulation [1737, 2344, 2285]. simulations [1233, 2468]. Simulink [1351]. Simulink/Stateflow [1351]. Simultaneous [1015, 228, 61, 1509, 212]. simultaneously [2164]. Single [1614, 1092, 937, 411, 2282, 833, 2012, 1164, 1649, 238, 681, 615, 894, 370, 279]. Single- [1614]. single-arm-gapped [238]. single-character [833]. Single-exponential [279]. single-minded [370]. single-processor [1649]. single-writer [411]. singleton [2099, 1488, 2328]. Singletons [1412]. singly [1749]. Singular [10]. sink [111, 1007, 1864]. sink-based [1864]. SINR [217, 1047, 1117, 1127]. siphons [2166]. site [2124]. site-directed [2124]. six [1875, 2021]. six-vertex [1875]. size [2229, 249, 1413, 208, 1251, 1252, 1883, 1254, 2126, 895, 1810, 875, 654, 2158, 282, 1731, 568, 603, 2091, 175]. sized [2466]. sizes [2267, 563]. skeleton [737]. Skew [378]. ski [1792]. skin [2321]. Skolem [2478]. skyline [1288]. Slanted [1255]. Sleep [1978]. Sleptsov [2356]. slide [1813]. sliding [1643, 1098, 303, 885]. slot [369]. slowdown [1021]. Small [863, 610, 361, 378, 671, 112, 1593, 1139]. Smallest [1775, 146, 288, 2272]. smart [488]. Smash [2398]. smooth [2256]. SMUAAKAP [1861]. snark [2374]. SNQ [863]. social [982, 1910, 2411, 32, 30, 1362, 2037, 1637, 1708, 1072, 31, 717, 560, 416, 425, 1534, 802, 1492, 949, 1636, 39, 517, 760, 855, 1746]. sofic [1771]. soft [2069, 1637, 2309]. Software [1912]. solid [96, 1217]. solid-metric [96]. Solitaire [1191]. Solitar [1249, 1487]. solution [1881, 1363, 2276, 1004, 335]. solutions [1423, 539, 2055, 270, 2404, 1469, 1571]. solve [1689]. solved [1074]. solves [2475]. Solving [2104, 1619, 618, 1900, 2207, 297, 2302, 1401, 1884]. Some [1107, 2419, 1402, 454, 1661, 1046, 1116, 1126, 1943, 1569, 1163, 1388, 2481, 768, 1601, 1373, 2070, 1106, 1899, 2377, 655, 2422, 1470, 2072]. sometimes [533]. sort [628, 1268, 1286, 2155, 1032, 1986, 1831]. Sorting [628, 2241, 2047, 2363, 940, 147, 1425, 1290, 1207, 1488]. sound [511]. source [1552, 1437, 1029]. source-wise [1029]. sources [610]. Space [2080, 2430, 2302, 637, 736, 192, 382, 777, 1562, 1539, 182, 235, 756, 1440, 2221, 2325, 2103, 1411, 2184, 513]. Space-efficient [2080, 637, 736, 192, 756]. spaces [159, 1909, 2071, 2231, 45, 393, 958, 1444, 1933, 720, 2166]. spacing [1969]. spanner [546, 1948, 1571]. spanners [1393, 1461, 2383, 1310, 1029, 765]. Spanning [474, 854, 1693, 450, 1971, 1328, 2094, 1904, 521, 2022, 146, 123, 1905, 1588, 1880, 1241, 1237, 2146]. Sparse [173, 2229, 1378, 2333, 1902, 294, 917, 1006, 981, 332]. sparsest [546]. Sparsification [1310]. spatial [1220, 2187]. spatio [1244]. spatio-temporal [1244]. Special [215, 1717, 157, 966, 1980, 1529, 1063, 2186, 742, 824, 342, 1935, 1998, 169, 655, 1628, 1530, 1716, 2400, 438]. Specification [777, 2198, 1355, 1785, 799]. specifications [1321]. Spectral [1250]. spectrum [248, 2086, 703]. Speech [1857]. speed [1981, 1791]. speed-up [1791]. Speeding [916, 444]. spherical [749]. spies [493]. spikes [863, 14]. Spiking [2132, 879, 2497, 14, 1607, 1846, 2109, 1235, 1631, 1522, 2287]. spined [1372]. spiro [1860]. spite [2102]. splicing [317]. split [174, 459, 2331, 624, 403, 1371]. split-star [403]. splits [607]. splittable [2257]. sports [1037]. spreading [210]. Spy [1598]. squad [1255]. Square [2219, 1256, 1558, 1862, 376, 911, 872, 1354, 86, 1727]. square-free [1558, 872]. squared [1105, 1884]. squared-exponential [1105]. Squarefree [889]. squares [93, 698, 1585, 2031, 3, 129]. stabbed [2447]. Stability [673, 2321, 768, 932]. stabilization [1874, 1748]. stabilizing [2102, 2011, 1885, 2022, 1784, 2170, 1010, 2384, 2320, 128, 1571]. Stable [1409, 2068, 1532, 1703, 1834, 1083, 618]. stably [2071, 393, 1844]. Stackelberg [1898]. Stacks [1040]. stage [2123, 2440, 1538, 2222, 51, 1031, 1782, 1072, 614, 1574, 119, 350, 748]. Stagnation [1945]. Stand [1804]. standard [1482, 1370, 1225, 2466, 1747, 913, 1674, 979]. Star [924, 2169, 401, 611, 403, 1603, 487, 1986, 1543, 2227]. Star-critical [924]. stars [1650]. State [2401, 2120, 2364, 211, 1440, 443, 873, 2142, 958, 2124, 1783, 2003].

State-based [2364]. state-dependent [1783]. state-space [1440]. Stateflow [1351]. static [1098]. stationary [2455, 2293]. stations [55, 1662, 1666]. Statistical [1997, 1244, 2082, 1440]. statistics [2145]. status [789]. stay [717]. steady [443]. Steiner [938, 544, 763, 1183, 480, 477, 704]. Step [658, 411, 1657]. Step-by-step [658]. Step-optimal [411]. Steps [2287, 1069]. Stick [1993]. Stochastic [11, 1783, 1270, 404, 1782, 1095, 153, 703]. stochastic-cascading [404]. stone [570]. stop [1507, 1294]. stopping [1469]. storage [1353, 1769, 1837, 1260, 1476]. stored [1220]. stories [886]. storing [524]. straight [740, 1221]. straight-line [740, 1221]. strangers [2160]. Strassen [22]. strategic [280]. strategically [223]. strategies [996, 2471, 442, 1872, 130, 250, 443, 541, 632, 180, 1928, 15, 1651]. strategy [441, 1637, 728, 800, 1072, 1655, 2246, 1931]. stream [1443]. Streamable [137]. Streaming [2077, 2249, 975, 1781, 1918, 2319]. streams [2197, 1917, 2325, 92, 2298]. strength [1786]. stretch [1902]. stretching [1789]. strict [2407]. strictly [992]. strikes [506]. String [1950, 1645, 2354, 1888, 656, 583, 244, 1900, 865, 2104, 927, 2202, 1737, 239, 718, 610, 2175, 708, 238, 236, 975, 1522]. string-processing [244]. Stringology [1133]. Strings [1798, 2286, 731, 1282, 2283, 761, 75, 601]. strip [457]. strips [101]. Strong [815, 1514, 1553, 1775, 1444, 1242, 2155, 657, 2292, 54]. Strongly [1565, 679, 1389, 241, 107]. Structural [1833, 1752, 617, 2275, 823, 206, 1742, 764, 572, 1919, 538, 1613, 2497, 14, 215, 2161]. Structure [697, 635, 152, 402, 1952, 1427, 1054, 382, 1372, 1567, 2078, 1751, 2418, 2460, 1920, 1174, 1258, 181, 1433, 284, 1141]. structures [1452, 1763, 619, 1984, 1950, 985, 2183, 455, 131, 1797, 796, 518, 874, 458, 883, 1220, 1092, 655, 2376]. study [817, 2407, 1440, 1298, 1849, 2073, 2379, 104]. sturdy [1640]. Sturmian [1795, 944]. style [266]. sub [1971, 1596, 2202, 801]. sub-exponential [1596]. sub-linear [801]. sub-quadratic [2202]. subbases [2071]. subclasses [439, 2139, 2315, 1601, 1106, 1899, 885]. Subcomputable [1204]. subcontinua [1222]. subcubic [58, 641]. subdivision [552]. subdivisions [2086]. subgraph [1670, 2094, 1904, 146, 33, 727, 520, 1138, 1903, 688]. subgraph-aware [33]. subgraphs [1990, 343, 123, 1309, 1588, 540, 1037]. subgroup [1550]. subgroups [2273]. subject [2077, 1709, 1186, 1774, 904]. Sublinear [2055, 1720, 679, 809, 2067]. Sublinear-time [1720]. submatrix [228]. Submodular [1917, 2123, 385, 2282, 1223, 2037, 1595, 1706, 553, 1913, 29, 2396, 1186, 2330, 1781, 709, 1774, 1537, 1783, 1918, 2328, 987, 748, 2261]. submodular/linear [1595]. submodularity [1568, 745, 1782, 709, 1774]. submonoids [1336]. Subnetwork [1286]. subposets [1445]. subregular [2176]. Subroutines [80]. subsequence [526, 1696, 1576, 2103, 296, 755]. subsequences [1510]. Subset [282, 2209, 2331, 1348, 108, 491, 1704, 349, 2219]. subset-sums [1704]. subsets [2273, 1887, 1761]. subshifts [719]. subspace [819]. substitutions [2230, 944]. substring [833, 871, 1433]. substring-substring [1433]. substrings [2451, 637, 2253]. substructure [1190, 697, 402, 1054]. subsystem [327]. subtraction [2343]. subtrees [2372, 2395, 1227]. Subversion [366, 827]. Subversion-resilient [366]. subwords [2337]. Succinct [1660, 1172, 1646]. such [336]. Sudoku [536]. suffice [191]. sufficient [954, 673]. sufficient-completeness [673]. Suffix [1587, 2286, 757, 1057, 1737, 1836, 269, 736]. Suffix-connected [1587]. suffixes [864]. sum [831, 734, 1379, 1443, 349, 707]. Sumcheck [1610]. Sumcheck-based [1610]. summaries [764]. summation [943]. sums [1374, 108, 1151, 1704, 1930]. Super [1174, 2204, 2217, 577, 1589]. Superfast [1961]. supergrid [2030]. superimposed [131]. supplier [2316, 1832]. supply [107]. support [2389, 1194]. supporting [1674]. sure [670]. surely [2446]. surfaces [1414, 481]. surroundings [81]. surveillance [1662]. survey [1254]. Survivability [1544]. sustainable [1688]. SVM [1755]. swap [787, 577, 1741, 1952, 1897, 1054]. swaps [1834]. swarm [1736]. swarms [2011]. Sweep [899, 1666, 484, 1663]. sweeping [2383]. switches [2091]. Sybil [2150]. Sybil-proof [2150]. symbol [1511, 2145]. symbolic [376, 86, 1441]. Symmetric [1202, 1285, 1511, 2406, 2183, 2206, 1757, 2029, 1157, 2318, 1790]. symmetric-persistent [2029]. symport [81, 1401]. symport/antiport [81]. synapses [1846, 2132, 1632, 1235]. synchronicity [683]. synchronisation [1255]. synchronism [1630]. Synchronization [124, 1889, 1147, 1193, 213, 1631]. synchronized [2441]. synchronous [2188]. Syntactic [1075, 169]. Syntax [2117, 558]. Synthesis [75, 601, 1321, 1188]. Synthesising [1406]. Synthesizing [1755]. system [2131, 2055, 261, 271, 1353, 1707, 1860, 800, 1623, 2436, 1774, 1424, 1301, 37, 1970, 511, 50]. system-level [271]. Systems [2367, 1063, 1406, 77, 76, 1524, 2152, 1236, 2102, 79, 2334, 1022, 1629, 1239, 1064, 1, 340, 326, 1065, 1850, 1020, 1066, 2198, 2104, 1399, 2393, 193, 1751, 1883, 2418, 716, 78, 1733, 446, 262, 2451, 1253, 1357, 1223, 775, 2111, 1067, 360, 502, 968, 1421, 2096, 337, 718, 1607, 2110, 1846, 1353, 1769, 788, 464, 1068, 1069, 80, 1400, 2132, 2365, 2477, 1714, 1906, 81, 82, 1401, 2032, 316, 2474, 317, 1731, 2184, 703, 334, 1260, 957, 1092, 1082, 2021, 992, 1608, 1454, 616, 879, 2497, 1351, 14, 1632, 2109, 335, 1235, 1631, 2378, 599, 1522, 1202, 2287, 863].

table [2103, 384]. tableau [266]. tableaux [1747]. tadpole [539]. tail [1268]. takes [219]. Taming [1342, 1973]. Target [517, 1981, 1892, 879, 39]. targeting [231]. targets [1010]. Tarski [339, 2]. task [40, 800, 1655, 1824, 2429]. task-based [1655]. tasks [1458]. tau [1075]. taxonomy [671]. TCS [1998]. Team [532, 1298, 497, 758]. technique [835, 373, 1707, 1425]. techniques [1552, 1910, 1672, 878, 334, 1261]. Tell [2375]. temperature [995, 770]. temperature-1 [995]. Temporal [98, 99, 2081, 799, 1244, 255, 260, 2407, 672, 2087, 2137, 2276, 497, 1321, 1622, 357, 509, 898, 1970, 511]. terminal [1228]. terminating [2446]. termination [673]. terms [23, 2097, 1897]. ternary [1814, 740, 986]. terrain [818]. terrains [659]. tessellation [16, 817, 1878]. test [19, 1370, 1219, 856, 1344, 1674]. tester [1979]. Testers [2016]. Testing [2488, 1014, 843, 295, 2475, 1607, 2110, 1015, 2348, 1506]. tests [295]. Tetris [2417]. text [777, 234, 1938, 1636]. th [1546, 513]. their [1046, 1116, 1126, 259, 942, 2476, 100, 1802, 309, 80, 1835, 2337, 2290, 2166, 1453]. Theor [591, 2212, 2094, 339, 2379, 1525, 129]. theorem [1957, 2263, 2478, 1678, 720, 1608, 1441, 2003, 2019, 22]. Theoret [340, 133]. theoretic [1987, 486, 2041, 92, 1803]. Theoretical [312, 905, 1816, 1980, 981, 2488, 2400, 404, 1462, 1823, 1717, 1935]. Theories [1628, 24]. Theory [824, 821, 323, 438, 1094, 1658, 1829, 1399, 671, 675, 2432, 2413, 629, 2178, 500, 117, 1697, 1932, 1906, 104, 1882, 1141, 2377, 2016]. theory-based [117]. thesis [1299]. thickness [1001]. thin [995]. things [1708, 1860]. Things-spiro [1860]. Three [2394, 386, 354, 565, 34, 2039, 2028]. three-machine [34]. Threshold [1919, 1805, 926, 2445, 2456, 1812, 718, 883, 1035, 1166, 1632, 1505]. Threshold-based [1919, 2456]. thresholds [2378]. Throttling [909]. Thue [1750, 2353]. tie [386]. tie-breaking [386]. Tight [757, 1943, 1596, 162, 1513, 1832, 663, 1796, 294, 1746, 1337, 537, 344, 1969]. tighter [1464]. Tightly [1225, 1304, 453, 1161]. tightly-secure [1161]. tiling [1731]. tilings [305]. TIM [1335]. Time [76, 729, 1241, 1663, 599, 382, 1853, 1393, 1874, 662, 79, 735, 1471, 937, 1562, 1734, 1436, 598, 300, 266, 255, 1339, 1044, 1114, 1124, 2392, 537, 585, 1817, 248, 1458, 2430, 2069, 2002, 354, 882, 322, 721, 1056, 2202, 2077, 1223, 2183, 272, 1720, 2221, 1218, 2366, 1007, 1221, 754, 567, 1216, 247, 288, 263, 445, 2012, 1012, 1580, 251, 1437, 1689, 2151, 44, 801, 1695, 240, 1599, 809, 3, 129, 929, 703, 302, 128, 107, 709, 1574, 644, 369, 335, 1657, 2364]. time-bounded [1853]. time-branching [703]. time-critical [1458]. Time-energy [729]. Time-free [599]. Time-freeness [76]. time-step [1657]. time-varying [567]. time-windows [882]. timed [254, 1673, 258]. Timeline [260, 1339, 309]. Timeline-based [260, 1339, 309]. times [2266, 563]. timestamps [1885]. Tipsy [2465]. tissue [2032]. Toeplitz [1542]. together [1294]. toggle [1421]. Token [216]. tokens [885]. tolerability [2058]. tolerance [2160, 1174, 635, 2357, 492, 853, 2169, 1084, 1952, 474, 1986, 1427, 2045, 1993]. tolerances [1779]. tolerant [1392, 2011, 1102, 1498, 338, 1885, 295, 2170, 762, 1491, 2029, 2155, 701, 657, 1168]. toll [2454]. tomography [1513, 1815]. tools [1261]. top [1645, 2253, 284, 1639]. top- [1645, 2253, 284]. top-down [1639]. topic [40]. Topics [1717]. Topological [1684, 2280, 932, 628, 684, 185, 1402, 776, 1761, 692]. topologies [836, 1265]. topology [850, 1571]. tori [706]. toroidal [808]. torus [61, 2216]. Total [1410, 1182, 1901, 912, 1056, 2077, 1036, 1977, 2297, 2237, 545, 132, 345, 1030, 696, 1495, 2036]. Toulmin’s [73]. tour [2312, 1342]. touring [925]. tournament [2404]. tournaments [2209]. tower [273]. trace [719, 2362, 1842, 1441]. trace-and-revoke [2362]. traceability [2013]. Tracking [606, 640, 1752]. Tracks [531, 1320]. Tractability [1893, 276, 523, 918, 1134, 102]. tractable [640, 1990, 500, 228]. tradeoff [325, 1201]. tradeoffs [729]. trading [784]. traffic [347]. trails [1580]. Train [1320]. trained [1708]. Trains [2136, 1320]. trajectories [1300]. Trajectory [1929, 746]. transaction [1653]. transactional [2242, 2427]. Transactions [1065]. transcriptional [2487]. transducers [967, 699, 2097, 2027, 2142, 975]. transductions [137]. transfer [1213, 1847]. transferable [1852]. transform [179, 1057, 235]. Transformation [1093, 1091, 699, 984, 582, 1717]. transformations [434, 1314, 1886, 1770, 1511]. transforms [514]. transients [2485, 1232]. transition [302]. transitive [1943]. Translating [198]. translation [967]. translational [481]. translations [2097]. translucent [2424]. transmission [1654, 1994]. transmitters [303]. transportation [2341]. Transporting [1445]. transposition [661]. transpositions [1290]. transversal [291, 348]. transversals [1348]. Trap [2166]. trapdoor [1088]. trapping [1469]. traps [374]. Traveling [1040, 1735, 1869]. Traveller [1872]. travelling [1675]. traversal [216]. traversals [23]. Treasure [726, 1547]. Tree [455, 35, 1392, 1087, 757, 2225, 1461, 967, 266, 1949, 1328, 1136, 2418, 851, 2238, 2022, 221, 2263, 1514, 699, 2126, 1780, 269, 850, 910, 2097, 608, 2302, 548, 611, 771, 686, 125, 797, 634, 1220, 1241, 1074, 681, 975, 435, 1639, 487, 102, 1845, 1134, 2004]. tree-coloring [771, 797]. tree-depth [1949]. tree-style [266]. tree-to-graph [967]. tree-walking [699]. treedepth [1499]. trees [789, 845, 105, 1684, 1206, 792, 1109, 2329, 1381, 1705, 1572, 938, 1337, 679, 564, 2242, 2254, 546, 183, 521, 740, 85, 1281, 2239, 1513, 218, 2294, 754, 458, 737, 1183, 1870, 713, 1880, 736, 909, 1772, 519, 107, 435, 615, 1195, 1496, 1227, 1142, 1651, 2146]. treewidth [279, 2395, 184, 938, 1377, 1003, 771]. triangle [700, 1044, 1114, 1124, 173, 818, 1922, 661, 1551]. triangle-free [1922]. triangles [1077]. triangular [2003]. triangulations [477, 547]. Trichotomy [788]. trie [1156]. tries [1268, 2126]. trigger [632]. trip [1029]. triples [170, 57, 468]. triviality [377]. tromino [305]. true [47, 1077]. truncated [385]. trusted [1856]. truth [915, 384]. truth-table [384]. truthful [2431, 2370]. Truthfulness [1726]. TSP [1463, 1849]. Tumble [1963]. tunable [283]. tuned [1714]. tunes [82]. Turing [2356, 1346, 1400, 1373, 2340, 1963, 1181]. Turing-Complete [1963, 2356]. turns [1342]. Twin [2127, 1459, 2457, 1264]. Twin-distance-hereditary [2127]. twisted [1667, 1174]. Two [230, 937, 122, 2123, 2440, 78, 901, 51, 2164, 1523, 1072, 973, 1951, 2049, 2409, 1368, 454, 524, 2225, 1343, 1538, 399, 729, 1056, 784, 2358, 916, 1031, 2493, 1794, 186, 970, 247, 1010, 643, 1400, 1005, 1782, 983, 1004, 2340, 150, 3, 129, 974, 755, 1574, 1828, 2347, 603, 400, 1243, 119, 350, 1228, 748, 1868, 1140]. Two-dimensional [230, 2164, 973, 150, 974]. Two-disjoint-cycle-cover [1951, 2049]. two-factor [2493]. two-interval [1343]. two-level [2340]. Two-machine [937, 1056]. two-party [983, 2347]. two-robot [643]. two-sided [186, 1004]. Two-stage [2123, 2440, 51, 1072, 1538, 1031, 1782, 1574, 119, 350, 748]. two-terminal [1228]. Two-way [901, 1523, 784, 970]. Type [247, 2041, 1684, 1, 340, 2413, 498, 394, 283, 2437, 1108, 650]. type-assignment [498]. Type-theoretic [2041]. Type-two [247]. types [249, 863, 143, 716, 2192, 2021, 2181, 1355, 283]. typical [432].

UAV [1675, 746, 1929]. UAVs [1663]. UFLP [513]. Ulam [191]. ultra [1374]. ultra-wide [1374]. ultrametric [419, 428]. UMU [2476]. unary [1313, 2142]. unbordered [731]. Unbounded [2491, 2449, 762]. uncertain [1409]. Uncertainty [1399, 1270, 940, 707]. unchained [410]. unconditionally [1994]. unconstrained [165]. uncooperative [121]. Uncut [688]. undecidability [1319, 532]. undirected [926]. unexpected [2172]. unicast [1591]. unicycle [1195]. unicyclic [661, 1588]. unidirectional [781, 654, 388, 192]. unification [675]. Unified [1351, 898, 800, 356, 1970, 511, 2182]. Uniform [534, 1224, 1448, 324, 439, 954, 2111, 2439, 1047, 1117, 1127, 328, 1826, 1103, 2142, 1844, 192, 720]. uniformity [1766]. uniformly [1228, 120]. Unifying [24, 1760, 1094]. uninitialized [722]. Union [1492, 1196]. Unique [60, 1555, 637, 90, 1283, 1861, 2375]. uniquely [1073, 58]. unit [1507, 651, 41, 2420, 182, 1920, 1921, 1267, 1721, 625, 490, 1430, 1354, 602, 2036, 1189]. unit-disk [1920]. univariate [721, 445]. Universal [239, 1049, 1119, 1129, 603, 513, 1473, 434, 2199, 1205, 2459]. universality [2178, 2468]. universally [351]. Universe [283]. unknown [875, 1573, 1148]. unless [2202]. unordered [1087]. unpaired [1432, 375]. Unprovability [1756]. unranked [967]. unrelated [814, 181, 372]. unreliable [2203, 1784, 1654, 118, 1837]. unsatisfiable [2476]. unsaturated [458]. unsupervised [1058]. Untangling [2137]. untyped [251, 650]. Updatable [2318, 2285, 50]. updates [754]. updating [882]. upgrades [2085]. upgrading [2347]. Upper [1329, 1268, 1818, 105, 757, 1995, 1150, 715, 1056, 2445, 56, 1038, 1290, 1596]. Upward [2475]. urban [488]. use [72, 2344]. user [971, 1709, 1861, 285, 476]. user-defined [971]. users [39, 1906, 517]. Using [1808, 124, 1912, 1552, 1544, 1000, 628, 1662, 1399, 1050, 1120, 1130, 994, 1076, 2420, 40, 1694, 659, 71, 73, 2111, 1619, 916, 895, 770, 2302, 600, 1707, 1857, 1860, 1911, 1708, 433, 1673, 717, 1623, 75, 1671, 334, 2437, 1636, 2327, 1710, 749, 2091]. utility [2089, 1852, 325]. UTP [2189]. UTP-CSP [2189]. UTVPI [1082].

validation [70]. valleys [1843]. Value [152, 2276, 10, 862]. Value-at-Risk [152]. valued [1747, 602]. values [2089, 844, 304]. vanilla [184]. variable [25, 1140, 319, 10, 1240, 1153, 1622, 1157]. variable-length [1140]. variables [2406, 2478]. variance [164]. variant [2152, 2131, 2350, 1346, 301, 1473, 1180]. Variants [1524, 1489, 1140, 387, 242, 2481, 2030, 1416, 2070, 974]. variational [164, 1923]. Varieties [1613]. Variety [2194]. various [2042]. varying [2077, 567]. vector [1194]. vectors [349]. vehicle [488, 333]. vehicles [980]. verifiable [1049, 1119, 1129, 2035, 1141, 1424]. Verification [1850, 2198, 2064, 2423, 829, 1443, 1141, 1300, 154, 1351, 335]. verified [716]. verifier [1209]. verifier-local [1209]. verifiers [2221]. Verify [356]. Verifying [600, 201, 2474]. version [1471, 1533]. versions [184]. versus [2366, 541, 323, 876]. Vertex [1139, 139, 624, 1459, 812, 1865, 1032, 459, 1190, 2209, 2331, 1548, 2306, 1582, 2317, 700, 2251, 2078, 1417, 1751, 1590, 489, 1076, 1749, 941, 89, 146, 360, 1364, 1499, 1875, 2457, 773, 2107, 936, 1577, 510, 948, 1420, 1155, 1001, 2448, 2006, 1951, 2249, 282, 440, 2049, 1551, 2005, 693, 1535, 1828, 581, 2061, 1162]. vertex-capturing [1582]. vertex-critical [2078]. vertex-disconnection [1590]. vertex-disjoint [1551]. Vertex-pancyclicity [1032, 773]. Vertex-weighted [139, 1548, 2006]. vertices [858, 1775, 686, 2028, 2244]. very [2295]. via [1581, 2332, 570, 546, 113, 9, 2449, 2445, 1887, 1891, 1196, 673, 1036, 631, 1961, 2265, 195, 1683, 1769, 216, 1035, 1755, 1501, 356, 1928, 1697, 213, 1485, 2377, 1956, 1505, 688, 2272]. viability [408]. Vickers [2071]. video [532]. view [866, 99, 708, 1404, 1075]. views [2293]. VIKOR [1671]. violation [2339]. Viral [30]. virus [2177]. visibility [1559, 2243, 2194, 727, 884, 701, 1142, 1651]. visible [858, 699]. vision [1076]. visit [1508]. visiting [1430]. visual [883, 1911]. Visualizing [307, 473]. voltage [355]. volume [454, 658]. volume-regular [658]. Voronoi [1079, 1047, 1117, 1127, 2147, 519]. vote [1843]. voting [1210, 276, 678, 2138, 1623, 2042]. vs [1539, 212, 2321, 1906]. vulnerabilities [2222]. vulnerable [732].

Wagner [2116]. Wait [1964]. Wait-free [1964]. waiting [567]. WALCOM [2186]. walk [240]. walking [699, 2252]. walks [2465, 302]. water [2241]. Watson [2451]. way [769, 901, 784, 970, 1523, 725, 2117]. wayness [769]. Weak [384, 1737, 958, 1154]. Weakly [2231, 2454]. web [1710, 1708]. Wegner [1633]. Weight [2068, 1044, 1114, 1124, 131, 1782, 520]. Weighted [1690, 1801, 1468, 2382, 441, 139, 1548, 1381, 1728, 2231, 546, 183, 738, 1335, 1514, 2239, 2193, 1085, 1685, 1313, 2006, 440, 705, 1818, 1866]. weighted-TIM [1335]. weighting [1379]. Weights [2313, 2132, 415, 424]. Weisfeiler [1271]. welfare [886, 1746]. well [1725]. wheel [1459, 697, 1652, 687]. Wheeler [235, 1057]. whether [1734, 2374]. Which [1165, 1524, 1794, 256]. while [1430]. Whither [136]. Who [533, 2375]. Whom [209]. wide [1374]. width [457, 378, 354, 272, 778, 2020, 2459]. Wiener [1108]. Wiener-type [1108]. Wildcarded [1345]. wildcards [1545]. will [2375]. win [1651]. windability [2448]. window [1643]. windowed [1412]. windows [1098, 882]. winning [2471, 1472]. wirelength [1372]. Wireless [220, 1544, 1532, 729, 100, 798, 1335, 1921, 850, 551, 216, 643, 746, 1929, 1655, 1856, 1620, 1571, 1688, 1557]. wise [833, 1029]. withheld [2411]. within [1563, 1941, 995, 809]. without [2219, 55, 827, 1276, 489, 1680, 1775, 1373, 1865, 388, 1077, 1371, 2339]. witness [2196, 533]. witnesses [533]. WK [1427]. WK-recursive [1427]. Wolfram [1944]. word [1562, 1734, 1448, 1374, 585, 1256, 270]. Words [1998, 1750, 689, 1643, 62, 1793, 1558, 287, 1644, 1642, 1814, 1862, 171, 585, 923, 834, 1285, 698, 196, 1542, 2199, 1200, 911, 872, 1497, 1835, 2353, 1795, 150, 1257, 889, 2326, 2337, 172, 804, 2031, 1283, 772, 1510, 467, 466]. work [1691]. Working [13]. Works [2222]. Workshop [1133, 1529]. Workshops [2186]. world [769, 1637]. worst [1874, 1852]. worst-case [1874, 1852]. worth [397]. WoT [1636]. write [494, 1137]. writer [411]. writes [1563]. wrt [2230]. WSN [1688]. WSNs [1672].

Xd [198]. Xd-C [198]. XML [699, 394]. XTR [1424].

Yao’s [46]. yield [2263]. Yuxi [1935]. zany [1570]. Zaslavsky [1592]. ZDDs [91]. Zeckendorf [831]. Zero [1486, 2491, 2128, 2449, 195, 1164, 2384, 1272, 536, 1997]. zerodegrading [2384]. Zero-freeness [1486]. zero-knowledge [2128, 1164, 1272, 536, 1997]. zeroing [2358, 335]. zeroness [19]. zeros [1267]. Zhang [336]. Ziv [1497]. ZKP [2171]. zoo [1570]. ZX [1294]. ZX-calculus [1294].

References

Blanqui:2002:IDT

[1] Frédéric Blanqui, Jean-Pierre Jouannaud, and Mitsuhiro Okada. Inductivedata-type systems. Theoretical Computer Science, 272(1–2):41–68, February 6, 2002. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                              //www.elsevier.com/gej-ng/10/41/
                                              16/222/27/29/abstract.html. See

corrigendum [340].

Dang:2018:CET

[2] Chuangyin Dang and Yinyu Ye. On the complexity of an expanded Tarski’s fixed point problem under the componentwise ordering. Theoretical Computer Science, 732(??): 26–45, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0304397518302433. See

erratum/correction [339].

Sadhu:2019:LTA

[3] Sanjib Sadhu, Sasanka Roy, Subhas C. Nandy, and Suchismita Roy. Linear time algorithm to cover and hit a set of line segments optimally by two axis-parallel squares. Theoretical Computer Science, 769(??): 63–74, May 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306303. See

corrigendum [129].

Lazard:2019:ELR

[4] Sylvain Lazard, William J. Lenhart, and Giuseppe Liotta. On the edgelength ratio of outerplanar graphs. Theoretical Computer Science, 770 (??):88–94, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306078. See

corrigendum [133].

Andres:2019:OCG

[5] Stephan Dominique Andres, Melissa Huggan, Fionn Mc Inerney, and Richard J. Nowakowski. The orthogonal colouring game. Theoretical Computer Science, 795(??):312– 325, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304499. See

corrigendum [591].

Anonymous:2020:PJa

[6] Anonymous. Pages 1–192 (1 January 2020). Theoretical Computer Science, 801(??):1–192, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBa

[7] Anonymous. Editorial Board. Theoretical Computer Science, 801(??): ii, January 1, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307339.

Kari:2020:Ea

[8] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 801(??):iii, January 1, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307364.

Doerr:2020:OPC

[9] Benjamin Doerr, Carola Doerr, and Jing Yang. Optimal parameter choices via precise black-box analysis. Theoretical Computer Science, 801(??): 1–34, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304013.

Gyongyosi:2020:SVD

[10] Laszlo Gyongyosi. Singular value decomposition assisted multicarrier continuous-variable quantum key distribution. Theoretical Computer Science, 801(??):35–63, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519304645.

Cappelletti:2020:SCR

[11] Daniele Cappelletti, Andrés Ortiz-Muñoz, David F. Anderson, and Erik Winfree. Stochastic chemical reaction networks for robustly approximating arbitrary probability distributions. Theoretical Computer Science, 801(??): 64–95, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930502X.

Boyer:2020:CSA

[12] Michel Boyer, Rotem Liss, and Tal Mor. Composable security against collective attacks of a modified BB84 QKD protocol with information only in one basis. Theoretical Computer Science, 801(??): 96–109, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305031.

Doerr:2020:WPB

[13] Benjamin Doerr and Weijie Zheng. Working principles of binary differential evolution. Theoretical Computer Science, 801(??):110–142, January 1, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305213.

Yang:2020:SNS

[14] Qian Yang, Bo Li, Yue Huang, Hong Peng, and Jun Wang. Spiking neural P systems with structural plasticity and anti-spikes. Theoretical Computer Science, 801(??):143– 156, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305304.

Shir:2020:CHR

[15] Ofer M. Shir and Amir Yehudayoff. On the covariance-Hessian relation in evolution strategies. Theoretical Computer Science, 801(??): 157–174, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305468.

Abreu:2020:GTC

[16] A. Abreu, L. Cunha, C. de Figueiredo, L. Kowada, F. Marquezino, D. Posner, and R. Portugal. The graph tessellation cover number: Chromatic bounds, efficient algorithms and hardness. Theoretical Computer Science, 801(??): 175–191, January 1, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305572.

Anonymous:2020:PJb

[17] Anonymous. Pages 1–196 (8 January 2020). Theoretical Computer Science, 802(??):1–196, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBb

[18] Anonymous. Editorial Board. Theoretical Computer Science, 802(??): ii, January 8, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307418.

Bergstra:2020:CCP

[19] J. A. Bergstra and C. A. Middelburg. On the complexity of the correctness problem for non-zeroness test instruction sequences. Theoretical Computer Science, 802(??): 1–18, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302130.

Corradini:2020:AGR

[20] Andrea Corradini, Dominique Duval, Rachid Echahed, Frédéric Prost, and Leila Ribeiro. Algebraic graph rewriting with controlled embedding. Theoretical Computer Science, 802(??): 19–37, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303871.

Cruz-Filipe:2020:CMC

[21] Luís Cruz-Filipe and Fabrizio Montesi. A core model for choreographic programming. Theoretical Computer Science, 802(??):38– 66, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304311.

Zhou:2020:STQ

[22] Li Zhou, Shenggang Ying, Nengkun Yu, and Mingsheng Ying. Strassen’s theorem for quantum couplings. Theoretical Computer Science, 802(??): 67–76, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305225.

Blum:2020:ELT

[23] William Blum. Evaluating lambda terms with traversals. Theoretical Computer Science, 802(??):77– 104, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305316.

Foster:2020:UTR

[24] Simon Foster, Ana Cavalcanti, Samuel Canham, Jim Woodcock, and Frank Zeyda. Unifying theories of reactive design contracts. Theoretical Computer Science, 802(??):105– 140, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305614.

Bednarczyk:2020:OVL

[25] Bartosz Bednarczyk. One-variable logic meets Presburger arithmetic. Theoretical Computer Science, 802(??): 141–146, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305791.

Castiglioni:2020:PDC

[26] Valentina Castiglioni and Simone Tini. Probabilistic divide and congruence: Branching bisimilarity. Theoretical Computer Science, 802(??):147– 196, January 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305882.

Anonymous:2020:PJc

[27] Anonymous. Pages 1–234 (10 January 2020). Theoretical Computer Science, 803(??):1–234, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBc

[28] Anonymous. Editorial Board. Theoretical Computer Science, 803(??): ii, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307741.

Shi:2020:BAM

[29] Yishuo Shi, Yingli Ran, Zhao Zhang, and Ding-Zhu Du. A bicriteria algorithm for the minimum submodular cost partial set multi-cover problem. Theoretical Computer Science, 803(??): 1–9, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301513.

Gao:2020:VMO

[30] Chuangen Gao, Hai Du, Weili Wu, and Hua Wang. Viral marketing of online game by DS decomposition in social networks. Theoretical Computer Science, 803(??):10– 21, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301537.

Liu:2020:PMP

[31] Bin Liu, Xiao Li, Huijuan Wang, Qizhi Fang, Junyu Dong, and Weili Wu. Profit Maximization problem with Coupons in social networks. Theoretical Computer Science, 803(??): 22–35, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301549.

Chen:2020:RAP

[32] Tiantian Chen, Bin Liu, Wenjing Liu, Qizhi Fang, Jing Yuan, and Weili Wu. A random algorithm for profit maximization in online social networks. Theoretical Computer Science, 803(??): 36–47, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930194X.

Han:2020:DRK

[33] Xiao Han, Chunhong Zhang, Chenchen Guo, Yang Ji, and Zheng Hu. Distributed representation of knowledge graphs with subgraph-aware proximity. Theoretical Computer Science, 803 (??):48–56, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301999.

Liu:2020:AAT

[34] Longcheng Liu, Yong Chen, Jianming Dong, Randy Goebel, Guohui Lin, Yue Luo, Guanqun Ni, Bing Su, Yao Xu, and An Zhang. Approximation algorithms for the three-machine proportionate mixed shop scheduling. Theoretical Computer Science, 803(??): 57–70, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303858.

Yu:2020:NLR

[35] Wei Yu, Zhaohui Liu, and Xiaoguang Bao. New LP relaxations for minimum cycle/path/tree cover problems. Theoretical Computer Science, 803(??): 71–81, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304086.

Fang:2020:GRB

[36] Qizhi Fang, Xin Chen, Qingqin Nong, Zongchao Zhang, Yongchang Cao, Yan Feng, Tao Sun, Suning Gong, and Dingzhu Du. General Rumor Blocking: an efficient random algorithm with martingale approach. Theoretical Computer Science, 803(??): 82–93, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304116.

Zhang:2020:PAA

[37] Chaoli Zhang, Jiapeng Xie, Fan Wu, Xiaofeng Gao, and Guihai Chen. Pricing and allocation algorithm designs in dynamic ridesharing system. Theoretical Computer Science, 803(??): 94–104, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304153.

Lv:2020:HAM

[38] Hongtao Lv, Fan Wu, Tie Luo, Xiaofeng Gao, and Guihai Chen. Hardness of and approximate mechanism design for the bike rebalancing problem. Theoretical Computer Science, 803(??): 105–115, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304657.

Yan:2020:CBA

[39] Ruidong Yan, Yuqing Zhu, Deying Li, and Yongcai Wang. Community based acceptance probability maximization for target users on social networks: Algorithms and analysis. Theoretical Computer Science, 803(??): 116–129, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304670.

Du:2020:QAO

[40] Yang Du, Yu-E Sun, He Huang, Liusheng Huang, Hongli Xu, and Xiaocan Wu. Quality-aware online task assignment mechanisms using latent topic model. Theoretical Computer Science, 803(??):130– 143, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304682.

Chen:2020:OSS

[41] Yong Chen, Randy Goebel, Guohui Lin, Bing Su, and An Zhang. Openshop scheduling for unit jobs under precedence constraints. Theoretical Computer Science, 803(??):144– 151, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306164.

Miao:2020:HRN

[42] Dongjing Miao, Jiguo Yu, and Zhipeng Cai. The hardness of resilience for nested aggregation query. Theoretical Computer Science, 803(??):152– 159, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930619X.

Kenter:2020:CBP

[43] Franklin Kenter, Daphne Skipper, and Dan Wilson. Computing bounds on product graph pebbling numbers. Theoretical Computer Science, 803(??): 160–177, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306206.

Ma:2020:QTA

[44] Hengzhao Ma and Jianzhong Li. An O(log n) query time algorithm for reducing ϵ-NN to (c, r)-NN. Theoretical Computer Science, 803(??):178– 195, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306279.

Hong:2020:BFA

[45] Yi Hong, Jiandong Liu, Deying Li, Chuanwen Luo, and Mengjie Chang. Balanced-flow algorithm for path network planning in hierarchical spaces. Theoretical Computer Science, 803(??): 196–206, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306577.

Miyahara:2020:PCB

[46] Daiki Miyahara, Yu ichi Hayashi, Takaaki Mizuki, and Hideaki Sone. Practical card-based implementations of Yao’s millionaire protocol. Theoretical Computer Science, 803(??): 207–221, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307042.

Belova:2020:AMP

[47] Tatiana Belova and Ivan Bliznets. Algorithms for (n, 3)-MAXSAT and parameterization above the all-true assignment. Theoretical Computer Science, 803(??):222–233, January 10, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519307595.

Anonymous:2020:PJd

[48] Anonymous. Pages 1–236 (12 January 2020). Theoretical Computer Science, 804(??):1–236, January 12, 2020.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBd

[49] Anonymous. Editorial Board. Theoretical Computer Science, 804(??): ii, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307807.

Zhou:2020:NUI

[50] Yanwei Zhou, Bo Yang, Tao Wang, and Yi Mu. Novel updatable identity-based hash proof system and its applications. Theoretical Computer Science, 804(??): 1–28, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306607.

Goerigk:2020:TSC

[51] Marc Goerigk, Adam Kasperski, and Paweł Zieliński. Two-stage combinatorial optimization problems under risk. Theoretical Computer Science, 804(??): 29–45, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306711.

Anand:2020:CEN

[52] Bijo S. Anand, Ullas Chandran S. V., Manoj Changat, Mitre C. Dourado, Ferdoos Hossein Nezhad, and Prasanth G. Narasimha-Shenoi. On the Carathéodory and exchange numbers of geodetic convexity in graphs. Theoretical Computer Science, 804(??): 46–57, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306917.

Stefanes:2020:SPA

[53] Marco A. Stefanes, Diego P. Rubert, and José Soares. Scalable parallel algorithms for maximum matching and Hamiltonian circuit in convex bipartite graphs. Theoretical Computer Science, 804(??):58– 71, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306929.

Ji:2020:SCF

[54] Meng Ji, Xueliang Li, and Xiaoyu Zhu. (Strong) conflict-free connectivity: Algorithm and complexity. Theoretical Computer Science, 804(??): 72–80, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306930.

Acharyya:2020:RAB

[55] Ankush Acharyya, Minati De, Subhas C. Nandy, and Bodhayan Roy. Range assignment of base-stations maximizing coverage area without interference. Theoretical Computer Science, 804(??):81–97, January 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306942.

Henning:2020:AAU

[56] Michael A. Henning and D. Pradhan. Algorithmic aspects of upper paired-domination in graphs. Theoretical Computer Science, 804(??): 98–114, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306954.

Lin:2020:PRT

[57] Zhicong Lin. Patterns of relation triples in inversion and ascent sequences. Theoretical Computer Science, 804(??): 115–125, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307066.

Furst:2020:EIM

[58] M. Fürst and D. Rautenbach. On the equality of the induced matching number and the uniquely restricted matching number for subcubic graphs. Theoretical Computer Science, 804(??): 126–138, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307467.

Wang:2020:GNC

[59] Zhao Wang, Yaping Mao, Sun-Yuan Hsieh, and Jichang Wu. On the ggood-neighbor connectivity of graphs. Theoretical Computer Science, 804(??): 139–148, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307479.

Bell:2020:UDF

[60] Paul C. Bell, Daniel Reidenbach, and Jeffrey O. Shallit. Unique decipherability in formal languages. Theoretical Computer Science, 804(??): 149–160, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307480.

Angelini:2020:BLP

[61] Patrizio Angelini, Giordano Da Lozzo, Giuseppe Di Battista, Fabrizio Frati, Maurizio Patrignani, and Ignaz Rutter. Beyond level planarity: Cyclic, torus, and simultaneous level planarity. Theoretical Computer Science, 804(??): 161–170, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307509.

Andrieu:2020:MWE

[62] Mélodie Andrieu and Anna E. Frid. Morphic words and equidistributed sequences. Theoretical Computer Science, 804(??):171–185, January 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930756X.

Moscarini:2020:CMB

[63] Marina Moscarini. Computing a metric basis of a 2-connected bipartite distance-hereditary graph. Theoretical Computer Science, 804(??):186– 206, January 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307571.

Meybodi:2020:PCD

[64] Mohsen Alambardar Meybodi, Fedor V. Fomin, Amer E. Mouawad, and Fahad Panolan. On the parameterized complexity of [1, j]- domination problems. Theoretical Computer Science, 804(??):207–218, January 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307583.

Le:2020:CCM

[65] Minh Ha Le, Vinh Duc Tran, Van Anh Trinh, and Viet Cuong Trinh. Compacting ciphertext in multi-channel broadcast encryption and attributebased encryption. Theoretical Computer Science, 804(??):219–235, January 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307601.

Anonymous:2020:PJe

[66] Anonymous. Pages 1–232 (18 January 2020). Theoretical Computer Science, 805(??):1–232, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBe

[67] Anonymous. Editorial Board. Theoretical Computer Science, 805(??): ii, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307996.

Kari:2020:Eb

[68] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 805(??):iii, January 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308023.

Xavier-Junior:2020:EAA

[69] João C. Xavier-Júnior, Alex A. Freitas, Teresa B. Ludermir, Antonino Feitosa-Neto, and Cephas A. S. Barreto. An evolutionary algorithm for automated machine learning focusing on classifier ensembles: an improved algorithm and extended results. Theoretical Computer Science, 805(??): 1–18, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307649.

Eustaquio:2020:ENI

[70] Fernanda Eustáquio and Tatiane Nogueira. Evaluating the numerical instability in fuzzy clustering validation of high-dimensional data. Theoretical Computer Science, 805(??): 19–36, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306826.

Freitas:2020:IQA

[71] Pedro Garcia Freitas, Luísa Peixoto da Eira, Samuel Soares Santos, and Mylène C. Q. Farias. Image quality assessment using BSIF, CLBP, LCP, and LPQ operators. Theoretical Computer Science, 805(??): 37–61, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306814.

Dantas:2020:UFL

[72] Augusto Dantas and Aurora Pozo. On the use of fitness landscape features in meta-learning based algorithm selection for the quadratic assignment problem. Theoretical Computer Science, 805(??):62–75, January 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306693.

Gabriel:2020:RBA

[73] Vágner de Oliveira Gabriel, Alison R. Panisson, Rafael H. Bordini, Diana Francisca Adamatti, and Cleo Zanella Billa. Reasoning in BDI agents using Toulmin’s argumentation model. Theoretical Computer Science, 805(??):76–91, January 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306553.

Maia:2020:RPA

[74] Artur Vidal Maia and Jaime Simão Sichman. Representing planning autonomy in agent organizational models. Theoretical Computer Science, 805(??): 92–108, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305456.

Rocha:2020:SDF

[75] Thiago Alves Rocha, Ana Teresa Martins, and Francicleber Martins Ferreira. Synthesis of a DNF formula from a sample of strings using Ehrenfeucht–Fraïssé games. Theoretical Computer Science, 805(??): 109–126, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305043.

Alhazov:2020:TFC

[76] Artiom Alhazov, Rudolf Freund, Sergiu Ivanov, Linqiang Pan, and Bosheng Song. Time-freeness and clock-freeness and related concepts in P systems. Theoretical Computer Science, 805(??): 127–143, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305747.

Alhazov:2020:SRR

[77] Artiom Alhazov, Rudolf Freund, and Sergiu Ivanov. P systems with randomized right-hand sides of rules. Theoretical Computer Science, 805(??): 144–160, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304973.

Ciencialova:2020:TNA

[78] Lucie Ciencialová and Ludek Cienciala. Two notes on APCol systems. Theoretical Computer Science, 805(??): 161–174, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304870.

Aman:2020:LTM

[79] Bogdan Aman, Péter Battyányi, Gabriel Ciobanu, and György Vaszil. Local time membrane systems and time Petri nets. Theoretical Computer Science, 805(??):175–192, January 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304213.

Leporati:2020:SSC

[80] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. Subroutines in P systems and closure properties of their complexity classes. Theoretical Computer Science, 805(??):193– 205, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304201.

Orellana-Martin:2020:SSA

[81] David Orellana-Martín, Miguel Á. Martínez del Amor, Luis Valencia-Cabrera, Bosheng Song, Linqiang Pan, and Mario J. Pérez-Jiménez. P systems with symport/antiport rules: When do the surroundings matter? Theoretical Computer Science, 805(??):

206–217, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303062.

Orellana-Martin:2020:WOP

[82] David Orellana-Martín, Miguel Á. Martínez del Amor, Ignacio Pérez-Hurtado, Agustín Riscos-Núñez, Luis Valencia-Cabrera, and Mario J. Pérez-Jiménez. When object production tunes the efficiency of membrane systems. Theoretical Computer Science, 805(??): 218–231, January 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302354.

Anonymous:2020:PFa

[83] Anonymous. Pages 1–690 (2 February 2020). Theoretical Computer Science, 806(??):1–690, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBf

[84] Anonymous. Editorial Board. Theoretical Computer Science, 806(??): ii, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300165.

Crochemore:2020:CLT

[85] Maxime Crochemore and Luís M. S. Russo. Cartesian and Lyndon trees. Theoretical Computer Science, 806(??): 1–9, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305279.

Peltomaki:2020:MDS

[86] Jarkko Peltomäki and Markus A. Whiteland. More on the dynamics of the symbolic square root map. Theoretical Computer Science, 806(??): 10–27, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305358.

Neraud:2020:EIC

[87] Jean Néraud and Carla Selmi. Embedding a θ-invariant code into a complete one. Theoretical Computer Science, 806 (??):28–41, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305383.

Estrada-Moreno:2020:PDG

[88] Alejandro Estrada-Moreno. On the kpartition dimension of graphs. Theoretical Computer Science, 806(??): 42–52, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305930.

Eroh:2020:CMD

[89] Linda Eroh, Cong X. Kang, and Eunjeong Yi. The connected metric dimension at a vertex of a graph. Theoretical Computer Science, 806(??): 53–69, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306674.

Qin:2020:IAH

[90] Tong Qin and Osamu Watanabe. An improvement of the algorithm of Hertli for the unique 3SAT problem. Theoretical Computer Science, 806(??): 70–80, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307138.

Fujita:2020:BND

[91] Takahiro Fujita, Kohei Hatano, and Eiji Takimoto. Boosting over nondeterministic ZDDs. Theoretical Computer Science, 806(??):81–89, February 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307175.

Lamarche-Perrin:2020:ITF

[92] Robin Lamarche-Perrin. An informationtheoretic framework for the lossy compression of link streams. Theoretical Computer Science, 806(??):90– 115, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307308.

Bulteau:2020:RBS

[93] Laurent Bulteau and Stéphane Vialette. Recognizing binary shuffle squares is NP-hard. Theoretical Computer Science, 806(??):116–132, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519300258.

Gutin:2020:APM

[94] Gregory Gutin, M. S. Ramanujan, Felix Reidl, and Magnus Wahlström. Alternative parameterizations of Metric Dimension. Theoretical Computer Science, 806(??):133–143, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519300532.

Alcon:2020:NLC

[95] Liliana Alcon, Marisa Gutierrez, Carmen Hernando, Mercè Mora, and Ignacio M. Pelayo. Neighbor-locating colorings in graphs. Theoretical Computer Science, 806(??):144–155, February 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300805.

Hakanen:2020:SMD

[96] Anni Hakanen, Ville Junnila, and Tero Laihonen. The solid-metric dimension. Theoretical Computer Science, 806(??): 156–170, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301197.

Yi:2020:BDN

[97] Eunjeong Yi. On the broadcast domination number of permutation graphs. Theoretical Computer Science, 806(??): 171–183, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301902.

Baste:2020:TM

[98] Julien Baste, Binh-Minh Bui-Xuan, and Antoine Roux. Temporal matching. Theoretical Computer Science, 806(??):184–196, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519301914.

Fluschnik:2020:TGC

[99] Till Fluschnik, Hendrik Molter, Rolf Niedermeier, Malte Renken, and Philipp Zschoche. Temporal graph classes: a view through temporal separators. Theoretical Computer Science, 806(??):197–218, February 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301975.

DeBonis:2020:NKS

[100] Annalisa De Bonis and Ugo Vaccaro. A new kind of selectors and their applications to conflict resolution in wireless multichannels networks. Theoretical Computer Science, 806(??): 219–235, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302014.

Bampas:2020:BSI

[101] Evangelos Bampas, Jurek Czyzowicz, David Ilcinkas, and Ralf Klasing.

Beachcombing on strips and islands. Theoretical Computer Science, 806(??): 236–255, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302312.

You:2020:FPT

[102] Jie You, Feng Shi, Jianxin Wang, and Qilong Feng. Fixed-parameter tractability for minimum tree cut/ paste distance and minimum common integer partition. Theoretical Computer Science, 806(??):256– 270, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302567.

Chen:2020:AAH

[103] Li-Hsuan Chen, Sun-Yuan Hsieh, Ling-Ju Hung, and Ralf Klasing. Approximation algorithms for the phub center routing problem in parameterized metric graphs. Theoretical Computer Science, 806(??):271– 280, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302877.

Neumann:2020:PSO

[104] Eike Neumann and Florian Steinberg. Parametrised second-order complexity theory with applications to the study of interval computation. Theoretical Computer Science, 806(??): 281–304, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302889.

Assiyatun:2020:CUB

[105] Hilda Assiyatun, Dian Kastika Syofyan, and Edy Tri Baskoro. Calculating an upper bound of the locatingchromatic number of trees. Theoretical Computer Science, 806(??): 305–309, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302993.

Yamazaki:2020:ENI

[106] Kazuaki Yamazaki, Toshiki Saitoh, Masashi Kiyomi, and Ryuhei Uehara. Enumeration of nonisomorphic interval graphs and nonisomorphic permutation graphs. Theoretical Computer Science, 806(??):310– 322, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303056.

Takayama:2020:SPT

[107] Koki Takayama and Yusuke Kobayashi. A strongly polynomial time algorithm for the maximum supply rate problem on trees. Theoretical Computer Science, 806(??):323– 331, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303226.

Cardinal:2020:RSA

[108] Jean Cardinal, Erik D. Demaine, David Eppstein, Robert A. Hearn, and Andrew Winslow. Reconfiguration of satisfying assignments and subset sums: Easy to find, hard to connect. Theoretical Computer Science, 806(??): 332–343, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303779.

Jager:2020:MD

[109] Gerold Jäger and Frank Drewes. The metric dimension of Z_{n} × Z_{n} × Z_{n} is ⌊3n/2⌋. Theoretical Computer Science, 806(??):344– 362, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304098.

Ahmadi:2020:CGB

[110] Mohamad Ahmadi, Abdolhamid Ghodselahi, Fabian Kuhn, and Anisur Rahaman Molla. The cost of global broadcast in dynamic radio networks. Theoretical Computer Science, 806(??): 363–387, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304414.

Benkoczi:2020:MSP

[111] Robert Benkoczi, Binay Bhattacharya, Yuya Higashikawa, Tsunehiko Kameda, and Naoki Katoh. Minsum k-sink problem on path networks. Theoretical Computer Science, 806(??): 388–401, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304451.

Chikhi:2020:BGS

[112] Rayan Chikhi, Vladan Jovičić, Stefan Kratsch, Paul Medvedev, Martin Milanič, Sofya Raskhodnikova, and Nithin Varma. Bipartite graphs of small readability. Theoretical Computer Science, 806(??):402–415, February 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304578.

Chan:2020:GHL

[113] T.-H. Hubert Chan and Zhibin Liang. Generalizing the hypergraph Laplacian via a diffusion process with mediators. Theoretical Computer Science, 806(??): 416–428, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304591.

Casteigts:2020:RNF

[114] Arnaud Casteigts, Swan Dubois, Franck Petit, and John M. Robson. Robustness: a new form of heredity motivated by dynamic networks. Theoretical Computer Science, 806(??): 429–445, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304979.

Lin:2020:AAP

[115] Mugang Lin, Qilong Feng, Bin Fu, and Jianxin Wang. An approximation algorithm for the lpseudoforest deletion problem. Theoretical Computer Science, 806(??): 446–454, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304980.

Ahmed:2020:OFA

[116] Abu Reyan Ahmed, Md. Saidur Rahman, and Stephen Kobourov. Online facility assignment. Theoretical Computer Science, 806(??):455– 467, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305006.

Liu:2020:GTB

[117] Gang Liu, Zheng Xiao, GuangHua Tan, Kenli Li, and Anthony Theodore Chronopoulos. Game theory-based optimization of distributed idle computing resources in cloud environments. Theoretical Computer Science, 806(??): 468–488, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305080.

Gilbert:2020:SBU

[118] Seth Gilbert, Nancy Lynch, Calvin Newport, and Dominik Pajak. On simple back-off in unreliable radio networks. Theoretical Computer Science, 806(??):489–508, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519305237.

Wu:2020:IAA

[119] Guangwei Wu, Jianer Chen, and Jianxin Wang. Improved approximation algorithms for two-stage flowshops scheduling problem. Theoretical Computer Science, 806(??):509– 515, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305249.

Xu:2020:DCP

[120] Chuan Xu, Joffroy Beauquier, Janna Burman, Shay Kutten, and Thomas Nowak. Data collection in population protocols with non-uniformly random scheduler. Theoretical Computer Science, 806(??):516–530, February 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305250.

Czyzowicz:2020:SNA

[121] Jurek Czyzowicz, Stefan Dobrev, Maxime Godon, Evangelos Kranakis, Toshinori Sakai, and Jorge Urrutia. Searching for a non-adversarial, uncooperative agent on a cycle. Theoretical Computer Science, 806(??): 531–542, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305274.

Baig:2020:TIS

[122] Mirza Galib Anwarul Husain Baig and Deepanjan Kesh. Two improved schemes in the bitprobe model. Theoretical Computer Science, 806(??):

543–552, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305298.

Hanaka:2020:RSI

[123] Tesshu Hanaka, Takehiro Ito, Haruka Mizuta, Benjamin Moore, Naomi Nishimura, Vijay Subramanya, Akira Suzuki, and Krishna Vaidyanathan. Reconfiguring spanning and induced subgraphs. Theoretical Computer Science, 806(??):553–566, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519305626.

Hounkanli:2020:GSC

[124] Kokouvi Hounkanli, Avery Miller, and Andrzej Pelc. Global synchronization and consensus using beeps in a faultprone multiple access channel. Theoretical Computer Science, 806(??): 567–576, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930564X.

Murano:2020:MCG

[125] Aniello Murano, Mimmo Parente, Sasha Rubin, and Loredana Sorrentino. Model-checking graded computationtree logic with finite path semantics. Theoretical Computer Science, 806(??): 577–586, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305651.

Tan:2020:NKS

[126] Guanlan Tan, Qilong Feng, Beilin Zhuo, Neng Huang, and Jianxin Wang. New kernels for several problems on planar graphs. Theoretical Computer Science, 806(??):587– 594, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305754.

Czyzowicz:2020:PEDa

[127] Jurek Czyzowicz, Konstantinos Georgiou, Ryan Killick, Evangelos Kranakis, Danny Krizanc, Lata Narayanan, Jaroslav Opatrny, and Sunil Shende. Priority evacuation from a disk: the case of n = 1, 2, 3. Theoretical Computer Science, 806(??):595– 616, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305778.

Sudo:2020:LSL

[128] Yuichi Sudo, Fukuhito Ooshita, Hirotsugu Kakugawa, Toshimitsu Masuzawa, Ajoy K. Datta, and Lawrence L. Larmore. Looselystabilizing leader election with polylogarithmic convergence time. Theoretical Computer Science, 806(??): 617–631, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305857.

Sadhu:2020:CLT

[129] Sanjib Sadhu, Xiaozhou He, Sasanka Roy, Subhas C. Nandy, and Suchismita Roy. Corrigendum to: “Linear time algorithm to cover and hit a set of line segments optimally by two axis-parallel squares” [Theor. Comput. Sci. 769 (2019) 63–74]. Theoretical Computer Science, 806(??): 632–640, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306140. See

[3].

Bilo:2020:PAA

[130] Vittorio Bilò and Cosimo Vinci. The price of anarchy of affine congestion games with similar strategies. Theoretical Computer Science, 806(??): 641–654, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306413.

Gargano:2020:LWS

[131] Luisa Gargano, Adele Anna Rescigno, and Ugo Vaccaro. Low-weight superimposed codes and related combinatorial structures: Bounds and applications. Theoretical Computer Science, 806(??): 655–672, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306619.

Kobayashi:2020:OIS

[132] Koji M. Kobayashi. Online interval scheduling to maximize total satisfaction. Theoretical Computer Science, 806(??):673–688, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519306966.

Lazard:2020:CEL

[133] Sylvain Lazard, William J. Lenhart, and Giuseppe Liotta. Corrigendum to “On the edge-length ratio of outerplanar graphs” [Theoret. Comput. Sci. 770 (2019) 88–94]. Theoretical Computer Science, 806(??): 689, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307443. See

[4].

Anonymous:2020:EBg

[134] Anonymous. Editorial Board. Theoretical Computer Science, 807(??): ii, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300256.

Ausiello:2020:P

[135] Giorgio Ausiello, Lila Kari, Grzegorz Rozenberg, Donald Sannella, Paul Spirakis, and Pierre-Louis Curien. Preface. Theoretical Computer Science, 807(??):1–2, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519308035.

Abramsky:2020:WS

[136] Samson Abramsky. Whither semantics? Theoretical Computer Science, 807(??):3–14, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930444X.

Alur:2020:SRT

[137] Rajeev Alur, Dana Fisman, Konstantinos Mamouras, Mukund Raghothaman, and Caleb Stanford. Streamable regular transductions. Theoretical Computer Science, 807(??):15– 41, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307376.

Avni:2020:DRA

[138] Guy Avni, Thomas A. Henzinger, and Orna Kupferman. Dynamic resource allocation games. Theoretical Computer Science, 807(??):42– 55, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304475.

Bar-Noy:2020:VWR

[139] Amotz Bar-Noy, David Peleg, and Dror Rawitz. Vertex-weighted realizations of graphs. Theoretical Computer Science, 807(??):56–72, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307959.

Blass:2020:BD

[140] Andreas Blass and Yuri Gurevich. Braided distributivity. Theoretical Computer Science, 807(??):73– 94, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307534.

Bruni:2020:BNS

[141] Roberto Bruni, Hernán Melgratti, and Ugo Montanari. Bayesian network semantics for Petri nets. Theoretical Computer Science, 807(??): 95–113, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304694.

Calude:2020:SSL

[142] Cristian S. Calude, Sanjay Jain, Wolfgang Merkle, and Frank Stephan. Searching for shortest and least programs. Theoretical Computer Science, 807(??):114–127, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519306401.

Castellani:2020:GTI

[143] Ilaria Castellani, Mariangiola Dezani-Ciancaglini, Paola Giannini, and Ross Horne. Global types with internal delegation. Theoretical Computer Science, 807(??):128–153, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930578X.

Durier:2020:TCR

[144] Adrien Durier, Daniel Hirschkoff, and Davide Sangiorgi. Towards ‘up to context’ reasoning about higher-order processes. Theoretical Computer Science, 807(??):154– 168, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305870.

Ehrhard:2020:CBP

[145] Thomas Ehrhard, Jean Krivine, and Ying Jiang. A calculus of branching processes. Theoretical Computer Science, 807(??):169– 184, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304438.

Georgiadis:2020:ASV

[146] Loukas Georgiadis, Giuseppe F. Italiano, and Aikaterini Karanasiou. Approximating the smallest 2-vertex connected spanning subgraph of a directed graph. Theoretical Computer Science, 807(??): 185–200, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306048.

Iwama:2020:IAC

[147] Kazuo Iwama and Junichi Teruyama. Improved average complexity for comparison-based sorting. Theoretical Computer Science, 807(??): 201–219, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304487.

Kozen:2020:LHC

[148] Dexter Kozen and Alexandra Silva. Left-handed completeness. Theoretical Computer Science, 807(??): 220–233, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306838.

Leone:2020:ESD

[149] Pierre Leone, Kasun Samarasinghe, and José D. P. Rolim. Every Schnyder drawing is a greedy embedding. Theoretical Computer Science, 807(??): 234–244, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304566.

Mahalingam:2020:LNP

[150] Kalpana Mahalingam, Palak Pandoh, and Kamala Krithivasan. On the least number of palindromes in two-dimensional words. Theoretical Computer Science, 807(??):245– 256, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304463.

Matiyasevich:2020:RHC

[151] Yu. Matiyasevich. The Riemann Hypothesis in computer science. Theoretical Computer Science, 807(??): 257–265, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304633.

Mavronicolas:2020:CVR

[152] Marios Mavronicolas and Burkhard Monien. Conditional value-at-risk: Structure and complexity of equilibria. Theoretical Computer Science, 807 (??):266–283, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305596.

Mislove:2020:DSP

[153] Michael Mislove. Domains and stochastic processes. Theoretical Computer Science, 807(??):284– 297, February 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302713.

Wu:2020:GBA

[154] Min Wu, Matthew Wicker, Wenjie Ruan, Xiaowei Huang, and Marta Kwiatkowska. A game-based approximate verification of deep neural networks with provable guarantees. Theoretical Computer Science, 807(??): 298–329, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304426.

Yu:2020:SIS

[155] Fang Yu, Daowen Qiu, Xiaoming Wang, Qin Li, Lvzhou Li, and Jozef Gruska. Security improvements of several basic quantum private query protocols with O(log N) communication complexity. Theoretical Computer Science, 807(??):

330–340, February 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307704.

Anonymous:2020:EBh

[156] Anonymous. Editorial Board. Theoretical Computer Science, 808(??): ii, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300396.

Hanneke:2020:SIA

[157] Steve Hanneke and Lev Reyzin. Special issue on ALT 2017: Guest Editors’ introduction. Theoretical Computer Science, 808(??):1, February 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307200.

Angluin:2020:PRC

[158] Dana Angluin and Tyler Dohrn. The power of random counterexamples. Theoretical Computer Science, 808(??): 2–13, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307212.

Balcan:2020:LLC

[159] Maria-Florina Balcan, Avrim Blum, and Vaishnavh Nagarajan. Lifelong learning in costly feature spaces. Theoretical Computer Science, 808(??): 14–37, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307224.

Bayeh:2020:FDE

[160] Fahimeh Bayeh, Ziyuan Gao, and Sandra Zilles. Finitely distinguishable erasing pattern languages. Theoretical Computer Science, 808(??): 38–73, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307236.

Ensign:2020:CEN

[161] Danielle Ensign, Scott Neville, Arnab Paul, and Suresh Venkatasubramanian. The complexity of explaining neural networks through (group) invariants. Theoretical Computer Science, 808(??): 74–85, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307248.

Feldman:2020:TBA

[162] Vitaly Feldman, Pravesh Kothari, and Jan Vondrák. Tight bounds on l _{1} approximation and learning of self-bounding functions. Theoretical Computer Science, 808(??):86– 98, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930725X.

Gupta:2020:PIM

[163] Arushi Gupta and Daniel Hsu. Parameter identification in Markov chain choice models. Theoretical Computer Science, 808(??):99–107, February 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307261.

Joulani:2020:MAA

[164] Pooria Joulani, András György, and Csaba Szepesvári. A modular analysis of adaptive (non-)convex optimization: Optimism, composite objectives, variance reduction, and variational bounds. Theoretical Computer Science, 808(??): 108–138, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307273.

Kotlowski:2020:SIU

[165] Wojciech Kotłowski. Scale-invariant unconstrained online learning. Theoretical Computer Science, 808(??): 139–158, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307285.

Long:2020:NBP

[166] Philip M. Long. New bounds on the price of bandit feedback for mistakebounded online multiclass learning. Theoretical Computer Science, 808(??): 159–163, February 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307297.

Anonymous:2020:PFb

[167] Anonymous. Pages 1–574 (24 February 2020). Theoretical Computer Science, 809(??):1–574, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBi

[168] Anonymous. Editorial Board. Theoretical Computer Science, 809(??): ii, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300566.

Ramya:2020:LBS

[169] C. Ramya and B. V. Raghavendra Rao. Lower bounds for special cases of syntactic multilinear ABPs. Theoretical Computer Science, 809(??): 1–20, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306978.

Abram:2020:PCM

[170] Antoine Abram, Mélodie Lapointe, and Christophe Reutenauer. Palindromization and construction of Markoff triples. Theoretical Computer Science, 809(??):21–29, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930698X.

Burcroff:2020:GLF

[171] Amanda Burcroff and Eric Winsor. Generalized Lyndon factorizations of infinite words. Theoretical Computer Science, 809(??):30– 38, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307029.

Postic:2020:LW

[172] Mickaël Postic and Luca Q. Zamboni. ω-Lyndon words. Theoretical Computer Science, 809(??):39– 44, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307030.

Censor-Hillel:2020:SMM

[173] Keren Censor-Hillel, Dean Leitersdorf, and Elia Turner. Sparse matrix multiplication and triangle listing in the Congested Clique model. Theoretical Computer Science, 809(??): 45–60, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307054.

Abas:2020:MDC

[174] Marcel Abas and Tomás Vetrík. Metric dimension of Cayley digraphs of split metacyclic groups. Theoretical Computer Science, 809(??):61– 72, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307510.

Zhao:2020:AAI

[175] Zhen Zhao, Fuchun Guo, Jianchang Lai, Willy Susilo, Baocang Wang, and Yupu Hu. Accountable authority identity-based broadcast encryption with constant-size private keys and ciphertexts. Theoretical Computer Science, 809(??):73– 87, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307613.

Lv:2020:ECE

[176] Mengjie Lv, Jianxi Fan, Jingya Zhou, Baolei Cheng, and Xiaohua Jia. The extra connectivity and extra diagnosability of regular interconnection networks. Theoretical Computer Science, 809(??):88– 102, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307637.

Katsumata:2020:LBR

[177] Shuichi Katsumata, Takahiro Matsuda, and Atsushi Takayasu. Lattice-based revocable (hierarchical) IBE with decryption key exposure resistance. Theoretical Computer Science, 809(??): 103–136, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307650.

Datta:2020:CPF

[178] Pratish Datta. Constrained pseudorandom functions from functional encryption. Theoretical Computer Science, 809(??):137–170, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307662.

Barbier:2020:CMT

[179] Morgan Barbier, Hayat Cheballah, and Jean-Marie Le Bars. On the computation of the Möbius transform. Theoretical Computer Science, 809(??): 171–188, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307674.

Li:2020:RAD

[180] Xiaoyan Li, Xiaohua Jia, Jianxi Fan, and Cheng-Kuan Lin. Reliability analysis of data center networks based on precise and imprecise diagnosis strategies. Theoretical Computer Science, 809(??):189– 203, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307686.

Page:2020:MMUa

[181] Daniel R. Page, Roberto Solis-Oba, and Marten Maack. Makespan minimization on unrelated parallel machines with simple job-intersection structure and bounded job assignments. Theoretical Computer Science, 809(??):204–217, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307844.

Dumitrescu:2020:OUC

[182] Adrian Dumitrescu, Anirban Ghosh, and Csaba D. Tóth. Online unit covering in Euclidean space. Theoretical Computer Science, 809(??):218– 230, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307856.

Cheng:2020:PDP

[183] Changjie Cheng, Changhong Lu, and Yu Zhou. The k-power domination problem in weighted trees. Theoretical Computer Science, 809(??):231– 238, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307881.

Bentz:2020:CMP

[184] Cédric Bentz and Pierre Le Bodic. Complexity of the multicut problem, in its vanilla, partial and generalized versions, in graphs of bounded treewidth. Theoretical Computer Science, 809(??): 239–249, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930790X.

DiGiacomo:2020:PDT

[185] Emilio Di Giacomo, Peter Eades, Giuseppe Liotta, Henk Meijer, and Fabrizio Montecchiani. Polyline drawings with topological constraints. Theoretical Computer Science, 809(??):

250–264, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307911.

Kamiyama:2020:PMT

[186] Naoyuki Kamiyama. Popular matchings with two-sided preference lists and matroid constraints. Theoretical Computer Science, 809(??):265– 276, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307923.

Zhou:2020:IBE

[187] Yanwei Zhou, Bo Yang, Zhe Xia, Mingwu Zhang, and Yi Mu. Identitybased encryption with leakage-amplified chosen-ciphertext attacks security. Theoretical Computer Science, 809(??): 277–295, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307935.

Wang:2020:PCA

[188] Di Wang and Jinhui Xu. Principal Component Analysis in the local differential privacy model. Theoretical Computer Science, 809(??):296– 312, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307947.

DeMarco:2020:DBC

[189] Gianluca De Marco, Mauro Leoncini, and Manuela Montangero. Distributed balanced color assignment on arbitrary networks. Theoretical Computer Science, 809(??):313– 326, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308059.

Czygrinow:2020:DAA

[190] Andrzej Czygrinow, Michał Hanćkowiak, Wojciech Wawrzyniak, and Marcin Witkowski. Distributed approximation algorithms for k-dominating set in graphs of bounded genus and linklessly embeddable graphs. Theoretical Computer Science, 809(??):327– 338, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308096.

Cicalese:2020:MIU

[191] Ferdinando Cicalese and Massimiliano Rossi. On the multi-interval Ulam–Rényi game: For 3 lies 4 intervals suffice. Theoretical Computer Science, 809(??):339–356, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308102.

Shibata:2020:SEU

[192] Masahiro Shibata, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Space-efficient uniform deployment of mobile agents in asynchronous unidirectional rings. Theoretical Computer Science, 809(??):357–371, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300013.

Chen:2020:NES

[193] Jiahui Chen, Jianting Ning, Jie Ling, Terry Shue Chien Lau, and Yacheng Wang. A new encryption scheme for multivariate quadratic systems. Theoretical Computer Science, 809(??): 372–383, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300025.

Hellmuth:2020:CMP

[194] Marc Hellmuth, Manuela Geiß, and Peter F. Stadler. Complexity of modification problems for reciprocal best match graphs. Theoretical Computer Science, 809(??):384– 393, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300037.

Hu:2020:CNP

[195] Wenyu Hu, Shenghao Li, Jinhong Huang, Tinghua Wang, and Gaohang Yu. Computing the nearest polynomial to multiple given polynomials with a given zero via l_{2,q}-norm minimization. Theoretical Computer Science, 809(??): 394–406, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300049.

Dekking:2020:MWB

[196] Michel Dekking. Morphic words, Beatty sequences and integer images of the Fibonacci language. Theoretical Computer Science, 809(??):407– 417, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300062.

DAngeli:2020:OEA

[197] Daniele D’Angeli, Emanuele Rodaro, and Jan Philipp Wächter. Orbit expandability of automaton semigroups and groups. Theoretical Computer Science, 809(??):418–429, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300074.

Wang:2020:TXC

[198] Meng Wang, Cong Tian, Nan Zhang, Zhenhua Duan, and Chenguang Yao. Translating Xd-C programs to MSVL programs. Theoretical Computer Science, 809(??):430– 465, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300086.

Dose:2020:OSC

[199] Titus Dose. An oracle separating conjectures about incompleteness in the finite domain. Theoretical Computer Science, 809(??):466– 481, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300116.

Mills:2020:AOR

[200] K. Alex Mills, R. Chandrasekaran, and Neeraj Mittal. Algorithms for optimal replica placement under correlated failure in hierarchical failure domains. Theoretical Computer Science, 809(??): 482–518, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300128.

Zhang:2020:NAV

[201] Nan Zhang, Zhenhua Duan, Cong Tian, and Hongwei Du. A novel approach to verifying context free properties of programs. Theoretical Computer Science, 809(??):519– 530, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300190.

Vereshchagin:2020:DCC

[202] Nikolay Vereshchagin. Descriptive complexity of computable sequences revisited. Theoretical Computer Science, 809(??):531–537, February 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300347.

Amrane:2020:LRL

[203] Amazigh Amrane and Nicolas Bedon. Logic and rational languages of scattered and countable series-parallel posets. Theoretical Computer Science, 809(??):538–562, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520300426.

Huang:2020:AAC

[204] Lingxiao Huang, Jian Li, and Qicai Shi. Approximation algorithms for the connected sensor cover problem. Theoretical Computer Science, 809(??): 563–574, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300487.

Anonymous:2020:EBj

[205] Anonymous. Editorial Board. Theoretical Computer Science, 810(??): ii, March 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300682.

Suomela:2020:SIC

[206] Jukka Suomela. Structural Information and Communication Complexity. Theoretical Computer Science, 810(??):1, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300645.

Bartschi:2020:CDE

[207] Andreas Bärtschi, Jérémie Chalopin, Shantanu Das, Yann Disser, Barbara Geissmann, Daniel Graf, Arnaud Labourel, and Matúš Mihalák.

Collaborative delivery with energyconstrained mobile robots. Theoretical Computer Science, 810(??): 2–14, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304656.

Brandes:2020:FSA

[208] Philipp Brandes, Marcin Kardas, Marek Klonowski, Dominik Pajak, and Roger Wattenhofer. Fast size approximation of a radio network in beeping model. Theoretical Computer Science, 810(??):15–25, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304620.

Cordasco:2020:WBI

[209] Gennaro Cordasco, Luisa Gargano, Manuel Lafond, Lata Narayanan, Adele A. Rescigno, Ugo Vaccaro, and Kangkang Wu. Whom to befriend to influence people. Theoretical Computer Science, 810(??): 26–42, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303682.

Daum:2020:RSB

[210] Sebastian Daum, Fabian Kuhn, and Yannic Maus. Rumor spreading with bounded in-degree. Theoretical Computer Science, 810(??):43– 57, March 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304080.

Emek:2020:DNF

[211] Yuval Emek and Jara Uitto. Dynamic networks of finite state machines. Theoretical Computer Science, 810(??):58–71, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304668.

Fischer:2020:PVP

[212] Orr Fischer, Rotem Oshman, and Uri Zwick. Public vs. private randomness in simultaneous multi-party communication complexity. Theoretical Computer Science, 810(??): 72–81, March 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302664.

Pandurangan:2020:MLB

[213] Gopal Pandurangan, David Peleg, and Michele Scquizzato. Message lower bounds via efficient network synchronization. Theoretical Computer Science, 810(??):82– 95, March 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307072.

Anonymous:2020:EBk

[214] Anonymous. Editorial Board. Theoretical Computer Science, 811(??): ii, April 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300839.

Das:2020:SIS

[215] Shantanu Das and Sebastien Tixeuil. Special issue on Structural Information and Communication Complexity. Theoretical Computer Science, 811(??):1–2, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300864.

Jurdzinski:2020:TTA

[216] Tomasz Jurdzinski, Dariusz R. Kowalski, Michal Rozanski, and Grzegorz Stachowiak. Token traversal in ad hoc wireless networks via implicit carrier sensing. Theoretical Computer Science, 811(??):3–20, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305602.

Halldorsson:2020:LES

[217] Magnús M. Halldórsson, Stephan Holzer, Evangelia Anna Markatou, and Nancy Lynch. Leader election in SINR model with arbitrary power control. Theoretical Computer Science, 811(??):21–28, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300490.

Halldorsson:2020:IDA

[218] Magnús M. Halldórsson and Christian Konrad. Improved distributed algorithms for coloring interval graphs with application to multicoloring trees. Theoretical Computer Science, 811 (??):29–41, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307187.

Feuilloley:2020:HLI

[219] Laurent Feuilloley. How long it takes for an ordinary node with an ordinary id to output? Theoretical Computer Science, 811(??): 42–55, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300489.

Brandt:2020:WER

[220] Sebastian Brandt, Klaus-Tycho Foerster, Benjamin Richner, and Roger Wattenhofer. Wireless evacuation on m rays with k searchers. Theoretical Computer Science, 811(??): 56–69, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306534.

Disser:2020:GLB

[221] Yann Disser, Frank Mousset, Andreas Noever, Nemanja Skorić, and Angelika Steger. A general lower bound for collaborative tree exploration. Theoretical Computer Science, 811(??):70–78, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301452.

DiLuna:2020:GDR

[222] Giuseppe Antonio Di Luna, Paola Flocchini, Linda Pagli, Giuseppe Prencipe, Nicola Santoro, and Giovanni Viglietta. Gathering in dynamic rings. Theoretical Computer Science, 811 (??):79–98, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830639X.

Narayanan:2020:HCF

[223] Lata Narayanan and Kangkang Wu. How to choose friends strategically. Theoretical Computer Science, 811 (??):99–111, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304948.

Censor-Hillel:2020:APL

[224] Keren Censor-Hillel, Ami Paz, and Mor Perry. Approximate prooflabeling schemes. Theoretical Computer Science, 811(??):112–124, April 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751830536X.

Anonymous:2020:EBl

[225] Anonymous. Editorial Board. Theoretical Computer Science, 812(??): ii, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301079.

Fici:2020:P

[226] Gabriele Fici and Giuseppe F. Italiano. Preface. Theoretical Computer Science, 812(??):1, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301031.

Alzamel:2020:FAM

[227] Mai Alzamel, Panagiotis Charalampopoulos, Costas S. Iliopoulos, Solon P. Pissis, Jakub Radoszewski, and Wing-Kin Sung. Faster algorithms for 1mappability of a sequence. Theoretical Computer Science, 812(??): 2–12, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303147.

R:2020:SCO

[228] Rani M. R., Subashini R., and Mohith Jagalmohanan. Simultaneous consecutive ones submatrix and editing problems: Classical complexity and fixed-parameter tractable results. Theoretical Computer Science, 812 (??):13–38, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304104.

Gawrychowski:2020:CRM

[229] Paweł Gawrychowski, Seungbum Jo, Shay Mozes, and Oren Weimann. Compressed range minimum queries. Theoretical Computer Science, 812 (??):39–48, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304281.

Amir:2020:TDM

[230] Amihood Amir, Gad M. Landau, Shoshana Marcus, and Dina Sokol. Two-dimensional maximal repetitions. Theoretical Computer Science, 812 (??):49–61, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304323.

Cordasco:2020:FFT

[231] Gennaro Cordasco, Luisa Gargano, Joseph G. Peters, Adele A. Rescigno, and Ugo Vaccaro. Fast and frugal targeting with incentives. Theoretical Computer Science, 812(??): 62–79, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304335.

Garay:2020:CHD

[232] Juan Garay, David Johnson, Aggelos Kiayias, and Moti Yung. The combinatorics of hidden diversity. Theoretical Computer Science, 812(??): 80–95, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304517.

Bannai:2020:RI

[233] Hideo Bannai, Travis Gagie, and Tomohiro I. Refining the r-index. Theoretical Computer Science, 812 (??):96–108, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304943.

Bernardini:2020:APM

[234] Giulia Bernardini, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Approximate pattern matching on elastic-degenerate text. Theoretical Computer Science, 812(??):109– 122, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305018.

Fuentes-Sepulveda:2020:PCB

[235] José Fuentes-Sepúlveda, Gonzalo Navarro, and Yakov Nekrich. Parallel computation of the Burrows Wheeler Transform in compact space. Theoretical Computer Science, 812 (??):123–136, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930581X.

Ryu:2020:FSM

[236] Cheol Ryu, Thierry Lecroq, and Kunsoo Park. Fast string matching for DNA sequences. Theoretical Computer Science, 812(??):137– 148, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305821.

Munro:2020:RDS

[237] J. Ian Munro, Gonzalo Navarro, Rahul Shah, and Sharma V. Thankachan.

Ranked document selection. Theoretical Computer Science, 812(??): 149–159, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306310.

Narisada:2020:ECL

[238] Shintaro Narisada, Diptarama Hendrian, Kazuyuki Narisawa, Shunsuke Inenaga, and Ayumi Shinohara. Efficient computation of longest singlearm-gapped palindromes in a string. Theoretical Computer Science, 812 (??):160–173, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306541.

Gawrychowski:2020:URS

[239] Paweł Gawrychowski, Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Universal reconstruction of a string. Theoretical Computer Science, 812 (??):174–186, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306565.

Narisada:2020:LTO

[240] Shintaro Narisada, Diptarama Hendrian, Ryo Yoshinaka, and Ayumi Shinohara. Linear-time online algorithm for inferring the shortest path graph from a walk label. Theoretical Computer Science, 812(??): 187–202, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306589.

Panahi:2020:API

[241] Ashkan Panahi, Morteza Haghir Chehreghani, and Devdatt Dubhashi. Accelerated proximal incremental algorithm schemes for nonstrongly convex functions. Theoretical Computer Science, 812(??): 203–213, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306590.

Egidi:2020:LMC

[242] Lavinia Egidi and Giovanni Manzini. Lightweight merging of compressed indices based on BWT variants. Theoretical Computer Science, 812(??): 214–229, April 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307005.

Giancarlo:2020:ABA

[243] Raffaele Giancarlo, Giovanni Manzini, Antonio Restivo, Giovanna Rosone, and Marinella Sciortino. The Alternating BWT: an algorithmic perspective. Theoretical Computer Science, 812(??):230–243, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307017.

Ayad:2020:LPP

[244] Lorraine A. K. Ayad, Giulia Bernardini, Roberto Grossi, Costas S. Iliopoulos, Nadia Pisanti, Solon P. Pissis, and Giovanna Rosone. Longest property-preserved common factor: a new string-processing framework. Theoretical Computer Science, 812 (??):244–251, April 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300967.

Anonymous:2020:PAa

[245] Anonymous. Pages 1–452 (12 April 2020). Theoretical Computer Science, 813(??):1–452, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBm

[246] Anonymous. Editorial Board. Theoretical Computer Science, 813(??): ii, April 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301201.

Kapron:2020:TTP

[247] Bruce M. Kapron and Florian Steinberg. Type-two polynomial-time and restricted lookahead. Theoretical Computer Science, 813(??): 1–19, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304293.

Castiglioni:2020:MLT

[248] Valentina Castiglioni, Michele Loreti, and Simone Tini. The metric lineartime branching-time spectrum on nondeterministic probabilistic processes. Theoretical Computer Science, 813 (??):20–69, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305638.

Baillot:2020:CLL

[249] Patrick Baillot and Alexis Ghyselen. Combining linear logic and size types for implicit complexity. Theoretical Computer Science, 813(??): 70–99, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305833.

DalLago:2020:RSC

[250] Ugo Dal Lago and Gabriele Vanoni. On randomised strategies in the λ-calculus. Theoretical Computer Science, 813(??): 100–116, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305845.

Laurent:2020:PTU

[251] Olivier Laurent. Polynomial time in untyped elementary linear logic. Theoretical Computer Science, 813(??): 117–142, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306255.

Matos:2020:FPP

[252] Armando B. Matos, Luca Paolini, and Luca Roversi. The fixed point problem of a simple reversible language. Theoretical Computer Science, 813(??): 143–154, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306280.

Giordano:2020:APS

[253] Laura Giordano and Alberto Policriti. Adding the power-set to description logics. Theoretical Computer Science, 813(??):155–174, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519306991.

Bersani:2020:ICT

[254] Marcello M. Bersani, Matteo Rossi, and Pierluigi San Pietro. On the initialization of clocks in timed formalisms. Theoretical Computer Science, 813(??): 175–198, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307492.

Bozzelli:2020:ATT

[255] Laura Bozzelli, Aniello Murano, and Loredana Sorrentino. Alternating-time temporal logics with linear past. Theoretical Computer Science, 813(??): 199–217, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307546.

Paolini:2020:CRP

[256] Luca Paolini, Mauro Piccolo, and Luca Roversi. A class of Recursive Permutations which is Primitive Recursive complete. Theoretical Computer Science, 813(??):

218–233, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307558.

DalLago:2020:EAS

[257] Ugo Dal Lago, Francesco Gavazzo, and Ryo Tanaka. Effectful applicative similarity for call-by-name lambda calculi. Theoretical Computer Science, 813(??): 234–247, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308072.

Spalazzi:2020:PMC

[258] Luca Spalazzi and Francesco Spegni. Parameterized model checking of networks of timed automata with Boolean guards. Theoretical Computer Science, 813(??):248–269, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519308084.

Barozzini:2020:BRL

[259] David Barozzini, David de Frutos-Escrig, Dario Della Monica, Angelo Montanari, and Pietro Sala. Beyond ω-regular languages: ωT - regular expressions and their automata and logic counterparts. Theoretical Computer Science, 813(??): 270–304, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308114.

Bozzelli:2020:TBP

[260] Laura Bozzelli, Alberto Molinari, Angelo Montanari, Adriano Peron, and Gerhard Woeginger. Timeline-based planning over dense temporal domains. Theoretical Computer Science, 813(??): 305–326, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519308126.

Fujita:2020:FSR

[261] Ken etsu Fujita. A formal system of reduction paths for parallel reduction. Theoretical Computer Science, 813(??): 327–340, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300104.

Couceiro:2020:ENF

[262] Miguel Couceiro, Erkko Lehtonen, Pierre Mercuriali, and Romain Péchoux. On the efficiency of normal form systems for representing Boolean functions. Theoretical Computer Science, 813(??):341–361, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752030030X.

Kernberger:2020:EPH

[263] Daniel Kernberger and Martin Lange. On the expressive power of hybrid branching-time logics. Theoretical Computer Science, 813(??): 362–374, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300359.

Aschieri:2020:CCC

[264] Federico Aschieri, Agata Ciabattoni, and Francesco A. Genco. On the concurrent computational content of intermediate logics. Theoretical Computer Science, 813(??): 375–409, April 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300505.

Wang:2020:LLC

[265] Longchun Wang and Qingguo Li. A logic for Lawson compact algebraic L-domains. Theoretical Computer Science, 813(??):410–427, April 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300591.

Bolotov:2020:BTL

[266] Alexander Bolotov, Montserrat Hermo, and Paqui Lucio. Branching-time logic ECTL# and its tree-style onepass tableau: Extending fairness expressibility of ECTL+. Theoretical Computer Science, 813(??):428– 451, April 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300992.

Anonymous:2020:PAb

[267] Anonymous. Pages 1–270 (24 April 2020). Theoretical Computer Science, 814(??):1–270, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBn

[268] Anonymous. Editorial Board. Theoretical Computer Science, 814(??): ii, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301353.

Gobonamang:2020:CBS

[269] Tshepo Kitso Gobonamang and Dimane Mpoeleng. Counter based suffix tree for DNA pattern repeats. Theoretical Computer Science, 814 (??):1–12, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307893.

Holub:2020:PSW

[270] Stepán Holub. Pseudo-solutions of word equations. Theoretical Computer Science, 814(??):13–18, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300050.

Guo:2020:CDN

[271] Chen Guo, Zhifang Xiao, Zhihong Liu, and Shuo Peng. R_{g} conditional diagnosability: a novel generalized measure of system-level diagnosis. Theoretical Computer Science, 814 (??):19–27, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300207.

Galby:2020:SDN

[272] Esther Galby, Andrea Munaro, and Bernard Ries. Semitotal domination: New hardness results and a polynomial-time algorithm for graphs of bounded mim-width. Theoretical Computer Science, 814(??):28– 48, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300219.

Zhu:2020:RAE

[273] Yuqing Zhu, Jiejing Wen, Jincheng Zhuang, Chang Lv, and Dongdai Lin. Refined analysis to the extended tower number field sieve. Theoretical Computer Science, 814(??): 49–68, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300311.

Bang-Jensen:2020:DLP

[274] J. Bang-Jensen, T. Bellitto, W. Lochet, and A. Yeo. The directed 2-linkage problem with length constraints. Theoretical Computer Science, 814(??):69–73, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300335.

Hajiaghayi:2020:AAC

[275] MohammadTaghi Hajiaghayi, Guy Kortsarz, Robert MacDavid, Manish Purohit, and Kanthi Sarpatwar. Approximation algorithms for connected maximum cut and related problems. Theoretical Computer Science, 814 (??):74–85, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300438.

Bredereck:2020:MIP

[276] Robert Bredereck, Piotr Faliszewski, Rolf Niedermeier, Piotr Skowron, and Nimrod Talmon. Mixed integer programming with convex/ concave constraints: Fixed-parameter tractability and applications to multicovering and voting. Theoretical Computer Science, 814(??):86– 105, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030044X.

Ohta:2020:NRC

[277] Satoru Ohta. The number of rearrangements for Clos networks — new results. Theoretical Computer Science, 814(??): 106–119, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300451.

Lee:2020:BDD

[278] Jia Lee, Ya-Hui Ye, Xin Huang, and Rui-Long Yang. Binarydecision-diagram-based decomposition of Boolean functions into reversible logic elements. Theoretical Computer Science, 814(??):120– 134, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300463.

Baste:2020:HMB

[279] Julien Baste, Ignasi Sau, and Dimitrios M. Thilikos. Hitting minors on bounded treewidth graphs. II. Single-exponential algorithms. Theoretical Computer Science, 814(??): 135–152, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300608.

Waniek:2020:CAO

[280] Marcin Waniek, Khaled Elbassioni, Flávio L. Pinheiro, César A. Hidalgo, and Aamena Alshamsi. Computational aspects of optimal strategic network diffusion. Theoretical Computer Science, 814(??):153– 168, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030061X.

Defrain:2020:DLG

[281] Oscar Defrain and Lhouari Nourine. Dualization in lattices given by implicational bases. Theoretical Computer Science, 814(??):169–176, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300621.

Papadopoulos:2020:SFV

[282] Charis Papadopoulos and Spyridon Tzimas. Subset feedback vertex set on graphs of bounded independent set size. Theoretical Computer Science, 814(??):177–188, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300633.

Juma:2020:CCA

[283] Nahid Juma, Werner Dietl, and Mahesh Tripunitara. A computational complexity analysis of tunable type inference for Generic Universe Types. Theoretical Computer Science, 814(??): 189–209, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300761.

Sanyal:2020:EMD

[284] Biswajit Sanyal, Subhashis Majumder, Wing-Kai Hon, and Prosenjit Gupta. Efficient meta-data structure in topk queries of combinations and multiitem procurement auctions. Theoretical Computer Science, 814(??): 210–222, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300773.

Tian:2020:LRB

[285] Yangguang Tian, Yingjiu Li, Binanda Sengupta, Nan Li, and Chunhua Su. Leakage-resilient biometricbased remote user authentication with fuzzy extractors. Theoretical Computer Science, 814(??):223– 233, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300785.

Ailon:2020:PME

[286] Nir Ailon. Paraunitary matrices, entropy, algebraic condition number and Fourier computation. Theoretical Computer Science, 814(??): 234–248, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300797.

Baril:2020:PPE

[287] Jean-Luc Baril and Vincent Vajnovszki. Popularity of patterns over d-equivalence classes of words and permutations. Theoretical Computer Science, 814(??):249–258, April 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030089X.

Keikha:2020:FPT

[288] Vahideh Keikha, Maarten Löffler, and Ali Mohades. A fully polynomial time approximation scheme for the smallest diameter of imprecise points. Theoretical Computer Science, 814(??): 259–270, April 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300906.

Anonymous:2020:PMa

[289] Anonymous. Pages 1–310 (2 May 2020). Theoretical Computer Science, 815(??):1–310, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBo

[290] Anonymous. Editorial Board. Theoretical Computer Science, 815(??): ii, May 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301444.

Cosmadakis:2020:RBA

[291] Stavros Cosmadakis, Dimitris Kavvadias, and Lina Panagopoulou. Resolution based algorithms for the transversal hypergraph generation problem. Theoretical Computer Science, 815 (??):1–10, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301316.

Sun:2020:RIB

[292] Yinxia Sun, Yi Mu, Willy Susilo, Futai Zhang, and Anmin Fu. Revocable identity-based encryption with server-aided ciphertext evolution. Theoretical Computer Science, 815(??): 11–24, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301298.

Ali:2020:FDH

[293] Mohammad Ali, Javad Mohajeri, Mohammad-Reza Sadeghi, and Ximeng Liu. A fully distributed hierarchical attribute-based encryption scheme. Theoretical Computer Science, 815(??):25–46, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301286.

Wang:2020:TLB

[294] Di Wang and Jinhui Xu. Tight lower bound of sparse covariance matrix estimation in the local differential privacy model. Theoretical Computer Science, 815(??):47–59, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301250.

Chang:2020:ETN

[295] Huilan Chang and Han-Min Chu. Error-tolerant nonadaptive interval group testing with density-based tests. Theoretical Computer Science, 815 (??):60–68, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301249.

Lo:2020:DBA

[296] Shou-Fu Lo, Kuo-Tsung Tseng, Chang-Biau Yang, and Kuo-Si Huang. A diagonal-based algorithm for the longest common increasing subsequence problem. Theoretical Computer Science, 815(??):69–78, May 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301158.

Hsu:2020:SGG

[297] Wei-Yuan Hsu, Chu-Ling Ko, Jr-Chang Chen, Ting-Han Wei, Chu-Hsuan Hsueh, and I-Chen Wu. On solving the 7, 7, 5-game and the 8, 8, 5game. Theoretical Computer Science, 815(??):79–94, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301146.

Navarrete:2020:PRE

[298] Lise Rommel Romero Navarrete and Guilherme P. Telles. Practical regular expression constrained sequence alignment. Theoretical Computer Science, 815(??):95–108, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301018.

Xiao:2020:IPA

[299] Mingyu Xiao and Zimo Sheng. Improved parameterized algorithms and kernels for mixed domination. Theoretical Computer Science, 815(??): 109–120, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520300980.

Bodewes:2020:RHG

[300] Jelco M. Bodewes, Hans L. Bodlaender, Gunther Cornelissen, and Marieke van der Wegen. Recognizing hyperelliptic graphs in polynomial time. Theoretical Computer Science, 815(??): 121–146, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300979.

Darmann:2020:SHV

[301] Andreas Darmann and Janosch Döcker. On a simple hard variant of Not-All-Equal 3-Sat. Theoretical Computer Science, 815(??):147–152, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520300943.

Shiraga:2020:CTD

[302] Takeharu Shiraga. The cover time of deterministic random walks for general transition probabilities. Theoretical Computer Science, 815(??): 153–162, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300931.

Mahdavi:2020:COPa

[303] Salma Sadat Mahdavi, Saeed Seddighin, and Mohammad Ghodsi. Covering orthogonal polygons with sliding k-transmitters. Theoretical Computer Science, 815(??):163– 181, May 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030092X.

Jaffke:2020:CDC

[304] Lars Jaffke and Paloma T. Lima. A complexity dichotomy for critical values of the b-chromatic number of graphs. Theoretical Computer Science, 815(??):182–196, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300918.

Akagi:2020:HEI

[305] Javier T. Akagi, Carlos F. Gaona, Fabricio Mendoza, Manjil P. Saikia, and Marcos Villagra. Hard and easy instances of L-tromino tilings. Theoretical Computer Science, 815(??): 197–212, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030116X.

Bose:2020:APF

[306] Kaustav Bose, Ranendu Adhikary, Manash Kumar Kundu, and Buddhadeb Sau. Arbitrary pattern formation on infinite grid by asynchronous oblivious robots. Theoretical Computer Science, 815(??): 213–227, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301006.

Calamoneri:2020:VCP

[307] Tiziana Calamoneri, Valentino Di Donato, Diego Mariottini, and Maurizio Patrignani. Visualizing cophylogenetic reconciliations. Theoretical Computer Science, 815(??): 228–245, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308060.

Murano:2020:P

[308] Aniello Murano and Sasha Rubin. Preface. Theoretical Computer Science, 815(??):246, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519308047.

Gigante:2020:TBG

[309] Nicola Gigante, Angelo Montanari, Andrea Orlandini, Marta Cialdea Mayer, and Mark Reynolds. On timeline-based games and their complexity. Theoretical Computer Science, 815(??): 247–269, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300955.

Barbay:2020:CCK

[310] Jérémy Barbay, Pablo Pérez-Lantero, and Javiel Rojas-Ledesma. Computing coverage kernels under restricted settings. Theoretical Computer Science, 815(??):270–288, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300499.

Castenow:2020:GAO

[311] Jannik Castenow, Matthias Fischer, Jonas Harbig, Daniel Jung, and Friedhelm Meyer auf der Heide. Gathering anonymous, oblivious robots on a grid. Theoretical Computer Science, 815(??):289–309, May 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030102X.

Aldini:2020:ICT

[312] Alessandro Aldini and Marco Bernardo. The Italian Conference on Theoretical Computer Science. Theoretical Computer Science, 815(??): 310, May 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301407.

Anonymous:2020:PMb

[313] Anonymous. Pages 1–270 (6 May 2020). Theoretical Computer Science, 816(??):1–270, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBp

[314] Anonymous. Editorial Board. Theoretical Computer Science, 816(??): ii, May 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030150X.

Kari:2020:Ec

[315] Lila Kari.  Editorial.  Theo-

retical Computer Science, 816(??): iii, May 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301535.

Pan:2020:CLS

[316] Linqiang Pan, David Orellana-Martín, Bosheng Song, and Mario J. Pérez-Jiménez. Cell-like P systems with polarizations and minimal rules. Theoretical Computer Science, 816(??):

1–18, May 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306231.

Paul:2020:DLD

[317] Prithwineel Paul and Kumar Sankar Ray. Derivation languages and descriptional complexity measures of restricted flat splicing systems. Theoretical Computer Science, 816(??): 19–36, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306267.

Gajardo:2020:PEC

[318] A. Gajardo, V. Nesme, and G. Theyssier. Pre-expansivity in cellular automata. Theoretical Computer Science, 816 (??):37–66, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930670X.

Gyongyosi:2020:MCV

[319] Laszlo Gyongyosi. Multicarrier continuous-variable quantum key distribution. Theoretical Computer Science, 816(??):67–95, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519307522.

Kotzing:2020:DLP

[320] Timo Kötzing, J. A. Gregor Lagodzinski, Johannes Lengler, and Anna Melnichenko. Destructiveness of lexicographic parsimony pressure and alleviation by a concatenation crossover in genetic programming. Theoretical Computer Science, 816(??):96– 113, May 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307625.

Klinge:2020:RBF

[321] Titus H. Klinge, James I. Lathrop, and Jack H. Lutz. Robust biomolecular finite automata. Theoretical Computer Science, 816(??): 114–143, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300293.

Doerr:2020:ILP

[322] Benjamin Doerr, Timo Kötzing, J. A. Gregor Lagodzinski, and Johannes Lengler. The impact of lexicographic parsimony pressure for OR-DER/MAJORITY on the run time. Theoretical Computer Science, 816 (??):144–168, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300323.

McGeoch:2020:TVP

[323] Catherine C. McGeoch. Theory versus practice in annealingbased quantum computing. Theoretical Computer Science, 816(??): 169–183, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300529.

Facca:2020:CNU

[324] Enrico Facca, Andreas Karrenbauer, Pavel Kolev, and Kurt Mehlhorn. Convergence of the non-uniform directed Physarum model. Theoretical Computer Science, 816(??):184– 194, May 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030075X.

Liu:2020:BPU

[325] Hai Liu, Zhenqiang Wu, Changgen Peng, Feng Tian, and Laifeng Lu. Bounded privacy-utility monotonicity indicating bounded tradeoff of differential privacy mechanisms. Theoretical Computer Science, 816(??): 195–220, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300888.

Bortolussi:2020:FAB

[326] Luca Bortolussi, Jane Hillston, and Michele Loreti. Fluid approximation of broadcasting systems. Theoretical Computer Science, 816(??): 221–248, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301110.

Zhang:2020:RAS

[327] Qifan Zhang, Liqiong Xu, Shuming Zhou, and Weihua Yang. Reliability analysis of subsystem in dual cubes. Theoretical Computer Science, 816 (??):249–259, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301262.

Karrenbauer:2020:CNU

[328] Andreas Karrenbauer, Pavel Kolev, and Kurt Mehlhorn. Convergence of the non-uniform Physarum dynamics. Theoretical Computer Science, 816 (??):260–269, May 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301304.

Anonymous:2020:PMc

[329] Anonymous. Pages 1–82 (12 May 2020). Theoretical Computer Science, 817(??):1–82, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBq

[330] Anonymous. Editorial Board. Theoretical Computer Science, 817(??): ii, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301730.

Kari:2020:Ed

[331] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 817(??):iii, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301766.

Zhu:2020:LDC

[332] Junlei Zhu and Yuehua Bu. List rdynamic coloring of sparse graphs. Theoretical Computer Science, 817 (??):1–11, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303810.

Yan:2020:IAG

[333] Chun Yan, Meixuan Li, Wei Liu, and Man Qi. Improved adaptive genetic algorithm for the vehicle Insurance Fraud Identification Model based on a BP Neural Network. Theoretical Computer Science, 817(??): 12–23, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304177.

Smiti:2020:DMC

[334] Abir Smiti and Zied Elouedi. Dynamic maintenance case base using knowledge discovery techniques for case based reasoning systems. Theoretical Computer Science, 817(??): 24–32, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304189.

Zhang:2020:DTF

[335] Yunong Zhang, Huanchang Huang, Min Yang, and Jian Li. Discretetime formulation, control, solution and verification of pendulum systems with zeroing neural dynamics. Theoretical Computer Science, 817(??):

33–43, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304190.

Zhang:2020:MES

[336] Yunong Zhang, Min Yang, Binbin Qiu, Jian Li, and Mingjie Zhu. From mathematical equivalence such as Ma equivalence to generalized Zhang equivalency including gradient equivalency. Theoretical Computer Science, 817 (??):44–54, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304621.

Hu:2020:KPK

[337] Juan Hu, Hong Peng, Jun Wang, and Wenping Yu. kNN-P: a kNN classifier optimized by P systems. Theoretical Computer Science, 817 (??):55–65, May 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300098.

Basak:2020:FTM

[338] Mithun Basak, Laxman Saha, Gour Kanta Das, and Kalishankar Tiwary. Faulttolerant metric dimension of circulant graphs C_{n}(1, 2, 3). Theoretical Computer Science, 817(??):66– 79, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300246.

Dang:2020:ECC

[339] Chuangyin Dang and Yinyu Ye. Erratum/correction to “On the complexity of an expanded Tarski’s fixed point problem under the componentwise ordering” [Theor. Comput. Sci. 732 (2018) 26–45]. Theoretical Computer Science, 817(??): 80, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301677. See

[2].

Blanqui:2020:CID

[340] Frédéric Blanqui, Jean-Pierre Jouannaud, and Mitsuhiro Okada. Corrigendum to “Inductive-data-type systems” [Theoret. Comput. Sci. 272 (1–2) (2002) 41–68]. Theoretical Computer Science, 817(??):81– 82, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300306. See

[1].

Anonymous:2020:EBr

[341] Anonymous. Editorial Board. Theoretical Computer Science, 818(??): ii, May 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301869.

Cao:2020:PSI

[342] Yixin Cao. Preface to the special issue on Computing and Combinatorics. Theoretical Computer Science, 818(??):1, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301821.

Conte:2020:EEM

[343] Alessio Conte, Mamadou Moustapha Kanté, Yota Otachi, Takeaki Uno, and Kunihiro Wasa. Efficient enumeration of maximal k-degenerate induced subgraphs of a chordal graph. Theoretical Computer Science, 818 (??):2–11, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305255.

Feng:2020:IKM

[344] Qilong Feng, Senmin Zhu, and Jianxin Wang. An improved kernel for Max-Bisection above tight lower bound. Theoretical Computer Science, 818 (??):12–21, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304432.

Koivisto:2020:NCR

[345] Mikko Koivisto, Petteri Laakkonen, and Juho Lauri. NP-completeness results for partitioning a graph into total dominating sets. Theoretical Computer Science, 818(??):22– 31, May 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302202.

Li:2020:MDC

[346] Qian Li and Xiaoming Sun. On the modulo degree complexity of Boolean functions. Theoretical Computer Science, 818(??):32–40, May 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303037.

Mehta:2020:ICE

[347] Ruta Mehta and Vijay V. Vazirani. An incentive compatible, efficient market for air traffic flow management. Theoretical Computer Science, 818 (??):41–50, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305711.

Misra:2020:LRT

[348] Pranabendu Misra, Fahad Panolan, M. S. Ramanujan, and Saket Saurabh. Linear representation of transversal matroids and gammoids parameterized by rank. Theoretical Computer Science, 818(??):51–59, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301269.

Shenmaier:2020:CAF

[349] Vladimir Shenmaier. Complexity and algorithms for finding a subset of vectors with the longest sum. Theoretical Computer Science, 818(??): 60–73, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302408.

Wu:2020:SMT

[350] Guangwei Wu, Jianer Chen, and Jianxin Wang. On scheduling multiple two-stage flowshops. Theoretical Computer Science, 818(??): 74–82, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302391.

Xu:2020:EUA

[351] Jian Xu, Soumya Banerjee, and Wenjing Rao. The existence of universally agreed fairest semi-matchings in any given bipartite graph. Theoretical Computer Science, 818(??): 83–91, May 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301816.

Anonymous:2020:EBs

[352] Anonymous. Editorial Board. Theoretical Computer Science, 819(??): ii, June 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301936.

Miao:2020:FDR

[353] Dongjing Miao, Zhipeng Cai, Xianmin Liu, and Jianzhong Li. Functional dependency restricted insertion propagation. Theoretical Computer Science, 819(??):1–8, June 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302827.

Das:2020:PTA

[354] Bireswar Das, Murali Krishna Enduri, and I. Vinod Reddy. Polynomialtime algorithm for isomorphism of graphs with clique-width at most three. Theoretical Computer Science, 819 (??):9–23, June 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306758.

Chau:2020:FSD

[355] Vincent Chau, Xin Chen, Ken C. K. Fong, Minming Li, and Kai Wang. Flow shop for dual CPUs in dynamic voltage scaling. Theoretical Computer Science, 819(??):24– 34, June 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730676X.

Lu:2020:VHU

[356] Xu Lu, Zhenhua Duan, Cong Tian, and Hongwei Du. Verify heaps via unified model checking. Theoretical Computer Science, 819(??): 35–49, June 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306874.

Shu:2020:DPC

[357] Xinfeng Shu, Zhenhua Duan, and Hongwei Du. A decision procedure and complete axiomatization for projection temporal logic. Theoretical Computer Science, 819(??): 50–84, June 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306886.

Anonymous:2020:PJf

[358] Anonymous. Pages 1–122 (8 June 2020). Theoretical Computer Science, 820(??):1–122, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBt

[359] Anonymous. Editorial Board. Theoretical Computer Science, 820(??): ii, June 8, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302061.

Georgiou:2020:LPS

[360] Konstantinos Georgiou, Andy (Jia) Jiang, Edward Lee, Astrid A. Olave, Ian Seong, and Twesh Upadhyaya. Lift and project systems performing on the partial-vertex-cover polytope. Theoretical Computer Science, 820(??):1–16, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301559.

Lingas:2020:SNC

[361] Andrzej Lingas. Small normalized circuits for semi-disjoint bilinear forms require logarithmic and-depth. Theoretical Computer Science, 820(??): 17–25, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301560.

Zhang:2020:MCA

[362] Yu an Zhang, Xiaofeng Qin, Qinglian Ma, Minghao Zhao, Satoru Hiwa, Tomoyuki Hiroyasu, and Hiroshi Furutani. Markov chain analysis of evolutionary algorithms on OneMax function — from coupon collector’s problem to (1 + 1) EA. Theoretical Computer Science, 820(??):26– 44, June 8, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301584.

Bamas:2020:LAM

[363] Étienne Bamas and Louis Esperet. Local approximation of the Maximum Cut in regular graphs. Theoretical Computer Science, 820(??): 45–59, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301596.

Yamakawa:2020:GHI

[364] Takashi Yamakawa, Shota Yamada, Goichiro Hanaoka, and Noboru Kunihiro. Generic hardness of inversion on ring and its relation to self-bilinear map. Theoretical Computer Science, 820(??):60–84, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301602.

Kuo:2020:FFC

[365] Che-Nan Kuo and Yu-Huei Cheng. Fault-free cycles embedding in folded hypercubes with F_{4}. Theoretical Computer Science, 820(??):85– 90, June 8, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301663.

Ateniese:2020:SRS

[366] Giuseppe Ateniese, Bernardo Magri, and Daniele Venturi. Subversionresilient signatures: Definitions, constructions and applications. Theoretical Computer Science, 820(??): 91–122, June 8, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301808.

Anonymous:2020:PJg

[367] Anonymous. Pages 1–124 (12 June 2020). Theoretical Computer Science, 821(??):1–124, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBu

[368] Anonymous. Editorial Board. Theoretical Computer Science, 821(??): ii, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030219X.

Wang:2020:CST

[369] Kai Wang. Calibration scheduling with time slot cost. Theoretical Computer Science, 821(??):

1–14, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301699.

Zhang:2020:OOA

[370] Yong Zhang, Francis Y. L. Chin, Sheung-Hung Poon, Hing-Fung Ting, Dachuan Xu, and Dongxiao Yu. Offline and online algorithms for singleminded selling problem. Theoretical Computer Science, 821(??):15– 22, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301687.

Gao:2020:IAI

[371] Chuangen Gao, Shuyang Gu, Ruiqi Yang, Jiguo Yu, Weili Wu, and Dachuan Xu. Interaction-aware influence maximization and iterated sandwich method. Theoretical Computer Science, 821(??):23– 33, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301675.

Page:2020:MMUb

[372] Daniel R. Page and Roberto Solis-Oba. Makespan minimization on unrelated parallel machines with a few bags. Theoretical Computer Science, 821(??):34–44, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030164X.

Farago:2020:NAD

[373] András Faragó and Rupei Xu. A new algorithm design technique for hard problems. Theoretical Computer Science, 821(??):45–56, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301638.

Lidbetter:2020:SGH

[374] Thomas Lidbetter and Kyle Y. Lin. A search game on a hypergraph with booby traps. Theoretical Computer Science, 821(??): 57–70, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301626.

Park:2020:CIG

[375] Jung-Heum Park and Hyeong-Seok Lim. Characterization of interval graphs that are unpaired 2disjoint path coverable. Theoretical Computer Science, 821(??):71– 86, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301614.

Gangloff:2020:AGR

[376] Silvère Gangloff and Alexandre Talon. Asymptotic growth rate of square grids dominating sets: a symbolic dynamics approach. Theoretical Computer Science, 821(??): 87–101, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301572.

Arora:2020:LCP

[377] Pranav Arora, Aritra Banik, Vijay Kumar Paliwal, and Venkatesh Raman. List-coloring — parameterizing from triviality. Theoretical Computer Science, 821(??):102– 110, June 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301134.

Balaji:2020:SCS

[378] Nikhil Balaji, Andreas Krebs, and Nutan Limaye. Skew circuits of small width. Theoretical Computer Science, 821(??):111–123, June 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517302074.

Anonymous:2020:PJh

[379] Anonymous. Pages 1–110 (24 June 2020). Theoretical Computer Science, 822(??):1–110, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBv

[380] Anonymous. Editorial Board. Theoretical Computer Science, 822(??): ii, June 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302371.

Banerjee:2020:FDA

[381] Niranka Banerjee, Venkatesh Raman, and Saket Saurabh. Fully dynamic arboricity maintenance. Theoretical Computer Science, 822(??): 1–14, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302152.

Abedin:2020:LSD

[382] Paniz Abedin, Arnab Ganguly, Wing-Kai Hon, Kotaro Matsuda, Yakov Nekrich, Kunihiko Sadakane, Rahul Shah, and Sharma V. Thankachan. A linear-space data structure for range-LCP queries in poly-logarithmic time. Theoretical Computer Science, 822 (??):15–22, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302139.

Nguyen:2020:CGK

[383] Viet-Ha Nguyen, Kévin Perrot, and Mathieu Vallet. NP-completeness of the game Kingdomino^{T M}. Theoretical Computer Science, 822(??): 23–35, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302115.

Zhang:2020:WMB

[384] Liyu Zhang, Mahmoud Quweider, Hansheng Lei, and Fitra Khan. Weak mitoticity of bounded disjunctive and conjunctive truth-table autoreducible sets. Theoretical Computer Science, 822(??):36–48, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302024.

Chen:2020:CCG

[385] Wei Chen, Xiaohan Shan, Xiaoming Sun, and Jialin Zhang. Coreness of cooperative games with truncated submodular profit functions. Theoretical Computer Science, 822 (??):49–60, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302012.

Anthony:2020:EDP

[386] Barbara M. Anthony and Christine Chung. Equilibria in Doodle polls under three tie-breaking rules. Theoretical Computer Science, 822(??): 61–71, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302000.

Docker:2020:PQV

[387] Janosch Döcker, Britta Dorn, Simone Linz, and Charles Semple. Placing quantified variants of 3-SAT and Not-All-Equal 3-SAT in the polynomial hierarchy. Theoretical Computer Science, 822(??):72–91, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301973.

Shibata:2020:MOP

[388] Masahiro Shibata, Norikazu Kawata, Yuichi Sudo, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Move-optimal partial gathering of mobile agents without identifiers or global knowledge in asynchronous unidirectional rings. Theoretical Computer Science, 822(??): 92–109, June 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301985.

Anonymous:2020:PJi

[389] Anonymous. Pages 1–100 (2 July 2020). Theoretical Computer Science, 823(??):1–100, July 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBw

[390] Anonymous. Editorial Board. Theoretical Computer Science, 823(??): ii, July 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302474.

Amato:2020:CSP

[391] Gianluca Amato, Maria Chiara Meo, and Francesca Scozzari. On collecting semantics for program analysis. Theoretical Computer Science, 823(??):1–25, July 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301122.

Blumensath:2020:BIM

[392] Achim Blumensath and Felix Wolf. Bisimulation invariant monadic-second order logic in the finite. Theoretical Computer Science, 823(??):26– 43, July 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301389.

Kawai:2020:PGD

[393] Tatsuji Kawai. Presenting de Groot duality of stably compact spaces. Theoretical Computer Science, 823 (??):44–68, July 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301390.

Im:2020:BTI

[394] Hyeonseung Im, Pierre Genevès, Nils Gesbert, and Nabil Layaïda. Backward type inference for XML queries. Theoretical Computer Science, 823 (??):69–99, July 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301791.

Anonymous:2020:PJj

[395] Anonymous. Pages 1–112 (6 July 2020). Theoretical Computer Science, 824–825(??):1–112, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBx

[396] Anonymous. Editorial Board. Theoretical Computer Science, 824–825 (??):ii, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302553.

Lopez-Ortiz:2020:PEW

[397] Alejandro López-Ortiz, Marc P. Renault, and Adi Rosén. Paid exchanges are worth the price. Theoretical Computer Science, 824–825 (??):1–10, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300724.

Barth:2020:EPI

[398] Dominik Barth, Moritz Beck, Titus Dose, Christian Glaßer, Larissa Michler, and Marc Technau. Emptiness problems for integer circuits. Theoretical Computer Science, 824–825 (??):11–35, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520301997.

Costa:2020:PCT

[399] Eurinardo Costa, Victor Lage Pessoa, Rudini Sampaio, and Ronan Soares. PSPACE-completeness of two graph coloring games. Theoretical Computer Science, 824–825 (??):36–45, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030181X.

Wang:2020:DTF

[400] Yihong Wang, Cheng-Kuan Lin, Xiaoyan Li, and Shuming Zhou. Diagnosability for two families of composition networks. Theoretical Computer Science, 824–825(??):46– 56, July 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302127.

Li:2020:LCF

[401] Pingshan Li and Min Xu. The largest component of faulty star graphs. Theoretical Computer Science, 824–825 (??):57–66, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302140.

Pan:2020:SCS

[402] Zhuowen Pan and Dongqin Cheng. Structure connectivity and substructure connectivity of the crossed cube. Theoretical Computer Science, 824–825 (??):67–80, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302255.

Lin:2020:RCG

[403] Limei Lin, Yanze Huang, Xiaoding Wang, and Li Xu. Restricted connectivity and good-neighbor diagnosability of split-star networks. Theoretical Computer Science, 824–825 (??):81–91, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302243.

Chen:2020:GTB

[404] Wei Chen, Shang-Hua Teng, and Hanrui Zhang. A graph-theoretical basis of stochastic-cascading network influence: Characterizations of influence-based centrality. Theoretical Computer Science, 824–825(??): 92–111, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302279.

Anonymous:2020:PJk

[405] Anonymous. Pages 1–112 (6 July 2020). Theoretical Computer Science, 824–825(??):1–112, July 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBy

[406] Anonymous. Editorial Board. Theoretical Computer Science, 826–827 (??):ii, July 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302668.

Fatourou:2020:P

[407] Panagiota Fatourou and Fernando Pedone. Preface. Theoretical Computer Science, 826–827(??):1, July 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302097.

Qadri:2020:VDC

[408] Ammar Qadri. On the viability of the deterministic consensus hierarchy. Theoretical Computer Science, 826–827 (??):2–11, July 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300811.

Chan:2020:BD

[409] David Yu Cheng Chan, Vassos Hadzilacos, and Sam Toueg. Bounded disagreement. Theoretical Computer Science, 826–827(??):12–24, July 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305978.

Comte:2020:KGU

[410] Céline Comte and Fabien Mathieu. Kleinberg’s grid unchained. Theoretical Computer Science, 826–827 (??):25–39, July 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305966.

Chen:2020:SOI

[411] Tian Ze Chen and Yuanhao Wei. Step-optimal implementations of large single-writer registers. Theoretical Computer Science, 826–827(??):40– 50, July 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302103.

Anonymous:2020:PAc

[412] Anonymous. Pages 1–80 (6 August 2020). Theoretical Computer Science, 828–829(??):1–80, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBz

[413] Anonymous. Editorial Board. Theoretical Computer Science, 828–829 (??):ii, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302784.

DaSilva:2020:FCPa

[414] Ian Albuquerque Raymundo Da Silva, Gruia Calinescu, and Nathan De Graaf. Faster compression of patterns to Rectangle Rule Lists. Theoretical Computer Science, 828–829 (??):1–18, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301651.

Rybarczyk:2020:GAGa

[415] Katarzyna Rybarczyk. GHS algorithm on a graph with random weights. Theoretical Computer Science, 828–829 (??):19–31, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302231.

Ni:2020:ICMa

[416] Qiufen Ni, Jianxiong Guo, Chuanhe Huang, and Weili Wu. Information coverage maximization for multiple products in social networks. Theoretical Computer Science, 828–829 (??):32–41, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302267.

Angelini:2020:RDGa

[417] Patrizio Angelini, Michael A. Bekos, Henry Förster, and Michael Kaufmann. On RAC drawings of graphs with one bend per edge. Theoretical Computer Science, 828–829(??): 42–54, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302280.

Chau:2020:MCBa

[418] Vincent Chau, Minming Li, Elaine Yinling Wang, Ruilong Zhang, and Yingchao Zhao. Minimizing the cost of batch calibrations. Theoretical Computer Science, 828–829(??): 55–64, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302309.

Chang:2020:UMSa

[419] Ching-Lueh Chang. On ultrametric 1-median selection. Theoretical Computer Science, 828–829(??): 65–69, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302334.

Huang:2020:NPEa

[420] Yanze Huang, Limei Lin, and Li Xu. A new proof for exact relationship between extra connectivity and extra diagnosability of regular connected graphs under MM* model. Theoretical Computer Science, 828–829(??): 70–80, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302310.

Anonymous:2020:PAd

[421] Anonymous. Pages 1–80 (6 August 2020). Theoretical Computer Science, 828–829(??):1–80, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBba

[422] Anonymous. Editorial Board. Theoretical Computer Science, 828–829 (??):ii, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302784.

DaSilva:2020:FCPb

[423] Ian Albuquerque Raymundo Da Silva, Gruia Calinescu, and Nathan De Graaf. Faster compression of patterns to Rectangle Rule Lists. Theoretical Computer Science, 828–829 (??):1–18, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301651.

Rybarczyk:2020:GAGb

[424] Katarzyna Rybarczyk. GHS algorithm on a graph with random weights. Theoretical Computer Science, 828–829 (??):19–31, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302231.

Ni:2020:ICMb

[425] Qiufen Ni, Jianxiong Guo, Chuanhe Huang, and Weili Wu. Information coverage maximization for multiple products in social networks. Theoretical Computer Science, 828–829 (??):32–41, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302267.

Angelini:2020:RDGb

[426] Patrizio Angelini, Michael A. Bekos, Henry Förster, and Michael Kaufmann. On RAC drawings of graphs with one bend per edge. Theoretical Computer Science, 828–829(??): 42–54, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302280.

Chau:2020:MCBb

[427] Vincent Chau, Minming Li, Elaine Yinling Wang, Ruilong Zhang, and Yingchao Zhao. Minimizing the cost of batch calibrations. Theoretical Computer Science, 828–829(??): 55–64, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302309.

Chang:2020:UMSb

[428] Ching-Lueh Chang. On ultrametric 1-median selection. Theoretical Computer Science, 828–829(??): 65–69, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302334.

Huang:2020:NPEb

[429] Yanze Huang, Limei Lin, and Li Xu. A new proof for exact relationship between extra connectivity and extra diagnosability of regular connected graphs under MM* model. Theoretical Computer Science, 828–829(??): 70–80, August 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302310.

Anonymous:2020:PAe

[430] Anonymous. Pages 1–90 (24 August 2020). Theoretical Computer Science, 830–831(??):1–90, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbb

[431] Anonymous. Editorial Board. Theoretical Computer Science, 830–831 (??):ii, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302930.

Kim:2020:FRR

[432] Gyu-Chol Kim, Su-Chol Li, and Hak-Chol Hwang. Fast rebalanced RSA signature scheme with typical prime generation. Theoretical Computer Science, 830–831(??): 1–19, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302413.

Mefenza:2020:ISP

[433] Thierry Mefenza and Damien Vergnaud. Inferring sequences produced by elliptic curve generators using Coppersmith’s methods. Theoretical Computer Science, 830–831(??):20– 42, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302401.

Almethen:2020:PLH

[434] Abdullah Almethen, Othon Michail, and Igor Potapov. Pushing lines helps: Efficient universal centralised transformations for programmable matter. Theoretical Computer Science, 830–831 (??):43–59, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302437.

Travnicek:2020:MBM

[435] Jan Trávníček, Jan Janoušek, Bořivoj Melichar, and Loek Cleophas. On modification of Boyer–Moore–Horspool’s algorithm for tree pattern matching in linearised trees. Theoretical Computer Science, 830–831(??):

60–90, August 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302425.

Anonymous:2020:EBbc

[436] Anonymous. Editorial Board. Theoretical Computer Science, 832(??): ii, September 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303066.

Kari:2020:Ee

[437] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 832(??):iii, September 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303091.

Oliveto:2020:TEC

[438] Pietro S. Oliveto and Andrew M. Sutton. Theory of evolutionary computation — Special Issue Editorial. Theoretical Computer Science, 832 (??):1–2, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303029.

Friedrich:2020:AES

[439] Tobias Friedrich, Timo Kötzing, J. A. Gregor Lagodzinski, Frank Neumann, and Martin Schirneck. Analysis of the (1 + 1) EA on subclasses of linear functions under uniform and linear constraints. Theoretical Computer Science, 832(??):

3–19, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303050.

Pourhassan:2020:RAR

[440] Mojgan Pourhassan, Vahid Roostapour, and Frank Neumann. Runtime analysis of RLS and (1 + 1) EA for the dynamic weighted vertex cover problem. Theoretical Computer Science, 832(??):20– 41, September 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301501.

Akimoto:2020:QGA

[441] Youhei Akimoto, Anne Auger, and Nikolaus Hansen. Quality gain analysis of the weighted recombination evolution strategy on general convex quadratic functions. Theoretical Computer Science, 832(??):42– 67, September 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303359.

Atamna:2020:ILC

[442] Asma Atamna, Anne Auger, and Nikolaus Hansen. On invariance and linear convergence of evolution strategies with augmented Lagrangian constraint handling. Theoretical Computer Science, 832(??): 68–97, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830611X.

Hellwig:2020:SSA

[443] Michael Hellwig and Hans-Georg Beyer. On the steady state analysis of covariance matrix self-adaptation evolution strategies on the noisy ellipsoid model. Theoretical Computer Science, 832(??):98–122, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397518303360.

Osuna:2020:DAD

[444] Edgar Covantes Osuna, Wanru Gao, Frank Neumann, and Dirk Sudholt. Design and analysis of diversitybased parent selection schemes for speeding up evolutionary multiobjective optimisation. Theoretical Computer Science, 832(??):123– 142, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304171.

Krejca:2020:LBR

[445] Martin S. Krejca and Carsten Witt. Lower bounds on the run time of the Univariate Marginal Distribution Algorithm on OneMax. Theoretical Computer Science, 832(??): 143–165, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304122.

Corus:2020:WHA

[446] Dogan Corus, Pietro S. Oliveto, and Donya Yazdani. When hypermutations and ageing enable artificial immune systems to outperform evolutionary algorithms. Theoretical Computer Science, 832(??): 166–185, September 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301495.

Anonymous:2020:PS

[447] Anonymous. Pages 1–172 (12 September 2020). Theoretical Computer Science, 833(??):1–172, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbd

[448] Anonymous. Editorial Board. Theoretical Computer Science, 833(??): ii, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303480.

Gu:2020:NMP

[449] Ran Gu, Yaping Mao, and Guoju Ye. Note on matching preclusion number of random graphs. Theoretical Computer Science, 833(??):1– 10, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307868.

Akitaya:2020:MCS

[450] Hugo A. Akitaya, Maarten Löffler, and Csaba D. Tóth. Multicolored spanning graphs. Theoretical Computer Science, 833(??):11– 25, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302322.

Dragan:2020:EFD

[451] Feodor F. Dragan and Heather M. Guarnera. Eccentricity function in distance-hereditary graphs. Theoretical Computer Science, 833(??):26– 40, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302607.

Hao:2020:RBE

[452] Rong-Xia Hao, Mei-Mei Gu, and Jou-Ming Chang. Relationship between extra edge connectivity and component edge connectivity for regular graphs. Theoretical Computer Science, 833(??): 41–55, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302693.

Tomida:2020:TSI

[453] Junichi Tomida. Tightly secure inner product functional encryption: Multi-input and functionhiding constructions. Theoretical Computer Science, 833(??):56– 86, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302711.

Ando:2020:FVS

[454] Ei Ando and Shuji Kijima. An FPTAS for the volume of some V-polytopes — It is hard to compute the volume of the intersection of two cross-polytopes. Theoretical Computer Science, 833(??): 87–106, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303157.

Gagie:2020:TPM

[455] Travis Gagie, Meng He, Gonzalo Navarro, and Carlos Ochoa. Tree path majority data structures. Theoretical Computer Science, 833(??):107– 119, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303248.

Grigoryan:2020:PAE

[456] Hayk A. Grigoryan and Samvel K. Shoukourian. Polynomial algorithm for equivalence problem of deterministic multitape finite automata. Theoretical Computer Science, 833(??):120– 132, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303297.

Bae:2020:MWD

[457] Sang Won Bae. Minimum-width double-strip and parallelogram annulus. Theoretical Computer Science, 833(??):133–146, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520303315.

Jedwab:2020:ICU

[458] Jonathan Jedwab, Tara Petrie, and Samuel Simon. An infinite class of unsaturated rooted trees corresponding to designable RNA secondary structures. Theoretical Computer Science, 833(??): 147–163, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303327.

Agrawal:2020:QVK

[459] Akanksha Agrawal, Sushmita Gupta, Pallavi Jain, and R. Krithika. Quadratic vertex kernel for split vertex deletion. Theoretical Computer Science, 833(??): 164–172, September 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303352.

Anonymous:2020:EBbe

[460] Anonymous. Editorial Board. Theoretical Computer Science, 834(??): ii, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303947.

DeFelice:2020:AL

[461] Clelia De Felice, Dominique Perrin, and Antonio Restivo. Aldo de Luca (1941–2018). Theoretical Computer Science, 834(??):1– 2, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303261.

Karhumaki:2020:CPA

[462] Juhani Karhumäki and Markus A. Whiteland. A compactness property of the k-abelian monoids. Theoretical Computer Science, 834(??):3– 13, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300517.

Carpi:2020:CML

[463] Arturo Carpi and Flavio D’Alessandro. Coding by minimal linear grammars. Theoretical Computer Science, 834(??): 14–25, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300736.

Kleijn:2020:PCP

[464] Jetty Kleijn, Maciej Koutny, and Grzegorz Rozenberg. Plug-in context providers for reaction systems. Theoretical Computer Science, 834(??):26– 42, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520300748.

Reghizzi:2020:DAL

[465] Stefano Crespi Reghizzi and Pierluigi San Pietro. Deque automata, languages, and planar graph representations. Theoretical Computer Science, 834(??):43–59, September 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301274.

Postic:2020:RLW

[466] Mickaël Postic and Luca Q. Zamboni. Reprint of: ω-Lyndon words. Theoretical Computer Science, 834(??):60– 65, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303224.

Burcroff:2020:RGL

[467] Amanda Burcroff and Eric Winsor. Reprint of: Generalized Lyndon factorizations of infinite words. Theoretical Computer Science, 834(??):66– 74, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030325X.

Abram:2020:RPC

[468] Antoine Abram, Mélodie Lapointe, and Christophe Reutenauer. Reprint of: Palindromization and construction of Markoff triples. Theoretical Computer Science, 834(??):75– 83, September 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303273.

Anonymous:2020:POa

[469] Anonymous. Pages 1–168 (2 October 2020). Theoretical Computer Science, 835(??):1–168, October 2, 2020.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbf

[470] Anonymous. Editorial Board. Theoretical Computer Science, 835(??): ii, October 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304382.

Lv:2020:RAA

[471] Mengjie Lv, Jianxi Fan, Guo Chen, Baolei Cheng, Jingya Zhou, and Jia Yu. The reliability analysis of k-ary n-cube networks. Theoretical Computer Science, 835(??): 1–14, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302619.

Song:2020:CBS

[472] Yongcheng Song, Xinyi Huang, Yi Mu, Wei Wu, and Huaxiong Wang. A code-based signature scheme from the Lyubashevsky framework. Theoretical Computer Science, 835(??): 15–30, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302723.

Tollis:2020:AVP

[473] Ioannis G. Tollis and Konstantinos G. Kakoulis. Algorithms for visualizing phylogenetic networks. Theoretical Computer Science, 835(??): 31–43, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303303.

Xu:2020:FTH

[474] Min Xu, Kshirasagar Naik, and Krishnaiyan Thulasiraman. Fault tolerance of hypercube like networks: Spanning laceability under edge faults. Theoretical Computer Science, 835(??): 44–57, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303339.

Agrawal:2020:PCH

[475] Akanksha Agrawal, N. R. Aravind, Subrahmanyam Kalyanasundaram, Anjeneya Swami Kare, Juho Lauri, Neeldhara Misra, and I. Vinod Reddy. Parameterized complexity of happy coloring problems. Theoretical Computer Science, 835(??): 58–81, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303364.

Wang:2020:PQH

[476] Luping Wang, Jie Chen, Kai Zhang, and Haifeng Qian. A post-quantum hybrid encryption based on QC-LDPC codes in the multi-user setting. Theoretical Computer Science, 835(??): 82–96, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303558.

Lubiw:2020:CTM

[477] Anna Lubiw and Debajyoti Mondal. On compatible triangulations with a minimum number of Steiner points. Theoretical Computer Science, 835(??): 97–107, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303546.

Ding:2020:CEC

[478] Tongtong Ding, Pingshan Li, and Min Xu. The component (edge) connectivity of shuffle-cubes. Theoretical Computer Science, 835(??): 108–119, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303583.

Cabello:2020:HMB

[479] Sergio Cabello and Éric Colin de Verdière. Hardness of Minimum Barrier Shrinkage and Minimum Installation Path. Theoretical Computer Science, 835(??): 120–133, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030356X.

Lintzmayer:2020:RAS

[480] Carla N. Lintzmayer, Flávio K. Miyazawa, Phablo F. S. Moura, and Eduardo C. Xavier. Randomized approximation scheme for Steiner Multi Cycle in the Euclidean plane. Theoretical Computer Science, 835(??): 134–155, October 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303649.

Perez-Diaz:2020:PRT

[481] Sonia Pérez-Díaz and Li-Yong Shen. Parameterization of rational translational surfaces. Theoretical Computer Science, 835(??):156–167, October 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730600X.

Anonymous:2020:POb

[482] Anonymous. Pages 1–124 (6 October 2020). Theoretical Computer Science, 836(??):1–124, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbg

[483] Anonymous. Editorial Board. Theoretical Computer Science, 836(??): ii, October 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304515.

Liu:2020:GSC

[484] Chuang Liu, Hongwei Du, Qiang Ye, and Wen Xu. Group sweep coverage with guaranteed approximation ratio. Theoretical Computer Science, 836 (??):1–15, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030284X.

Lee:2020:CDC

[485] Chia-Wei Lee. Conditional diagnosability of component-composition graphs under the PMC model. Theoretical Computer Science, 836(??): 16–28, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302838.

Chen:2020:GTA

[486] Xiuyang Chen and Zhao Zhang. A game theoretic approach for minimal connected dominating set. Theoretical Computer Science, 836(??): 29–36, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302966.

Wang:2020:MSD

[487] Ye Wang, Yusheng Li, and Yan Li. Maximum star deleted from Ramsey graphs of book and tree. Theoretical Computer Science, 836(??): 37–41, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303145.

Luo:2020:DRO

[488] Chuanwen Luo, Deying Li, Xingjian Ding, and Weili Wu. Delivery Route Optimization with automated vehicle in smart urban environment. Theoretical Computer Science, 836(??): 42–52, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303376.

Cui:2020:NVA

[489] Xuyang Cui, Wenshun Teng, Xing Liu, and Huijuan Wang. A note of vertex arboricity of planar graphs without 4-cycles intersecting with 6-cycles. Theoretical Computer Science, 836(??): 53–58, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303388.

Liu:2020:AAC

[490] Pengcheng Liu, Zhao Zhang, and Xiaohui Huang. Approximation algorithm for (connected) bounded-degree deletion problem on unit disk graphs. Theoretical Computer Science, 836(??): 59–64, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303613.

Dong:2020:SRP

[491] Luobing Dong, Qiumin Guo, Weili Wu, and Meghana N. Satpute. A semantic relatedness preserved subset extraction method for language corpora based on pseudo-Boolean optimization. Theoretical Computer Science, 836(??): 65–75, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304023.

Navon:2020:MFT

[492] Tal Navon and David Peleg. Mixed fault tolerance in server assignment:

Combining reinforcement and backup. Theoretical Computer Science, 836(??): 76–93, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303753.

Czyzowicz:2020:GPL

[493] Jurek Czyzowicz, Ryan Killick, Evangelos Kranakis, Danny Krizanc, and Oscar Morales-Ponce. Gathering in the plane of location-aware robots in the presence of spies. Theoretical Computer Science, 836(??):94– 109, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303741.

Godard:2020:LBA

[494] Emmanuel Godard, Damien Imbs, Michel Raynal, and Gadi Taubenfeld. Leader-based de-anonymization of an anonymous read/write memory. Theoretical Computer Science, 836(??): 110–123, October 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304114.

Anonymous:2020:POc

[495] Anonymous. Pages 1–246 (12 October 2020). Theoretical Computer Science, 837(??):1–246, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbh

[496] Anonymous. Editorial Board. Theoretical Computer Science, 837(??): ii, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304576.

Luck:2020:CLT

[497] Martin Lück. On the complexity of linear temporal logic with team semantics. Theoretical Computer Science, 837 (??):1–25, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302292.

Curzi:2020:TAL

[498] Gianluca Curzi and Luca Roversi. A type-assignment of linear erasure and duplication. Theoretical Computer Science, 837(??):26– 53, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302516.

Cristescu:2020:ECI

[499] Ioana Cristescu, Thomas Given-Wilson, and Axel Legay. Expressiveness of concurrent intensionality. Theoretical Computer Science, 837(??): 54–83, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030270X.

Liaw:2020:GTC

[500] Tsong-Ming Liaw and Simon C. Lin. A general theory of concept lattice with tractable implication exploration. Theoretical Computer Science, 837(??):

84–114, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302826.

Kimelfeld:2020:CEP

[501] Benny Kimelfeld, Ester Livshits, and Liat Peterfreund. Counting and enumerating preferred database repairs. Theoretical Computer Science, 837(??): 115–157, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302899.

Gossler:2020:CAF

[502] Gregor Gössler and Jean-Bernard Stefani. Causality analysis and fault ascription in component-based systems. Theoretical Computer Science, 837(??): 158–180, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303510.

Dimovski:2020:GAR

[503] Aleksandar S. Dimovski, Axel Legay, and Andrzej Wasowski. Generalized abstraction-refinement for game-based CTL lifted model checking. Theoretical Computer Science, 837(??): 181–206, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303522.

Abdala:2020:NPH

[504] Jarafe Abdala and Fagner Santana. A note on “Posets having continuous interval”. Theoretical Computer Science, 837(??):207–208, October 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303662.

Baaz:2020:FOI

[505] Matthias Baaz and Anela Lolic. First-order interpolation derived from propositional interpolation. Theoretical Computer Science, 837(??):209– 222, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304266.

Aceto:2020:API

[506] Luca Aceto, Elli Anastasiadi, Valentina Castiglioni, Anna Ingólfsdóttir, Bas Luttik, and Mathias Ruggaard Pedersen. On the axiomatisability of priority III: Priority strikes again. Theoretical Computer Science, 837(??): 223–246, October 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304278.

Anonymous:2020:POd

[507] Anonymous. Pages 1–250 (24 October 2020). Theoretical Computer Science, 838(??):1–250, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbi

[508] Anonymous. Editorial Board. Theoretical Computer Science, 838(??): ii, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304795.

Shu:2020:EDP

[509] Xinfeng Shu, Nan Zhang, Xiaobing Wang, and Liang Zhao. Efficient decision procedure for propositional projection temporal logic. Theoretical Computer Science, 838(??): 1–16, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302735.

Le:2020:LSP

[510] Hung Le and Baigong Zheng. Local search is a PTAS for feedback vertex set in minor-free graphs. Theoretical Computer Science, 838(??): 17–24, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302747.

Zhao:2020:SCP

[511] Liang Zhao, Xiaobing Wang, Xinfeng Shu, and Nan Zhang. A sound and complete proof system for a unified temporal logic. Theoretical Computer Science, 838(??): 25–44, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302851.

Bonamy:2020:DCU

[512] Marthe Bonamy, Marc Heinrich, Takehiro Ito, Yusuke Kobayashi, Haruka Mizuta, Moritz Mühlenthaler, Akira Suzuki, and Kunihiro Wasa. Diameter of colorings under Kempe changes. Theoretical Computer Science, 838(??): 45–57, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303169.

Xu:2020:UUF

[513] Yicheng Xu, Dachuan Xu, Yong Zhang, and Juan Zou. M^{p} UFLP: Universal facility location problem in the p-th power of metric space. Theoretical Computer Science, 838(??): 58–67, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303236.

Dinesh:2020:SAT

[514] Krishnamoorthy Dinesh and Jayalal Sarma. Sensitivity, affine transforms and quantum communication complexity. Theoretical Computer Science, 838 (??):68–80, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303340.

Jin:2020:FBK

[515] Kai Jin. On 1-factorizations of bipartite Kneser graphs. Theoretical Computer Science, 838(??):81– 93, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303406.

Bliznets:2020:LBH

[516] Ivan Bliznets and Danil Sagunov. Lower bounds for the happy coloring problems. Theoretical Computer Science, 838(??):94–110, October 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030339X.

Yan:2020:TUA

[517] Ruidong Yan, Hongwei Du, Yi Li, Wenping Chen, Yongcai Wang, Yuqing Zhu, and Deying Li. Target users’ activation probability maximization with different seed set constraints in social networks. Theoretical Computer Science, 838(??): 111–125, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303418.

J:2020:DDS

[518] Girish Raguvir J., Manas Jyoti Kashyop, and N. S. Narayanaswamy. Dynamic data structures for interval coloring. Theoretical Computer Science, 838(??):126–142, October 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303650.

Sun:2020:ORM

[519] Xiaoming Sun, Yuan Sun, Zhiyu Xia, and Jialin Zhang. The oneround multi-player discrete Voronoi game on grids and trees. Theoretical Computer Science, 838(??):143– 159, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303704.

Liu:2020:AAM

[520] Pengcheng Liu, Zhao Zhang, and Xiaohui Huang. Approximation algorithm for minimum weight connectedk-subgraph cover. Theoretical Computer Science, 838(??):160–167, October 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303285.

Couto:2020:HEM

[521] Fernanda Couto and Luís Felipe I. Cunha. Hardness and efficiency on minimizing maximum distances in spanning trees. Theoretical Computer Science, 838(??):168–179, October 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303534.

Chen:2020:ENE

[522] Xujin Chen, Xiaodong Hu, Chenhao Wang, and Xiaoying Wu. The efficiency of Nash equilibria in the load balancing game with a randomizing scheduler. Theoretical Computer Science, 838(??): 180–194, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304059.

Gao:2020:RTB

[523] Xiangyu Gao, Jianzhong Li, Dongjing Miao, and Xianmin Liu. Recognizing the tractability in big data computing. Theoretical Computer Science, 838(??): 195–207, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304072.

Baig:2020:IBT

[524] Mirza Galib Anwarul Husain Baig and Deepanjan Kesh. Improved bounds for two query adaptive bitprobe schemes storing five elements. Theoretical Computer Science, 838(??): 208–230, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304229.

Gao:2020:AAP

[525] Guichen Gao, Li Ning, Hing-Fung Ting, Yicheng Xu, Yong Zhang, and Yifei Zou. Approximation algorithms for the partial assignment problem. Theoretical Computer Science, 838(??): 231–237, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304242.

Asahiro:2020:EAR

[526] Yuichi Asahiro, Jesper Jansson, Guohui Lin, Eiji Miyano, Hirotaka Ono, and Tadatoshi Utashima. Exact algorithms for the repetition-bounded longest common subsequence problem. Theoretical Computer Science, 838(??):

238–249, October 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304230.

Anonymous:2020:PNa

[527] Anonymous. Pages 1–206 (2 November 2020). Theoretical Computer Science, 839(??):1–206, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbj

[528] Anonymous. Editorial Board. Theoretical Computer Science, 839(??): ii, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304904.

Tamir:2020:POE

[529] Tami Tamir. The power of one evil secret agent. Theoretical Computer Science, 839(??):1–12, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520303005.

Bodlaender:2020:ECP

[530] Hans L. Bodlaender and Tom C. van der Zanden. On the exact complexity of polyomino packing. Theoretical Computer Science, 839(??): 13–20, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520302991.

Almanza:2020:THW

[531] Matteo Almanza, Stefano Leucci, and Alessandro Panconesi. Tracks from hell — when finding a proof may be easier than checking it. Theoretical Computer Science, 839(??): 21–29, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030311X.

Coulombe:2020:CVG

[532] Michael J. Coulombe and Jayson Lynch. Cooperating in video games? Impossible! Undecidability of team multiplayer games. Theoretical Computer Science, 839(??): 30–40, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303108.

Abel:2020:WWW

[533] Zachary Abel, Jeffrey Bosboom, Michael Coulombe, Erik D. Demaine, Linus Hamilton, Adam Hesterberg, Justin Kopinsky, Jayson Lynch, Mikhail Rudoy, and Clemens Thielen. Who witnesses The Witness? Finding witnesses in The Witness is hard and sometimes impossible. Theoretical Computer Science, 839(??): 41–102, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303133.

Kitamura:2020:UDP

[534] Naoki Kitamura, Yuya Kawabata, and Taisuke Izumi. Uniform distribution for Pachinko. Theoretical Computer Science, 839(??):103– 121, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303121.

Bultel:2020:FCC

[535] Xavier Bultel, Jannik Dreier, Jean-Guillaume Dumas, and Pascal Lafourcade. A faster cryptographer’s Conspiracy Santa. Theoretical Computer Science, 839(??):122–134, November 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303170.

Sasaki:2020:ECB

[536] Tatsuya Sasaki, Daiki Miyahara, Takaaki Mizuki, and Hideaki Sone. Efficient card-based zero-knowledge proof for Sudoku. Theoretical Computer Science, 839(??):135–142, November 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303200.

Brandt:2020:TLB

[537] Sebastian Brandt, Yuval Emek, Jara Uitto, and Roger Wattenhofer. A tight lower bound for the capture time of the Cops and Robbers game. Theoretical Computer Science, 839(??): 143–163, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303443.

Limnios:2020:EDA

[538] Stratis Limnios, Christophe Paul, Joanny Perret, and Dimitrios M. Thilikos. Edge degeneracy: Algorithmic and structural results. Theoretical Computer Science, 839(??):164– 175, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030342X.

Brandt:2020:OGE

[539] Sebastian Brandt, Klaus-Tycho Foerster, Jonathan Maurer, and Roger Wattenhofer. Online graph exploration on a restricted graph class: Optimal solutions for tadpole graphs. Theoretical Computer Science, 839(??): 176–185, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303431.

Masjoody:2020:CRG

[540] Masood Masjoody and Ladislav Stacho. Cops and robbers on graphs with a set of forbidden induced subgraphs. Theoretical Computer Science, 839(??): 186–194, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030373X.

Kawamura:2020:SSV

[541] Akitoshi Kawamura and Makoto Soejima. Simple strategies versus optimal schedules in multi-agent patrolling. Theoretical Computer Science, 839(??): 195–206, November 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304163.

Anonymous:2020:PNb

[542] Anonymous. Pages 1–270 (6 November 2020). Theoretical Computer Science, 840(??):1–270, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbk

[543] Anonymous. Editorial Board. Theoretical Computer Science, 840(??): ii, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030534X.

Ding:2020:AAB

[544] Wei Ding and Ke Qiu. A 2approximation algorithm and beyond for the minimum diameter k-Steiner forest problem. Theoretical Computer Science, 840(??):1– 15, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930787X.

Jha:2020:AHR

[545] Anupriya Jha, D. Pradhan, and S. Banerjee. Algorithm and hardness results on neighborhood total domination in graphs. Theoretical Computer Science, 840(??):16– 32, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302590.

Censor-Hillel:2020:SAS

[546] Keren Censor-Hillel, Ami Paz, and Noam Ravid. The sparsest additive spanner via multiple weighted BFS trees. Theoretical Computer Science, 840(??):33–44, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520303194.

Salam:2020:LCP

[547] Sameera M. Salam, Jasine Babu, and K. Murali Krishnan. A local characterization for perfect plane near-triangulations. Theoretical Computer Science, 840(??):45– 58, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303728.

Ko:2020:AAN

[548] Jeremy Ko. The amortized analysis of a non-blocking chromatic tree. Theoretical Computer Science, 840(??): 59–121, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303844.

Blahoudek:2020:LSL

[549] František Blahoudek, Juraj Major, and Jan Strejček. LTL to self-loop alternating automata with generic acceptance and back. Theoretical Computer Science, 840(??):122– 142, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030390X.

Gandhi:2020:RAS

[550] Rajiv Gandhi, Magnús M. Halldórsson, Christian Konrad, Guy Kortsarz, and Hoon Oh. Radio aggregation scheduling. Theoretical Computer Science, 840(??):143–153, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752030414X.

Halldorsson:2020:LCW

[551] Magnús M. Halldórsson, Christian Konrad, and Tigran Tonoyan. Limitations of current wireless link scheduling algorithms. Theoretical Computer Science, 840(??):154– 165, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304151.

Jana:2020:CPR

[552] Satyabrata Jana and Supantha Pandit. Covering and packing of rectilinear subdivision. Theoretical Computer Science, 840(??):166–176, November 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304175.

Nong:2020:FAM

[553] Qingqin Nong, Jiazhu Fang, Suning Gong, Yan Feng, and Xiaoying Qu. A fast algorithm for maximizing a non-monotone DR-submodular integer lattice function. Theoretical Computer Science, 840(??):177– 186, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304643.

Miao:2020:MRP

[554] Dongjing Miao, Jianzhong Li, and Zhipeng Cai. Maximum reachability preserved graph cut. Theoretical Computer Science, 840(??):187– 198, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304692.

Miao:2020:PCK

[555] Dongjing Miao, Jianzhong Li, and Zhipeng Cai. The parameterized complexity and kernelization of resilience for database queries. Theoretical Computer Science, 840(??): 199–211, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304710.

Li:2020:CBG

[556] Yan Li, Yusheng Li, and Ye Wang. Complete bipartite graphs deleted in Ramsey graphs. Theoretical Computer Science, 840(??):212–218, November 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304746.

Ding:2020:ESM

[557] Xingjian Ding, Wenping Chen, Yongcai Wang, Deying Li, and Yi Hong. Efficient scheduling of a mobile charger in large-scale sensor networks. Theoretical Computer Science, 840(??): 219–233, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304734.

Buro:2020:SEL

[558] Samuele Buro and Isabella Mastroeni. On the semantic equivalence of language syntax formalisms. Theoretical Computer Science, 840(??):234– 248, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304758.

Bhattacharyya:2020:ILP

[559] Arnab Bhattacharyya, Ameet Gadekar, and Ninad Rajgopal. Improved learning of k-parities. Theoretical Computer Science, 840(??):249– 256, November 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304850.

Ni:2020:CBR

[560] Qiufen Ni, Jianxiong Guo, Chuanhe Huang, and Weili Wu. Communitybased rumor blocking maximization in social networks: Algorithms and analysis. Theoretical Computer Science, 840(??):257–269, November 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304953.

Anonymous:2020:PNc

[561] Anonymous. Pages 1–198 (12 November 2020). Theoretical Computer Science, 841(??):1–198, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbl

[562] Anonymous. Editorial Board. Theoretical Computer Science, 841(??): ii, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305405.

Li:2020:SJS

[563] Yijie Li and Shuguang Li. Scheduling jobs with sizes and delivery times on identical parallel batch machines. Theoretical Computer Science, 841(??): 1–9, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303637.

Brodal:2020:FPB

[564] Gerth Stølting Brodal, Spyros Sioutas, Konstantinos Tsakalidis, and Kostas Tsichlas. Fully persistent B-trees. Theoretical Computer Science, 841(??): 10–26, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303686.

Huang:2020:HRT

[565] Zhong Huang and Xueliang Li. Hardness results for three kinds of colored connections of graphs. Theoretical Computer Science, 841(??):27– 38, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303698.

Brandenburg:2020:FCG

[566] Franz J. Brandenburg. On fancrossing graphs. Theoretical Computer Science, 841(??):39–49, November 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303777.

Ilcinkas:2020:ECB

[567] David Ilcinkas and Ahmed M. Wade. Exploration of carrier-based timevarying networks: the power of waiting. Theoretical Computer Science, 841(??): 50–61, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303789.

Suzuki:2020:EPK

[568] Kaichi Suzuki, Atsushi Takayasu, and Noboru Kunihiro. Extended partial key exposure attacks on RSA: Improvement up to full size decryption exponents. Theoretical Computer Science, 841(??):62– 83, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303820.

Gilboa-Freedman:2020:BID

[569] Gail Gilboa-Freedman and Rann Smorodinsky. On the behavioral implications of differential privacy. Theoretical Computer Science, 841(??): 84–93, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303819.

Amanatidis:2020:MBO

[570] Georgios Amanatidis, Evangelos Markakis, and Apostolos Ntokos. Multiple birds with one stone: Beating 1/2 for EFX and GMMS via envy cycle elimination. Theoretical Computer Science, 841(??): 94–109, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303790.

Kyropoulou:2020:AEF

[571] Maria Kyropoulou, Warut Suksompong, and Alexandros A. Voudouris. Almost envy-freeness in group resource allocation. Theoretical Computer Science, 841(??):110–123, November 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303807.

Jansen:2020:PKH

[572] Bart M. P. Jansen and Astrid Pieterse. Polynomial kernels for hitting forbidden minors under structural parameterizations. Theoretical Computer Science, 841(??):124– 166, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303856.

Emek:2020:BGN

[573] Yuval Emek, Shay Kutten, Ron Lavi, and Yangguang Shi. Bayesian generalized network design. Theoretical Computer Science, 841(??):167– 185, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303832.

Yan:2020:PAI

[574] Sherry H. F. Yan and Yao Yu. Pattern-avoiding inversion sequences and open partition diagrams. Theoretical Computer Science, 841(??):186– 197, November 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303868.

Anonymous:2020:PNd

[575] Anonymous. Pages 1–136 (24 November 2020). Theoretical Computer Science, 842(??):1–136, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbm

[576] Anonymous. Editorial Board. Theoretical Computer Science, 842(??): ii, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305594.

Ning:2020:CSC

[577] Wantao Ning. Connectivity and super connectivity of the divideand-swap cube. Theoretical Computer Science, 842(??):1–5, November 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303571.

Voudouris:2020:SCA

[578] Alexandros A. Voudouris. Simple combinatorial auctions with budget constraints. Theoretical Computer Science, 842(??):6–17, November 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304035.

Ito:2020:CMS

[579] Takehiro Ito, Naonori Kakimura, and Yusuke Kobayashi. Complexity of the multi-service center problem. Theoretical Computer Science, 842(??): 18–27, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304060.

Ding:2020:FBC

[580] Hu Ding. Faster balanced clusterings in high dimension. Theoretical Computer Science, 842(??):28– 40, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030400X.

Xiao:2020:PMA

[581] Han Xiao, Qizhi Fang, and Ding-Zhu Du. Population monotonic allocation schemes for vertex cover games. Theoretical Computer Science, 842(??): 41–49, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304047.

M-Alizadeh:2020:DNP

[582] Benyamin M.-Alizadeh and Amir Hashemi. Deterministic normal position transformation and its applications. Theoretical Computer Science, 842(??):50–64, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520304011.

Monti:2020:SFM

[583] A. Monti and B. Sinaimeri. String factorisations with maximum or minimum dimension. Theoretical Computer Science, 842(??):65–73, November 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304102.

Yuster:2020:ACM

[584] Raphael Yuster. A 2^{O(k)}n algorithm for k-cycle in minor-closed graph families. Theoretical Computer Science, 842(??): 74–85, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304205.

Burcsi:2020:GGC

[585] Péter Burcsi, Gabriele Fici, Zsuzsanna Lipták, Rajeev Raman, and Joe Sawada. Generating a Gray code for prefix normal words in amortized polylogarithmic time per word. Theoretical Computer Science, 842(??):86– 99, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304217.

Hosoyamada:2020:QAM

[586] Akinori Hosoyamada, Yu Sasaki, Seiichiro Tani, and Keita Xagawa. Quantum algorithm for the multicollision problem. Theoretical Computer Science, 842(??):100–117, November 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304187.

Cui:2020:GMD

[587] Yong Cui, Qian Liu, Guo Chen, and Hujun Zhang. A general method for decomposing self-intersecting polygon to normal based on self-intersection points. Theoretical Computer Science, 842(??):118–129, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520304199.

Ellen:2020:P

[588] Faith Ellen and Luís Rodrigues. Preface. Theoretical Computer Science, 842 (??):130, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520301778.

Ito:2020:FE

[589] Hiro Ito, Stefano Leonardi, Linda Pagli, and Giuseppe Prencipe. FUN editorial. Theoretical Computer Science, 842(??):131, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520303674.

Anonymous:2020:P

[590] Anonymous.  Preface.  Theoreti-

cal Computer Science, 842(??):132, November 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303984.

Andres:2020:COC

[591] Stephan Dominique Andres, Melissa Huggan, Fionn Mc Inerney, and Richard J. Nowakowski. Corrigendum to “The orthogonal colouring game” [Theor. Comput. Sci. 795 (2019) 312–325]. Theoretical Computer Science, 842(??):133– 135, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307698. See

[5].

Anonymous:2020:PDa

[592] Anonymous. Pages 1–164 (2 December 2020). Theoretical Computer Science, 843(??):1–164, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbn

[593] Anonymous. Editorial Board. Theoretical Computer Science, 843(??): ii, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305715.

Kari:2020:Ef

[594] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 843(??):iii, December 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305764.

Perrot:2020:ERO

[595] Kévin Perrot and Éric Rémila. On the emergence of regularities on onedimensional decreasing sandpiles. Theoretical Computer Science, 843(??): 1–24, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303601.

Bridoux:2020:EAN

[596] Florian Bridoux, Maximilien Gadouleau, and Guillaume Theyssier. Expansive automata networks. Theoretical Computer Science, 843(??):25– 44, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303595.

Dennunzio:2020:DBA

[597] Alberto Dennunzio, Enrico Formenti, Darij Grinberg, and Luciano Margara. Dynamical behavior of additive cellular automata over finite abelian groups. Theoretical Computer Science, 843(??): 45–56, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303625.

Bian:2020:RTA

[598] Chao Bian, Chao Qian, Ke Tang, and Yang Yu. Running time analysis of the (1 + 1)-EA for robust linear optimization. Theoretical Computer Science, 843(??):57– 72, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303765.

Zhao:2020:TFC

[599] Yuzhen Zhao, Xiyu Liu, Minghe Sun, Jianhua Qu, and Yuanjie Zheng. Timefree cell-like P systems with multiple promoters/inhibitors. Theoretical Computer Science, 843(??):73– 83, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303996.

Johnson:2020:VPR

[600] Robert F. Johnson and Erik Winfree. Verifying polymer reaction networks using bisimulation. Theoretical Computer Science, 843(??):84– 114, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304473.

Rocha:2020:SQF

[601] Thiago Alves Rocha, Ana Teresa Martins, and Francicleber Martins Ferreira. Synthesis of quantifier-free DNF sentences from inconsistent samples of strings with EF games and SAT. Theoretical Computer Science, 843(??): 115–135, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304990.

Valle:2020:QVR

[602] Marcos Eduardo Valle and Rodolfo Anibal Lobo. Quaternion-valued recurrent projection neural networks on unit quaternions. Theoretical Computer Science, 843(??):136–152, December 2, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304989.

Verlan:2020:UIG

[603] Sergey Verlan, Henning Fernau, and Lakshmanan Kuppusamy. Universal insertion grammars of size two. Theoretical Computer Science, 843(??): 153–163, December 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305016.

Anonymous:2020:PDb

[604] Anonymous. Pages 1–226 (6 December 2020). Theoretical Computer Science, 844(??):1–226, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbo

[605] Anonymous. Editorial Board. Theoretical Computer Science, 844(??): ii, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305934.

Bilo:2020:TRC

[606] Davide Bilò, Luciano Gualà, Stefano Leucci, and Guido Proietti. Tracking routes in communication networks. Theoretical Computer Science, 844(??): 1–15, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303881.

Asahiro:2020:GOS

[607] Yuichi Asahiro, Jesper Jansson, Eiji Miyano, Hesam Nikpey, and Hirotaka Ono. Graph orientation with splits. Theoretical Computer Science, 844(??): 16–25, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030387X.

Hou:2020:AAP

[608] Xin Hou, Wen Liu, and Bo Hou. An approximation algorithm for the k-prize-collecting multicut on a tree problem. Theoretical Computer Science, 844(??): 26–33, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303893.

Drgas-Burchardt:2020:EIC

[609] Ewa Drgas-Burchardt, Hanna Furmańczyk, and Elżbieta Sidorowicz. Equitable improper choosability of graphs. Theoretical Computer Science, 844(??): 34–45, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030428X.

Jacquet:2020:JSC

[610] Philippe Jacquet, Dimitris Milioris, and Wojciech Szpankowski. Joint string complexity for Markov sources: Small data matters. Theoretical Computer Science, 844(??):46–80, December 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304424.

Li:2020:STC

[611] Chunfang Li, Shangwei Lin, and Shengjia Li. The 4-set tree connectivity of (n, k)-star networks. Theoretical Computer Science, 844(??): 81–86, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304436.

Cui:2020:RDS

[612] Hui Cui, Baodong Qin, Willy Susilo, and Surya Nepal. Robust digital signature revisited. Theoretical Computer Science, 844(??):87– 96, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030445X.

Andersen:2020:PCP

[613] J. B. Andersen, J. Bang-Jensen, and A. Yeo. On the parameterized complexity of 2-partitions. Theoretical Computer Science, 844(??):97– 105, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304606.

Mankowski:2020:EDP

[614] Michal Mankowski and Mikhail Moshkov. Extensions of dynamic programming for multi-stage combinatorial optimization. Theoretical Computer Science, 844(??):106–132, December 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304618.

Van:2020:CAI

[615] Le Huynh My Van, Kien Trung Nguyen, and Nguyen Thanh Hung. Computational aspects of the inverse single facility location problem on trees under l_{k}-norm. Theoretical Computer Science, 844(??):133– 141, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304631.

Wojciechowski:2020:AFH

[616] Piotr Wojciechowski, R. Chandrasekaran, and K. Subramani. Analyzing fractional Horn constraint systems. Theoretical Computer Science, 844(??): 142–153, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030462X.

Jansen:2020:SPS

[617] Klaus Jansen, Marten Maack, and Roberto Solis-Oba. Structural parameters for scheduling with assignment restrictions. Theoretical Computer Science, 844(??):154–170, December 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304680.

Meeks:2020:SHS

[618] Kitty Meeks and Baharak Rastegari. Solving hard stable matching problems involving groups of similar agents. Theoretical Computer Science, 844(??): 171–194, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304709.

Bazhenov:2020:OPF

[619] Nikolay Bazhenov, Iskander Kalimullin, Alexander Melnikov, and Keng Meng Ng. Online presentations of finitely generated structures. Theoretical Computer Science, 844(??):195–216, December 6, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304722.

Anand:2020:CHN

[620] Bijo S. Anand, Arun Anil, Manoj Changat, Mitre C. Dourado, and Sabeer S. Ramla. Computing the hull number in Δ-convexity. Theoretical Computer Science, 844(??): 217–226, December 6, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304849.

Anonymous:2020:PDc

[621] Anonymous. Pages 1–242 (12 December 2020). Theoretical Computer Science, 845(??):1–242, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbp

[622] Anonymous. Editorial Board. Theoretical Computer Science, 845(??): ii, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305995.

Demko:2020:NNG

[623] Christophe Demko, Karell Bertet, Cyril Faucher, Jean-François Viaud, and Sergei O. Kuznetsov. NextPriorityConcept: a new and generic algorithm computing concepts from complex and heterogeneous data. Theoretical Computer Science, 845(??): 1–20, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304837.

Choudhary:2020:VDS

[624] Pratibha Choudhary, Pallavi Jain, R. Krithika, and Vibha Sahlot. Vertex deletion on split graphs: Beyond 4-hitting set. Theoretical Computer Science, 845(??):21–37, December 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030493X.

Jallu:2020:LDU

[625] Ramesh K. Jallu, Sangram K. Jena, and Gautam K. Das. Liar’s domination in unit disk graphs. Theoretical Computer Science, 845(??): 38–49, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304965.

Singh:2020:PCD

[626] Ranveer Singh. Parameterized complexity of determinant and permanent. Theoretical Computer Science, 845(??): 50–58, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304941.

Dinesh:2020:NBE

[627] Krishnamoorthy Dinesh, Samir Otiv, and Jayalal Sarma. New bounds for energy complexity of Boolean functions. Theoretical Computer Science, 845(??): 59–75, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305041.

Behera:2020:SAU

[628] Balaram D. Behera. Sorting an array using the topological sort of a corresponding comparison graph. Theoretical Computer Science, 845(??): 76–97, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305028.

Gonzalez:2020:TTM

[629] Miguel Coviello Gonzalez and Marek Chrobak. Towards a theory of mixing graphs: a characterization of perfect mixability. Theoretical Computer Science, 845(??):98–121, December 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305065.

Takashima:2020:CBP

[630] Ken Takashima, Yuta Abe, Tatsuya Sasaki, Daiki Miyahara, Kazumasa Shinagawa, Takaaki Mizuki, and Hideaki Sone. Card-based protocols for secure ranking computations. Theoretical Computer Science, 845(??):122– 135, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305089.

Gorgi:2020:ASM

[631] Abbass Gorgi, Mourad El Ouali, Anand Srivastav, and Mohamed Hachimi. Approximation of set multicover via hypergraph matching. Theoretical Computer Science, 845(??):136– 143, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305077.

Konstantinidis:2020:PTS

[632] G. Konstantinidis and Ath. Kehagias. On positionality of trigger strategies Nash equilibria in SCAR. Theoretical Computer Science, 845(??):144– 158, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305120.

Belot:2020:ANF

[633] Gordon Belot. Absolutely no free lunches! Theoretical Computer Science, 845(??):159–180, December 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305107.

Park:2020:FPP

[634] Sung Gwan Park, Magsarjav Bataa, Amihood Amir, Gad M. Landau, and Kunsoo Park. Finding patterns and periods in Cartesian tree matching.

Theoretical Computer Science, 845(??): 181–197, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305132.

Liu:2020:SFT

[635] Heqin Liu and Dongqin Cheng. Structure fault tolerance of balanced hypercubes. Theoretical Computer Science, 845(??):198–207, December 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305119.

Levy:2020:OPD

[636] Avivit Levy and B. Riva Shalom. Online parameterized dictionary matching with one gap. Theoretical Computer Science, 845(??):208–229, December 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030517X.

Mieno:2020:SEA

[637] Takuya Mieno, Dominik Köppl, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Spaceefficient algorithms for computing minimal/shortest unique substrings. Theoretical Computer Science, 845(??):230– 242, December 12, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305181.

Anonymous:2020:PDd

[638] Anonymous. Pages 1–172 (18 December 2020). Theoretical Computer Science, 846(??):1–172, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbq

[639] Anonymous. Editorial Board. Theoretical Computer Science, 846(??): ii, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306204.

Banik:2020:FPT

[640] Aritra Banik, Pratibha Choudhary, Venkatesh Raman, and Saket Saurabh. Fixed-parameter tractable algorithms for Tracking Shortest Paths. Theoretical Computer Science, 846(??): 1–13, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305053.

Harutyunyan:2020:MIS

[641] A. Harutyunyan, M. Lampis, V. Lozin, and J. Monnot. Maximum independent sets in subcubic graphs: New results. Theoretical Computer Science, 846(??): 14–26, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305090.

Boppana:2020:SLI

[642] Ravi B. Boppana, Magnús M. Halldórsson, and Dror Rawitz. Simple and local independent set approximation. Theoretical Computer Science, 846(??):27–37, December 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305193.

Lamprou:2020:FTR

[643] Ioannis Lamprou, Russell Martin, and Sven Schewe. Fast two-robot disk evacuation with wireless communication. Theoretical Computer Science, 846(??): 38–60, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030520X.

Turau:2020:DAF

[644] Volker Turau. A distributed algorithm for finding Hamiltonian cycles in random graphs in O(log n) time. Theoretical Computer Science, 846(??): 61–74, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305211.

Wang:2020:FMP

[645] Jinling Wang. Fractional matching preclusion of product networks. Theoretical Computer Science, 846(??): 75–81, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305168.

Bujtas:2020:CFG

[646] Csilla Bujtás, Stanislav Jendrol’, and Zsolt Tuza. On caterpillar factors in graphs. Theoretical Computer Science, 846(??):82–90, December 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305156.

Czyzowicz:2020:PEDb

[647] J. Czyzowicz, K. Georgiou, R. Killick, E. Kranakis, D. Krizanc, L. Narayanan, J. Opatrny, and S. Shende. Priority evacuation from a disk: the case of n ≥ 4. Theoretical Computer Science, 846(??):91–102, December 18, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305223.

Xiao:2020:PAK

[648] Mingyu Xiao and Shaowei Kou. Parameterized algorithms and kernels for almost induced matching. Theoretical Computer Science, 846(??):103– 113, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305284.

DiGiacomo:2020:CCC

[649] Emilio Di Giacomo, Leszek Ga̧sieniec, Giuseppe Liotta, and Alfredo Navarra. On the curve complexity of 3-colored point-set embeddings. Theoretical Computer Science, 846(??):114– 140, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305260.

deLiguoro:2020:UCC

[650] Ugo de’Liguoro and Riccardo Treglia. The untyped computational λ-calculus and its intersection type discipline. Theoretical Computer Science, 846(??): 141–159, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305259.

Bhattacharya:2020:PFM

[651] Sayan Bhattacharya, Elias Koutsoupias, Janardhan Kulkarni, Stefano Leonardi, Tim Roughgarden, and Xiaoming Xu. Prior-free multi-unit auctions with ordered bidders. Theoretical Computer Science, 846(??):160– 171, December 18, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305296.

Anonymous:2020:PDe

[652] Anonymous. Pages 1–198 (22 December 2020). Theoretical Computer Science, 847(??):1–198, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbr

[653] Anonymous. Editorial Board. Theoretical Computer Science, 847(??): ii, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306307.

Lai:2020:CSC

[654] Junzuo Lai, Zhengan Huang, Man Ho Au, and Xianping Mao. Constantsize CCA-secure multi-hop unidirectional proxy re-encryption from indistinguishability obfuscation. Theoretical Computer Science, 847(??): 1–16, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305302.

Wang:2020:OCM

[655] Lusheng Wang and Boting Yang. The one-cop-moves game on graphs with some special structures. Theoretical Computer Science, 847(??):17– 26, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305466.

Komusiewicz:2020:RPC

[656] Christian Komusiewicz, Mateus de Oliveira Oliveira, and Meirav Zehavi. Revisiting the parameterized complexity of Maximum-Duo Preservation String Mapping. Theoretical Computer Science, 847(??): 27–38, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305454.

Xu:2020:EFT

[657] Min Xu and Pingshan Li. Edgefault-tolerant strong Menger edge connectivity on regular graphs. Theoretical Computer Science, 847(??):39– 48, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305442.

Becchetti:2020:SSC

[658] Luca Becchetti, Emilio Cruciani, Francesco Pasquale, and Sara Rizzo. Step-by-step community detection in volume-regular graphs. Theoretical Computer Science, 847(??):49– 67, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030548X.

Dutta:2020:OFL

[659] Binayak Dutta, Arindam Karmakar, and Sasanka Roy. Optimal facility location problem on polyhedral terrains using descending paths. Theoretical Computer Science, 847(??): 68–75, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305478.

Dose:2020:FOS

[660] Titus Dose. Further oracles separating conjectures about incompleteness in the finite domain. Theoretical Computer Science, 847(??):76– 94, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030551X.

Li:2020:LMF

[661] Peiheng Li and Jixiang Meng. Linearly many faults in Cayley graphs generated by transposition triangle free unicyclic graphs. Theoretical Computer Science, 847(??):95– 102, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305545.

Aman:2020:LTC

[662] Md A. A. A. Aman, Apurba Sarkar, Mousumi Dutt, and Arindam Biswas. A linear time combinatorial algorithm to compute the relative orthogonal convex hull of digital objects. Theoretical Computer Science, 847(??):103– 121, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305533.

Lee:2020:TSG

[663] Youngkyung Lee, Jong Hwan Park, Kwangsu Lee, and Dong Hoon Lee. Tight security for the generic construction of identity-based signature (in the multi-instance setting). Theoretical Computer Science, 847(??):122– 133, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305557.

Fahrenberg:2020:CBD

[664] Uli Fahrenberg, Axel Legay, and Karin Quaas. Computing branching distances with quantitative games. Theoretical Computer Science, 847(??):134– 146, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305636.

Bozzelli:2020:HCP

[665] Laura Bozzelli, Aniello Murano, Giuseppe Perelli, and Loredana Sorrentino. Hierarchical cost-parity games. Theoretical Computer Science, 847(??): 147–174, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030565X.

Mahdavi:2020:COPb

[666] Salma Sadat Mahdavi and Mohammad Ghodsi. Clearing an orthogonal polygon to find the evaders. Theoretical Computer Science, 847(??):175– 184, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305648.

Chen:2020:FLG

[667] Zhihuai Chen, Ken C. K. Fong, Minming Li, Kai Wang, Hongning Yuan, and Yong Zhang. Facility location games with optional preference. Theoretical Computer Science, 847(??):185– 197, December 22, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305661.

Anonymous:2020:PDf

[668] Anonymous. Pages 1–140 (24 December 2020). Theoretical Computer Science, 848(??):1–140, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2020:EBbs

[669] Anonymous. Editorial Board. Theoretical Computer Science, 848(??): ii, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306538.

Koponen:2020:CPL

[670] Vera Koponen. Conditional probability logic, lifted Bayesian networks, and almost sure quantifier elimination. Theoretical Computer Science, 848(??): 1–27, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304461.

Cantone:2020:CAD

[671] Domenico Cantone, Pietro Maugeri, and Eugenio G. Omodeo. Complexity assessments for decidable fragments of set theory. II: a taxonomy for ‘small’ languages involving membership. Theoretical Computer Science, 848(??): 28–46, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304825.

Chiari:2020:OPT

[672] Michele Chiari, Dino Mandrioli, and Matteo Pradella. Operator precedence temporal logic and model checking. Theoretical Computer Science, 848(??): 47–81, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304977.

Gaina:2020:STS

[673] Daniel Găină, Masaki Nakamura, Kazuhiro Ogata, and Kokichi Futatsugi. Stability of termination and sufficient-completeness under pushouts via amalgamation. Theoretical Computer Science, 848(??):82–105, December 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305235.

delaTour:2020:PRA

[674] Thierry Boy de la Tour and Rachid Echahed. Parallel rewriting of attributed graphs. Theoretical Computer Science, 848(??):106–132, December 24, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305247.

Cerna:2020:AUT

[675] David M. Cerna. Anti-unification and the theory of semirings. Theoretical Computer Science, 848(??):133– 139, December 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306034.

Anonymous:2021:PJa

[676] Anonymous. Pages 1–236 (6 January 2021). Theoretical Computer Science, 849(??):1–236, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBa

[677] Anonymous. Editorial Board. Theoretical Computer Science, 849(??): ii, January 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306824.

Dey:2021:LDC

[678] Palash Dey. Local distance constrained bribery in voting. Theoretical Computer Science, 849(??): 1–21, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305673.

Brandt:2021:BLM

[679] Sebastian Brandt, Manuela Fischer, and Jara Uitto. Breaking the linearmemory barrier in MPC: Fast MIS on trees with strongly sublinear memory. Theoretical Computer Science, 849 (??):22–34, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305788.

Bartschi:2021:NGE

[680] Andreas Bärtschi, Evangelos Bampas, Jérémie Chalopin, Shantanu Das, Christina Karousatou, and Matúš Mihalák. Near-gathering of energyconstrained mobile agents. Theoretical Computer Science, 849(??): 35–46, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305818.

Song:2021:FAS

[681] Siwoo Song, Geonmo Gu, Cheol Ryu, Simone Faro, Thierry Lecroq, and Kunsoo Park. Fast algorithms for single and multiple pattern Cartesian tree matching. Theoretical Computer Science, 849(??): 47–63, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305806.

Takayasu:2021:RIB

[682] Atsushi Takayasu and Yohei Watanabe. Revocable identity-based encryption with bounded decryption key exposure resistance: Latticebased construction and more. Theoretical Computer Science, 849(??): 64–98, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030579X.

Cicerone:2021:GRG

[683] Serafino Cicerone, Gabriele Di Stefano, and Alfredo Navarra. Gathering robots in graphs: the central role of synchronicity. Theoretical Computer Science, 849(??):99– 120, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030582X.

Castaneda:2021:TPD

[684] Armando Castañeda, Pierre Fraigniaud, Ami Paz, Sergio Rajsbaum, Matthieu Roy, and Corentin Travers. A topological perspective on distributed network algorithms. Theoretical Computer Science, 849(??): 121–137, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305831.

Bose:2021:APF

[685] Kaustav Bose, Manash Kumar Kundu, Ranendu Adhikary, and Buddhadeb Sau. Arbitrary pattern formation by asynchronous opaque robots with lights. Theoretical Computer Science, 849(??):138–158, January 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305867.

Mitchell:2021:PTD

[686] Lon Mitchell. Propagation tree decompositions and linearly independent vertices. Theoretical Computer Science, 849(??):159–162, January 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305880.

Wei:2021:CDC

[687] Yulong Wei and Min Xu. Conditional diagnosability of Cayley graphs generated by wheel graphs under the PMC model. Theoretical Computer Science, 849(??):163– 172, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305879.

Zhang:2021:AMU

[688] Peng Zhang and Zhendong Liu. Approximating Max k-Uncut via LProunding plus greed, with applications to Densest k-Subgraph. Theoretical Computer Science, 849(??): 173–183, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305892.

Ago:2021:GHP

[689] Kristina Ago, Bojan Bašić, Stefan Hačko, and Danijela Mitrović. On generalized highly potential words. Theoretical Computer Science, 849(??): 184–196, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306058.

Wang:2021:NCD

[690] Yihong Wang, Cheng-Kuan Lin, Qianru Zhou, and Shuming Zhou. Note on R_{g}-conditional diagnosability of hypercube. Theoretical Computer Science, 849(??):197– 201, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030606X.

Zhang:2021:RHD

[691] Nianpeng Zhang and Qiang Zhu. Reliability and hybrid diagnosis of exchanged hypercube. Theoretical Computer Science, 849(??):202– 209, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306071.

Starke:2021:ETE

[692] Florian Starke. Exploring the topological entropy of formal languages. Theoretical Computer Science, 849(??): 210–226, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306083.

Tamura:2021:AIF

[693] Yuma Tamura, Takehiro Ito, and Xiao Zhou. Approximability of the independent feedback vertex set problem for bipartite graphs. Theoretical Computer Science, 849(??):227– 236, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306101.

Anonymous:2021:PJb

[694] Anonymous. Pages 1–262 (4 January 2021). Theoretical Computer Science, 850(??):1–262, January 4, 2021.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBb

[695] Anonymous. Editorial Board. Theoretical Computer Science, 850(??): ii, January 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307003.

Panda:2021:GTD

[696] B. S. Panda and Pooja Goyal. Global total k-domination: Approximation and hardness results. Theoretical Computer Science, 850(??): 1–19, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306095.

Feng:2021:SCS

[697] Wei Feng and Shiying Wang. Structure connectivity and substructure connectivity of wheel networks. Theoretical Computer Science, 850(??): 20–29, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306113.

Currie:2021:CLB

[698] James D. Currie and Jesse T. Johnson. Characterization of the lengths of binary circular words containing no squares other than 00, 11, and 0101. Theoretical Computer Science, 850(??): 30–39, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306125.

Engelfriet:2021:XNT

[699] Joost Engelfriet, Hendrik Jan Hoogeboom, and Bart Samwel. XML navigation and transformation by treewalking automata and transducers with visible and invisible pebbles. Theoretical Computer Science, 850(??): 40–97, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306137.

Bonomo-Braberman:2021:BCA

[700] Flavia Bonomo-Braberman, Maria Chudnovsky, Jan Goedgebeur, Peter Maceli, Oliver Schaudt, Maya Stein, and Mingxian Zhong. Better 3coloring algorithms: Excluding a triangle and a seven vertex path. Theoretical Computer Science, 850(??): 98–115, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306150.

Poudel:2021:FTC

[701] Pavan Poudel, Aisha Aljohani, and Gokarna Sharma. Fault-tolerant complete visibility for asynchronous robots with lights under one-axis agreement. Theoretical Computer Science, 850(??): 116–134, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306162.

Li:2021:RCG

[702] Xiaowang Li, Shuming Zhou, Xia Guo, and Tianlong Ma. The h-restricted connectivity of the generalized hypercubes. Theoretical Computer Science, 850(??): 135–147, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306265.

Sharma:2021:LTB

[703] Arpit Sharma. The linear timebranching time spectrum of equivalences for stochastic systems with non-determinism. Theoretical Computer Science, 850(??):148–167, January 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306356.

Ras:2021:CCC

[704] C. J. Ras, M. Brazil, and D. A. Thomas. Computational complexity of the 2-connected Steiner network problem in the ℓ_{p} plane. Theoretical Computer Science, 850(??): 168–184, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306344.

Seki:2021:ORP

[705] Hiroyuki Seki, Reo Yoshimura, and Yoshiaki Takata. Optimal run problem for weighted register automata. Theoretical Computer Science, 850(??): 185–201, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306368.

Gotoh:2021:EDT

[706] Tsuyoshi Gotoh, Yuichi Sudo, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Exploration of dynamic tori by multiple agents. Theoretical Computer Science, 850(??): 202–220, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306332.

Yu:2021:FMS

[707] Hung-I Yu, Tien-Ching Lin, and D. T. Lee. Finding maximum sum segments in sequences with uncertainty. Theoretical Computer Science, 850(??): 221–235, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306381.

Mantaci:2021:CVS

[708] Sabrina Mantaci, Antonio Restivo, Giuseppe Romana, Giovanna Rosone, and Marinella Sciortino. A combinatorial view on string attractors. Theoretical Computer Science, 850(??): 236–248, January 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030637X.

Tang:2021:BPS

[709] Shaojie Tang. Beyond pointwise submodularity: Non-monotone adaptive submodular maximization in linear time. Theoretical Computer Science, 850(??):249–261, January 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306393.

Anonymous:2021:PJc

[710] Anonymous. Pages 1–154 (6 January 2021). Theoretical Computer Science, 851(??):1–154, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBc

[711] Anonymous. Editorial Board. Theoretical Computer Science, 851(??): ii, January 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307301.

Kari:2021:Ea

[712] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 851(??):iii, January 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307337.

Liu:2021:ILB

[713] Daphne Der-Fen Liu, Laxman Saha, and Satyabrata Das. Improved lower bounds for the radio number of trees. Theoretical Computer Science, 851(??): 1–13, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303017.

Liu:2021:FDB

[714] Jiafei Liu, Shuming Zhou, Zhendong Gu, Qianru Zhou, and Dajin Wang. Fault diagnosability of Bicube networks under the PMC diagnostic model. Theoretical Computer Science, 851(??): 14–23, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305144.

Doerr:2021:EUB

[715] Benjamin Doerr. Exponential upper bounds for the runtime of randomized search heuristics. Theoretical Computer Science, 851(??): 24–38, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305430.

Christensen:2021:SFS

[716] Michael Christensen, Joseph McMahan, Lawton Nichols, Jared Roesch, Timothy Sherwood, and Ben Hardekopf. Safe functional systems through integrity types and verified assembly. Theoretical Computer Science, 851(??): 39–61, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305508.

Minatel:2021:LER

[717] Diego Minatel, Vinícius Ferreira, and Alneu de Andrade Lopes. Local-entity resolution for building location-based social networks by using stay points. Theoretical Computer Science, 851(??):

62–76, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305843.

Huang:2021:CPD

[718] Yue Huang, Wenmei Yi, Hong Peng, Jun Wang, Xiaohui Luo, and Qian Yang. Computational power of dynamic threshold neural P systems for generating string languages. Theoretical Computer Science, 851(??): 77–91, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306046.

Kopra:2021:TSF

[719] Johan Kopra. On the trace subshifts of fractional multiplication automata. Theoretical Computer Science, 851(??): 92–110, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306423.

Singh:2021:FPT

[720] Pankaj Kumar Singh and Surabhi Tiwari. A fixed point theorem in rough semi-linear uniform spaces. Theoretical Computer Science, 851(??): 111–120, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306447.

Doerr:2021:SRT

[721] Benjamin Doerr and Martin S. Krejca. A simplified run time analysis of the univariate marginal distribution algorithm on LeadingOnes. Theoretical Computer Science, 851(??): 121–128, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306691.

Takahashi:2021:PUQ

[722] Yasuhiro Takahashi and Seiichiro Tani. Power of uninitialized qubits in shallow quantum circuits. Theoretical Computer Science, 851(??):129– 153, January 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306861.

Anonymous:2021:PJd

[723] Anonymous. Pages 1–212 (8 January 2021). Theoretical Computer Science, 852(??):1–212, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBd

[724] Anonymous. Editorial Board. Theoretical Computer Science, 852(??): ii, January 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030743X.

Ko:2021:COW

[725] Ker-I Ko and Lidong Wu. On continuous one-way functions. Theoretical Computer Science, 852(??): 1–17, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306149.

Georgiou:2021:TEO

[726] Konstantinos Georgiou, George Karakostas, and Evangelos Kranakis. Treasure evacuation with one robot on a disk. Theoretical Computer Science, 852(??): 18–28, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030640X.

Li:2021:IGS

[727] Shouwei Li, Friedhelm Meyer auf der Heide, and Pavel Podlipyan. The impact of the Gabriel subgraph of the visibility graph on the gathering of mobile autonomous robots. Theoretical Computer Science, 852(??): 29–40, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306411.

Li:2021:CSC

[728] Shouwei Li, Christine Markarian, Friedhelm Meyer auf der Heide, and Pavel Podlipyan. A continuous strategy for collisionless gathering. Theoretical Computer Science, 852(??): 41–60, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306435.

Czyzowicz:2021:TET

[729] Jurek Czyzowicz, Konstantinos Georgiou, Ryan Killick, Evangelos Kranakis, Danny Krizanc, Manuel Lafond, Lata Narayanan, Jaroslav Opatrny, and Sunil Shende. Time-energy tradeoffs for evacuation by two robots in the wireless model. Theoretical Computer Science, 852(??):61– 72, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306459.

Alhakim:2021:RPS

[730] Abbas Alhakim, Evan Sala, and Joe Sawada. Revisiting the Prefersame and Prefer-opposite de Bruijn sequence constructions. Theoretical Computer Science, 852(??):73– 77, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306629.

Cording:2021:MUF

[731] Patrick Hagge Cording, Travis Gagie, Mathias Bæk Tejs Knudsen, and Tomasz Kociumaka. Maximal unbordered factors of random strings. Theoretical Computer Science, 852(??): 78–83, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306630.

Chen:2021:IMP

[732] Huiping Chen, Grigorios Loukides, Solon P. Pissis, and Hau Chan. Influence maximization in the presence of vulnerable nodes: a ratio perspective. Theoretical Computer Science, 852(??): 84–103, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306575.

Fertin:2021:MCA

[733] Guillaume Fertin, Julien Fradin, and Géraldine Jean. The Maximum Colorful Arborescence problem: How (computationally) hard can it be? Theoretical Computer Science, 852(??): 104–120, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306617.

Desharnais:2021:ASP

[734] Charles Desharnais and Sylvie Hamel. Automedian sets of permutations: direct sum and shuffle. Theoretical Computer Science, 852(??):121– 131, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306654.

Amir:2021:TRT

[735] Amihood Amir and Eitan Kondratovsky. Towards a real time algorithm for parameterized longest common prefix computation. Theoretical Computer Science, 852(??):132– 137, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306642.

Prezza:2021:SEC

[736] Nicola Prezza and Giovanna Rosone. Space-efficient construction of compressed suffix trees. Theoretical Computer Science, 852(??):138–156, January 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306599.

Klein:2021:OSR

[737] Shmuel T. Klein, Jakub Radoszewski, Tamar C. Serebro, and Dana Shapira. Optimal skeleton and reduced Huffman trees. Theoretical Computer Science, 852(??):157–171, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520306605.

Dahn:2021:FPA

[738] Christine Dahn, Nils M. Kriege, Petra Mutzel, and Julian Schilling. Fixed-parameter algorithms for the weighted Max-Cut problem on embedded 1-planar graphs. Theoretical Computer Science, 852(??):172– 184, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306721.

Carpi:2021:BLC

[739] Arturo Carpi and Flavio D’Alessandro. On bounded linear codes and the commutative equivalence. Theoretical Computer Science, 852(??):185– 196, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306708.

Covella:2021:ARP

[740] Barbara Covella, Fabrizio Frati, and Maurizio Patrignani. On the area requirements of planar straight-line orthogonal drawings of ternary trees. Theoretical Computer Science, 852(??): 197–211, January 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030671X.

Anonymous:2021:EBe

[741] Anonymous. Editorial Board. Theoretical Computer Science, 853(??): ii, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307635.

Sun:2021:SII

[742] Xiaoming Sun and Jialin Zhang. Special issue on the International Conference on Algorithmic Aspects in Information and Management 2019 (AAIM’19). Theoretical Computer Science, 853(??):1, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397521000086.

Huang:2021:AAE

[743] Yufei Huang, Bo Hou, Wen Liu, Lidong Wu, Stephen Rainwater, and Suogang Gao. On approximation algorithm for the edge metric dimension problem. Theoretical Computer Science, 853(??): 2–6, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302620.

Tan:2021:NAA

[744] Guanlan Tan, Qilong Feng, Xiangzhong Meng, and Jianxin Wang. A new approximation algorithm for contig-based genomic scaffold filling. Theoretical Computer Science, 853(??): 7–15, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302863.

Gong:2021:MMF

[745] Suning Gong, Qingqin Nong, Tao Sun, Qizhi Fang, Dingzhu Du, and Xiaoyu Shao. Maximize a monotone function with a generic submodularity ratio. Theoretical Computer Science, 853 (??):16–24, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302875.

Luo:2021:OFT

[746] Chuanwen Luo, Wenping Chen, Deying Li, Yongcai Wang, Hongwei Du, Lidong Wu, and Weili Wu. Optimizing flight trajectory of UAV for efficient data collection in wireless sensor networks. Theoretical Computer Science, 853(??): 25–42, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302887.

Wang:2021:AAD

[747] Limin Wang, Zhao Zhang, Chenchen Wu, Dachuan Xu, and Xiaoyan Zhang. Approximation algorithms for the dynamic k-level facility location problems. Theoretical Computer Science, 853(??):43–56, January 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520302978.

Yang:2021:CTS

[748] Ruiqi Yang, Shuyang Gu, Chuangen Gao, Weili Wu, Hua Wang, and Dachuan Xu. A constrained twostage submodular maximization. Theoretical Computer Science, 853(??): 57–64, January 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030298X.

Zhang:2021:AAS

[749] Dongmei Zhang, Yukun Cheng, Min Li, Yishui Wang, and Dachuan Xu. Approximation algorithms for spherical k-means problem using local search scheme. Theoretical Computer Science, 853(??):65–77, January 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520303716.

Anonymous:2021:PJe

[750] Anonymous. Pages 1–174 (16 January 2021). Theoretical Computer Science, 854(??):1–174, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBf

[751] Anonymous. Editorial Board. Theoretical Computer Science, 854(??): ii, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000050.

Shalom:2021:PDM

[752] B. Riva Shalom. Parameterized dictionary matching and recognition with one gap. Theoretical Computer Science, 854(??):1–16, January 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306587.

Czajka:2021:EEA

[753] Patryk Czajka and Jakub Radoszewski. Experimental evaluation of algorithms for computing quasiperiods. Theoretical Computer Science, 854(??): 17–29, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306757.

Ibtehaz:2021:MST

[754] Nabil Ibtehaz, M. Kaykobad, and M. Sohel Rahman. Multidimensional segment trees can do range updates in poly-logarithmic time. Theoretical Computer Science, 854(??):30– 43, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306745.

Ta:2021:CLC

[755] Toan Thang Ta, Yi-Kung Shieh, and Chin Lung Lu. Computing a longest common almost-increasing subsequence of two sequences. Theoretical Computer Science, 854(??): 44–51, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306733.

Ganguly:2021:FDS

[756] Arnab Ganguly, Wing-Kai Hon, Kunihiko Sadakane, Rahul Shah, Sharma V. Thankachan, and Yilin Yang. A framework for designing space-efficient dictionaries for parameterized and order-preserving matching. Theoretical Computer Science, 854(??):52– 62, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306769.

Badkobeh:2021:TUL

[757] Golnaz Badkobeh, Paweł Gawrychowski, Juha Kärkkäinen, Simon J. Puglisi, and Bella Zhukova. Tight upper and lower bounds on suffix tree breadth. Theoretical Computer Science, 854(??): 63–67, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306770.

Wang:2021:BAB

[758] Yijing Wang, Dachuan Xu, Donglei Du, and Ran Ma. Bicriteria algorithms to balance coverage and cost in team formation under online model. Theoretical Computer Science, 854(??): 68–76, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307076.

Liu:2021:MMN

[759] Jin Liu, Zhenhua Duan, and Cong Tian. Multi-matching nested relations. Theoretical Computer Science, 854(??): 77–93, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307064.

Zhang:2021:MCC

[760] Yapu Zhang, Jianxiong Guo, Wenguo Yang, and Weili Wu. Mixed-case community detection problem in social networks: Algorithms and analysis. Theoretical Computer Science, 854(??): 94–104, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307052.

Louza:2021:NAR

[761] Felipe A. Louza, Neerja Mhaskar, and W. F. Smyth. A new approach to regular & indeterminate strings. Theoretical Computer Science, 854(??): 105–115, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307118.

Idalino:2021:NCF

[762] Thais Bardini Idalino and Lucia Moura. Nested cover-free families for unbounded fault-tolerant aggregate signatures. Theoretical Computer Science, 854(??):116–130, January 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307106.

Hon:2021:NGN

[763] Wing-Kai Hon, Ton Kloks, Hsiang-Hsuan Liu, Hung-Lung Wang, and Yue-Li Wang. A note on the geodetic number and the Steiner number of AT-free graphs. Theoretical Computer Science, 854(??):131– 135, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030712X.

Blume:2021:FCM

[764] Till Blume, David Richerby, and Ansgar Scherp. FLUID: a common model for semantic structural graph summaries based on equivalence relations. Theoretical Computer Science, 854(??): 136–158, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307234.

vanRenssen:2021:BDS

[765] André van Renssen and Gladys Wong. Bounded-degree spanners in the presence of polygonal obstacle. Theoretical Computer Science, 854(??):159– 173, January 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307507.

Anonymous:2021:PFa

[766] Anonymous. Pages 1–186 (6 February 2021). Theoretical Computer Science, 855(??):1–186, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBg

[767] Anonymous. Editorial Board. Theoretical Computer Science, 855(??): ii, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000128.

Kaklamanis:2021:PSS

[768] Christos Kaklamanis, Panagiotis Kanellopoulos, Konstantinos Papaioannou, and Dimitris Patouchas. On the price of stability of some simple graph-based hedonic games. Theoretical Computer Science, 855(??): 1–15, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306460.

Cao:2021:BPO

[769] Shujiao Cao and Rui Xue. Being a permutation is also orthogonal to onewayness in quantum world: Impossibilities of quantum one-way permutations from one-wayness primitives. Theoretical Computer Science, 855(??): 16–42, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306472.

Huntemann:2021:BGT

[770] Svenja Huntemann, Richard J. Nowakowski, and Carlos Pereira dos Santos. Bounding game temperature using confusion intervals. Theoretical Computer Science, 855(??):43– 60, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306484.

Li:2021:ELT

[771] Yan Li and Xin Zhang. Equitable list tree-coloring of bounded treewidth graphs. Theoretical Computer Science, 855(??):61–67, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520306496.

Wei:2021:PCI

[772] Jianxin Wei. Proof of a conjecture on 2-isometric words. Theoretical Computer Science, 855(??): 68–73, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306666.

Hao:2021:VPD

[773] Rong-Xia Hao and Zengxian Tian. The vertex-pancyclicity of data center networks. Theoretical Computer Science, 855(??):74–89, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520306678.

Vijayalakshmi:2021:SGM

[774] Vipin Ravindran Vijayalakshmi, Marc Schröder, and Tami Tamir. Scheduling games with machine-dependent priority lists. Theoretical Computer Science, 855(??):90–103, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520306915.

Ganesh:2021:CRF

[775] Chaya Ganesh, Bernardo Magri, and Daniele Venturi. Cryptographic reverse firewalls for interactive proof systems. Theoretical Computer Science, 855(??): 104–132, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306927.

Matsumoto:2021:CEO

[776] Kengo Matsumoto. Coded equivalence of one-sided topological Markov shifts. Theoretical Computer Science, 855(??): 133–140, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306903.

Attiya:2021:SSC

[777] Hagit Attiya, Sebastian Burckhardt, Alexey Gotsman, Adam Morrison, Hongseok Yang, and Marek Zawirski. Specification and space complexity of collaborative text editing. Theoretical Computer Science, 855(??):141– 160, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306952.

Gurski:2021:HCD

[778] Frank Gurski, Dominique Komander, and Carolin Rehs. How to compute digraph width measures on directed co-graphs. Theoretical Computer Science, 855(??):161– 185, February 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306940.

Anonymous:2021:PFb

[779] Anonymous. Pages 1–172 (8 February 2021). Theoretical Computer Science, 856(??):1–172, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBh

[780] Anonymous. Editorial Board. Theoretical Computer Science, 856(??): ii, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000396.

Deppe:2021:BCE

[781] Christian Deppe, Vladimir Lebedev, and Georg Maringer. Bounds for the capacity error function for unidirectional channels with noiseless feedback. Theoretical Computer Science, 856(??): 1–13, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306964.

Vereshchagin:2021:PCI

[782] Nikolay Vereshchagin. Proofs of conservation inequalities for Levin’s notion of mutual information of 1974. Theoretical Computer Science, 856(??): 14–20, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030709X.

Schrock:2021:CCB

[783] Tyler Schrock and Rafael Frongillo. Computational complexity of k-block conjugacy. Theoretical Computer Science, 856(??):21–40, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520307143.

Fung:2021:OTW

[784] Stanley P. Y. Fung. Online two-way trading: Randomization and advice. Theoretical Computer Science, 856(??): 41–50, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307192.

Feuilloley:2021:HLD

[785] Laurent Feuilloley, Pierre Fraigniaud, and Juho Hirvonen. A hierarchy of local decision. Theoretical Computer Science, 856(??):51– 67, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307210.

Berczi:2021:GCS

[786] Kristóf Bérczi, Endre Boros, Ondřej Čepek, Khaled Elbassioni, Petr Kučera, and Kazuhisa Makino. Generating clause sequences of a CNF formula. Theoretical Computer Science, 856(??):68–74, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397520307246.

Chang:2021:CDC

[787] Yu-Huei Chang, Kung-Jui Pai, Chiun-Chieh Hsu, Jinn-Shyong Yang, and Jou-Ming Chang. Constructing dual-CISTs of folded divide-and-swap cubes. Theoretical Computer Science, 856(??): 75–87, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307489.

Kimura:2021:TRP

[788] Kei Kimura and Akira Suzuki. Trichotomy for the reconfiguration problem of integer linear systems. Theoretical Computer Science, 856(??): 88–109, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307519.

Abiad:2021:SST

[789] Aida Abiad, Boris Brimkov, and Alexander Grigoriev. On the status sequences of trees. Theoretical Computer Science, 856(??):110– 120, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307581.

Janssen:2021:CPN

[790] Remie Janssen and Yukihiro Murakami. On cherry-picking and network containment. Theoretical Computer Science, 856(??):121–150, February 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307593.

Prokhorchuk:2021:GAF

[791] Veronika Prokhorchuk. Generation of amalgamated free products of cyclic groups by finite automata over minimal alphabet. Theoretical Computer Science, 856(??):151– 164, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307684.

Baskoro:2021:IAL

[792] Edy Tri Baskoro and Devi Imulia Dian Primaskun. Improved algorithm for the locating-chromatic number of trees. Theoretical Computer Science, 856(??): 165–168, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307696.

Li:2021:CCL

[793] Yaqiao Li. Conflict complexity is lower bounded by block sensitivity. Theoretical Computer Science, 856(??): 169–172, February 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307726.

Anonymous:2021:PFc

[794] Anonymous. Pages 1–192 (12 February 2021). Theoretical Computer Science, 857(??):1–192, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBi

[795] Anonymous. Editorial Board. Theoretical Computer Science, 857(??): ii, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000566.

Hooshmand:2021:EDS

[796] Sahar Hooshmand, Paniz Abedin, M. Oğuzhan Külekci, and Sharma V. Thankachan. I/O-efficient data structures for non-overlapping indexing. Theoretical Computer Science, 857(??): 1–7, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307131.

Niu:2021:HAE

[797] Bei Niu, Bi Li, and Xin Zhang. Hardness and algorithms of equitable treecoloring problem in chordal graphs. Theoretical Computer Science, 857(??): 8–15, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307258.

Ding:2021:OWC

[798] Xingjian Ding, Jianxiong Guo, Deying Li, and Weili Wu. Optimal wireless charger placement with individual energy requirement. Theoretical Computer Science, 857(??): 16–28, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307520.

Zhang:2021:TLS

[799] Nan Zhang, Bin Yu, Cong Tian, Zhenhua Duan, and Xiaoshuai Yuan. Temporal logic specification mining of programs. Theoretical Computer Science, 857(??):29–42, February 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307568.

Li:2021:UTR

[800] Zhiyao Li, Bosen Cheng, Xiaofeng Gao, Huai Chen, and Guihai Chen. A unified task recommendation strategy for realistic mobile crowdsourcing system. Theoretical Computer Science, 857(??): 43–58, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030757X.

Ma:2021:SLT

[801] Hengzhao Ma and Jianzhong Li. A sub-linear time algorithm for approximating k-nearest-neighbor with full quality guarantee. Theoretical Computer Science, 857(??):59– 70, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307702.

Rao:2021:MIM

[802] Guoyao Rao, Yongcai Wang, Wenping Chen, Deying Li, and Weili Wu. Matching influence maximization in social networks. Theoretical Computer Science, 857(??):71– 86, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307714.

Chen:2021:NAM

[803] Shengminjie Chen, Wenguo Yang, Suixiang Gao, and Rong Jin. Novel algorithms for maximum DS decomposition. Theoretical Computer Science, 857(??):87–96, February 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307738.

Richomme:2021:SID

[804] Gwenaël Richomme. On sets of indefinitely desubstitutable words. Theoretical Computer Science, 857(??): 97–113, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000177.

Bu:2021:ICP

[805] Yuehua Bu, Chentao Qi, Junlei Zhu, and Ting Xu. Injective coloring of planar graphs. Theoretical Computer Science, 857(??):114– 122, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000189.

Giuliani:2021:WDM

[806] Sara Giuliani, Zsuzsanna Lipták, Francesco Masillo, and Romeo Rizzi. When a dollar makes a BWT. Theoretical Computer Science, 857(??): 123–146, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000207.

Berglund:2021:FIB

[807] Martin Berglund, Willem Bester, and Brink van der Merwe. Formalising and implementing Boost POSIX regular expression matching. Theoretical Computer Science, 857(??):147– 165, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000232.

delaMaza:2021:CRO

[808] Sebastián González Hermosillo de la Maza, Seyyed Aliasghar Hosseini, Fiachra Knox, Bojan Mohar, and Bruce Reed. Cops and robbers on oriented toroidal grids. Theoretical Computer Science, 857(??):166– 176, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000281.

Ning:2021:RPW

[809] Baoling Ning, Jianzhong Li, and Shouxu Jiang. Range partitioning within sublinear time: Algorithms and lower bounds. Theoretical Computer Science, 857(??):177– 191, February 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100030X.

Anonymous:2021:PFd

[810] Anonymous. Pages 1–146 (16 February 2021). Theoretical Computer Science, 858(??):1–146, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBj

[811] Anonymous. Editorial Board. Theoretical Computer Science, 858(??): ii, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000682.

Lamprou:2021:IBC

[812] Ioannis Lamprou, Ioannis Sigalas, and Vassilis Zissimopoulos. Improved budgeted connected domination and budgeted edge-vertex domination. Theoretical Computer Science, 858(??): 1–12, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000529.

Fotakis:2021:RMF

[813] Dimitris Fotakis, Loukas Kavouras, Panagiotis Kostopanagiotis, Philip Lazos, Stratis Skoulakis, and Nikos Zarifis. Reallocating multiple facilities on the line. Theoretical Computer Science, 858(??):13– 34, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000517.

Nguyen:2021:IBC

[814] Trung Thanh Nguyen and Jörg Rothe. Improved bi-criteria approximation schemes for load balancing on unrelated machines with cost constraints. Theoretical Computer Science, 858(??): 35–48, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030726X.

Chiarelli:2021:SCD

[815] Nina Chiarelli, Berenice Martínez-Barona, Martin Milanič, Jérôme Monnot, and Peter Muršič. Strong cliques in diamond-free graphs. Theoretical Computer Science, 858(??):49– 63, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307088.

Pattanayak:2021:RGA

[816] Debasish Pattanayak, John Augustine, and Partha Sarathi Mandal. Randomized gathering of asynchronous mobile robots. Theoretical Computer Science, 858(??):64– 80, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306939.

Abreu:2021:CCC

[817] Alexandre Abreu, Luís Cunha, Celina de Figueiredo, Luis Kowada, Franklin Marquezino, Renato Portugal, and Daniel Posner. A computational complexity comparative study of graph tessellation problems. Theoretical Computer Science, 858(??):81– 89, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306897.

Das:2021:LTI

[818] Arun Kumar Das, Sandip Das, and Joydeep Mukherjee. Largest triangle inside a terrain. Theoretical Computer Science, 858(??):90– 99, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307222.

Deshpande:2021:SBD

[819] Amit Deshpande and Rameshwar Pratap. Sampling-based dimension reduction for subspace approximation with outliers. Theoretical Computer Science, 858(??):100– 113, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000323.

Liu:2021:OAB

[820] Zhicheng Liu, Ling Chen, Hong Chang, Donglei Du, and Xiaoyan Zhang. Online algorithms for BP functions maximization. Theoretical Computer Science, 858(??):114– 121, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000335.

Baranwal:2021:OAS

[821] Aseem Baranwal, Luke Schaeffer, and Jeffrey Shallit. Ostrowski-automatic sequences: Theory and applications. Theoretical Computer Science, 858(??): 122–142, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000311.

Du:2021:ECA

[822] Ding-Zhu Du and Jie Wang. Editorial: Complexity and approximation: In honor of Ker-I Ko. Theoretical Computer Science, 858(??): 143, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306782.

Lotker:2021:SAI

[823] Zvi Lotker and Boaz Patt-Shamir. Selected articles from the 25th International Colloquium on Structural Information and Communication Complexity. Theoretical Computer Science, 858 (??):144, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305491.

Angelopoulos:2021:PSI

[824] Spyros Angelopoulos, Nancy E. Clarke, Fedor V. Fomin, Archontia C. Giannopoulou, and Roman Rabinovich. Preface to the special issue on Graph Searching: Theory and Applications. Theoretical Computer Science, 858(??): 145–146, February 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000645.

Anonymous:2021:PMa

[825] Anonymous. Pages 1–174 (6 March 2021). Theoretical Computer Science, 859(??):1–174, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBk

[826] Anonymous. Editorial Board. Theoretical Computer Science, 859(??): ii, March 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000979.

Ateniese:2021:IAC

[827] Giuseppe Ateniese, Danilo Francati, Bernardo Magri, and Daniele Venturi. Immunization against complete subversion without random oracles. Theoretical Computer Science, 859 (??):1–36, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000190.

Shang:2021:LCP

[828] Jianhui Shang, Peng Li, and Yi Shi. The longest cycle problem is polynomial on interval graphs. Theoretical Computer Science, 859(??): 37–47, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000165.

Rong:2021:RVC

[829] Guozhen Rong, Wenjun Li, Yongjie Yang, and Jianxin Wang. Reconstruction and verification of chordal graphs with a distance oracle. Theoretical Computer Science, 859(??): 48–56, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000153.

Yamanaka:2021:ORO

[830] Katsuhisa Yamanaka, Takashi Horiyama, and Kunihiro Wasa. Optimal reconfiguration of optimal ladder lotteries. Theoretical Computer Science, 859 (??):57–69, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000219.

Dekking:2021:SDF

[831] F. Michel Dekking. The sum of digits functions of the Zeckendorf and the base phi expansions. Theoretical Computer Science, 859(??):

70–79, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000244.

Morales:2021:MDH

[832] Pedro Ramaciotti Morales, Robin Lamarche-Perrin, Raphaël Fournier-Sńiehotta, Rémy Poulain, Lionel Tabourier, and Fabien Tarissan. Measuring diversity in heterogeneous information networks. Theoretical Computer Science, 859(??):80–115, March 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000256.

Funakoshi:2021:CLP

[833] Mitsuru Funakoshi, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Computing longest palindromic substring after singlecharacter or block-wise edits. Theoretical Computer Science, 859(??): 116–133, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100027X.

Cicalese:2021:IPN

[834] Ferdinando Cicalese, Zsuzsanna Lipták, and Massimiliano Rossi. On infinite prefix normal words. Theoretical Computer Science, 859(??): 134–148, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000268.

DArco:2021:SSS

[835] Paolo D’Arco, Roberto De Prisco, and Alfredo De Santis. Secret sharing schemes for infinite sets of participants: a new design technique. Theoretical Computer Science, 859(??): 149–161, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000347.

Barth:2021:OEN

[836] Dominique Barth, Thierry Mautor, Arnaud de Moissac, Dimitri Watel, and Marc-Antoine Weisser. Optimisation of electrical network configuration: Complexity and algorithms for ring topologies. Theoretical Computer Science, 859(??): 162–173, March 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100044X.

Anonymous:2021:PMb

[837] Anonymous. Pages 1–134 (8 March 2021). Theoretical Computer Science, 860(??):1–134, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBl

[838] Anonymous. Editorial Board. Theoretical Computer Science, 860(??): ii, March 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001171.

Zhou:2021:CLR

[839] Yanwei Zhou, Yuan Xu, Zirui Qiao, Bo Yang, and Mingwu Zhang. Continuous leakage-resilient certificatebased signcryption scheme and application in cloud computing. Theoretical Computer Science, 860(??): 1–22, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000451.

Dick:2021:RAP

[840] Jeffery Dick, Laura K. Hutchinson, Robert Mercaş, and Daniel Reidenbach. Reducing the ambiguity of Parikh matrices. Theoretical Computer Science, 860(??): 23–40, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000487.

Al-Najjar:2021:ARN

[841] Yacine Al-Najjar, Walid Ben-Ameur, and Jérémie Leguay. On the approximability of robust network design. Theoretical Computer Science, 860(??):41–50, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000463.

Klavik:2021:GIR

[842] Pavel Klavík, Dušan Knop, and Peter Zeman. Graph isomorphism restricted by lists. Theoretical Computer Science, 860(??):51–71, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000475.

Calderoni:2021:DPP

[843] Luca Calderoni, Luciano Margara, and Moreno Marzolla. Direct product primality testing of graphs is GI-hard. Theoretical Computer Science, 860 (??):72–83, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000505.

Georgiou:2021:CEE

[844] Chryssis Georgiou, Marios Mavronicolas, and Burkhard Monien. The complexity of (E + Var)-equilibria, ESRequilibria, and SuperE-equilibria for 2-players games with few cost values. Theoretical Computer Science, 860 (??):84–97, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000499.

Agrawal:2021:PTC

[845] Akanksha Agrawal, Lawqueen Kanesh, Saket Saurabh, and Prafullkumar Tale. Paths to trees and cacti. Theoretical Computer Science, 860(??): 98–116, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000621.

Bar-Noy:2021:GBC

[846] Amotz Bar-Noy, Thomas Erlebach, Dror Rawitz, and Peter Terlecky.

“Green” barrier coverage with mobile sensors. Theoretical Computer Science, 860(??):117–134, March 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439752100061X.

Anonymous:2021:PMc

[847] Anonymous. Pages 1–154 (12 March 2021). Theoretical Computer Science, 861(??):1–154, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBm

[848] Anonymous. Editorial Board. Theoretical Computer Science, 861(??): ii, March 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001298.

Alam:2021:DBE

[849] Jawaherul Md. Alam, Michael A. Bekos, Vida Dujmović, Martin Gronemann, Michael Kaufmann, and Sergey Pupyrev. On dispersable book embeddings. Theoretical Computer Science, 861(??):1–22, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000633.

Gorain:2021:SLS

[850] Barun Gorain and Andrzej Pelc. Short labeling schemes for topology recognition in wireless tree networks. Theoretical Computer Science, 861(??): 23–44, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000876.

Czyzowicz:2021:GEC

[851] Jurek Czyzowicz, Dariusz Dereniowski, Robert Ostrowski, and Wojciech Rytter. Gossiping by energy-constrained mobile agents in tree networks. Theoretical Computer Science, 861(??): 45–65, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000852.

Jansen:2021:ESF

[852] Klaus Jansen, Oliver Sinnen, and Huijun Wang. An EPTAS for scheduling fork-join graphs with communication delay. Theoretical Computer Science, 861(??):66– 79, March 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000827.

Ning:2021:GME

[853] Wantao Ning and Hao Li. The generalized measure of edge fault tolerance in exchanged crossed cube. Theoretical Computer Science, 861(??): 80–84, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000803.

Acharyya:2021:CSL

[854] Ankush Acharyya, Anil Maheshwari, and Subhas C. Nandy. Colorspanning localized query. Theoretical Computer Science, 861(??):85– 101, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000864.

Zhang:2021:RCM

[855] Yapu Zhang, Wenguo Yang, and Ding-Zhu Du. Rumor correction maximization problem in social networks. Theoretical Computer Science, 861(??): 102–116, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000839.

Lin:2021:IBE

[856] Xi-Jun Lin, Qihui Wang, Lin Sun, and Haipeng Qu. Identity-based encryption with equality test and datestampbased authorization mechanism. Theoretical Computer Science, 861(??): 117–132, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000815.

Bienkowski:2021:NRM

[857] Marcin Bienkowski, Martin Böhm, Jaroslaw Byrka, Marek Chrobak, Christoph Dürr, Lukáš Folwarczný, Łukasz Jeż, Jiří Sgall, Nguyen Kim Thang, and Pavel Veselý. New results on multi-level aggregation. Theoretical Computer Science, 861(??): 133–143, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000888.

Bose:2021:CRB

[858] Prosenjit Bose, Matias Korman, André van Renssen, and Sander Verdonschot. Constrained routing between non-visible vertices. Theoretical Computer Science, 861(??): 144–154, March 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100089X.

Anonymous:2021:EBn

[859] Anonymous. Editorial Board. Theoretical Computer Science, 862(??): ii, March 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001353.

Kari:2021:Eb

[860] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 862(??):iii, March 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001389.

Kari:2021:FRC

[861] Lila Kari, Ion Petre, Grzegorz Rozenberg, and Arto Salomaa. A fascinating rainbow of computation — Honoring Gheorghe Păun on the occasion of his 70th birthday. Theoretical Computer Science, 862(??): 1–2, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001250.

Trejo:2021:NQR

[862] José Manuel Agüero Trejo and Cristian S. Calude. A new quantum random number generator certified by value indefiniteness. Theoretical Computer Science, 862(??): 3–13, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304679.

Bilbie:2021:SSP

[863] Florin-Daniel Bîlbîe and Andrei Păun. Small SNQ P Systems with multiple types of spikes. Theoretical Computer Science, 862(??):14– 23, March 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305855.

Bonizzoni:2021:LCP

[864] Paola Bonizzoni, Clelia De Felice, Rocco Zaccagnino, and Rosalba Zizza. On the longest common prefix of suffixes in an inverse Lyndon factorization and other properties. Theoretical Computer Science, 862(??): 24–41, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306253.

Bonizzoni:2021:CMS

[865] Paola Bonizzoni, Gianluca Della Vedova, Yuri Pirola, Marco Previtali, and Raffaella Rizzi. Computing the multi-string BWT and LCP array in external memory. Theoretical Computer Science, 862(??):

42–58, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306885.

Bottoni:2021:CVD

[866] Paolo Bottoni and Anna Labella. A compositional view of derivations as interactive processes with applications to regulated and distributed rewriting. Theoretical Computer Science, 862 (??):59–80, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306241.

Bockenhauer:2021:ACO

[867] Hans-Joachim Böckenhauer, Juraj Hromkovič, Sacha Krug, and Walter Unger. On the advice complexity of the online dominating set problem. Theoretical Computer Science, 862(??): 81–96, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000438.

Carpi:2021:RBB

[868] Arturo Carpi, Flavio D’Alessandro, Oscar H. Ibarra, and Ian McQuillan. Relationships between bounded languages, counter machines, finiteindex grammars, ambiguity, and commutative regularity. Theoretical Computer Science, 862(??):97– 118, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305776.

Dassow:2021:REC

[869] Jürgen Dassow. Remarks on external contextual grammars with selection. Theoretical Computer Science, 862(??): 119–129, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304126.

Franco:2021:ERS

[870] Giuditta Franco, Vincenzo Manca, Marco Andreolli, and Silvia Lampis. Emergence of random selections in evolution of biological populations. Theoretical Computer Science, 862(??): 130–143, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304862.

Han:2021:CSP

[871] Yo-Sub Han, Sang-Ki Ko, Timothy Ng, and Kai Salomaa. Closest substring problems for regular languages. Theoretical Computer Science, 862(??): 144–154, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030503X.

Harju:2021:DSF

[872] Tero Harju. Disposability in squarefree words. Theoretical Computer Science, 862(??):155–159, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520304096.

Ipate:2021:FRL

[873] Florentin Ipate, Marian Gheorghe, and Raluca Lefticaru. Fundamental results for learning deterministic extended finite state machines from queries. Theoretical Computer Science, 862(??): 160–173, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520305272.

Janicki:2021:RSC

[874] Ryszard Janicki, Jetty Kleijn, Maciej Koutny, and Łukasz Mikulski. Relational structures for concurrent behaviours. Theoretical Computer Science, 862(??):174–192, March 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520306022.

Kimoto:2021:RCA

[875] Nobuya Kimoto, Shigetaka Nakamura, Ken Komiya, Kenzo Fujimoto, and Satoshi Kobayashi. Reducing control alphabet size for the control of right linear grammars with unknown behaviors. Theoretical Computer Science, 862(??): 193–213, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307040.

Mitrana:2021:FDR

[876] Victor Mitrana and Maria C. Negru. Filters defined by random contexts versus polarization in networks of evolutionary processors. Theoretical Computer Science, 862(??):

214–223, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520304448.

Okubo:2021:CPR

[877] Fumiya Okubo and Takashi Yokomori. On the computing powers of L-reductions of insertion languages. Theoretical Computer Science, 862(??): 224–235, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752030668X. See

corrigendum [1525].

Orellana-Martin:2021:PTM

[878] David Orellana-Martín, Luis Valencia-Cabrera, and Mario J. Pérez-Jiménez. Proof techniques in Membrane Computing. Theoretical Computer Science, 862(??):236–249, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397520305521.

Wu:2021:SNP

[879] Tingfang Wu, Luping Zhang, and Linqiang Pan. Spiking neural P systems with target indications. Theoretical Computer Science, 862(??): 250–261, March 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520303972.

Anonymous:2021:PAa

[880] Anonymous. Pages 1–156 (8 April 2021). Theoretical Computer Science, 863(??):1–156, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBo

[881] Anonymous. Editorial Board. Theoretical Computer Science, 863(??): ii, April 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001596.

Demange:2021:OPT

[882] Marc Demange, David Ellison, and Bertrand Jouve. Orienteering problem with time-windows and updating delay. Theoretical Computer Science, 863(??):1–18, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000220.

Ju:2021:RNC

[883] GwangSu Ju and UnGwang Ko. Research on a novel construction of probabilistic visual cryptography scheme (k, n, 0, 1, 1)-PVCS for threshold access structures. Theoretical Computer Science, 863(??):19– 39, April 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000840.

Liotta:2021:OPV

[884] Giuseppe Liotta, Fabrizio Montecchiani, and Alessandra Tappini. Orthopolygon visibility representations of 3-connected 1-plane graphs. Theoretical Computer Science, 863(??):

40–52, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000906.

Yamada:2021:SRS

[885] Takeshi Yamada and Ryuhei Uehara. Shortest reconfiguration of sliding tokens on subclasses of interval graphs. Theoretical Computer Science, 863 (??):53–68, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000918.

Amanatidis:2021:MNW

[886] Georgios Amanatidis, Georgios Birmpas, Aris Filos-Ratsikas, Alexandros Hollender, and Alexandros A. Voudouris. Maximum Nash welfare and other stories about EFX. Theoretical Computer Science, 863 (??):69–85, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000931.

McCollum:2021:QFL

[887] Joey McCollum and Thomas Krauss. QUBO formulations of the longest path problem. Theoretical Computer Science, 863(??):86–101, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752100092X.

Albers:2021:NRS

[888] Susanne Albers and Leon Ladewig. New results for the k-secretary problem. Theoretical Computer Science, 863(??):102–119, April 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521001122.

Milosevic:2021:SWI

[889] Marko Milosevic and Narad Rampersad. Squarefree words with interior disposable factors. Theoretical Computer Science, 863(??):120– 126, April 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001109.

Emura:2021:ERI

[890] Keita Emura, Jae Hong Seo, and Yohei Watanabe. Efficient revocable identity-based encryption with short public parameters. Theoretical Computer Science, 863(??):127– 155, April 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001134.

Anonymous:2021:PAb

[891] Anonymous. Pages 1–140 (10 April 2021). Theoretical Computer Science, 864(??):1–140, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBp

[892] Anonymous. Editorial Board. Theoretical Computer Science, 864(??): ii, April 10, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001730.

Liu:2021:MFL

[893] Wenjing Liu, Yuan Ding, Xin Chen, Qizhi Fang, and Qingqin Nong. Multiple facility location games with envy ratio. Theoretical Computer Science, 864(??):1–9, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000293.

Wang:2021:IHA

[894] Xing Wang, Guangting Chen, Yong Chen, Guohui Lin, Yonghao Wang, and An Zhang. Improved hardness and approximation results for single allocation hub location problems. Theoretical Computer Science, 864(??): 10–19, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000761.

Gorain:2021:FSD

[895] Barun Gorain and Andrzej Pelc. Finding the size and the diameter of a radio network using short labels. Theoretical Computer Science, 864(??): 20–33, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000797.

Liu:2021:PCC

[896] Xianmin Liu, Jianzhong Li, Yingshu Li, and Yuqiang Feng. Parameterized complexity of completeness reasoning for conjunctive queries. Theoretical Computer Science, 864(??): 34–49, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000785.

Wang:2021:IPD

[897] Xiaomeng Wang, Shou-Jun Xu, and Xianyue Li. Independent perfect dominating sets in semi-Cayley graphs. Theoretical Computer Science, 864 (??):50–57, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100075X.

Zhang:2021:UTL

[898] Nan Zhang, Zhenhua Duan, and Cong Tian. Unified temporal logic. Theoretical Computer Science, 864(??): 58–69, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000773.

Nie:2021:AAG

[899] Zixiong Nie and Hongwei Du. An approximation algorithm for General Energy Restricted Sweep Coverage problem. Theoretical Computer Science, 864(??):70–79, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100102X.

Cameron:2021:CGF

[900] Kathie Cameron, Jan Goedgebeur, Shenwei Huang, and Yongtang Shi. k-Critical graphs in P _{5}-free graphs. Theoretical Computer Science, 864 (??):80–91, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001067.

Fazekas:2021:TWD

[901] Szilárd Zsolt Fazekas, Kaito Hoshi, and Akihiro Yamamura. Two-way deterministic automata with jumping mode. Theoretical Computer Science, 864(??): 92–102, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001079.

Durand:2021:EEM

[902] Martin Durand and Fanny Pascual. Efficiency and equity in the multi organization scheduling problem. Theoretical Computer Science, 864(??): 103–117, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001055.

Liang:2021:MPP

[903] Wei Liang, Menghong Li, Zhao Zhang, and Xiaohui Huang. Minimum power partial multi-cover on a line. Theoretical Computer Science, 864(??): 118–128, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001031.

Zhang:2021:PMN

[904] Hongxiang Zhang, Dachuan Xu, Longkun Guo, and Jingjing Tan. Parallelized maximization of nonsubmodular function subject to a cardinality constraint. Theoretical Computer Science, 864(??):129– 137, April 10, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001006.

Cherubini:2021:TCS

[905] Alessandra Cherubini, Nicoletta Sabadini, and Simone Tini. Theoretical computer science in Italy. Theoretical Computer Science, 864(??): 138, April 10, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100164X.

Das:2021:ADM

[906] Gautam Kumar Das, Subhas C. Nandy, and M. Sohel Rahman. Algorithms and discrete mathematics — celebrating the silver jubilee of IITG Guwahati. Theoretical Computer Science, 864(??): 139–140, April 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001651.

Anonymous:2021:PAc

[907] Anonymous. Pages 1–140 (14 April 2021). Theoretical Computer Science, 865(??):1–140, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBq

[908] Anonymous. Editorial Board. Theoretical Computer Science, 865(??): ii, April 14, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001948.

Ross:2021:TPE

[909] Michael S. Ross. Throttling processes equivalent to full throttling on trees. Theoretical Computer Science, 865(??):1–11, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001110.

Hasanzadeh:2021:CCI

[910] Mehran Hasanzadeh, Behrooz Alizadeh, and Fahimeh Baroughi. The cardinality constrained inverse center location problems on tree networks with edge length augmentation. Theoretical Computer Science, 865(??):12–33, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001080.

Ghareghani:2021:SFC

[911] N. Ghareghani and P. Sharifani. On square factors and critical factors of k-bonacci words on infinite alphabet. Theoretical Computer Science, 865 (??):34–43, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001092.

Ciftci:2021:CBD

[912] Canan Çiftçi. Complexity and bounds for disjunctive total bondage. Theoretical Computer Science, 865(??): 44–51, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001043.

Luo:2021:ABP

[913] Fucai Luo, Saif Al-Kuwari, Fuqun Wang, and Kefei Chen. Attributebased proxy re-encryption from standard lattices. Theoretical Computer Science, 865(??):52–62, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521001201.

Schaller:2021:CMP

[914] David Schaller, Peter F. Stadler, and Marc Hellmuth. Complexity of modification problems for best match graphs. Theoretical Computer Science, 865 (??):63–84, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001213.

Wang:2021:IGT

[915] Di Wang and Jinhui Xu. Inferring ground truth from crowdsourced data under local attribute differential privacy. Theoretical Computer Science, 865(??):85–98, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001237.

Golin:2021:SAD

[916] Mordecai Golin and Elfarouk Harb. Speeding up the AIFV-2 dynamic programs by two orders of magnitude using Range Minimum Queries. Theoretical Computer Science, 865(??): 99–118, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001249.

Wang:2021:DPH

[917] Di Wang and Jinhui Xu. Differentially private high dimensional sparse covariance matrix estimation. Theoretical Computer Science, 865(??): 119–130, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001407.

Hanaka:2021:FMM

[918] Tesshu Hanaka, Yasuaki Kobayashi, Yusuke Kobayashi, and Tsuyoshi Yagita. Finding a maximum minimal separator: Graph classes and fixed-parameter tractability. Theoretical Computer Science, 865(??): 131–140, April 14, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001444.

Anonymous:2021:PAd

[919] Anonymous. Pages 1–170 (18 April 2021). Theoretical Computer Science, 866(??):1–170, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBr

[920] Anonymous. Editorial Board. Theoretical Computer Science, 866(??): ii, April 18, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002036.

Dippel:2021:RPR

[921] Matthew Dippel, Ravi Sundaram, and Akshar Varma. Realization problems on reachability sequences. Theoretical Computer Science, 866(??): 1–13, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001018.

Holub:2021:BIF

[922] Štěpán Holub and Štěpán Starosta. Binary intersection formalized. Theoretical Computer Science, 866(??): 14–24, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001481.

Castiglione:2021:PSW

[923] Giuseppa Castiglione, Gabriele Fici, and Antonio Restivo. Primitive sets of words. Theoretical Computer Science, 866(??):25–36, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001468.

Li:2021:SCR

[924] Yan Li, Yusheng Li, and Ye Wang. Star-critical Ramsey number of large cycle and book of different orders. Theoretical Computer Science, 866 (??):37–42, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001511.

Tan:2021:TRR

[925] Xuehou Tan. The touring rays and related problems. Theoretical Computer Science, 866(??):43– 55, April 18, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100150X.

Currie:2021:URT

[926] James D. Currie and Lucas Mol. The undirected repetition threshold and undirected pattern avoidance. Theoretical Computer Science, 866(??): 56–69, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100147X.

Crochemore:2021:SCA

[927] Maxime Crochemore, Costas S. Iliopoulos, Jakub Radoszewski, Wojciech Rytter, Juliusz Straszyński, Tomasz Waleń, and Wiktor Zuba. Shortest covers of all cyclic shifts of a string. Theoretical Computer Science, 866(??):70–81, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001493.

Yang:2021:MSL

[928] Hyeyun Yang and Antoine Vigneron. Matching sets of line segments. Theoretical Computer Science, 866(??): 82–95, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001547.

Sarker:2021:LTA

[929] Anik Sarker, Wing-Kin Sung, and M. Sohel Rahman. A linear time algorithm for the r-gathering problem on the line. Theoretical Computer Science, 866(??):96– 106, April 18, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001535.

Klootwijk:2021:PAO

[930] Stefan Klootwijk, Bodo Manthey, and Sander K. Visser. Probabilistic analysis of optimization problems on generalized random shortest path metrics. Theoretical Computer Science, 866 (??):107–122, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001559.

Asahiro:2021:CAH

[931] Yuichi Asahiro, Hiroshi Eto, Tesshu Hanaka, Guohui Lin, Eiji Miyano, and Ippei Terabaru. Complexity and approximability of the happy set problem. Theoretical Computer Science, 866(??): 123–144, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001699. See

corrigendum [2212].

vanderHoog:2021:TSK

[932] Ivor van der Hoog, Marc van Kreveld, Wouter Meulemans, Kevin Verbeek, and Jules Wulms. Topological stability of kinetic k-centers. Theoretical Computer Science, 866(??): 145–159, April 18, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001791.

Chau:2021:MEH

[933] Vincent Chau, Chi Kit Ken Fong, Shengxin Liu, Elaine Yinling Wang, and Yong Zhang. Minimizing energy on homogeneous processors with shared memory. Theoretical Computer Science, 866(??):160– 170, April 18, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001833.

Anonymous:2021:PMd

[934] Anonymous. Pages 1–140 (6 May 2021). Theoretical Computer Science, 867(??):1–140, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBs

[935] Anonymous. Editorial Board. Theoretical Computer Science, 867(??): ii, May 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100219X.

Kanesh:2021:PCF

[936] Lawqueen Kanesh, Soumen Maity, Komal Muluk, and Saket Saurabh. Parameterized complexity of fair feedback vertex set problem. Theoretical Computer Science, 867(??):1– 12, May 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001432.

Babou:2021:TMO

[937] Nadia Babou, Djamal Rebaine, and Mourad Boudhar. Two-machine open shop problem with a single server and set-up time considerations. Theoretical Computer Science, 867(??): 13–29, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001419.

Bodlaender:2021:STH

[938] Hans L. Bodlaender, Nick Brettell, Matthew Johnson, Giacomo Paesani, Daniël Paulusma, and Erik Jan van Leeuwen. Steiner trees for hereditary graph classes: a treewidth perspective. Theoretical Computer Science, 867(??):30–39, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001420.

Cheng:2021:FGO

[939] Siu-Wing Cheng, Otfried Cheong, Taegyoung Lee, and Zhengtong Ren. Fitting a graph to one-dimensional data. Theoretical Computer Science, 867(??):40–49, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001663.

Halldorsson:2021:QCS

[940] Magnús M. Halldórsson and Murilo Santos de Lima. Query-competitive sorting with uncertainty. Theoretical Computer Science, 867(??):50– 67, May 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001675.

Dragan:2021:HGG

[941] Feodor F. Dragan and Heather M. Guarnera. Helly-gap of a graph and vertex eccentricities. Theoretical Computer Science, 867(??):68– 84, May 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001687.

Brandenburg:2021:FCF

[942] Franz J. Brandenburg. Fan-crossing free graphs and their relationship to other beyond-planar graphs. Theoretical Computer Science, 867(??): 85–100, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001869.

Kostolanyi:2021:RES

[943] Peter Kostolányi. Rational elements of summation semirings. Theoretical Computer Science, 867(??): 101–127, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001845.

Pelantova:2021:SSC

[944] Edita Pelantová and Štěpán Starosta. On Sturmian substitutions closed under derivation. Theoretical Computer Science, 867(??):128–139, May 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521001857.

Anonymous:2021:PMe

[945] Anonymous. Pages 1–112 (8 May 2021). Theoretical Computer Science, 868(??):1–112, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBt

[946] Anonymous. Editorial Board. Theoretical Computer Science, 868(??): ii, May 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002267.

Lehner:2021:CCL

[947] Florian Lehner and Christian Lindorfer. Comparing consecutive letter counts in multiple context-free languages. Theoretical Computer Science, 868(??):1–5, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001870.

Li:2021:VKD

[948] Wenjun Li, Yang Ding, Yongjie Yang, and Guozhen Rong. A (2 + ϵ)kvertex kernel for the dual coloring problem. Theoretical Computer Science, 868(??):6–11, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001882.

Shi:2021:PMC

[949] Qihao Shi, Can Wang, Deshi Ye, Jiawei Chen, Sheng Zhou, Yan Feng, Chun Chen, and Yanhao Huang. Profit maximization for competitive social advertising. Theoretical Computer Science, 868(??):12–29, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001900.

Ito:2021:AGG

[950] Takehiro Ito, Naoyuki Kamiyama, Yusuke Kobayashi, and Yoshio Okamoto. Algorithms for gerrymandering over graphs. Theoretical Computer Science, 868(??):30–45, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001894.

Heeger:2021:MGP

[951] Klaus Heeger, Anne-Sophie Himmel, Frank Kammer, Rolf Niedermeier, Malte Renken, and Andrej Sajenko.

Multistage graph problems on a global budget. Theoretical Computer Science, 868(??):46–64, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002061.

Birx:2021:ILB

[952] Alexander Birx, Yann Disser, Alexander V. Hopp, and Christina Karousatou. An improved lower bound for competitive graph exploration. Theoretical Computer Science, 868(??): 65–86, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002073.

Chalopin:2021:CDF

[953] Jérémie Chalopin, Shantanu Das, Yann Disser, Arnaud Labourel, and Matúš Mihalák. Collaborative delivery on a fixed path with homogeneous energy-constrained agents. Theoretical Computer Science, 868 (??):87–96, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002097.

Frosini:2021:NSC

[954] Andrea Frosini, Christophe Picouleau, and Simone Rinaldi. New sufficient conditions on the degree sequences of uniform hypergraphs. Theoretical Computer Science, 868(??): 97–111, May 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002103.

Anonymous:2021:PMf

[955] Anonymous. Pages 1–194 (12 May 2021). Theoretical Computer Science, 869(??):1–194, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBu

[956] Anonymous. Editorial Board. Theoretical Computer Science, 869(??): ii, May 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002474.

Spreen:2021:GIS

[957] Dieter Spreen. Generalised information systems capture L-domains. Theoretical Computer Science, 869 (??):1–28, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000013.

Lanotte:2021:WSA

[958] Ruggero Lanotte and Simone Tini. A weak semantic approach to bisimulation metrics in models with nondeterminism and continuous state spaces. Theoretical Computer Science, 869 (??):29–61, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000359.

Cousot:2021:CDR

[959] Patrick Cousot. Calculational design of a regular model checker by abstract interpretation. Theoretical Computer Science, 869(??):62– 84, May 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000712.

Bice:2021:DDC

[960] Tristan Bice. Distance domains: Continuity. Theoretical Computer Science, 869(??):85–107, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000724.

Marsik:2021:ICE

[961] Jirka Maršík, Maxime Amblard, and Philippe de Groote. Introducing |(λ)|, a λ-calculus for effectful computation. Theoretical Computer Science, 869(??): 108–155, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001225.

Auletta:2021:OMD

[962] Vincenzo Auletta, Diodato Ferraioli, and Gianluigi Greco. Optimal majority dynamics for the diffusion of an opinion when multiple alternatives are available. Theoretical Computer Science, 869(??):156– 180, May 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001638.

Fu:2021:MIA

[963] Yuxi Fu. Model independent approach to probabilistic models. Theoretical Computer Science, 869(??): 181–194, May 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001997.

Anonymous:2021:EBv

[964] Anonymous. Editorial Board. Theoretical Computer Science, 870(??): ii, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002565.

Kari:2021:Ec

[965] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 870(??):iii, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002590.

Jiraskova:2021:SII

[966] Galina Jirásková and Cezar Câmpeanu. Special issue on implementation and application of automata. Theoretical Computer Science, 870(??):1– 2, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002528.

Bjorklund:2021:BUT

[967] Johanna Björklund, Shay B. Cohen, Frank Drewes, and Giorgio Satta. Bottom-up unranked treeto-graph transducers for translation into semantic graphs. Theoretical Computer Science, 870(??):3– 28, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307167.

Han:2021:GAO

[968] Yo-Sub Han, Hwee Kim, Yusei Masuda, and Shinnosuke Seki. A general architecture of oritatami systems for simulating arbitrary finite automata. Theoretical Computer Science, 870 (??):29–52, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307155.

Hospodar:2021:PPC

[969] Michal Hospodár. Power, positive closure, and quotients on convex languages. Theoretical Computer Science, 870(??):53–74, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000748.

Kapoutsis:2021:ATW

[970] Christos Kapoutsis and Mohammad Zakzok. Alternation in twoway finite automata. Theoretical Computer Science, 870(??):75– 102, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307179.

Konstantinidis:2021:PDR

[971] Stavros Konstantinidis, Nelma Moreira, and Rogério Reis. Partial derivatives of regular expressions over alphabet-invariant and user-defined labels. Theoretical Computer Science, 870(??):103–120, May 16, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307544.

Kutrib:2021:IDM

[972] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Inputdriven multi-counter automata. Theoretical Computer Science, 870(??): 121–136, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521000608.

Mraz:2021:TDP

[973] František Mráz, Daniel Pruša, and Michael Wehar. Two-dimensional pattern matching against local and regular-like picture languages. Theoretical Computer Science, 870(??): 137–152, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307490.

Smith:2021:DPP

[974] Taylor J. Smith and Kai Salomaa. Decision problems and projection languages for restricted variants of two-dimensional automata. Theoretical Computer Science, 870(??):

153–164, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307532.

Takahashi:2021:SRT

[975] Yuta Takahashi, Kazuyuki Asada, and Keisuke Nakano. Streaming ranked-tree-to-string transducers. Theoretical Computer Science, 870(??): 165–187, May 16, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307556.

Anonymous:2021:PJf

[976] Anonymous. Pages 1–146 (6 June 2021). Theoretical Computer Science, 871(??):1–146, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBw

[977] Anonymous. Editorial Board. Theoretical Computer Science, 871(??): ii, June 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002784.

Frohlich:2021:HCI

[978] Nicolas Fröhlich and Stefan Ruzika. On the hardness of covering-interdiction problems. Theoretical Computer Science, 871(??):1–15, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521002115.

Dutta:2021:CRI

[979] Priyanka Dutta, Willy Susilo, Dung Hoang Duong, and Partha Sarathi Roy. Collusion-resistant identity-based Proxy Re-encryption: Lattice-based constructions in Standard Model. Theoretical Computer Science, 871(??): 16–29, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002127.

Gu:2021:ARP

[980] Qian-Ping Gu, Jiajian Leo Liang, and Guochuan Zhang. Approximate ridesharing of personal vehicles problem. Theoretical Computer Science, 871(??):30–50, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002140.

Zhang:2021:TAE

[981] Huaqiang Zhang and Xin Zhang. Theoretical aspects of equitable partition of networks into sparse modules. Theoretical Computer Science, 871(??):51–61, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002139.

Amanatidis:2021:IAP

[982] Georgios Amanatidis, Peter Fulla, Evangelos Markakis, and Krzysztof Sornat. Inequity aversion pricing over social networks: Approximation algorithms and hardness results. Theoretical Computer Science, 871 (??):62–78, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002346.

Lin:2021:HGT

[983] Chuang-Chieh Lin, Chi-Jen Lu, and Po-An Chen. How good is a twoparty election game? Theoretical Computer Science, 871(??):79– 93, June 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002322.

Kumar:2021:TAC

[984] Vinod Kumar and Krishnendra Shekhawat. A transformation algorithm to construct a rectangular floorplan. Theoretical Computer Science, 871(??): 94–106, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002310.

Ferragina:2021:PLD

[985] Paolo Ferragina, Fabrizio Lillo, and Giorgio Vinciguerra. On the performance of learned data structures. Theoretical Computer Science, 871(??): 107–120, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002334.

Yang:2021:ECT

[986] Yuxing Yang. Embedded connectivity of ternary n-cubes. Theoretical Computer Science, 871(??):

121–125, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002358.

Wang:2021:AAS

[987] Xin Wang, Suogang Gao, Bo Hou, Lidong Wu, and Wen Liu. Approximation algorithms for the submodular edge cover problem with submodular penalties. Theoretical Computer Science, 871 (??):126–133, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100236X.

Egecioglu:2021:NSC

[988] Ömer Eğecioğlu, Elif Saygı, and Zülfükar Saygı. The number of short cycles in Fibonacci cubes. Theoretical Computer Science, 871(??): 134–146, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002395.

Anonymous:2021:PJg

[989] Anonymous. Pages 1–116 (8 June 2021). Theoretical Computer Science, 872(??):1–116, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBx

[990] Anonymous. Editorial Board. Theoretical Computer Science, 872(??): ii, June 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002851.

Kari:2021:Ed

[991] Lila Kari.  Editorial.  Theo-

retical Computer Science, 872(??): iii, June 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002899.

Teh:2021:RSM

[992] Wen Chean Teh, Kien Trung Nguyen, and Chuei Yee Chen. Ranks of strictly minimal reaction systems induced by permutations. Theoretical Computer Science, 872(??):1– 14, June 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307209.

Souza:2021:DPL

[993] Marlo Souza, Renata Vieira, and Álvaro Moreira. Dynamic Preference Logic meets iterated belief change: Representation results and postulates characterization. Theoretical Computer Science, 872(??): 15–40, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307751.

Coimbra:2021:AQL

[994] Danilo B. Coimbra, Rafael M. Martins, Edson Mota, Tacito Tiburtino, Pedro Diamantino, and Maycon L. M. Peixoto. Analyzing the quality of local and global multidimensional projections using performance evaluation planning. Theoretical Computer Science, 872(??): 41–54, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752030774X.

Furcy:2021:SAO

[995] David Furcy, Scott M. Summers, and Christian Wendlandt. Self-assembly of and optimal encoding within thin rectangles at temperature-1 in 3D. Theoretical Computer Science, 872 (??):55–78, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521000736.

Adak:2021:ISG

[996] Sumit Adak and Sukanta Das. (Imperfect) strategies to generate primitive polynomials over GF (2). Theoretical Computer Science, 872(??): 79–96, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001456.

Pardo:2021:SRB

[997] Jérémie Pardo, Sergiu Ivanov, and Franck Delaplace. Sequential reprogramming of biological network fate. Theoretical Computer Science, 872 (??):97–116, June 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001523.

Anonymous:2021:PJh

[998] Anonymous. Pages 1–114 (10 June 2021). Theoretical Computer Science, 873(??):1–114, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBy

[999] Anonymous. Editorial Board. Theoretical Computer Science, 873(??): ii, June 10, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002930.

Bazgan:2021:DAU

[1000] Cristina Bazgan, Pierre Cazals, and Janka Chlebíková. Degreeanonymization using edge rotations. Theoretical Computer Science, 873 (??):1–15, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002413.

Liu:2021:PAF

[1001] Yunlong Liu, Jie Chen, Jingui Huang, and Jianxin Wang. On parameterized algorithms for fixed-order book thickness with respect to the pathwidth of the vertex ordering. Theoretical Computer Science, 873(??): 16–24, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002425.

Lamprou:2021:MRC

[1002] Ioannis Lamprou, Russell Martin, Sven Schewe, Ioannis Sigalas, and Vassilis Zissimopoulos. Maximum rooted connected expansion. Theoretical Computer Science, 873(??): 25–37, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002401.

Hanaka:2021:POA

[1003] Tesshu Hanaka, Yasuaki Kobayashi, and Taiga Sone. A (probably) optimal algorithm for Bisection on bounded-treewidth graphs. Theoretical Computer Science, 873(??): 38–46, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002437.

Ma:2021:SBT

[1004] Jingjing Ma, Daming Zhu, Haitao Jiang, and Binhai Zhu. On the solution bound of two-sided scaffold filling. Theoretical Computer Science, 873 (??):47–63, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002371.

Li:2021:DPC

[1005] Luyi Li and Xueliang Li. Digraphs with proper connection number two. Theoretical Computer Science, 873 (??):64–75, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002504.

Yan:2021:IAL

[1006] Di Yan, Yu Yu, Hanlin Liu, Shuoyao Zhao, and Jiang Zhang. An improved algorithm for learning sparse parities in the presence of noise. Theoretical Computer Science, 873(??): 76–86, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002516.

Higashikawa:2021:ALT

[1007] Yuya Higashikawa, Naoki Katoh, Junichi Teruyama, and Koji Watase. Almost linear time algorithms for minsum k-sink problems on dynamic flow path networks. Theoretical Computer Science, 873(??): 87–113, June 10, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002620.

Anonymous:2021:PJi

[1008] Anonymous. Pages 1–94 (12 June 2021). Theoretical Computer Science, 874(??):1–94, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBz

[1009] Anonymous. Editorial Board. Theoretical Computer Science, 874(??): ii, June 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002991.

Kim:2021:SSA

[1010] Yonghwan Kim, Masahiro Shibata, Yuichi Sudo, Junya Nakamura, Yoshiaki Katayama, and Toshimitsu Masuzawa. A self-stabilizing algorithm for constructing a minimal reachable directed acyclic graph with two senders and two targets. Theoretical Computer Science, 874(??): 1–14, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002632.

Dey:2021:PPP

[1011] Palash Dey, Neeldhara Misra, Swaprava Nath, and Garima Shakya. A parameterized perspective on protecting elections. Theoretical Computer Science, 874(??):15–31, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002656.

Kurita:2021:CAT

[1012] Kazuhiro Kurita, Kunihiro Wasa, Takeaki Uno, and Hiroki Arimura. A constant amortized time enumeration algorithm for independent sets in graphs with bounded clique number. Theoretical Computer Science, 874 (??):32–41, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002668.

Geneson:2021:NPB

[1013] Jesse Geneson. A note on the price of bandit feedback for mistakebounded online learning. Theoretical Computer Science, 874(??):42– 45, June 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002711.

Bshouty:2021:ODG

[1014] Nader H. Bshouty and Catherine A.

Haddad-Zaknoon. Optimal deterministic group testing algorithms to estimate the number of defectives. Theoretical Computer Science, 874(??): 46–58, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002723.

Liotta:2021:SFO

[1015] Giuseppe Liotta, Ignaz Rutter, and Alessandra Tappini. Simultaneous FPQ-ordering and hybrid planarity testing. Theoretical Computer Science, 874(??):59–79, June 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002693.

Bonato:2021:GCE

[1016] Anthony Bonato, Melissa A. Huggan, Trent G. Marbach, and Fionn Mc Inerney. The game of Cops and Eternal Robbers. Theoretical Computer Science, 874(??):80– 93, June 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002681.

Anonymous:2021:PJj

[1017] Anonymous. Pages 1–80 (6 July 2021). Theoretical Computer Science, 875(??): 1–80, July 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBba

[1018] Anonymous. Editorial Board. Theoretical Computer Science, 875(??): ii, July 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003170.

Kari:2021:Ee

[1019] Lila Kari.  Editorial.  Theo-

retical Computer Science, 875(??): iii, July 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003200.

Brodo:2021:LGF

[1020] Linda Brodo, Roberto Bruni, and Moreno Falaschi. A logical and graphical framework for reaction systems. Theoretical Computer Science, 875 (??):1–27, July 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001778.

Lengler:2021:ESL

[1021] Johannes Lengler and Xun Zou. Exponential slowdown for larger populations: the (μ + 1)-EA on monotone functions. Theoretical Computer Science, 875(??):28– 51, July 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100178X.

Assaf:2021:CFS

[1022] George Assaf, Monika Heiner, and Fei Liu. Colouring fuzziness for systems biology. Theoretical Computer Science, 875(??):52–64, July 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002152.

Bhojraj:2021:PFQ

[1023] Tejas Bhojraj. Prefix-free quantum Kolmogorov complexity. Theoretical Computer Science, 875(??): 65–80, July 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002814.

Anonymous:2021:PJk

[1024] Anonymous. Pages 1–78 (12 July 2021). Theoretical Computer Science, 876(??):1–78, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbb

[1025] Anonymous. Editorial Board. Theoretical Computer Science, 876(??): ii, July 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003248.

Mitchell:2021:MMC

[1026] Joseph S. B. Mitchell and Supantha Pandit. Minimum membership covering and hitting. Theoretical Computer Science, 876(??):1– 11, July 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002619.

Purcell:2021:RCG

[1027] Christopher Purcell and Puck Rombach. Role colouring graphs in hereditary classes. Theoretical Computer Science, 876(??):12–24, July 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100270X.

Hohne:2021:BMC

[1028] Felix Höhne and Rob van Stee. Buffer minimization with conflicts on a line. Theoretical Computer Science, 876 (??):25–33, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002735.

Zhu:2021:FAS

[1029] Chun Jiang Zhu, Song Han, and Kam-Yiu Lam. A fast algorithm for source-wise round-trip spanners. Theoretical Computer Science, 876 (??):34–44, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003029.

Panda:2021:DTD

[1030] B. S. Panda, Pooja Goyal, and D. Pradhan. Differentiating-total domination: Approximation and hardness results. Theoretical Computer Science, 876(??):45–58, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003066.

Jiang:2021:IAT

[1031] Xiaojuan Jiang, An Zhang, Yong Chen, Guangting Chen, and Kangbok Lee. An improved algorithm for a two-stage production scheduling problem with an outsourcing option. Theoretical Computer Science, 876(??):59–69, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003054.

Wang:2021:VPB

[1032] Xin Wang, Chaoqun Ma, and Jia Guo. Vertex-pancyclicity of the (n, k)-bubble-sort networks. Theoretical Computer Science, 876(??): 70–78, July 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003078.

Anonymous:2021:PJl

[1033] Anonymous. Pages 1–90 (20 July 2021). Theoretical Computer Science, 877(??):1–90, July 20, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbc

[1034] Anonymous. Editorial Board. Theoretical Computer Science, 877(??): ii, July 20, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003303.

Kerkmann:2021:LFH

[1035] Anna Maria Kerkmann, Nhan-Tam Nguyen, and Jörg Rothe. Local fairness in hedonic games via individual threshold coalitions. Theoretical Computer Science, 877(??):1– 17, July 20, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001808.

Galby:2021:BTD

[1036] E. Galby, F. Mann, and B. Ries. Blocking total dominating sets via edge contractions. Theoretical Computer Science, 877(??):18–35, July 20, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752100181X.

Urrutia:2021:RSS

[1037] Sebastián Urrutia, Dominique de Werra, and Tiago Januario. Recoloring subgraphs of K_{2n} for sports scheduling. Theoretical Computer Science, 877 (??):36–45, July 20, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521001821.

Monnot:2021:AAU

[1038] Jérôme Monnot, Henning Fernau, and David Manlove. Algorithmic aspects of upper edge domination. Theoretical Computer Science, 877(??): 46–57, July 20, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521001985.

Demange:2021:GOC

[1039] Marc Demange and Martin Olsen.

Generalised online colouring problems in overlap graphs. Theoretical Computer Science, 877(??):58– 73, July 20, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002607.

Alfandari:2021:ADT

[1040] Laurent Alfandari and Sophie Toulouse.

Approximation of the Double Traveling Salesman Problem with Multiple Stacks. Theoretical Computer Science, 877(??):74–89, July 20, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002747.

Anonymous:2021:PJm

[1041] Anonymous. Pages 1–112 (22 July 2021). Theoretical Computer Science, 878–879(??):1–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbd

[1042] Anonymous. Editorial Board. Theoretical Computer Science, 878–879 (??):ii, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003777.

Bi:2021:ICMa

[1043] Huimin Bi and Xin Zhang. Incidence coloring of Mycielskians with fast algorithm. Theoretical Computer Science, 878–879(??):1– 10, July 22, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003108.

Brandstadt:2021:MWIa

[1044] Andreas Brandstädt and Raffaele Mosca. Maximum weight independent sets for (S_{1,2,4}, triangle)-free graphs in polynomial time. Theoretical Computer Science, 878–879 (??):11–25, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100308X.

Duchene:2021:IPSa

[1045] Eric Duchêne, Stéphane Gonzalez, Aline Parreau, Eric Rémila, and Philippe Solal. influence: a partizan scoring game on graphs. Theoretical Computer Science, 878–879 (??):26–46, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003091.

Barcucci:2021:EGSa

[1046] Elena Barcucci, Antonio Bernini, and Renzo Pinzani. Exhaustive generation of some lattice paths and their prefixes. Theoretical Computer Science, 878–879 (??):47–52, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307180.

Kantor:2021:NSVa

[1047] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. Nonuniform SINR + Voronoi diagrams are effectively uniform. Theoretical Computer Science, 878–879(??): 53–66, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002383.

Zhang:2021:IAPa

[1048] Zhen Zhang, Yutian Guo, Junyu Huang, Jianxin Wang, and Feng Shi. Improved approximation for prizecollecting red-blue median. Theoretical Computer Science, 878–879 (??):67–82, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003133.

Goldreich:2021:ULVa

[1049] Oded Goldreich and Tom Gur. Universal locally verifiable codes and 3-round interactive proofs of proximity for CSP. Theoretical Computer Science, 878–879 (??):83–101, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003339.

Chen:2021:CDCa

[1050] Yu-Han Chen, Shyue-Ming Tang, Kung-Jui Pai, and Jou-Ming Chang. Constructing dual-CISTs with short diameters using a generic adjustment scheme on bicubes. Theoretical Computer Science, 878–879(??): 102–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003352.

Anonymous:2021:PAe

[1051] Anonymous. Pages 1–138 (3 August 2021). Theoretical Computer Science, 880(??):1–138, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbe

[1052] Anonymous. Editorial Board. Theoretical Computer Science, 880(??): ii, August 3, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003832.

Kanellopoulos:2021:MSG

[1053] Panagiotis Kanellopoulos, Maria Kyropoulou, and Alexandros A. Voudouris. Modified Schelling games. Theoretical Computer Science, 880(??): 1–19, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003340.

Zhou:2021:SSC

[1054] Qianru Zhou, Shuming Zhou, Jiafei Liu, and Xiaoqing Liu. Structure and substructure connectivity of divide-and-swap cube. Theoretical Computer Science, 880(??): 20–36, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003364.

Lee:2021:RHI

[1055] Kwangsu Lee. Revocable hierarchical identity-based encryption with adaptive security. Theoretical Computer Science, 880(??):37– 68, August 3, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003376.

Dosa:2021:IPA

[1056] Gyorgy Dosa, Hans Kellerer, Tomas Olaj, and Zsolt Tuza. An improved parametric algorithm on two-machine scheduling with given lower and upper bounds for the total processing time. Theoretical Computer Science, 880(??): 69–81, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003479.

Daykin:2021:CSA

[1057] Jacqueline W. Daykin, Neerja Mhaskar, and W. F. Smyth. Computation of the suffix array, Burrows–Wheeler transform and FM-index in V -order. Theoretical Computer Science, 880(??):

82–96, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003492.

Zhu:2021:DUM

[1058] Yazhou Zhu, Xiang Pan, Tianxu Lv, Yuan Liu, and Lihua Li. DESN: an unsupervised MR image denoising network with deep image prior. Theoretical Computer Science, 880(??): 97–110, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003522.

Goncharov:2021:MGI

[1059] Sergey Goncharov, Christoph Rauch, and Lutz Schröder. A metalanguage for guarded iteration. Theoretical Computer Science, 880(??): 111–137, August 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521002085.

Hierons:2021:P

[1060] Robert M. Hierons and Mohamed Mosbah. Preface. Theoretical Computer Science, 880(??):138, August 3, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004047.

Anonymous:2021:EBbf

[1061] Anonymous. Editorial Board. Theoretical Computer Science, 881(??): ii, August 15, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004515.

Kari:2021:Ef

[1062] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 881(??):iii, August 15, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004552.

Mikulski:2021:PSI

[1063] Łukasz Mikulski and Ion Petre. Preface: Special issue on reaction systems. Theoretical Computer Science, 881(??):1–2, August 15, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100431X.

Barbuti:2021:EBN

[1064] Roberto Barbuti, Roberta Gori, and Paolo Milazzo. Encoding Boolean networks into reaction systems for investigating causal dependencies in gene regulation. Theoretical Computer Science, 881(??):3–24, August 15, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304138.

Bottoni:2021:TCB

[1065] Paolo Bottoni and Anna Labella.

Transactions and contracts based on reaction systems. Theoretical Computer Science, 881(??):25– 61, August 15, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004163.

Brodo:2021:PAA

[1066] Linda Brodo, Roberto Bruni, and Moreno Falaschi. A process algebraic approach to reaction systems. Theoretical Computer Science, 881(??): 62–82, August 15, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520305004.

Genova:2021:CRR

[1067] Daniela Genova, Hendrik Jan Hoogeboom, and Jetty Kleijn. Comparing reactions in reaction systems. Theoretical Computer Science, 881(??): 83–96, August 15, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520307039.

Koutny:2021:APR

[1068] Maciej Koutny, Marta Pietkiewicz-Koutny, and Alex Yakovlev. Asynchrony and persistence in reaction systems. Theoretical Computer Science, 881(??):97–110, August 15, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397520306873.

Kreowski:2021:CAR

[1069] Hans-Jörg Kreowski and Aaron Lye. A categorial approach to reaction systems: First steps. Theoretical Computer Science, 881(??):111– 128, August 15, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520304667.

Anonymous:2021:PAf

[1070] Anonymous. Pages 1–142 (23 August 2021). Theoretical Computer Science, 882(??):1–142, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbg

[1071] Anonymous. Editorial Board. Theoretical Computer Science, 882(??): ii, August 23, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100459X.

Liang:2021:TSP

[1072] Ziwei Liang, He Yuan, and Hongwei Du. Two-stage pricing strategy with price discount in online social networks. Theoretical Computer Science, 882(??): 1–14, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100339X.

Chaudhary:2021:CMM

[1073] Juhi Chaudhary and B. S. Panda. On the complexity of minimum maximal uniquely restricted matching. Theoretical Computer Science, 882(??): 15–28, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003388.

Shieh:2021:EMP

[1074] Yi-Kung Shieh, Shyong Jian Shyu, Chin Lung Lu, and Richard Chia-Tung Lee. The exact multiple pattern matching problem solved by a reference tree approach. Theoretical Computer Science, 882(??): 29–48, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003480.

Rytter:2021:SVS

[1075] Wojciech Rytter and Wiktor Zuba. Syntactic view of sigma-tau generation of permutations. Theoretical Computer Science, 882(??):49– 62, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003534.

Daescu:2021:NBG

[1076] Ovidiu Daescu and Hemant Malik. New bounds on guarding problems for orthogonal polygons in the plane using vertex guards with halfplane vision. Theoretical Computer Science, 882(??): 63–76, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003583.

Shu:2021:AEC

[1077] Qiaojun Shu, Yong Chen, Shuguang Han, Guohui Lin, Eiji Miyano, and An Zhang. Acyclic edge coloring conjecture is true on planar graphs without intersecting triangles. Theoretical Computer Science, 882(??): 77–108, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003595.

Nguyen:2021:MEP

[1078] Cam Ly Nguyen, Vorapong Suppakitpaisarn, Athasit Surarerks, and Phanu Vajanopath. On the maximum edgepair embedding bipartite matching. Theoretical Computer Science, 882(??): 109–124, August 23, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003698.

Banik:2021:VGP

[1079] Aritra Banik, Arun Kumar Das, Sandip Das, Anil Maheshwari, and Sarvottamananda. Voronoi game on polygons. Theoretical Computer Science, 882(??):125–142, August 23, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003686.

Anonymous:2021:PSa

[1080] Anonymous. Pages 1–122 (3 September 2021). Theoretical Computer Science, 883(??):1–122, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbh

[1081] Anonymous. Editorial Board. Theoretical Computer Science, 883(??): ii, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004837.

Subramani:2021:PCR

[1082] K. Subramani and P. Wojciechowski. On the parametrized complexity of read-once refutations in UTVPI+ constraint systems. Theoretical Computer Science, 883(??):1–18, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521002644.

Gupta:2021:BSM

[1083] Sushmita Gupta, Sanjukta Roy, Saket Saurabh, and Meirav Zehavi. Balanced stable marriage: How close is close enough? Theoretical Computer Science, 883(??):19– 43, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100267X.

Sun:2021:CCF

[1084] Xueli Sun, Jianxi Fan, Baolei Cheng, Zhao Liu, and Jia Yu. Component conditional fault tolerance of hierarchical folded cubic networks. Theoretical Computer Science, 883(??): 44–58, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003467.

Jiang:2021:CIM

[1085] Yong Jiang, Weifeng Lin, Longcheng Liu, and Anzhen Peng. Constrained inverse minimum flow problems under the weighted Hamming distance. Theoretical Computer Science, 883(??): 59–68, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003510.

Barbay:2021:CDD

[1086] Jérémy Barbay, Pablo Pérez-Lantero, and Javiel Rojas-Ledesma. Computing the depth distribution of a set of boxes. Theoretical Computer Science, 883(??): 69–82, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003509.

Akutsu:2021:NIA

[1087] Tatsuya Akutsu, Jesper Jansson, Ruiming Li, Atsuhiro Takasu, and Takeyuki Tamura. New and improved algorithms for unordered tree inclusion. Theoretical Computer Science, 883(??): 83–98, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003625.

Li:2021:CSA

[1088] Qinyi Li, Xavier Boyen, and Ernest Foo. CCA-security from adaptive allbut-one lossy trapdoor functions. Theoretical Computer Science, 883(??):99– 121, September 3, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003601.

Anonymous:2021:PSb

[1089] Anonymous. Pages 1–136 (7 September 2021). Theoretical Computer Science, 884(??):1–136, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbi

[1090] Anonymous. Editorial Board. Theoretical Computer Science, 884(??): ii, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004904.

Campbell:2021:CGG

[1091] Graham Campbell and Detlef Plump. Confluence up to garbage in graph transformation. Theoretical Computer Science, 884(??):1–22, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397521003546.

Stunkel:2021:SPR

[1092] Patrick Stünkel and Harald König. Single pushout rewriting in comprehensive systems of graph-like structures. Theoretical Computer Science, 884(??): 23–43, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004059.

Lambers:2021:TRN

[1093] Leen Lambers and Fernando Orejas.

Transformation rules with nested application conditions: Critical pairs, initial conflicts & minimality. Theoretical Computer Science, 884(??): 44–67, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004278.

Behr:2021:RTL

[1094] Nicolas Behr, Jean Krivine, Jakob L.

Andersen, and Daniel Merkle. Rewriting theory for the life sciences: a unifying theory of CTMC semantics. Theoretical Computer Science, 884(??):68– 115, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100428X.

Marin:2021:DPS

[1095] Andrea Marin, Carla Piazza, and Sabina Rossi. D_{PSNI} : Delimited persistent stochastic non-interference. Theoretical Computer Science, 884(??):116– 135, September 7, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004643.

Anonymous:2021:PSc

[1096] Anonymous. Pages 1–158 (11 September 2021). Theoretical Computer Science, 885(??):1–158, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbj

[1097] Anonymous. Editorial Board. Theoretical Computer Science, 885(??): ii, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004989.

BenBasat:2021:AQP

[1098] Ran Ben Basat, Seungbum Jo, Srinivasa Rao Satti, and Shubham Ugare. Approximate query processing over static sets and sliding windows. Theoretical Computer Science, 885(??):1– 14, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003571.

Hollender:2021:CPE

[1099] Alexandros Hollender. The classes PPA-k: Existence from arguments modulo k. Theoretical Computer Science, 885(??):15–29, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521003613.

Sood:2021:CCD

[1100] Gaurav Sood and K. Murali Krishnan. On the computational complexity of Data Flow Analysis over finite bounded meet semilattices. Theoretical Computer Science, 885(??):30– 40, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003637.

Zhang:2021:FEC

[1101] Zheng Zhang and Fangguo Zhang.

Functional encryption for cubic polynomials and implementation. Theoretical Computer Science, 885(??):41– 54, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003649.

Auvolat:2021:BTC

[1102] Alex Auvolat, Davide Frey, Michel Raynal, and François Taïani. Byzantinetolerant causal broadcast. Theoretical Computer Science, 885(??):55– 68, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003662.

Koerkamp:2021:RFC

[1103] Ragnar Groot Koerkamp and Stanislav Živný. On rainbow-free colourings of uniform hypergraphs. Theoretical Computer Science, 885(??):69– 76, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003704.

Li:2021:DBM

[1104] Yang Li, Lloyd Allison, and Kevin B.

Korb. The difficulty of being moral. Theoretical Computer Science, 885(??): 77–90, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003650.

Xu:2021:CCS

[1105] Shengfeng Xu, Xiangxue Li, Haifeng Qian, and Kefei Chen. CPA/CCA2secure PKE with squared-exponential DFR from low-noise LPN. Theoretical Computer Science, 885(??):91– 103, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003674.

Panda:2021:DIM

[1106] B. S. Panda and Juhi Chaudhary. Dominating induced matching in some subclasses of bipartite graphs. Theoretical Computer Science, 885(??):104– 115, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003881.

Friman:2021:SMG

[1107] Oren Friman and Gabriel Nivasch. Some i-Mark games. Theoretical Computer Science, 885(??):116–124, September 11, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003893.

Susilo:2021:WTA

[1108] Willy Susilo and Joseph Tonien. A Wiener-type attack on an RSAlike cryptosystem constructed from cubic Pell equations. Theoretical Computer Science, 885(??):125– 130, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100390X.

Basunia:2021:ANF

[1109] Alamgir Rahaman Basunia, Satyabrata Das, Laxman Saha, and Kalishankar Tiwary. Antipodal number of full m-ary trees. Theoretical Computer Science, 885(??):131– 145, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003935.

Liu:2021:AAF

[1110] Qian Liu, Jianxin Liu, Min Li, and Yang Zhou. Approximation algorithms for fuzzy C-means problem based on seeding method. Theoretical Computer Science, 885(??):146– 158, September 11, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003911.

Anonymous:2021:PJn

[1111] Anonymous. Pages 1–112 (22 July 2021). Theoretical Computer Science, 878–879(??):1–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbk

[1112] Anonymous. Editorial Board. Theoretical Computer Science, 878–879 (??):ii, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003777.

Bi:2021:ICMb

[1113] Huimin Bi and Xin Zhang. Incidence coloring of Mycielskians with fast algorithm. Theoretical Computer Science, 878–879(??):1– 10, July 22, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003108.

Brandstadt:2021:MWIb

[1114] Andreas Brandstädt and Raffaele Mosca. Maximum weight independent sets for (S_{1,2,4}, triangle)-free graphs in polynomial time. Theoretical Computer Science, 878–879 (??):11–25, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100308X.

Duchene:2021:IPSb

[1115] Eric Duchêne, Stéphane Gonzalez, Aline Parreau, Eric Rémila, and Philippe Solal. Influence: a partizan scoring game on graphs. Theoretical Computer Science, 878–879 (??):26–46, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003091.

Barcucci:2021:EGSb

[1116] Elena Barcucci, Antonio Bernini, and Renzo Pinzani. Exhaustive generation of some lattice paths and their prefixes. Theoretical Computer Science, 878–879 (??):47–52, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307180.

Kantor:2021:NSVb

[1117] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. Nonuniform SINR + Voronoi diagrams are effectively uniform. Theoretical Computer Science, 878–879(??): 53–66, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002383.

Zhang:2021:IAPb

[1118] Zhen Zhang, Yutian Guo, Junyu Huang, Jianxin Wang, and Feng Shi. Improved approximation for prizecollecting red-blue median. Theoretical Computer Science, 878–879 (??):67–82, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003133.

Goldreich:2021:ULVb

[1119] Oded Goldreich and Tom Gur. Universal locally verifiable codes and 3-round interactive proofs of proximity for CSP. Theoretical Computer Science, 878–879 (??):83–101, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003339.

Chen:2021:CDCb

[1120] Yu-Han Chen, Shyue-Ming Tang, Kung-Jui Pai, and Jou-Ming Chang. Constructing dual-CISTs with short diameters using a generic adjustment scheme on bicubes. Theoretical Computer Science, 878–879(??): 102–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003352.

Anonymous:2021:PJo

[1121] Anonymous. Pages 1–112 (22 July 2021). Theoretical Computer Science, 878–879(??):1–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbl

[1122] Anonymous. Editorial Board. Theoretical Computer Science, 878–879 (??):ii, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003777.

Bi:2021:ICMc

[1123] Huimin Bi and Xin Zhang. Incidence coloring of Mycielskians with fast algorithm. Theoretical Computer Science, 878–879(??):1– 10, July 22, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003108.

Brandstadt:2021:MWIc

[1124] Andreas Brandstädt and Raffaele Mosca. Maximum weight independent sets for (S_{1,2,4}, triangle)-free graphs in polynomial time. Theoretical Computer Science, 878–879 (??):11–25, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100308X.

Duchene:2021:IPSc

[1125] Eric Duchêne, Stéphane Gonzalez, Aline Parreau, Eric Rémila, and Philippe Solal. Influence: a partizan scoring game on graphs. Theoretical Computer Science, 878–879 (??):26–46, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003091.

Barcucci:2021:EGSc

[1126] Elena Barcucci, Antonio Bernini, and Renzo Pinzani. Exhaustive generation of some lattice paths and their prefixes. Theoretical Computer Science, 878–879 (??):47–52, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397520307180.

Kantor:2021:NSVc

[1127] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. Nonuniform SINR + Voronoi diagrams are effectively uniform. Theoretical Computer Science, 878–879(??): 53–66, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521002383.

Zhang:2021:IAPc

[1128] Zhen Zhang, Yutian Guo, Junyu Huang, Jianxin Wang, and Feng Shi.

Improved approximation for prizecollecting red-blue median. Theoretical Computer Science, 878–879 (??):67–82, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003133.

Goldreich:2021:ULVc

[1129] Oded Goldreich and Tom Gur. Universal locally verifiable codes and 3-round interactive proofs of proximity for CSP. Theoretical Computer Science, 878–879 (??):83–101, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003339.

Chen:2021:CDCc

[1130] Yu-Han Chen, Shyue-Ming Tang, Kung-Jui Pai, and Jou-Ming Chang. Constructing dual-CISTs with short diameters using a generic adjustment scheme on bicubes. Theoretical Computer Science, 878–879(??): 102–112, July 22, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003352.

Anonymous:2021:PSd

[1131] Anonymous. Pages 1–178 (13 September 2021). Theoretical Computer Science, 886(??):1–178, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbm

[1132] Anonymous. Editorial Board. Theoretical Computer Science, 886(??): ii, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005296.

Ayad:2021:ELS

[1133] Lorraine A. K. Ayad, Carl Barton, and Costas Iliopoulos. Editorial: the 27th London Stringology Days & London Algorithmic Workshop (LSD & LAW) 2019. Theoretical Computer Science, 886(??):1– 2, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005016.

Shi:2021:FPT

[1134] Feng Shi, Jie You, Zhen Zhang, Jingyi Liu, and Jianxin Wang. Fixed-parameter tractability for the Tree Assembly problem. Theoretical Computer Science, 886(??):3– 12, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003923.

Delporte-Gallet:2021:AP

[1135] Carole Delporte-Gallet, Hugues Fauconnier, Eli Gafni, and Giuliano Losa. The assignment problem. Theoretical Computer Science, 886(??):13– 26, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003972.

Chatterjee:2021:CLN

[1136] Bapi Chatterjee, Ivan Walulya, and Philippas Tsigas. Concurrent linearizable nearest neighbour search in LockFree-kD-tree. Theoretical Computer Science, 886(??):27–48, September 13, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003984.

Imbs:2021:SCD

[1137] Damien Imbs, Achour Mostéfaoui, Matthieu Perrin, and Michel Raynal. Set-constrained delivery broadcast: a communication abstraction for read/write implementable distributed objects. Theoretical Computer Science, 886(??):49–68, September 13, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004011.

Martinod:2021:CIR

[1138] T. Martinod, V. Pollet, B. Darties, R. Giroudeau, and J.-C. König. Complexity and inapproximability results for balanced connected subgraph problem. Theoretical Computer Science, 886(??):69–83, September 13, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004114.

Agrawal:2021:SSV

[1139] Garima Agrawal and Soumen Maity. The Small Set Vertex Expansion Problem. Theoretical Computer Science, 886(??):84–93, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521004217.

Chen:2021:SVL

[1140] Hua Chen, Lin Chen, Guochuan Zhang, and Vincent Chau. Scheduling with variable-length calibrations: Two agreeable variants. Theoretical Computer Science, 886(??):94– 105, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004254.

Sehrawat:2021:EST

[1141] Vipin Singh Sehrawat, Foo Yee Yeo, and Yvo Desmedt. Extremal set theory and LWE based access structure hiding verifiable secret sharing with maliciousmajority and free verification. Theoretical Computer Science, 886(??):106– 138, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004266.

Yang:2021:OVC

[1142] Boting Yang and Tanzina Akter. Onevisibility cops and robber on trees. Theoretical Computer Science, 886(??): 139–156, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004436.

Deng:2021:FMD

[1143] Yunyun Deng, Longkun Guo, Kewen Liao, and Yi Chen. On finding maximum disjoint paths with different colors: Computational complexity and practical LP-based algorithms. Theoretical Computer Science, 886(??):157– 168, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004680.

He:2021:DGP

[1144] Xiaozhou He, Wenfeng Lai, Binhai Zhu, and Peng Zou. Dispersing and grouping points on planar segments. Theoretical Computer Science, 886(??): 169–177, September 13, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004692.

Anonymous:2021:POa

[1145] Anonymous. Pages 1–164 (2 October 2021). Theoretical Computer Science, 887(??):1–164, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbn

[1146] Anonymous. Editorial Board. Theoretical Computer Science, 887(??): ii, October 2, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005405.

Eftekhari:2021:SOD

[1147] Mohsen Eftekhari, Paola Flocchini, Lata Narayanan, Jaroslav Opatrny, and Nicola Santoro. On synchronization and orientation in distributed barrier coverage with relocatable sensors. Theoretical Computer Science, 887(??): 1–10, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003959.

Yakami:2021:SEU

[1148] Takahiro Yakami, Yukiko Yamauchi, Shuji Kijima, and Masafumi Yamashita. Searching for an evader in an unknown dark cave by an optimal number of asynchronous searchers. Theoretical Computer Science, 887(??): 11–29, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003996.

Simard:2021:GCR

[1149] Frédéric Simard, Josée Desharnais, and François Laviolette. General Cops and Robbers games with randomness. Theoretical Computer Science, 887(??): 30–50, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100400X.

Chu:2021:IUB

[1150] Huairui Chu, Mingyu Xiao, and Zhe Zhang. An improved upper bound for SAT. Theoretical Computer Science, 887(??):51– 62, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004023.

Koiran:2021:DAR

[1151] Pascal Koiran and Mateusz Skomra. Derandomization and absolute reconstruction for sums of powers of linear forms. Theoretical Computer Science, 887(??):63–84, October 2, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004151.

Lin:2021:DDA

[1152] Yuhang Lin, Limei Lin, Yanze Huang, and Jiaru Wang. The t/sdiagnosability and t/s-diagnosis algorithm of folded hypercube under the PMC/MM* model. Theoretical Computer Science, 887(??): 85–98, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004138.

Ruangwises:2021:SCV

[1153] Suthee Ruangwises and Toshiya Itoh. Securely computing the n-variable equality function with 2n cards. Theoretical Computer Science, 887(??): 99–110, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004126.

Molla:2021:ODA

[1154] Anisur Rahaman Molla, Kaushik Mondal, and William K. Moses. Optimal dispersion on an anonymous ring in the presence of weak Byzantine robots. Theoretical Computer Science, 887(??): 111–121, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004096.

Lima:2021:ARV

[1155] Paloma T. Lima, Erik Jan van Leeuwen, and Marieke van der Wegen. Algorithms for the rainbow vertex coloring problem on graph classes. Theoretical Computer Science, 887(??): 122–142, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004102.

Sugahara:2021:ECR

[1156] Ryo Sugahara, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Efficiently computing runs on a trie. Theoretical Computer Science, 887(??):143– 151, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004084.

Su:2021:NCO

[1157] Sihong Su, Jingjing Li, and Bingxin Wang. A new construction of oddvariable rotation symmetric Boolean functions with optimal algebraic immunity and higher nonlinearity. Theoretical Computer Science, 887(??): 152–163, October 2, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004230.

Anonymous:2021:POb

[1158] Anonymous. Pages 1–132 (4 October 2021). Theoretical Computer Science, 888(??):1–132, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbo

[1159] Anonymous. Editorial Board. Theoretical Computer Science, 888(??): ii, October 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005673.

Blair:2021:EIA

[1160] Jean R. S. Blair and Christophe Crespelle. On the effectiveness of the incremental approach to minimal chordal edge modification. Theoretical Computer Science, 888(??): 1–12, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004175.

Lai:2021:GCT

[1161] Jianchang Lai, Ge Wu, Peng Jiang, Zhen Zhao, Willy Susilo, and Fuchun Guo. Generic construction for tightlysecure signatures from discrete log. Theoretical Computer Science, 888(??): 13–21, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004205.

Zhou:2021:AAM

[1162] Peiyan Zhou, Haitao Jiang, Daming Zhu, and Binhai Zhu. Approximation algorithms for the maximum vertex coverage problem on bounded degree graphs. Theoretical Computer Science, 888(??):22– 30, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004199.

Das:2021:CRS

[1163] Sandip Das, Harmender Gahlawat, Uma Kant Sahoo, and Sagnik Sen. Cops and Robber on some families of oriented graphs. Theoretical Computer Science, 888(??):31– 40, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004187.

Lafourcade:2021:HCP

[1164] Pascal Lafourcade, Daiki Miyahara, Takaaki Mizuki, Léo Robert, Tatsuya Sasaki, and Hideaki Sone. How to construct physical zeroknowledge proofs for puzzles with a “single loop” condition. Theoretical Computer Science, 888(??): 41–55, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004242.

Chen:2021:WFA

[1165] Zheng Chen, Zhaoquan Gu, and Yuexuan Wang. Which is the fairest allocation in the max-min fairnessbased coalitional game? Theoretical Computer Science, 888(??): 56–69, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004229.

Omelchenko:2021:STP

[1166] Oleksii Omelchenko and Andrei A. Bulatov. Satisfiability threshold for power law random 2-SAT in configuration model. Theoretical Computer Science, 888(??):70–94, October 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004321.

Liu:2021:RAC

[1167] Jiafei Liu, Shuming Zhou, Eddie Cheng, Qianru Zhou, and Xiaoqing Liu. Reliability analysis of the cactus-based networks. Theoretical Computer Science, 888(??): 95–107, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004345.

Yang:2021:HFT

[1168] Yuxing Yang and Jing Li. Hybrid fault-tolerant prescribed hyper-Hamiltonian laceability of hypercubes. Theoretical Computer Science, 888(??): 108–116, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004333.

Araujo:2021:PON

[1169] J. Araujo, A. Cezar, C. V. G. C.

Lima, V. F. dos Santos, and A. Silva. On the proper orientation number of chordal graphs. Theoretical Computer Science, 888(??):117– 132, October 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004369.

Anonymous:2021:POc

[1170] Anonymous. Pages 1–188 (8 October 2021). Theoretical Computer Science, 889(??):1–188, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbp

[1171] Anonymous. Editorial Board. Theoretical Computer Science, 889(??): ii, October 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005843.

Alves:2021:SCM

[1172] Mateus Rodrigues Alves, Mateus de Oliveira Oliveira, Janio Carlos Nascimento Silva, and Uéverton dos Santos Souza. Succinct certification of monotone circuits. Theoretical Computer Science, 889(??): 1–13, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004357.

Cai:2021:CCE

[1173] Jin-Yi Cai and Artem Govorov. The complexity of counting edge colorings for simple graphs. Theoretical Computer Science, 889(??):14– 24, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004382.

Kung:2021:SFT

[1174] Tzu-Liang Kung, Yuan-Hsiang Teng, and Cheng-Kuan Lin. Super faulttolerance assessment of locally twisted cubes based on the structure connectivity. Theoretical Computer Science, 889 (??):25–40, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004400.

Gartner:2021:MRC

[1175] Bernd Gärtner and Ahad N. Zehmakan. Majority rule cellular automata. Theoretical Computer Science, 889(??): 41–59, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004370.

Hoi:2021:IAG

[1176] Gordon Hoi and Frank Stephan. Improved algorithms for the general exact satisfiability problem. Theoretical Computer Science, 889(??): 60–84, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004394.

Tellache:2021:NCR

[1177] Nour ElHouda Tellache. New complexity results for shop scheduling problems with agreement graphs. Theoretical Computer Science, 889(??): 85–95, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004412.

Li:2021:CCG

[1178] He Li, Shumin Zhang, Chengfu Ye, and Shuming Zhou. The g-component connectivity of graphs. Theoretical Computer Science, 889(??):96– 104, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004424.

Soulignac:2021:CDA

[1179] Francisco J. Soulignac. A certifying and dynamic algorithm for the recognition of proper circular-arc graphs. Theoretical Computer Science, 889(??): 105–134, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004448.

Zheng:2021:CRV

[1180] Mengce Zheng, Noboru Kunihiro, and Yuanzhi Yao. Cryptanalysis of the RSA variant based on cubic Pell equation. Theoretical Computer Science, 889(??): 135–144, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100445X.

Samarin:2021:CMK

[1181] Somayeh Dolatnezhad Samarin, Dario Fiore, Daniele Venturi, and Morteza Amini. A compiler for multi-key homomorphic signatures for Turing machines. Theoretical Computer Science, 889(??):145–170, October 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004473.

Abu-Affash:2021:MTI

[1182] A. Karim Abu-Affash, Paz Carmi, and Matthew J. Katz. Minimizing total interference in asymmetric sensor networks. Theoretical Computer Science, 889(??):171–181, October 8, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004461.

Li:2021:SET

[1183] Xingfu Li, Guihai Yu, Sandi Klavžar, Jie Hu, and Bo Li. The Steiner k-eccentricity on trees. Theoretical Computer Science, 889(??):182– 188, October 8, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004667.

Anonymous:2021:POd

[1184] Anonymous. Pages 1–240 (12 October 2021). Theoretical Computer Science, 890(??):1–240, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbq

[1185] Anonymous. Editorial Board. Theoretical Computer Science, 890(??): ii, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005934.

Sun:2021:DAA

[1186] Xin Sun, Dachuan Xu, Longkun Guo, and Min Li. Deterministic approximation algorithm for submodular maximization subject to a matroid constraint. Theoretical Computer Science, 890(??):1–15, October 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004709.

Gurski:2021:ECO

[1187] Frank Gurski, Dominique Komander, and Marvin Lindemann. Efficient computation of the oriented chromatic number of recursively defined digraphs. Theoretical Computer Science, 890(??): 16–35, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004710.

Tredup:2021:PCS

[1188] Ronny Tredup and Evgeny Erofeev.

On the parameterized complexity of the synthesis of Boolean nets with restricted place environments. Theoretical Computer Science, 890(??): 36–69, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004746.

Wojciechowski:2021:CCH

[1189] Piotr Wojciechowski and K. Subramani. Copy complexity of Horn formulas with respect to unit readonce resolution. Theoretical Computer Science, 890(??):70–86, October 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004758.

Babu:2021:SBL

[1190] Jasine Babu, Veena Prabhakaran, and Arko Sharma. A substructure based lower bound for eternal vertex cover number. Theoretical Computer Science, 890(??):87– 104, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004771.

Arena:2021:CSS

[1191] Francesco Arena and Miriam Di Ianni. Complexity of Scorpion Solitaire and applications to Klondike. Theoretical Computer Science, 890(??):105– 124, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100476X.

Lanotte:2021:PCA

[1192] Ruggero Lanotte, Massimo Merro, and Andrei Munteanu. A process calculus approach to detection and mitigation of PLC malware. Theoretical Computer Science, 890(??):125– 146, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004795.

Hoffmann:2021:CSC

[1193] Stefan Hoffmann. Constrained synchronization and commutativity. Theoretical Computer Science, 890(??): 147–170, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005077.

Tukan:2021:CSV

[1194] Murad Tukan, Cenk Baykal, Dan Feldman, and Daniela Rus. On coresets for support vector machines. Theoretical Computer Science, 890(??): 171–191, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005247.

Wang:2021:ADU

[1195] Haitao Wang and Yiming Zhao. Algorithms for diameters of unicycle graphs and diameter-optimally augmenting trees. Theoretical Computer Science, 890(??):192–209, October 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005351.

Fu:2021:ASU

[1196] Bin Fu, Pengfei Gu, and Yuming Zhao. Approximate set union via approximate randomization. Theoretical Computer Science, 890(??):210– 239, October 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005454.

Anonymous:2021:PNa

[1197] Anonymous. Pages 1–230 (4 November 2021). Theoretical Computer Science, 891(??):1–230, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbr

[1198] Anonymous. Editorial Board. Theoretical Computer Science, 891(??): ii, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006113.

Yuan:2021:PKP

[1199] Hanchun Yuan, Yuping Ke, and Yixin Cao. Polynomial kernels for pawfree edge modification problems. Theoretical Computer Science, 891(??): 1–12, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004722.

Frid:2021:PPL

[1200] Anna E. Frid, Enzo Laborde, and Jarkko Peltomäki. On prefix palindromic length of automatic words. Theoretical Computer Science, 891(??):

13–23, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004734.

Zhao:2021:STB

[1201] Qiulan Zhao and Jinjiang Yuan.

Scheduling to tradeoff between the number and the length of accepted jobs. Theoretical Computer Science, 891(??):24–34, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521004783.

Zhu:2021:SPM

[1202] Qiang Zhu, Krishnaiyan Thulasiraman, Kshirasagar Naik, Sridhar Radhakrishnan, and Min Xu. Symmetric PMC model of diagnosis, bmatchings in graphs and fault identification in t-diagnosable systems. Theoretical Computer Science, 891(??): 35–49, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005028.

Larsson:2021:GG

[1203] Urban Larsson and Yakov Babichenko.

Golden games. Theoretical Computer Science, 891(??):50–58, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439752100503X.

Sureson:2021:SHF

[1204] Claude Sureson. Subcomputable Hausdorff function dimension. Theoretical Computer Science, 891(??):59– 83, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005041.

Honkala:2021:QUR

[1205] Juha Honkala. Quasi-universal k-regular sequences. Theoretical Computer Science, 891(??):84–89, November 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005065.

Bantva:2021:ORL

[1206] Devsi Bantva and Daphne Der-Fen Liu. Optimal radio labellings of block graphs and line graphs of trees. Theoretical Computer Science, 891(??): 90–104, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005053.

Pai:2021:AAS

[1207] Jayakumar Pai and Bhadrachalam Chitturi. Approximation algorithms for sorting permutations by extreme block-interchanges. Theoretical Computer Science, 891(??):105– 115, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005089.

Kaboli:2021:GCH

[1208] Reza Kaboli, Shahram Khazaei, and Maghsoud Parviz. On groupcharacterizability of homomorphic secret sharing schemes. Theoretical Computer Science, 891(??):116– 130, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005090.

Perera:2021:AFA

[1209] Maharage Nisansala Sevwandi Perera, Toru Nakamura, Masayuki Hashimoto, Hiroyuki Yokoyama, and Kouichi Sakurai. Almost fully anonymous attributebased group signatures with verifierlocal revocation and member registration from lattice assumptions. Theoretical Computer Science, 891(??): 131–148, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005107.

Anand:2021:DRM

[1210] Aditya Anand and Palash Dey. Distance restricted manipulation in voting. Theoretical Computer Science, 891(??):149–165, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521005119.

Alon:2021:PSO

[1211] Bar Alon and Anat Paskin-Cherniavsky.

On perfectly secure 2PC in the OThybrid model. Theoretical Computer Science, 891(??):166–188, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521005120.

Chen:2021:CCC

[1212] Lin Chen, Ahmed Imtiaz Sunny, Lei Xu, Shouhuai Xu, Zhimin Gao, Yang Lu, Weidong Shi, and Nolan Shah. Computational complexity characterization of protecting elections from bribery. Theoretical Computer Science, 891(??):189–209, November 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005132.

Libert:2021:AOT

[1213] Benoît Libert, San Ling, Fabrice Mouhartem, Khoa Nguyen, and Huaxiong Wang. Adaptive oblivious transfer with access control from lattice assumptions. Theoretical Computer Science, 891(??):210– 229, November 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005168.

Anonymous:2021:PNb

[1214] Anonymous. Pages 1–288 (12 November 2021). Theoretical Computer Science, 892(??):1–288, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbs

[1215] Anonymous. Editorial Board. Theoretical Computer Science, 892(??): ii, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006186.

Jonsson:2021:ETH

[1216] Peter Jonsson, Victor Lagerkvist, Johannes Schmidt, and Hannes Uppman. The exponential-time hypothesis and the relative complexity of optimization and logical reasoning problems. Theoretical Computer Science, 892(??): 1–24, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005211.

Jurgensen:2021:ASC

[1217] Helmut Jürgensen and Ludwig Staiger.

Automata for solid codes. Theoretical Computer Science, 892(??):25– 47, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005223.

Goldberg:2021:FET

[1218] Leslie Ann Goldberg, John Lapinskas, and David Richerby. Faster exponential-time algorithms for approximately counting independent sets. Theoretical Computer Science, 892(??): 48–84, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005235.

Li:2021:EPK

[1219] Qinyi Li and Xavier Boyen. Efficient public-key encryption with equality test from lattices. Theoretical Computer Science, 892(??):85– 107, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005259.

Santolaya:2021:ECS

[1220] Fernando Santolaya, Mónica Caniupán, Luis Gajardo, Miguel Romero, and Rodrigo Torres-Avilés. Efficient computation of spatial queries over points stored in k^{2}-tree compact data structures. Theoretical Computer Science, 892(??):108– 131, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005338.

Hong:2021:REP

[1221] Seok-Hee Hong and Hiroshi Nagamochi. Re-embedding a 1-plane graph for a straight-line drawing in linear time. Theoretical Computer Science, 892(??):132–154, November 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005363.

Cacic:2021:CSS

[1222] Vedran Čačić, Marko Horvat, and Zvonko Iljazović. Computable subcontinua of semicomputable chainable Hausdorff continua. Theoretical Computer Science, 892(??):155– 169, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005478.

Fujii:2021:PTA

[1223] Kaito Fujii, Tasuku Soma, and Yuichi Yoshida. Polynomial-time algorithms for submodular Laplacian systems. Theoretical Computer Science, 892(??): 170–186, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005442.

Yasumi:2021:UBP

[1224] Hiroto Yasumi, Fukuhito Ooshita, Michiko Inoue, and Sébastien Tixeuil. Uniform bipartition in the population protocol model with arbitrary graphs. Theoretical Computer Science, 892(??): 187–207, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100548X.

Hara:2021:TSR

[1225] Keisuke Hara and Keisuke Tanaka.

Tightly secure ring signatures in the standard model. Theoretical Computer Science, 892(??):208–237, November 12, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005508.

Bensmail:2021:RC

[1226] Julien Bensmail, Foivos Fioravantes, and Fionn Mc Inerney. On the role of 3s for the 1-2-3 Conjecture. Theoretical Computer Science, 892(??):238– 257, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005533.

Yang:2021:ESB

[1227] Yu Yang, Xiao xiao Li, Meng yuan Jin, Long Li, Hua Wang, and Xiao-Dong Zhang. Enumeration of subtrees and BC-subtrees with maximum degree no more than k in trees. Theoretical Computer Science, 892(??):258– 278, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005521.

Xie:2021:NUM

[1228] Sun Xie, Haixing Zhao, and Jun Yin. Nonexistence of uniformly most reliable two-terminal graphs. Theoretical Computer Science, 892(??):279– 288, November 12, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100551X.

Anonymous:2021:PNc

[1229] Anonymous. Pages 1–220 (21 November 2021). Theoretical Computer Science, 893(??):1–220, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbt

[1230] Anonymous. Editorial Board. Theoretical Computer Science, 893(??): ii, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006307.

Kari:2021:Eg

[1231] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 893(??):iii, November 21, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006332.

Petrov:2021:LLT

[1232] Tatjana Petrov, Claudia Igler, Ali Sezgin, Thomas A. Henzinger, and Calin C. Guet. Long lived transients in gene regulation. Theoretical Computer Science, 893(??):1– 16, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003121.

Innocentini:2021:PFM

[1233] Guilherme C. P. Innocentini, Arran Hodgkinson, Fernando Antoneli, Arnaud Debussche, and Ovidiu Radulescu. Push-forward method for piecewise deterministic biochemical simulations. Theoretical Computer Science, 893(??):17–40, November 21, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100311X.

Perez-Verona:2021:LSA

[1234] Isabel Cristina Perez-Verona, Mirco Tribastone, and Andrea Vandin. A large-scale assessment of exact lumping of quantitative models in the BioModels repository. Theoretical Computer Science, 893(??):41– 59, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003716.

Zhang:2021:CLA

[1235] Luping Zhang, Fei Xu, Ajeesh Ramanujan, and K. G. Subramanian. Control languages accepted by labeled spiking neural P systems with rules on synapses. Theoretical Computer Science, 893(??):60– 71, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003728.

Allart:2021:CDA

[1236] Emilie Allart, Joachim Niehren, and Cristian Versari. Computing difference abstractions of linear equation systems. Theoretical Computer Science, 893(??): 72–104, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100387X.

Su:2021:MAS

[1237] Lirun Su and Fusheng Yu. Matrix approach to spanning matroids of rough sets and its application to attribute reduction. Theoretical Computer Science, 893(??):105– 116, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003947.

Takahashi:2021:CSQ

[1238] Yasuhiro Takahashi, Yuki Takeuchi, and Seiichiro Tani. Classically simulating quantum circuits with local depolarizing noise. Theoretical Computer Science, 893(??):117– 132, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004291.

Bao:2021:CPS

[1239] Tingting Bao, Qian Yang, Hong Peng, Xiaohui Luo, Jun Wang, and Xiaoxiao Song. Computational power of sequential dendrite P systems. Theoretical Computer Science, 893(??):133– 145, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004679.

Gyongyosi:2021:PPO

[1240] Laszlo Gyongyosi. Post-processing optimization for continuous-variable quantum key distribution. Theoretical Computer Science, 893(??):146– 158, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004941.

Shi:2021:TCA

[1241] Feng Shi, Frank Neumann, and Jianxin Wang. Time complexity analysis of evolutionary algorithms for 2-hop (1,2)-minimum spanning tree problem. Theoretical Computer Science, 893(??): 159–175, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005193.

Luan:2021:FSM

[1242] Yu Luan, Mei Lu, and Yi Zhang. The fractional (strong) matching preclusion number of complete k-partite graph. Theoretical Computer Science, 893(??): 176–182, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100520X.

Wang:2021:GCT

[1243] Jing Wang. The generalized 3connectivity of two kinds of regular networks. Theoretical Computer Science, 893(??):183–190, November 21, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100534X.

Bentriou:2021:AAS

[1244] Mahmoud Bentriou, Paolo Ballarini, and Paul-Henry Cournède. Automaton-ABC: a statistical method to estimate the probability of spatiotemporal properties for parametric Markov population models. Theoretical Computer Science, 893(??):191– 219, November 21, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005739.

Anonymous:2021:EBbu

[1245] Anonymous. Editorial Board. Theoretical Computer Science, 894(??): ii, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100640X.

Kari:2021:Eh

[1246] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 894(??):iii, November 26, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006435.

Bonizzoni:2021:BBH

[1247] Paola Bonizzoni, Lila Kari, Ion Petre, and Grzegorz Rozenberg. Building bridges — Honoring Nataša Jonoska on the occasion of her 60th birthday. Theoretical Computer Science, 894(??): 1–2, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006368.

Angeleska:2021:CNP

[1248] Angela Angeleska, Sara Omranian, and Zoran Nikoloski. Coherent network partitions: Characterizations with cographs and prime graphs. Theoretical Computer Science, 894(??): 3–11, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005892.

Aubrun:2021:DPB

[1249] Nathalie Aubrun and Jarkko Kari. On the domino problem of the Baumslag–Solitar groups. Theoretical Computer Science, 894(??):12– 22, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100517X.

Bonnici:2021:SCG

[1250] V. Bonnici, G. Franco, and V. Manca. Spectral concepts in genome informational analysis. Theoretical Computer Science, 894(??):23– 30, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003960.

Calude:2021:BID

[1251] Cristian S. Calude, Karen Frilya Celine, Ziyuan Gao, Sanjay Jain, Ludwig Staiger, and Frank Stephan. Biimmunity over different size alphabets. Theoretical Computer Science, 894(??): 31–49, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005181.

Cannon:2021:EHS

[1252] Sarah Cannon, Erik D. Demaine, Martin L. Demaine, Sarah Eisenstat, David Furcy, Matthew J. Patitz, Robert Schweller, Scott M. Summers, and Andrew Winslow. On the effects of hierarchical self-assembly for reducing program-size complexity. Theoretical Computer Science, 894(??):50– 78, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005326.

DeFelice:2021:HGM

[1253] Clelia De Felice, Rocco Zaccagnino, and Rosalba Zizza. Hybrid and generalized marked systems. Theoretical Computer Science, 894(??):79– 90, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005144.

Doty:2021:SSC

[1254] David Doty and Mahsa Eftekhari.

A survey of size counting in population protocols. Theoretical Computer Science, 894(??):91–102, November 26, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005156.

Durand-Lose:2021:AGC

[1255] Jérôme Durand-Lose and Aurélien Emmanuel. Abstract geometrical computation 11: Slanted firing squad synchronisation on signal machines. Theoretical Computer Science, 894(??):103– 120, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003558.

Fazekas:2021:SNW

[1256] Szilárd Zsolt Fazekas and Shinnosuke Seki. Square network on a word. Theoretical Computer Science, 894(??):121– 134, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100462X.

Mahalingam:2021:BRF

[1257] Kalpana Mahalingam, Anuran Maity, Palak Pandoh, and Rama Raghavan. Block reversal on finite words. Theoretical Computer Science, 894(??):135– 151, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521004035.

Manca:2021:EDS

[1258] V. Manca and G. Scollo. Explaining DNA structure. Theoretical Computer Science, 894(??):152– 171, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100356X.

McColm:2021:CPG

[1259] Gregory L. McColm. Cut-and-project graphs and other complexes. Theoretical Computer Science, 894(??):172– 189, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005594.

Song:2021:MDO

[1260] Xin Song, Shalin Shah, and John Reif. Multidimensional data organization and random access in largescale DNA storage systems. Theoretical Computer Science, 894(??):190– 202, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005491.

Xu:2021:RRE

[1261] Penghao Xu and Francesca Storici.

RESCOT: Restriction enzyme set and combination optimization tools for rNMP capture techniques. Theoretical Computer Science, 894(??):203– 213, November 26, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004655.

Anonymous:2021:PDa

[1262] Anonymous. Pages 1–212 (4 December 2021). Theoretical Computer Science, 895(??):1–212, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbv

[1263] Anonymous. Editorial Board. Theoretical Computer Science, 895(??): ii, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006587.

Misra:2021:IPT

[1264] Neeldhara Misra and Harshil Mittal. Imbalance parameterized by twin cover revisited. Theoretical Computer Science, 895(??):1–15, December 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005466.

Clementi:2021:PLB

[1265] Andrea Clementi, Emanuele Natale, and Isabella Ziccardi. Parallel Load Balancing on constrained client-server topologies. Theoretical Computer Science, 895(??):16– 33, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005570.

DalLago:2021:DLR

[1266] Ugo Dal Lago and Francesco Gavazzo. Differential logical relations, part II increments and derivatives. Theoretical Computer Science, 895(??): 34–47, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005569.

Graca:2021:SHE

[1267] Daniel S. Graça and Ning Zhong. The set of hyperbolic equilibria and of invertible zeros on the unit ball is computable. Theoretical Computer Science, 895(??):48–54, December 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005600.

Bercea:2021:UTA

[1268] Ioana O. Bercea and Guy Even.

Upper tail analysis of bucket sort and random tries. Theoretical Computer Science, 895(??):55–67, December 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005582.

Li:2021:ECC

[1269] Bi Li, Jingfen Lan, Wantao Ning, Yongcui Tian, Xin Zhang, and Qiang Zhu. h-extra r-component connectivity of interconnection networks with application to hypercubes. Theoretical Computer Science, 895(??):68– 74, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005624.

Chaplick:2021:QMU

[1270] Steven Chaplick, Magnús M. Halldórsson, Murilo S. de Lima, and Tigran Tonoyan. Query minimization under stochastic uncertainty. Theoretical Computer Science, 895(??):75– 95, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005612.

Fuhlbruck:2021:WLA

[1271] Frank Fuhlbrück, Johannes Köbler, Ilia Ponomarenko, and Oleg Verbitsky. The Weisfeiler–Leman algorithm and recognition of graph properties. Theoretical Computer Science, 895(??):

96–114, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005636.

Ruangwises:2021:PZK

[1272] Suthee Ruangwises and Toshiya Itoh. Physical zero-knowledge proof for Ripple Effect. Theoretical Computer Science, 895(??):115–123, December 4, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005557.

Bandyapadhyay:2021:GPN

[1273] Sayan Bandyapadhyay, Aritra Banik, Sujoy Bhore, and Martin Nöllenburg. Geometric planar networks on bichromatic collinear points. Theoretical Computer Science, 895(??):124– 136, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005545.

Chlebikova:2021:CCC

[1274] Janka Chlebíková and Clément Dallard. Colourful components in kcaterpillars and planar graphs. Theoretical Computer Science, 895(??): 137–150, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521005752.

Auletta:2021:BIQ

[1275] Vincenzo Auletta, Diodato Ferraioli, Ashutosh Rai, Giannicola Scarpa, and Andreas Winter. Belief-invariant and quantum equilibria in games of incomplete information. Theoretical Computer Science, 895(??):151– 177, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005764.

Blazy:2021:HSS

[1276] Olivier Blazy, Laura Brouilhet, Celine Chevalier, Patrick Towa, Ida Tucker, and Damien Vergnaud. Hardware security without secure hardware: How to decrypt with a password and a server. Theoretical Computer Science, 895(??): 178–211, December 4, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005776.

Anonymous:2021:PDb

[1277] Anonymous. Pages 1–170 (6 December 2021). Theoretical Computer Science, 896(??):1–170, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2021:EBbw

[1278] Anonymous. Editorial Board. Theoretical Computer Science, 896(??): ii, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006642.

Nakamura:2021:NCE

[1279] Shuhei Nakamura, Yasuhiko Ikematsu, Yacheng Wang, Jintai Ding, and Tsuyoshi Takagi. New complexity estimation on the Rainbow-Band-Separation attack. Theoretical Computer Science, 896(??):1– 18, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005806.

DiGiacomo:2021:PRH

[1280] Emilio Di Giacomo, William J. Lenhart, Giuseppe Liotta, Timothy W. Randolph, and Alessandra Tappini. (k, p)-planarity: a relaxation of hybrid planarity. Theoretical Computer Science, 896(??):19– 30, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100579X.

DiGiacomo:2021:CPS

[1281] Emilio Di Giacomo, Jaroslav Hančl, and Giuseppe Liotta. 2-colored pointset embeddings of partial 2-trees. Theoretical Computer Science, 896(??): 31–45, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005788.

Fici:2021:ALC

[1282] Gabriele Fici, Nicola Prezza, and Rossano Venturini. Adaptive learning of compressible strings. Theoretical Computer Science, 896(??): 46–52, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006022.

Rukavicka:2021:UER

[1283] Josef Rukavicka. A unique extension of rich words. Theoretical Computer Science, 896(??):53– 64, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005971.

Kobayashi:2021:OAB

[1284] Koji M. Kobayashi. An optimal algorithm for 2-bounded delay buffer management with lookahead. Theoretical Computer Science, 896(??): 65–78, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005983.

Cori:2021:DSD

[1285] Robert Cori, Andrea Frosini, Giulia Palma, Elisa Pergola, and Simone Rinaldi. On doubly symmetric Dyck words. Theoretical Computer Science, 896(??):79–97, December 6, 2021. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005995.

Feng:2021:SRA

[1286] Kai Feng, Xinyu Ma, and Wei Wei.

Subnetwork reliability analysis of bubble-sort graph networks. Theoretical Computer Science, 896(??):98– 110, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006010.

Albers:2021:AEC

[1287] Susanne Albers and Jens Quedenfeld. Algorithms for energy conservation in heterogeneous data centers. Theoretical Computer Science, 896(??):111– 131, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006034.

Ganguly:2021:OCS

[1288] Arnab Ganguly, Daniel Gibney, Sharma V. Thankachan, and Rahul Shah. I/O-optimal categorical 3sided skyline queries. Theoretical Computer Science, 896(??):132– 144, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006046.

Zhuang:2021:CCC

[1289] Hongbin Zhuang, Wenzhong Guo, Xiao-Yan Li, Ximeng Liu, and Cheng-Kuan Lin. The component connectivity, component diagnosability, and t/k-diagnosability of Bicube networks. Theoretical Computer Science, 896(??): 145–157, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100606X.

Nair:2021:IUB

[1290] Pramod P. Nair, Rajan Sundaravaradhan, and Bhadrachalam Chitturi. Improved upper bound for sorting permutations by prefix transpositions. Theoretical Computer Science, 896(??): 158–167, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006058.

Ochem:2021:FFR

[1291] Pascal Ochem. A family of formulas with reversal of arbitrarily high avoidability index. Theoretical Computer Science, 896(??):168– 170, December 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006071.

Anonymous:2022:PJa

[1292] Anonymous. Pages 1–184 (2 January 2022). Theoretical Computer Science, 897(??):1–184, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBa

[1293] Anonymous. Editorial Board. Theoretical Computer Science, 897(??): ii, January 2, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006769.

Coecke:2022:KQM

[1294] Bob Coecke, Dominic Horsman, Aleks Kissinger, and Quanlong Wang. Kindergarden quantum mechanics graduates … or how I learned to stop gluing LEGO together and love the ZX-calculus. Theoretical Computer Science, 897(??):1–22, January 2, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004308.

Casal:2022:MS

[1295] Filipe Casal, Andreia Mordido, and Vasco T. Vasconcelos. Mixed sessions. Theoretical Computer Science, 897(??): 23–48, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004631.

Yang:2022:CNC

[1296] Qizhe Yang and Yuxi Fu. Counting nondeterministic computations. Theoretical Computer Science, 897(??): 49–63, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521004862.

Zhang:2022:ESC

[1297] Yan Zhang, Zhaohui Zhu, and Jinjin Zhang. Encoding safety in CLL_{R}. Theoretical Computer Science, 897(??): 64–82, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005715.

Gorrieri:2022:STB

[1298] Roberto Gorrieri. A study on team bisimulation and H-team bisimulation for BPP nets. Theoretical Computer Science, 897(??):83– 113, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005727.

Fu:2022:TI

[1299] Yuxi Fu. A thesis for interaction.

Theoretical Computer Science, 897(??): 114–134, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005880.

Selvaratnam:2022:SPT

[1300] Daniel Selvaratnam, Michael Cantoni, J. M. Davoren, and Iman Shames. Sampling polynomial trajectories for LTL verification. Theoretical Computer Science, 897(??): 135–163, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006459.

Ying:2022:PSD

[1301] Mingsheng Ying, Li Zhou, Yangjia Li, and Yuan Feng. A proof system for disjoint parallel quantum programs. Theoretical Computer Science, 897(??): 164–184, January 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006484.

Anonymous:2022:PJb

[1302] Anonymous. Pages 1–180 (4 January 2022). Theoretical Computer Science, 898(??):1–180, January 4, 2022.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBb

[1303] Anonymous. Editorial Board. Theoretical Computer Science, 898(??): ii, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006836.

Liu:2022:TCS

[1304] Xiangyu Liu, Shengli Liu, Shuai Han, and Dawu Gu. Tightly CCAsecure inner product functional encryption scheme. Theoretical Computer Science, 898(??):1–19, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006009.

Boyaci:2022:MCC

[1305] Arman Boyacı, Tınaz Ekim, and Mordechai Shalom. On the maximum cardinality cut problem in proper interval graphs and related graph classes. Theoretical Computer Science, 898(??): 20–29, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100623X.

Yuan:2022:CDI

[1306] Jun Yuan, Huijuan Qiao, and Aixia Liu. The R_{g}-conditional diagnosability of international networks. Theoretical Computer Science, 898(??): 30–43, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006241.

Chimani:2022:CNB

[1307] Markus Chimani, Philipp Kindermann, Fabrizio Montecchiani, and Pavel Valtr. Crossing numbers of beyond-planar graphs. Theoretical Computer Science, 898(??):44– 49, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006228.

Reischuk:2022:KP

[1308] Rüdiger Reischuk. The kangaroo problem. Theoretical Computer Science, 898(??):50–58, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006265.

Kern:2022:DPC

[1309] Walter Kern, Barnaby Martin, Daniël Paulusma, Siani Smith, and Erik Jan van Leeuwen. Disjoint paths and connected subgraphs for H-free graphs. Theoretical Computer Science, 898(??): 59–68, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006344.

Tale:2022:SLB

[1310] Prafullkumar Tale. Sparsification lower bound for linear spanners in directed graphs. Theoretical Computer Science, 898(??):69–74, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006472.

Lopes:2022:RDD

[1311] Raul Lopes and Ignasi Sau. A relaxation of the Directed Disjoint Paths problem: a global congestion metric helps. Theoretical Computer Science, 898(??):75–91, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006460.

Bridoux:2022:SPC

[1312] Florian Bridoux. Sequentialization and procedural complexity in automata networks. Theoretical Computer Science, 898(??):92–109, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006496.

Kostolanyi:2022:DUW

[1313] Peter Kostolányi. Determinisability of unary weighted automata over the rational numbers. Theoretical Computer Science, 898(??):110– 131, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006502.

Almethen:2022:ECP

[1314] Abdullah Almethen, Othon Michail, and Igor Potapov. On efficient connectivity-preserving transformations in a grid. Theoretical Computer Science, 898(??):132–148, January 4, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006514.

Hosseini:2022:PPC

[1315] S. N. Hosseini and L. Yeganeh. Partial pullback complement rewriting along admissible matches. Theoretical Computer Science, 898(??):149– 179, January 4, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100668X.

Anonymous:2022:PJc

[1316] Anonymous. Pages 1–98 (6 January 2022). Theoretical Computer Science, 899(??):1–98, January 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBc

[1317] Anonymous. Editorial Board. Theoretical Computer Science, 899(??): ii, January 6, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007131.

Conte:2022:ODC

[1318] Alessio Conte and Etsuji Tomita.

On the overall and delay complexity of the CLIQUES and Bron– Kerbosch algorithms. Theoretical Computer Science, 899(??):1–24, January 6, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006538.

Wolf:2022:DUC

[1319] Petra Wolf. From decidability to undecidability by considering regular sets of instances. Theoretical Computer Science, 899(??):25– 38, January 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100654X.

Kuszmaul:2022:TTG

[1320] William Kuszmaul. Train tracks with gaps: Applying the probabilistic method to trains. Theoretical Computer Science, 899(??): 39–47, January 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006927.

Montanari:2022:RSI

[1321] Angelo Montanari and Pietro Sala. Reactive synthesis from interval temporal logic specifications. Theoretical Computer Science, 899(??): 48–79, January 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006940.

Bandyapadhyay:2022:NOC

[1322] Sayan Bandyapadhyay, Tanmay Inamdar, Shreyas Pai, and Sriram V.

Pemmaraju. Near-optimal clustering in the k-machine model. Theoretical Computer Science, 899(??): 80–97, January 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007015.

Anonymous:2022:PJd

[1323] Anonymous. Pages 1–120 (8 January 2022). Theoretical Computer Science, 900(??):1–120, January 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBd

[1324] Anonymous. Editorial Board. Theoretical Computer Science, 900(??): ii, January 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007301.

Nies:2022:RIS

[1325] André Nies and Frank Stephan. Randomness and initial segment complexity for measures. Theoretical Computer Science, 900(??):1–19, January 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006526.

Moscarini:2022:CMB

[1326] Marina Moscarini. Computing a metric basis of a bipartite distancehereditary graph. Theoretical Computer Science, 900(??):20–24, January 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006903.

Oitavem:2022:PHF

[1327] Isabel Oitavem. The polynomial hierarchy of functions and its levels. Theoretical Computer Science, 900(??): 25–34, January 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006915.

Chakraborty:2022:DCB

[1328] Maumita Chakraborty, Ranjan Mehera, and Rajat Kumar Pal. Divide-and-conquer based all spanning tree generation algorithm of a simple connected graph. Theoretical Computer Science, 900(??):35–52, January 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006952.

Asahiro:2022:ULD

[1329] Yuichi Asahiro, Jesper Jansson, Eiji Miyano, and Hirotaka Ono. Upper and lower degree-constrained graph orientation with minimum penalty. Theoretical Computer Science, 900(??): 53–78, January 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006964.

Blasius:2022:CDD

[1330] Thomas Bläsius, Tobias Friedrich, and Martin Schirneck. The complexity of dependency detection and discovery in relational databases. Theoretical Computer Science, 900(??): 79–96, January 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006939.

Emura:2022:IBE

[1331] Keita Emura, Shuichi Katsumata, and Yohei Watanabe. Identitybased encryption with security against the KGC: a formal model and its instantiations. Theoretical Computer Science, 900(??):97–119, January 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100699X.

Anonymous:2022:PJe

[1332] Anonymous. Pages 1–114 (12 January 2022). Theoretical Computer Science, 901(??):1–114, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBe

[1333] Anonymous. Editorial Board. Theoretical Computer Science, 901(??): ii, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007428.

Shen:2022:IBA

[1334] Shiyu Shen, Hongbing Wang, and Yunlei Zhao. Identity-based authenticated encryption with identity confidentiality. Theoretical Computer Science, 901 (??):1–18, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006897.

Dong:2022:IDW

[1335] Lu Dong, Dariusz R. Kowalski, Harshita Kudaravalli, and Miguel A. Mosteiro. Information dissemination in wireless ad-hoc networks under the weighted-TIM framework. Theoretical Computer Science, 901(??): 19–34, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006988.

Castiglione:2022:IMS

[1336] Giuseppa Castiglione and Štěpán Holub. The intersection of 3maximal submonoids. Theoretical Computer Science, 901(??):35– 44, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006976.

Bosek:2022:TBS

[1337] Bartłomiej Bosek, Dariusz Leniowski, Piotr Sankowski, and Anna Zych-Pawlewicz. A tight bound for shortest augmenting paths on trees. Theoretical Computer Science, 901(??): 45–61, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007003.

Chen:2022:OPP

[1338] Mengjing Chen, Haoqiang Huang, Weiran Shen, Pingzhong Tang, Zihe Wang, and Jie Zhang. Optimal pricing policy design for selling cost-reducing innovation in Cournot games. Theoretical Computer Science, 901(??): 62–86, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007040.

Bozzelli:2022:CIT

[1339] Laura Bozzelli, Angelo Montanari, and Adriano Peron. Complexity issues for timeline-based planning over dense time under future and minimal semantics. Theoretical Computer Science, 901(??):87– 113, January 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007076.

Anonymous:2022:PJf

[1340] Anonymous. Pages 1–92 (18 January 2022). Theoretical Computer Science, 902(??):1–92, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBf

[1341] Anonymous. Editorial Board. Theoretical Computer Science, 902(??): ii, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000044.

Besa:2022:TKT

[1342] Juan Jose Besa, Timothy Johnson, Nil Mamano, Martha C. Osegueda, and Parker Williams. Taming the knight’s tour: Minimizing turns and crossings. Theoretical Computer Science, 902(??): 1–20, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007052.

Bose:2022:PCT

[1343] Prosenjit Bose, Saeed Mehrabi, and Debajyoti Mondal. Parameterized complexity of two-interval pattern problem. Theoretical Computer Science, 902(??):21–28, January 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100709X.

Matos-Rodriguez:2022:ICE

[1344] Yuniel Matos-Rodríguez, Evaristo José Madarro-Capó, Carlos Miguel Legón-Pérez, Omar Rojas, and Guillermo Sosa-Gómez. Improvements in the computing efficiency of the probabilities of the LIL test for the PRNG evaluation. Theoretical Computer Science, 902(??):29–40, January 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007088.

Pham:2022:WIB

[1345] Minh Thuy Truc Pham, Ngoc Ai Van Nguyen, Mei Jiang, Dung Hoang Duong, and Willy Susilo. Wildcarded identity-based encryption from lattices. Theoretical Computer Science, 902(??): 41–53, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007167.

Colombo:2022:NTM

[1346] João Guilherme Fritsche Colombo and Jerusa Marchi. A nondeterministic Turing machine variant to compute functions. Theoretical Computer Science, 902(??):54– 63, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007179.

Elbassioni:2022:QPA

[1347] Khaled Elbassioni. Quasi-polynomial algorithms for list-coloring of nearly intersecting hypergraphs. Theoretical Computer Science, 902(??):64– 75, January 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007234.

Brettell:2022:CST

[1348] Nick Brettell, Matthew Johnson, Giacomo Paesani, and Daniël Paulusma. Computing subset transversals in H-free graphs. Theoretical Computer Science, 902(??):76–92, January 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007222.

Anonymous:2022:PFa

[1349] Anonymous. Pages 1–144 (8 February 2022). Theoretical Computer Science, 903(??):1–144, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBg

[1350] Anonymous. Editorial Board. Theoretical Computer Science, 903(??): ii, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200024X.

Xu:2022:UGC

[1351] Xiong Xu, Shuling Wang, Bohua Zhan, Xiangyu Jin, Jean-Pierre Talpin, and Naijun Zhan. Unified graphical comodeling, analysis and verification of cyber-physical systems by combining AADL and Simulink/Stateflow. Theoretical Computer Science, 903(??): 1–25, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521006691.

Jancar:2022:BBP

[1352] Petr Jančar. Bisimilarity on Basic Parallel Processes. Theoretical Computer Science, 903(??):26– 38, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007027.

Jin:2022:ACP

[1353] Zhao Jin, Bowen Zhang, Lei Zhang, Yongzhi Cao, and Hanpin Wang.

An adaptation-complete proof system for local reasoning about cloud storage systems. Theoretical Computer Science, 903(??):39–73, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521007350.

Neumann:2022:CSA

[1354] Eike Neumann. On the computability of the set of automorphisms of the unit square. Theoretical Computer Science, 903(??): 74–83, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007349.

Gadducci:2022:CSI

[1355] Fabio Gadducci, Hernán Melgratti, Christian Roldán, and Matteo Sammartino. Categorical specification and implementation of Replicated Data Types. Theoretical Computer Science, 903(??):84–112, February 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007362.

Stankovic:2022:MBA

[1356] Miroslav Stankovič, Ezio Bartocci, and Laura Kovács. Moment-based analysis of Bayesian network properties. Theoretical Computer Science, 903(??): 113–133, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007374.

Elder:2022:RSP

[1357] Murray Elder and Adam Piggott. Rewriting systems, plain groups, and geodetic graphs. Theoretical Computer Science, 903(??):134– 144, February 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007386.

Anonymous:2022:PFb

[1358] Anonymous. Pages 1–80 (15 February 2022). Theoretical Computer Science, 904(??):1–80, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBh

[1359] Anonymous. Editorial Board. Theoretical Computer Science, 904(??): ii, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000585.

Laforest:2022:CID

[1360] Christian Laforest and Timothée Martinod. On the complexity of independent dominating set with obligations in graphs. Theoretical Computer Science, 904(??):1– 14, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521003030.

Kim:2022:CAC

[1361] Eun Jung Kim, Martin Milanič, Jérôme Monnot, and Christophe Picouleau.

Complexity and algorithms for constant diameter augmentation problems. Theoretical Computer Science, 904(??): 15–26, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521003042.

Gilbert:2022:BPC

[1362] Hugo Gilbert, Tom Portoleau, and Olivier Spanjaard. Beyond pairwise comparisons in social choice: a setwise Kemeny aggregation problem. Theoretical Computer Science, 904(??): 27–47, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100414X.

Casel:2022:CSE

[1363] Katrin Casel, Henning Fernau, Mehdi Khosravian Ghadikolaei, Jérôme Monnot, and Florian Sikora. On the complexity of solution extension of optimization problems. Theoretical Computer Science, 904(??):48– 65, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006253.

Ghadikolaei:2022:EPC

[1364] Mehdi Khosravian Ghadikolaei, Nikolaos Melissinos, Jérôme Monnot, and Aris Pagourtzis. Extension and its price for the connected vertex cover problem. Theoretical Computer Science, 904(??):66– 80, February 15, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007039.

Anonymous:2022:PFc

[1365] Anonymous. Pages 1–106 (22 February 2022). Theoretical Computer Science, 905(??):1–106, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBi

[1366] Anonymous. Editorial Board. Theoretical Computer Science, 905(??): ii, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000640.

Li:2022:RRM

[1367] Yalan Li, Jichang Wu, Shumin Zhang, and Chengfu Ye. Reliability of the round matching composition networks based on g-extra conditional fault. Theoretical Computer Science, 905(??): 1–9, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007064.

Zhang:2022:TIP

[1368] Nianpeng Zhang and Qiang Zhu. Two important parameters of HPMC model. Theoretical Computer Science, 905(??): 10–17, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007192.

Zhang:2022:FAE

[1369] Xin Zhang, Huaqiang Zhang, Bei Niu, and Bi Li. Fast algorithm of equitably partitioning degenerate graphs into graphs with lower degeneracy. Theoretical Computer Science, 905(??): 18–30, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007180.

Duong:2022:CCL

[1370] Dung Hoang Duong, Partha Sarathi Roy, Willy Susilo, Kazuhide Fukushima, Shinsaku Kiyomoto, and Arnaud Sipasseuth. Chosen-ciphertext latticebased public key encryption with equality test in standard model. Theoretical Computer Science, 905(??): 31–53, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007210.

Wang:2022:MRC

[1371] Jiabing Wang and Jiaye Chen. Maximizing the ratio of cluster split to cluster diameter without and with cardinality constraints. Theoretical Computer Science, 905(??):54– 68, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007246.

Arockiaraj:2022:EAS

[1372] Micheal Arockiaraj, Arul Jeya Shalini, and J. Nancy Delaila. Embedding algorithm of spined cube into grid structure and its wirelength computation.

Theoretical Computer Science, 905(??): 69–86, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100726X.

Luo:2022:SFP

[1373] Weidong Luo. On some FPT problems without polynomial Turing compressions. Theoretical Computer Science, 905(??):87–98, February 22, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397521007258.

Bille:2022:PSU

[1374] Philip Bille, Inge Li Gørtz, and Frederik Rye Skjoldjensen. Partial sums on the ultra-wide word RAM. Theoretical Computer Science, 905(??): 99–105, February 22, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000111.

Anonymous:2022:PMa

[1375] Anonymous. Pages 1–114 (5 March 2022). Theoretical Computer Science, 906(??):1–114, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBj

[1376] Anonymous. Editorial Board. Theoretical Computer Science, 906(??): ii, March 5, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000780.

Eppstein:2022:THG

[1377] David Eppstein, Daniel Frishberg, and William Maxwell. On the treewidth of Hanoi graphs. Theoretical Computer Science, 906(??): 1–17, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521007209.

Amiri:2022:DDC

[1378] Saeed Akhoondian Amiri and Ben Wiederhake. Distributed distance-r covering problems on sparse high-girth graphs. Theoretical Computer Science, 906(??):18–31, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000081.

Panda:2022:CAN

[1379] B. S. Panda and Priyamvada. Complexity and algorithms for neighborsum-2-distinguishing {1, 3}-edge-weighting of graphs. Theoretical Computer Science, 906(??):32–51, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439752200010X.

Dourado:2022:HNC

[1380] Mitre C. Dourado, Lucia D. Penso, and Dieter Rautenbach. The hull number in the convexity of induced paths of order

3. Theoretical Computer Science, 906 (??):52–63, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000093.

Bhattacharya:2022:WCP

[1381] Binay Bhattacharya, Sandip Das, and Subhadeep Ranjan Dev. The weighted k-center problem in trees for fixed k. Theoretical Computer Science, 906 (??):64–75, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000135.

Liang:2022:CLE

[1382] Zuosong Liang, Juan Wang, Junqing Cai, and Xinxin Yang. On the complexity of local-equitable coloring of graphs. Theoretical Computer Science, 906(??):76–82, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000147.

Li:2022:IKA

[1383] Wenjun Li, Huan Peng, and Yongjie Yang. Improved kernel and algorithm for claw and diamond free edge deletion based on refined observations. Theoretical Computer Science, 906 (??):83–93, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000160.

Bilo:2022:RAP

[1384] Vittorio Bilò. On the robustness of the approximate price of anarchy in generalized congestion games. Theoretical Computer Science, 906(??):

94–113, March 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000172.

Anonymous:2022:PMb

[1385] Anonymous. Pages 1–128 (12 March 2022). Theoretical Computer Science, 907(??):1–128, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBk

[1386] Anonymous. Editorial Board. Theoretical Computer Science, 907(??): ii, March 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001062.

Bampis:2022:SRS

[1387] Evripidis Bampis, Dimitris Christou, Bruno Escoffier, Alexander Kononov, and Kim Thang Nguyen. A simple rounding scheme for multistage optimization. Theoretical Computer Science, 907(??):1–10, March 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000159.

Farina:2022:ECR

[1388] Antonio Fariña, Travis Gagie, Szymon Grabowski, Giovanni Manzini, Gonzalo Navarro, and Alberto Ordóñez. Efficient and compact representations of some non-canonical prefix-free codes. Theoretical Computer Science, 907(??): 11–25, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000196.

Li:2022:REP

[1389] Luyi Li and Xueliang Li. Rainbow edge-pancyclicity of strongly edge-colored graphs. Theoretical Computer Science, 907(??):26–33, March 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000123.

Krenn:2022:DRS

[1390] Daniel Krenn and Jeffrey Shallit. Decidability and k-regular sequences. Theoretical Computer Science, 907(??): 34–44, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200041X.

Feldkord:2022:OFL

[1391] Björn Feldkord, Till Knollmann, and Friedhelm Meyer auf der Heide. Online facility location with mobile facilities. Theoretical Computer Science, 907 (??):45–61, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000378.

Abdallah:2022:FTH

[1392] Mohamad Abdallah and Eddie Cheng. Fault-tolerant Hamiltonian connectivity of 2-tree-generated networks. Theoretical Computer Science, 907(??): 62–81, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000342.

Alstrup:2022:CLS

[1393] Stephen Alstrup, Søren Dahlgaard, Arnold Filtser, Morten Stöckel, and Christian Wulff-Nilsen. Constructing light spanners deterministically in near-linear time. Theoretical Computer Science, 907(??):82–112, March 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000408.

Gaikwad:2022:PCS

[1394] Ajinkya Gaikwad, Soumen Maity, and Shuvam Kant Tripathi. Parameterized complexity of satisfactory partition problem. Theoretical Computer Science, 907(??):113– 127, March 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000391.

Anonymous:2022:PMc

[1395] Anonymous. Pages 1–140 (24 March 2022). Theoretical Computer Science, 908(??):1–140, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBl

[1396] Anonymous. Editorial Board. Theoretical Computer Science, 908(??): ii, March 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001153.

Kari:2022:Ea

[1397] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 908(??):1, March 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001190.

Combe:2022:GIC

[1398] Noémie Combe, Yuri I. Manin, and Matilde Marcolli. Geometry of information: Classical and quantum aspects. Theoretical Computer Science, 908(??):2–27, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006356.

Campagner:2022:URD

[1399] Andrea Campagner, Davide Ciucci, and Valentina Dorigatti. Uncertainty representation in dynamical systems using rough set theory. Theoretical Computer Science, 908(??): 28–42, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006708.

Leporati:2022:DTP

[1400] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, and Claudio Zandron. Depth-two P systems can simulate Turing machines with NP oracles. Theoretical Computer Science, 908(??): 43–55, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752100671X.

Orellana-Martin:2022:SES

[1401] David Orellana-Martín, Luis Valencia-Cabrera, and Mario J. Pérez-Jiménez. P systems with evolutional symport and membrane creation rules solving QSAT. Theoretical Computer Science, 908(??):56–63, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397521006873.

Liu:2022:SGT

[1402] Bo Liu and Mengya Shen. Some geometrical and topological properties of DNNs’ decision boundaries. Theoretical Computer Science, 908(??): 64–75, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006885.

Wang:2022:DMC

[1403] Shiying Wang. The diagnosability of Möbius cubes for the g-extra condition. Theoretical Computer Science, 908(??): 76–88, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000184.

Melgratti:2022:PNV

[1404] Hernán Melgratti, Claudio Antares Mezzina, and G. Michele Pinna. A Petri net view of covalent bonds. Theoretical Computer Science, 908(??):

89–119, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000202.

Shen:2022:IC

[1405] Alexander Shen. Individual codewords. Theoretical Computer Science, 908(??): 120–122, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000688.

Ahmed:2022:SEN

[1406] Aishah Ahmed, Maciej Koutny, and Marta Pietkiewicz-Koutny. Synthesising elementary net systems with localities. Theoretical Computer Science, 908 (??):123–140, March 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200069X.

Anonymous:2022:PMd

[1407] Anonymous. Pages 1-110 (28 March 2022). Theoretical Computer Science, 909(??):1–110, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBm

[1408] Anonymous. Editorial Board. Theoretical Computer Science, 909(??): ii, March 28, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001323.

Aziz:2022:SMU

[1409] Haris Aziz, Péter Biró, Tamás Fleiner, Serge Gaspers, Ronald de Haan, Nicholas Mattei, and Baharak Rastegari. Stable matching with uncertain pairwise preferences. Theoretical Computer Science, 909(??): 1–11, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000421.

Zhou:2022:TCR

[1410] Yangyang Zhou, Dongyang Zhao, Mingyuan Ma, and Jin Xu. Total coloring of recursive maximal planar graphs. Theoretical Computer Science, 909(??): 12–18, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000366.

Russo:2022:RMQ

[1411] Luís M. S. Russo. Range minimum queries in minimal space. Theoretical Computer Science, 909(??): 19–38, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200038X.

Zhou:2022:SSR

[1412] Qian M. Zhou, Alice Calvert, and Maxwell Young. Singletons for simpletons revisiting windowed backoff with Chernoff bounds. Theoretical Computer Science, 909(??): 39–53, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000329.

Bilo:2022:CBS

[1413] Davide Bilò, Luciano Gualà, Stefano Leucci, Guido Proietti, and Giacomo Scornavacca. Cutting bamboo down to size. Theoretical Computer Science, 909(??):54– 67, March 28, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000330.

Esperet:2022:LCG

[1414] Louis Esperet and Benjamin Lévêque. Local certification of graphs on surfaces. Theoretical Computer Science, 909(??):68–75, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397522000354.

Durand:2022:CRC

[1415] Anaïs Durand, Michel Raynal, and Gadi Taubenfeld. Contention-related crash failures: Definitions, agreement algorithms, and impossibility results. Theoretical Computer Science, 909(??): 76–86, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000433.

Lima:2022:PHV

[1416] Carlos V. G. C. Lima, Thiago Marcilon, Nicolas Martins, and Rudini Sampaio. PSPACE-hardness of variants of the graph coloring game. Theoretical Computer Science, 909(??): 87–96, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000494.

Carneiro:2022:KFV

[1417] Alan Diêgo Aurélio Carneiro, Fábio Protti, and Uéverton dos Santos Souza. On knot-free vertex deletion: Fine-grained parameterized complexity analysis of a deadlock resolution graph problem. Theoretical Computer Science, 909(??):97– 109, March 28, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000469.

Anonymous:2022:PAa

[1418] Anonymous. Pages 1–120 (2 April 2022). Theoretical Computer Science, 910(??):1–120, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBn

[1419] Anonymous. Editorial Board. Theoretical Computer Science, 910(??): ii, April 2, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001396.

Li:2022:VKP

[1420] Wenjun Li, Junjie Ye, and Yixin Cao.

A 5k-vertex kernel for P_{2}-packing. Theoretical Computer Science, 910 (??):1–13, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000470.

Hanaoka:2022:GBR

[1421] Erika Hanaoka and Taizo Sadahiro. Generalized bitstring rotation and its applications to toggle dynamical systems. Theoretical Computer Science, 910(??):14–33, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000482.

Nguyen:2022:RNC

[1422] Viet-Ha Nguyen and Kévin Perrot. Rikudo is NP-complete. Theoretical Computer Science, 910(??):34– 47, April 2, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000457.

Berczi:2022:ALM

[1423] Kristóf Bérczi, Tamás Király, Yutaro Yamaguchi, and Yu Yokoi. Approximation by lexicographically maximal solutions in matching and matroid intersection problems. Theoretical Computer Science, 910(??): 48–53, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000445.

Yang:2022:NVM

[1424] Jing Yang and Fang-Wei Fu. New (k, l, m)-verifiable multi-secret sharing schemes based on XTR public key system. Theoretical Computer Science, 910(??):54–67, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000500.

Mubarak:2022:MNT

[1425] Abbas Mubarak, Sajid Iqbal, Tariq Naeem, and Shafiq Hussain. 2 mm: a new technique for sorting data. Theoretical Computer Science, 910 (??):68–90, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000512.

Barak-Pelleg:2022:MRI

[1426] D. Barak-Pelleg, D. Berend, and J. C. Saunders. A model of random industrial SAT. Theoretical Computer Science, 910(??):91– 112, April 2, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000524.

You:2022:SFT

[1427] Lantao You, Yuejuan Han, Jianxi Fan, and Dong Ye. Structure fault tolerance of WK-recursive networks. Theoretical Computer Science, 910(??): 113–120, April 2, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000536.

Anonymous:2022:PAb

[1428] Anonymous. Pages 1–124 (8 April 2022). Theoretical Computer Science, 911(??):1–124, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBo

[1429] Anonymous. Editorial Board. Theoretical Computer Science, 911(??): ii, April 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001736.

Morales-Ponce:2022:OPH

[1430] Oscar Morales-Ponce. Optimal patrolling of high priority segments while visiting the unit interval with a set of mobile robots. Theoretical Computer Science, 911(??):1– 18, April 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000731.

Liu:2022:CEC

[1431] Dong Liu, Pingshan Li, and Bicheng Zhang. Component edge connectivity of hypercube-like networks. Theoretical Computer Science, 911(??): 19–25, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000743.

Niu:2022:UMM

[1432] Ruichao Niu and Min Xu. The unpaired many-to-many k-disjoint paths in bipartite hypercube-like networks. Theoretical Computer Science, 911 (??):26–40, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000846.

Sakai:2022:DSS

[1433] Yoshifumi Sakai. A data structure for substring-substring LCS length queries. Theoretical Computer Science, 911(??):41–54, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000858.

Xu:2022:ROL

[1434] Jun Xu, Santanu Sarkar, and Lei Hu.

Revisiting orthogonal lattice attacks on approximate common divisor problems. Theoretical Computer Science, 911(??):55–69, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200086X.

Poureidi:2022:AHR

[1435] Abolfazl Poureidi. Algorithm and hardness results in double Roman domination of graphs. Theoretical Computer Science, 911(??):70– 79, April 8, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000834.

Behague:2022:LCT

[1436] Natalie C. Behague, Anthony Bonato, Melissa A. Huggan, Trent G. Marbach, and Brittany Pittman. The localization capture time of a graph. Theoretical Computer Science, 911 (??):80–91, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000871.

Lecomte:2022:PAS

[1437] Victor Lecomte, Gergely Ódor, and Patrick Thiran. The power of adaptivity in source identification with time queries on the path. Theoretical Computer Science, 911(??): 92–123, April 8, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000901.

Anonymous:2022:PAc

[1438] Anonymous. Pages 1–132 (12 April 2022). Theoretical Computer Science, 912(??):1–132, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBp

[1439] Anonymous. Editorial Board. Theoretical Computer Science, 912(??): ii, April 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001918.

Genitrini:2022:QSF

[1440] Antoine Genitrini, Martin Pépin, and Frédéric Peschanski. A quantitative study of fork–join processes with nondeterministic choice: Application to the statistical exploration of the statespace. Theoretical Computer Science, 912(??):1–36, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000299.

Vaandrager:2022:MNT

[1441] Frits Vaandrager and Abhisek Midya.

A Myhill–Nerode theorem for register automata and symbolic trace languages. Theoretical Computer Science, 912(??):37–55, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000287.

Damiani:2022:LEC

[1442] Ferruccio Damiani, Michael Lienhardt, and Luca Paolini. On logical and extensional characterizations of attributed feature models. Theoretical Computer Science, 912(??): 56–80, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000305.

Safari:2022:FVP

[1443] Mohsen Safari and Marieke Huisman. Formal verification of parallel prefix sum and stream compaction algorithms in CUDA. Theoretical Computer Science, 912(??): 81–98, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001232.

Lu:2022:MSQ

[1444] Jing Lu. The monad on strong quasi-metric spaces. Theoretical Computer Science, 912(??):99–108, April 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001025.

Poncet:2022:TCP

[1445] Paul Poncet. Transporting continuity properties from a poset to its subposets. Theoretical Computer Science, 912(??):109–132, April 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397522001013.

Anonymous:2022:PAd

[1446] Anonymous. Pages 1–174 (20 April 2022). Theoretical Computer Science, 913(??):1–174, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBq

[1447] Anonymous. Editorial Board. Theoretical Computer Science, 913(??): ii, April 20, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002018.

Bezem:2022:LCU

[1448] Marc Bezem and Thierry Coquand. Loop-checking and the uniform word problem for join-semilattices with an inflationary endomorphism. Theoretical Computer Science, 913(??):1– 7, April 20, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000317.

Durier:2022:EFP

[1449] Adrien Durier, Daniel Hirschkoff, and Davide Sangiorgi. Eager functions as processes. Theoretical Computer Science, 913(??):8– 42, April 20, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000718.

Xu:2022:MCQ

[1450] Ming Xu, Jianling Fu, Jingyi Mei, and Yuxin Deng. Model checking QCTL plus on quantum Markov chains. Theoretical Computer Science, 913(??): 43–72, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000706.

Deng:2022:FSC

[1451] Yuxin Deng and Yuan Feng. Formal semantics of a classical-quantum language. Theoretical Computer Science, 913(??):73–93, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000986.

Baldan:2022:BLC

[1452] Paolo Baldan, Daniele Gorla, Tommaso Padoan, and Ivano Salvo. Behavioural logics for configuration structures. Theoretical Computer Science, 913(??): 94–112, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000998.

Xie:2022:GRL

[1453] Ruifeng Xie and Zhenjiang Hu. Generic recursive lens combinators and their calculation laws. Theoretical Computer Science, 913(??): 113–137, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001001.

Wang:2022:ISC

[1454] Longchun Wang, Xiangnan Zhou, and Qingguo Li. Information systems for continuous semi-lattices. Theoretical Computer Science, 913(??): 138–150, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000974.

Peters:2022:D

[1455] Kirstin Peters, Uwe Nestmann, and Anna Schmitt. On distributability. Theoretical Computer Science, 913(??): 151–173, April 20, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001773.

Anonymous:2022:PMe

[1456] Anonymous. Pages 1–114 (7 May 2022). Theoretical Computer Science, 914(??):1–114, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBr

[1457] Anonymous. Editorial Board. Theoretical Computer Science, 914(??): ii, May 7, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002213.

Chen:2022:OST

[1458] Zuzhi Chen and Jialin Zhang. Online scheduling of time-critical tasks to minimize the number of calibrations. Theoretical Computer Science, 914(??):1–13, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000548.

Dai:2022:VCH

[1459] Yingjun Dai, Angèle M. Foley, and Chính T. Hoàng. Vertex coloring (4K_{1}, hole-twin, 5-wheel)-free graphs. Theoretical Computer Science, 914 (??):14–22, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000937.

Haitner:2022:CFC

[1460] Iftach Haitner, Nikolaos Makriyannis, and Eran Omri. On the complexity of fair coin flipping. Theoretical Computer Science, 914(??): 23–38, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000949.

Bendele:2022:ATS

[1461] Oliver Bendele and Dieter Rautenbach. Additive tree O(ρ log n)- spanners from tree breadth ρ. Theoretical Computer Science, 914(??): 39–46, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000962.

Demange:2022:GTA

[1462] Marc Demange, Alessia Di Fonso, Gabriele Di Stefano, and Pierpaolo Vittorini. A graph theoretical approach to the firebreak locating problem. Theoretical Computer Science, 914(??):47–72, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000895.

Gottlieb:2022:FAO

[1463] Lee-Ad Gottlieb, Robert Krauthgamer, and Havana Rika. Faster algorithms for orienteering and k-TSP. Theoretical Computer Science, 914 (??):73–83, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000950.

Castagnos:2022:TPC

[1464] Guilhem Castagnos, Fabien Laguillaumie, and Ida Tucker. A tighter proof for CCA secure inner product functional encryption: Genericity meets efficiency. Theoretical Computer Science, 914(??):84–113, May 7, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000913.

Anonymous:2022:PMf

[1465] Anonymous. Pages 1–100 (14 May 2022). Theoretical Computer Science, 915(??):1–100, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBs

[1466] Anonymous. Editorial Board. Theoretical Computer Science, 915(??): ii, May 14, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002298.

Zdanowski:2022:ENN

[1467] Konrad Zdanowski. On efficiency of notations for natural numbers. Theoretical Computer Science, 915 (??):1–10, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000925.

Janczewski:2022:WSH

[1468] Robert Janczewski, Paweł Obszarski, and Krzysztof Turowski. Weighted 2sections and hypergraph reconstruction. Theoretical Computer Science, 915(??):11–25, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522000883.

Velasquez:2022:CSM

[1469] Alvaro Velasquez, K. Subramani, and Piotr Wojciechowski. On the complexity of and solutions to the minimum stopping and trapping set problems. Theoretical Computer Science, 915(??):26–44, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200127X.

Xu:2022:DHD

[1470] Liqiong Xu, Shanshan Yin, and Liyang Zhai. Diagnosability and hybrid diagnosability of some classes of graphs under the BPMC model. Theoretical Computer Science, 915(??): 45–54, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001244.

Antonopoulos:2022:CAE

[1471] Antonis Antonopoulos, Eleni Bakali, Aggeliki Chalki, Aris Pagourtzis, Petros Pantavos, and Stathis Zachos. Completeness, approximability and exponential time results for counting problems with easy decision version. Theoretical Computer Science, 915 (??):55–73, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001256.

Gupta:2022:RCF

[1472] Sushmita Gupta, Sanjukta Roy, Saket Saurabh, and Meirav Zehavi. Resolute control: Forbidding candidates from winning an election is hard. Theoretical Computer Science, 915 (??):74–89, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001281.

Yang:2022:UPL

[1473] Shaojun Yang and Xinyi Huang.

Universal product learning with errors: a new variant of LWE for lattice-based cryptography. Theoretical Computer Science, 915(??): 90–100, May 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001268.

Anonymous:2022:PMg

[1474] Anonymous. Pages 1–110 (19 May 2022). Theoretical Computer Science, 916(??):1–110, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBt

[1475] Anonymous. Editorial Board. Theoretical Computer Science, 916(??): ii, May 19, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002353.

Thorncharoensri:2022:PPF

[1476] Pairat Thorncharoensri, Willy Susilo, and Yang-Wai Chow. Privacypreserving file sharing on cloud storage with certificateless signcryption. Theoretical Computer Science, 916 (??):1–21, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001359.

Czygrinow:2022:DDD

[1477] Andrzej Czygrinow, Michał Hanćkowiak, and Marcin Witkowski. Distributed distance domination in graphs with no K_{2,t}-minor. Theoretical Computer Science, 916(??):22–30, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397522001463.

Wang:2022:NCP

[1478] Na Wang, Jixiang Meng, and Yingzhi Tian. Neighbor-connectivity of pancake networks and burnt pancake networks. Theoretical Computer Science, 916(??):31–39, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001438.

Das:2022:CFD

[1479] Bibhuti Das, Abhinav Chakraborty, Subhash Bhagat, and Krishnendu Mukhopadhyaya. k-Circle formation by disoriented asynchronous robots. Theoretical Computer Science, 916 (??):40–61, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001451.

Katoh:2022:PIA

[1480] Takashi Katoh and Keisuke Goto. In-place initializable arrays. Theoretical Computer Science, 916(??): 62–69, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200144X.

Rusu:2022:HPD

[1481] Irena Rusu. Hamiltonian problems in directed graphs with simple row patterns. Theoretical Computer Science, 916(??):70–85, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001487.

Baghery:2022:SKB

[1482] Karim Baghery, Alonso González, Zaira Pindado, and Carla Ràfols. Signatures of knowledge for Boolean circuits under standard assumptions. Theoretical Computer Science, 916 (??):86–110, May 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001475.

Anonymous:2022:PMh

[1483] Anonymous. Pages 1–122 (25 May 2022). Theoretical Computer Science, 917(??):1–122, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBu

[1484] Anonymous. Editorial Board. Theoretical Computer Science, 917(??): ii, May 25, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002663.

Rong:2022:DCA

[1485] Guozhen Rong, Yongjie Yang, Wenjun Li, and Jianxin Wang. A divideand-conquer approach for reconstruction of C_{≥5}-free graphs via betweenness queries. Theoretical Computer Science, 917(??):1–11, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001566.

Bai:2022:ZFA

[1486] Zonglei Bai, Yongzhi Cao, and Hanpin Wang. Zero-freeness and approximation of real Boolean Holant problems. Theoretical Computer Science, 917(??):12–30, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001554.

Esnay:2022:ASB

[1487] Solène J. Esnay and Etienne Moutot. Aperiodic SFTs on Baumslag–Solitar groups. Theoretical Computer Science, 917(??):31–50, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001529.

Xie:2022:AAS

[1488] Shengjun Xie, Haodi Feng, Haitao Jiang, and Daming Zhu. Approximation algorithms for sorting by bounded singleton moves. Theoretical Computer Science, 917(??): 51–65, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001542.

Bordewich:2022:COV

[1489] Magnus Bordewich, Charles Semple, and Kristina Wicke. On the complexity of optimising variants of phylogenetic diversity on phylogenetic networks. Theoretical Computer Science, 917(??):66–80, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001517.

Hasan:2022:PPS

[1490] Md. Manzurul Hasan, Debajyoti Mondal, and Md. Saidur Rahman. Positive planar satisfiability problems under 3-connectivity constraints. Theoretical Computer Science, 917(??): 81–93, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001578.

Jean:2022:ECG

[1491] Devin C. Jean and Suk J. Seo.

Extremal cubic graphs for faulttolerant locating domination. Theoretical Computer Science, 917(??): 94–106, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001591.

Rao:2022:UAP

[1492] Guoyao Rao, Yongcai Wang, Wenping Chen, Deying Li, and Weili Wu. Union acceptable profit maximization in social networks. Theoretical Computer Science, 917(??): 107–121, May 25, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001530.

Anonymous:2022:PMi

[1493] Anonymous. Pages 1–122 (29 May 2022). Theoretical Computer Science, 918(??):1–122, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBv

[1494] Anonymous. Editorial Board. Theoretical Computer Science, 918(??): ii, May 29, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200295X.

Poureidi:2022:ARS

[1495] Abolfazl Poureidi. Algorithmic results in secure total dominating sets on graphs. Theoretical Computer Science, 918(??):1–17, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001505.

Wang:2022:EAM

[1496] Biing-Feng Wang. Efficient algorithms for the minmax regret path center problem with length constraint on trees. Theoretical Computer Science, 918(??):18–31, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200158X.

Jahannia:2022:CZL

[1497] Marieh Jahannia, Morteza Mohammad-Noori, Narad Rampersad, and Manon Stipulanti. Closed Ziv–Lempel factorization of the m-bonacci words.

Theoretical Computer Science, 918 (??):32–47, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001621.

Bar-Natan:2022:FTD

[1498] Aviv Bar-Natan, Panagiotis Charalampopoulos, Paweł Gawrychowski, Shay Mozes, and Oren Weimann. Faulttolerant distance labeling for planar graphs. Theoretical Computer Science, 918(??):48–59, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001633.

Gima:2022:EGB

[1499] Tatsuya Gima, Tesshu Hanaka, Masashi Kiyomi, Yasuaki Kobayashi, and Yota Otachi. Exploring the gap between treedepth and vertex cover through vertex integrity. Theoretical Computer Science, 918(??): 60–76, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200161X.

Bockenhauer:2022:CAPa

[1500] Hans-Joachim Böckenhauer, Dennis Komm, and Raphael Wegner. Call admission problems on grids with advice. Theoretical Computer Science, 918(??):77–93, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001682.

London:2022:GCG

[1501] András London, Ryan R. Martin, and András Pluhár. Graph clustering via generalized colorings. Theoretical Computer Science, 918(??): 94–104, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001645.

Geffert:2022:IDP

[1502] Viliam Geffert, Zuzana Bednárová, and Alexander Szabari. Inputdriven pushdown automata for edit distance neighborhood. Theoretical Computer Science, 918(??):105– 122, May 29, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001694.

Anonymous:2022:PJg

[1503] Anonymous. Pages 1–144 (5 June 2022). Theoretical Computer Science, 919(??):1–144, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBw

[1504] Anonymous. Editorial Board. Theoretical Computer Science, 919(??): ii, June 5, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003012.

Zamaraeva:2022:CRF

[1505] Elena Zamaraeva and Joviša Žunić. A characterization of 2-threshold functions via pairs of prime segments.

Theoretical Computer Science, 919 (??):1–17, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001670.

Klonowski:2022:GFG

[1506] Marek Klonowski, Dariusz R. Kowalski, and Dominik Paja̧k. Generalized framework for Group Testing: Queries, feedbacks and adversaries. Theoretical Computer Science, 919 (??):18–35, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001669.

Baiou:2022:CUS

[1507] Mourad Baïou, Rafael Colares, and Hervé Kerivin. The complexity of the unit stop number problem and its implications to other related problems. Theoretical Computer Science, 919(??):36–46, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001657.

Raynal:2022:VME

[1508] Michel Raynal and Gadi Taubenfeld.

A visit to mutual exclusion in seven dates. Theoretical Computer Science, 919(??):47–65, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001815.

Chaplick:2022:SAP

[1509] Steven Chaplick, Stefan Felsner, Philipp Kindermann, Jonathan Klawitter, Ignaz Rutter, and Alexander Wolff. Simple algorithms for partial and simultaneous rectangular duals with given contact orientations. Theoretical Computer Science, 919 (??):66–74, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001797.

Zak:2022:FBW

[1510] Radosław Żak. Finding binary words with a given number of subsequences. Theoretical Computer Science, 919 (??):75–79, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001803.

Du:2022:CRR

[1511] Jiao Du, Ziyu Chen, Shaojing Fu, Longjiang Qu, and Chao Li. Constructions of 2-resilient rotation symmetric Boolean functions through symbol transformations of cyclic Hadamard matrix. Theoretical Computer Science, 919(??):80–91, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001839.

Bose:2022:FCA

[1512] Prosenjit Bose, Pilar Cano, Rolf Fagerberg, John Iacono, Riko Jacob, and Stefan Langerman. Fragile complexity of adaptive algorithms. Theoretical Computer Science, 919(??):

92–102, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001840.

Galesi:2022:TBL

[1513] Nicola Galesi and Fariba Ranjbar.

Tight bounds to localize failure nodes on trees, grids and through embeddings under boolean network tomography. Theoretical Computer Science, 919 (??):103–117, June 5, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001876.

Droste:2022:FIP

[1514] Manfred Droste, Zoltán Fülöp, Dávid Kószó, and Heiko Vogler. Finiteimage property of weighted tree automata over past-finite monotonic strong bimonoids. Theoretical Computer Science, 919(??):118– 143, June 5, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001827.

Anonymous:2022:PJh

[1515] Anonymous. Pages 1–114 (12 June 2022). Theoretical Computer Science, 920(??):1–114, June 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBx

[1516] Anonymous. Editorial Board. Theoretical Computer Science, 920(??): ii, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003164.

Kari:2022:Eb

[1517] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 920(??):iii, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003218.

Bonifaci:2022:PIM

[1518] Vincenzo Bonifaci, Enrico Facca, Frederic Folz, Andreas Karrenbauer, Pavel Kolev, Kurt Mehlhorn, Giovanna Morigi, Golnoosh Shahkarami, and Quentin Vermande. Physaruminspired multi-commodity flow dynamics. Theoretical Computer Science, 920 (??):1–20, June 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522000822.

Gurevich:2022:QCC

[1519] Yuri Gurevich and Andreas Blass. Quantum circuits with classical channels and the principle of deferred measurements. Theoretical Computer Science, 920(??):21–32, June 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397522000810.

Tamm:2022:QAR

[1520] Hellis Tamm. Quotients and atoms of reversible languages. Theoretical Computer Science, 920(??):33– 45, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001104.

Gluck:2022:RPL

[1521] Robert Glück, Robin Kaarsgaard, and Tetsuo Yokoyama. From reversible programming languages to reversible metalanguages. Theoretical Computer Science, 920(??):46– 63, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001116.

Zhou:2022:CCS

[1522] Nan Zhou, Hong Peng, Jun Wang, Qian Yang, and Xiaohui Luo. Computational completeness of spiking neural P systems with inhibitory rules for generating string languages. Theoretical Computer Science, 920(??): 64–75, June 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001219.

Khadiev:2022:TWO

[1523] Kamil Khadiev, Aliya Khadieva, Mansur Ziatdinov, Ilnaz Mannapov, Dmitry Kravchenko, Alexander Rivosh, and Ramis Yamilov. Two-way and one-way quantum and classical automata with advice for online minimization problems. Theoretical Computer Science, 920(??): 76–94, June 12, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001220.

Alhazov:2022:VDM

[1524] Artiom Alhazov, Rudolf Freund, Sergiu Ivanov, and Marion Oswald. Variants of derivation modes for which purely catalytic P systems are computationally complete. Theoretical Computer Science, 920(??):95– 112, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001499.

Okubo:2022:CCP

[1525] Fumiya Okubo and Takashi Yokomori. Corrigendum to “On the computing powers of L-reductions of insertion languages” [Theor. Comput. Sci. 862 (2021) 224–235]. Theoretical Computer Science, 920(??): 113, June 12, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752100373X. See

[877].

Anonymous:2022:PJi

[1526] Anonymous. Pages 1–126 (19 June 2022). Theoretical Computer Science, 921(??):1–126, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBy

[1527] Anonymous. Editorial Board. Theoretical Computer Science, 921(??): ii, June 19, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003279.

Escoffier:2022:MJM

[1528] Bruno Escoffier, Laurent Gourvès, and Vangelis Paschos. In memory of Jérôme Monnot. Theoretical Computer Science, 921(??):1– 3, June 19, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200072X.

Li:2022:SII

[1529] Minming Li. Special Issue on the 14th International Frontiers of Algorithmics Workshop. Theoretical Computer Science, 921(??):4, June 19, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521006721.

Anonymous:2022:ESIa

[1530] Anonymous. Editorial: Special issue Algorithms and Computation. Theoretical Computer Science, 921 (??):5, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397521005740.

Cheng:2022:EDE

[1531] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the g-extra diagnosability of enhanced hypercubes. Theoretical Computer Science, 921(??): 6–19, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001864.

Cholvi:2022:SRS

[1532] Vicent Cholvi, P. Garncarek, T. Jurdziński, and Dariusz R. Kowalski. Stable routing scheduling algorithms in multi-hop wireless networks. Theoretical Computer Science, 921(??): 20–35, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001852.

Nguyen:2022:NVH

[1533] Minh Hieu Nguyen, Canh Ngoc Hoang, A. A. Moldovyan, N. A. Moldovyan, Huy Quang Vu, and Danh Kim Le Tran. A novel version of the hidden logarithm problem for postquantum signature algorithms. Theoretical Computer Science, 921(??): 36–49, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001967.

Ni:2022:MAB

[1534] Qiufen Ni, Jianxiong Guo, Hongmin W. Du, and Huan Wang. Multi-attribute based influence maximization in social networks: Algorithms and analysis. Theoretical Computer Science, 921(??):50– 62, June 19, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001979.

Tsur:2022:FAP

[1535] Dekel Tsur. Faster algorithm for pathwidth one vertex deletion. Theoretical Computer Science, 921(??): 63–74, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002067.

Gong:2022:EOC

[1536] Ye Gong and Qian-Ping Gu. An efficient oracle for counting shortest paths in planar graphs. Theoretical Computer Science, 921(??): 75–85, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002055.

Tang:2022:MSS

[1537] Zhongzheng Tang, Chenhao Wang, and Hau Chan. Monotone ksubmodular secretary problems: Cardinality and knapsack constraints. Theoretical Computer Science, 921(??): 86–99, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002079.

Chen:2022:SMT

[1538] Jianer Chen, Minjie Huang, and Yin Guo. Scheduling multiple two-stage flowshops with a deadline. Theoretical Computer Science, 921(??): 100–111, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002043.

Bisoyi:2022:PSV

[1539] Sagar Bisoyi, Krishnamoorthy Dinesh, and Jayalal Sarma. On pure space vs catalytic space. Theoretical Computer Science, 921(??): 112–126, June 19, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002080.

Anonymous:2022:PJj

[1540] Anonymous. Pages 1–502 (24 June 2022). Theoretical Computer Science, 922(??):1–502, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBz

[1541] Anonymous. Editorial Board. Theoretical Computer Science, 922(??): ii, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003425.

Fici:2022:PCT

[1542] Gabriele Fici and Jeffrey Shallit. Properties of a class of Toeplitz words. Theoretical Computer Science, 922 (??):1–12, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002109.

Yeh:2022:HAS

[1543] Hao-Ping Yeh, Wei Lu, Li-Hsuan Chen, Ling-Ju Hung, Ralf Klasing, and Sun-Yuan Hsieh. Hardness and approximation for the star p-Hub Routing Cost Problem in metric graphs.

Theoretical Computer Science, 922 (??):13–24, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002122.

Alshudukhi:2022:SDW

[1544] Jalawi Alshudukhi and Kusum Yadav. Survivability development of wireless sensor networks using neuro fuzzyclonal selection optimization. Theoretical Computer Science, 922(??): 25–36, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002110.

Barton:2022:ACC

[1545] Carl Barton. On the average-case complexity of pattern matching with wildcards. Theoretical Computer Science, 922(??):37–45, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002092.

Tian:2022:RMC

[1546] Zhaoxia Tian, Mingzu Zhang, and Xing Feng. Reliability measure of the n-th cartesian product of complete graph K_{4} on h-extra edge-connectivity. Theoretical Computer Science, 922 (??):46–60, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002134.

Gorain:2022:PGO

[1547] Barun Gorain, Kaushik Mondal, Himadri Nayak, and Supantha Pandit. Pebble guided optimal treasure hunt in anonymous graphs. Theoretical Computer Science, 922(??): 61–80, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002158.

Bar-Noy:2022:VWR

[1548] Amotz Bar-Noy, Toni Böhnlein, David Peleg, and Dror Rawitz. On vertex-weighted realizations of acyclic and general graphs. Theoretical Computer Science, 922(??): 81–95, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002171.

Avin:2022:HGF

[1549] Chen Avin, Avi Cohen, Zvi Lotker, and David Peleg. Hotelling games in fault-prone settings. Theoretical Computer Science, 922(??):96– 107, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200216X.

Ye:2022:SCH

[1550] Zekun Ye and Lvzhou Li. Sample complexity of hidden subgroup problem. Theoretical Computer Science, 922(??): 108–121, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002146.

Sheng:2022:IKP

[1551] Zimo Sheng and Mingyu Xiao. An improved kernel for planar vertexdisjoint triangle packing. Theoretical Computer Science, 922(??): 122–130, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002456.

Ahire:2022:SCM

[1552] Pallavi Ahire and Jibi Abraham. Secure cloud model for intellectual privacy protection of arithmetic expressions in source codes using data obfuscation techniques. Theoretical Computer Science, 922(??):131– 149, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002390.

Elkin:2022:DSD

[1553] Michael Elkin and Ofer Neiman.

Distributed strong diameter network decomposition. Theoretical Computer Science, 922(??):150–157, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002407.

Dvorakova:2022:MCE

[1554] L’ubomíra Dvořáková, Edita Pelantová, Daniela Opočenská, and Arseny M. Shur. On minimal critical exponent of balanced sequences. Theoretical Computer Science, 922(??): 158–169, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002432.

Berczi:2022:UKH

[1555] Kristóf Bérczi, Endre Boros, Ondřej Čepek, Petr Kučera, and Kazuhisa Makino. Unique key Horn functions. Theoretical Computer Science, 922(??): 170–178, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002419.

Yuan:2022:NIG

[1556] Jun Yuan, Ying Li, Aixia Liu, and Huijuan Qiao. The non-inclusive ggood-neighbor diagnosability of interconnection networks. Theoretical Computer Science, 922(??):179– 192, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002444.

Hong:2022:EEO

[1557] Yi Hong, Chuanwen Luo, Deying Li, Zhibo Chen, Xiyun Wang, and Xiao Li. Energy efficiency optimization for multiple chargers in Wireless Rechargeable Sensor Networks. Theoretical Computer Science, 922(??): 193–205, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002468.

Badkobeh:2022:ASF

[1558] Golnaz Badkobeh, Tero Harju, Pascal Ochem, and Matthieu Rosenfeld. Avoiding square-free words on free groups. Theoretical Computer Science, 922(??):206–217, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397522002493.

Adhikary:2022:MVG

[1559] Ranendu Adhikary, Kaustav Bose, Manash Kumar Kundu, and Buddhadeb Sau. Mutual visibility on grid by asynchronous luminous robots. Theoretical Computer Science, 922(??): 218–247, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002481.

Shen:2022:PAD

[1560] Zhao Shen. Prime automata do not exist. Theoretical Computer Science, 922 (??):248–256, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002511.

Wu:2022:CCA

[1561] Suying Wu, Jianxi Fan, Baolei Cheng, Jia Yu, and Yan Wang. Connectivity and constructive algorithms of disjoint paths in dragonfly networks. Theoretical Computer Science, 922(??): 257–270, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200247X.

Badkobeh:2022:ISA

[1562] Golnaz Badkobeh, Panagiotis Charalampopoulos, Dmitry Kosolobov, and Solon P. Pissis. Internal shortest absent word queries in constant time and linear space. Theoretical Computer Science, 922(??): 271–282, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002523.

Bouzid:2022:AWF

[1563] Zohir Bouzid, Pierre Sutra, and Corentin Travers. Agreeing within a few writes. Theoretical Computer Science, 922(??):283–299, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752200250X.

Mojdeh:2022:CNL

[1564] Doost Ali Mojdeh. On the conjectures of neighbor locating coloring of graphs. Theoretical Computer Science, 922(??): 300–307, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002559.

Liu:2022:SME

[1565] Xuemei Liu and Jixiang Meng. Strongly Menger-edge-connectedness of DCell networks. Theoretical Computer Science, 922(??):308–311, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002560.

Ingels:2022:EIF

[1566] Florian Ingels and Romain Azaïs. Enumeration of irredundant forests. Theoretical Computer Science, 922(??): 312–334, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002584.

Ba:2022:CSC

[1567] Lina Ba and Heping Zhang. The cyclestructure connectivity of crossed cubes. Theoretical Computer Science, 922(??): 335–345, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002535.

Chierichetti:2022:AAS

[1568] Flavio Chierichetti, Anirban Dasgupta, and Ravi Kumar. On additive approximate submodularity. Theoretical Computer Science, 922(??): 346–360, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002547.

Chen:2022:CSF

[1569] Hong Chen, Meirun Chen, Michel Habib, and Cheng-Kuan Lin. Connectivity for some families of composition networks. Theoretical Computer Science, 922(??):361–367, June 24, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002572.

Harland:2022:GCB

[1570] James Harland. Generating candidate busy beaver machines (or how to build the zany zoo). Theoretical Computer Science, 922(??): 368–394, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002596.

Zhang:2022:SSS

[1571] Xiujuan Zhang, Yongcai Wang, Deying Li, Wenping Chen, and Xingjian Ding. Self-stabilizing spanner topology control solutions in wireless ad hoc networks. Theoretical Computer Science, 922(??): 395–409, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002717.

Bockenhauer:2022:CAPb

[1572] Hans-Joachim Böckenhauer, Nina Corvelo Benz, and Dennis Komm. Call admission problems on trees. Theoretical Computer Science, 922(??): 410–423, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200278X.

Tan:2022:IEU

[1573] Xuehou Tan and Qi Wei. Improved exploration of unknown polygons. Theoretical Computer Science, 922(??): 424–437, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002778.

Tong:2022:PTA

[1574] Weitian Tong, Yao Xu, and Huili Zhang. A polynomial-time approximation scheme for parallel two-stage flowshops under makespan constraint. Theoretical Computer Science, 922(??): 438–446, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002699.

Crawford-Kahrl:2022:JRM

[1575] Peter Crawford-Kahrl, Bree Cummins, and Tomáš Gedeon. Joint realizability of monotone Boolean functions. Theoretical Computer Science, 922(??): 447–474, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002766.

Liang:2022:EAL

[1576] Ting-Wei Liang, Chang-Biau Yang, and Kuo-Si Huang. An efficient algorithm for the longest common palindromic subsequence problem. Theoretical Computer Science, 922(??): 475–485, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002705.

Konrad:2022:DMV

[1577] Christian Konrad and Viktor Zamaraev. Distributed minimum vertex coloring and maximum independent set in chordal graphs. Theoretical Computer Science, 922(??): 486–502, June 24, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002729.

Anonymous:2022:PJk

[1578] Anonymous. Pages 1–376 (26 June 2022). Theoretical Computer Science, 923(??):1–376, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBba

[1579] Anonymous. Editorial Board. Theoretical Computer Science, 923(??): ii, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003747.

Kurita:2022:CAT

[1580] Kazuhiro Kurita and Kunihiro Wasa. Constant amortized time enumeration of Eulerian trails. Theoretical Computer Science, 923(??):1– 12, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002754.

Akitaya:2022:RCG

[1581] Hugo A. Akitaya, Matias Korman, Oliver Korten, Diane L. Souvaine, and Csaba D. Tóth. Reconfiguration of connected graph partitions via recombination. Theoretical Computer Science, 923(??):13–26, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002730.

Bensmail:2022:VCG

[1582] Julien Bensmail and Fionn Mc Inerney. On a vertex-capturing game. Theoretical Computer Science, 923 (??):27–46, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002742.

El-Mhamdi:2022:RAD

[1583] El-Mahdi El-Mhamdi, Rachid Guerraoui, Lê Nguyen Hoang, and Alexandre Maurer. Removing algorithmic discrimination (with minimal individual error). Theoretical Computer Science, 923(??):47–55, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002791.

Habib:2022:GAS

[1584] Michel Habib, Fabien de Montgolfier, Lalla Mouatadid, and Mengchuan Zou. A general algorithmic scheme for combinatorial decompositions with application to modular decompositions of hypergraphs. Theoretical Computer Science, 923(??):56– 73, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002808.

Gudmundsson:2022:CSL

[1585] Joachim Gudmundsson, Mees van de Kerkhof, André van Renssen, Frank Staals, Lionov Wiratma, and Sampson Wong. Covering a set of line segments with a few squares. Theoretical Computer Science, 923(??): 74–98, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200281X.

Senda:2022:CRR

[1586] Ryoma Senda, Yoshiaki Takata, and Hiroyuki Seki. Complexity results on register context-free grammars and related formalisms. Theoretical Computer Science, 923(??):99– 125, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002833.

Goulet-Ouellet:2022:SCL

[1587] Herman Goulet-Ouellet. Suffixconnected languages. Theoretical Computer Science, 923(??):126–143, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002857.

Liang:2022:MCN

[1588] Jing Liang, Haixing Zhao, and Jun Yin. A method to calculate the number of spanning connected unicyclic(bicyclic) subgraphs in 2-separable networks. Theoretical Computer Science, 923(??): 144–159, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002894.

Ning:2022:CSC

[1589] Wantao Ning and Litao Guo. Connectivity and super connectivity of the exchanged 3-ary n-cube. Theoretical Computer Science, 923(??): 160–166, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002870.

Chen:2022:PVD

[1590] You Chen and Xueliang Li. The proper vertex-disconnection of graphs. Theoretical Computer Science, 923(??): 167–178, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002882.

Patt-Shamir:2022:PLS

[1591] Boaz Patt-Shamir and Mor Perry.

Proof-labeling schemes: Broadcast, unicast and in between. Theoretical Computer Science, 923(??): 179–195, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002900.

Bensmail:2022:GAN

[1592] Julien Bensmail, François Dross, Nacim Oijid, and Éric Sopena. Generalising the achromatic number to Zaslavsky’s colourings of signed graphs. Theoretical Computer Science, 923(??): 196–221, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003103.

Nassr:2022:ISP

[1593] Dieaa I. Nassr, M. Anwar, and Hatem M. Bahig. Improving small private exponent attack on the Murru– Saettone cryptosystem. Theoretical Computer Science, 923(??):222– 234, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003085.

Hosteins:2022:CCN

[1594] Pierre Hosteins, Rosario Scatamacchia, Andrea Grosso, and Roberto Aringhieri. The Connected Critical Node Problem. Theoretical Computer Science, 923(??):235–255, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003115.

Liu:2022:AAM

[1595] Xiaofei Liu, Weidong Li, and Han Dai. Approximation algorithms for the minimum power cover problem with submodular/linear penalties. Theoretical Computer Science, 923(??): 256–270, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003073.

Dublois:2022:UDS

[1596] Louis Dublois, Michael Lampis, and Vangelis Th. Paschos. Upper Dominating Set: Tight algorithms for pathwidth and subexponential approximation. Theoretical Computer Science, 923(??):

271–291, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003127.

Zhang:2022:BGM

[1597] Zhen Zhang, Yipeng Zhou, and Shaoqian Yu. Better guarantees for kmedian with service installation costs. Theoretical Computer Science, 923(??): 292–303, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003097.

Costa:2022:SGF

[1598] Eurinardo Rodrigues Costa, Nicolas Almeida Martins, and Rudini Sampaio. Spy game: FPT-algorithm, hardness and graph products. Theoretical Computer Science, 923(??): 304–317, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003061.

Nguyen:2022:LTA

[1599] Kien Trung Nguyen, Wen Chean Teh, Nguyen Thanh Hung, and Huong Nguyen-Thu. A linear time algorithm for connected p-centdian problem on block graphs. Theoretical Computer Science, 923(??):318– 326, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003310.

Tang:2022:OEH

[1600] Zhiyi Tang. Optimal embedding of hypercube into cylinder. Theoretical Computer Science, 923(??): 327–336, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003346.

Lin:2022:CDS

[1601] Min-Sheng Lin. Counting dominating sets in some subclasses of bipartite graphs. Theoretical Computer Science, 923(??):337–347, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200336X.

Alimi:2022:MIC

[1602] Morteza Alimi, Amir Daneshgar, and Mohammad-Hadi Foroughmand-Araabi. Mean isoperimetry with control on outliers: Exact and approximation algorithms. Theoretical Computer Science, 923(??):348– 365, June 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003358.

Shu:2022:NRS

[1603] Chai Shu, Xiang-Jun Li, and Meijie Ma. Note on reliability of star graphs. Theoretical Computer Science, 923(??): 366–375, June 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003504.

Anonymous:2022:PJl

[1604] Anonymous. Pages 1–186 (30 July 2022). Theoretical Computer Science, 924(??):1–186, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbb

[1605] Anonymous. Editorial Board. Theoretical Computer Science, 924(??): ii, July 30, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200367X.

Kari:2022:Ec

[1606] Lila Kari.  Editorial.  Theo-

retical Computer Science, 924(??): iii, July 30, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200370X.

Ipate:2022:MLB

[1607] Florentin Ipate and Marian Gheorghe. A model learning based testing approach for spiking neural P systems. Theoretical Computer Science, 924(??):1–16, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001608.

Teh:2022:ERS

[1608] Wen Chean Teh and Johnny Lim.

Evolvability of reaction systems and the invisibility theorem. Theoretical Computer Science, 924(??):17– 33, July 30, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522001955.

Salo:2022:CAB

[1609] Ville Salo, Guillaume Theyssier, and Ilkka Törmä. Cellular automata and bootstrap percolation. Theoretical Computer Science, 924(??): 34–45, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002249.

Takeuchi:2022:SBD

[1610] Yuki Takeuchi, Tomoyuki Morimae, and Seiichiro Tani. Sumcheckbased delegation of quantum computing to rational server. Theoretical Computer Science, 924(??): 46–67, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002250.

Skapinakis:2022:MPM

[1611] Eduardo Skapinakis and José Félix Costa. Machines that perform measurements. Theoretical Computer Science, 924(??):68–93, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397522002602.

Gao:2022:CMA

[1612] Ziyuan Gao, Sanjay Jain, Zeyong Li, Ammar Fathin Sabili, and Frank Stephan. A computation model with automatic functions and relations as primitive operations. Theoretical Computer Science, 924(??): 94–116, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002626.

Svozil:2022:VCB

[1613] Karl Svozil. Varieties of contextuality based on probability and structural nonembeddability. Theoretical Computer Science, 924(??):117– 128, July 30, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002614.

Roostapour:2022:SMO

[1614] Vahid Roostapour, Aneta Neumann, and Frank Neumann. Single- and multi-objective evolutionary algorithms for the knapsack problem with dynamically changing constraints. Theoretical Computer Science, 924(??): 129–147, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200305X.

Philippou:2022:CIS

[1615] Anna Philippou and Kyriaki Psara.

A collective interpretation semantics for reversing Petri nets. Theoretical Computer Science, 924(??): 148–170, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003231.

Wu:2022:QCQ

[1616] Zhenggang Wu, Daowen Qiu, Jiawei Tan, Hao Li, and Guangya Cai. Quantum and classical query complexities for generalized Simon’s problem. Theoretical Computer Science, 924(??): 171–186, July 30, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003474.

Anonymous:2022:PAe

[1617] Anonymous. Pages 1–86 (10 August 2022). Theoretical Computer Science, 925(??):1–86, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbc

[1618] Anonymous. Editorial Board. Theoretical Computer Science, 925(??): ii, August 10, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004029.

Ge-Ernst:2022:SDQ

[1619] Aile Ge-Ernst, Christoph Scholl, Juraj Síč, and Ralf Wimmer. Solving dependency quantified Boolean formulas using quantifier localization. Theoretical Computer Science, 925(??): 1–24, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522001785.

Sudha:2022:ADA

[1620] C. Sudha, D. Suresh, and A. Nagesh. Accurate data aggregation created by neural network and data classification processed through machine learning in wireless sensor networks. Theoretical Computer Science, 925(??): 25–36, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002420.

Barendregt:2022:PCA

[1621] Henk Barendregt and Sebastiaan A.

Terwijn. Partial combinatory algebra and generalized numberings. Theoretical Computer Science, 925(??): 37–44, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002821.

Rybakov:2022:CFV

[1622] Mikhail Rybakov and Dmitry Shkatov. Complexity of finite-variable fragments of propositional temporal and modal logics of computation. Theoretical Computer Science, 925(??): 45–60, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002845.

Revathy:2022:IVS

[1623] G. Revathy, K. Bhavana Raj, Anil Kumar, Spurthi Adibatti, Priyanka Dahiya, and T. M. Latha. Investigation of E-voting system using face recognition using convolutional neural network (CNN). Theoretical Computer Science, 925(??): 61–67, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522002869.

Kuhn:2022:MDM

[1624] Stefan Kuhn and Irek Ulidowski. Modelling of DNA mismatch repair with a reversible process calculus. Theoretical Computer Science, 925(??): 68–86, August 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003632.

Anonymous:2022:PAf

[1625] Anonymous. Pages 1–74 (18 August 2022). Theoretical Computer Science, 926(??):1–74, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbd

[1626] Anonymous. Editorial Board. Theoretical Computer Science, 926(??): ii, August 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004200.

Kari:2022:Ed

[1627] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 926(??):iii, August 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004236.

Alexandru:2022:TLC

[1628] Andrei Alexandru, Bogdan Aman, and Ross Horne. Theories of life and computation: Special issue on the occasion of the 65th birthday of Professor Gabriel Ciobanu. Theoretical Computer Science, 926(??): 1–2, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004157.

Bagossy:2022:CRC

[1629] Attila Bagossy and György Vaszil.

Controlled reversibility in communicating reaction systems. Theoretical Computer Science, 926(??): 3–20, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003528.

Balbi:2022:NMS

[1630] Pedro Paulo Balbi, Enrico Formenti, Kévin Perrot, Sara Riva, and Eurico L. P. Ruivo. Non-maximal sensitivity to synchronism in elementary cellular automata: Exact asymptotic measures. Theoretical Computer Science, 926(??): 21–50, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003383.

Zhang:2022:AHS

[1631] Luping Zhang and Fei Xu. Asynchronous homogeneous spiking neural P systems with local rule synchronization. Theoretical Computer Science, 926(??):51–61, August 18, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003371.

Yang:2022:DTP

[1632] Silu Yang, Dong Li, Hong Peng, Wenjie Zhou, Xiaohui Luo, Qian Yang, Jun Wang, and Xiaoxiao Song. Dynamic threshold P systems with delay on synapses for shortest path problems. Theoretical Computer Science, 926(??): 62–70, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003620.

Zhu:2022:WCD

[1633] Junlei Zhu, Yuehua Bu, and Hongguo Zhu. Wegner’s Conjecture on 2distance coloring for planar graphs. Theoretical Computer Science, 926(??): 71–74, August 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003814.

Anonymous:2022:PAg

[1634] Anonymous. Pages 1-162 (26 August 2022). Theoretical Computer Science, 927(??):1–162, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbe

[1635] Anonymous. Editorial Board. Theoretical Computer Science, 927(??): ii, August 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004662.

Surekha:2022:DMF

[1636] T. Lakshmi Surekha, N. Chandra Sekhara Rao, C. K. Shahnazeer, Syed Mufassir Yaseen, Surendra Kumar Shukla, Singh Bharat, and Mahendran Arumugam. Digital misinformation and fake news detection using WoT integration with Asian social networks fusion based feature extraction with text and image classification by machine learning architectures. Theoretical Computer Science, 927(??): 1–14, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003322.

Kumar:2022:HSC

[1637] A. Satish Kumar and S. Revathy.

A hybrid soft computing with big data analytics based protection and recovery strategy for security enhancement in large scale real world online social networks. Theoretical Computer Science, 927(??):15– 30, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003334.

Lemus:2022:CMD

[1638] Vladimir Lemus, Eduardo Acuña-Yeomans, Víctor Zamora, Francisco Hernández-Quiroz, and Hector Zenil. Computable model discovery and high-level-programming approximations to algorithmic complexity. Theoretical Computer Science, 927(??): 31–41, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003486.

Vagvolgyi:2022:DTT

[1639] Sándor Vágvölgyi. Deterministic topdown tree automata with Boolean deterministic look-ahead. Theoretical Computer Science, 927(??):42– 64, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003498.

Clokie:2022:CAS

[1640] Trevor Clokie, Thomas F. Lidbetter, Antonio Molina Lovett, Jeffrey Shallit, and Leon Witzman. Computational aspects of sturdy and flimsy numbers. Theoretical Computer Science, 927(??): 65–86, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003516.

Khan:2022:OHM

[1641] Mohd Abdul Rahim Khan, Shavkatov Navruzbek Shavkatovich, Bharti Nagpal, Anil Kumar, Mohd Anul Haq, V. Jeevika Tharini, Sathishkumar Karupusamy, and Malik Bader Alazzam. Optimizing hybrid metaheuristic algorithm with cluster head to improve performance metrics on the IoT. Theoretical Computer Science, 927(??): 87–97, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200353X.

Bell:2022:LCW

[1642] Jason P. Bell and Jeffrey Shallit.

Lie complexity of words. Theoretical Computer Science, 927(??):98– 108, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003541.

Akagi:2022:CMA

[1643] Tooru Akagi, Yuki Kuhara, Takuya Mieno, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Combinatorics of minimal absent words for a sliding window. Theoretical Computer Science, 927(??): 109–119, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003553.

Baril:2022:GCF

[1644] Jean-Luc Baril, Sergey Kirgizov, and Vincent Vajnovszki. Gray codes for Fibonacci q-decreasing words. Theoretical Computer Science, 927(??): 120–132, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003577.

Bille:2022:SIT

[1645] Philip Bille, Inge Li Gørtz, Max Rishøj Pedersen, Eva Rotenberg, and Teresa Anna Steiner. String indexing for top-k close consecutive occurrences. Theoretical Computer Science, 927(??): 133–147, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003565.

Zhang:2022:PSN

[1646] Shuangjun Zhang, Haibin Kan, and Liguan Wang. Preprocessing succinct non-interactive arguments for rank-1 constraint satisfiability from holographic proofs. Theoretical Computer Science, 927(??):148–161, August 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003590.

Anonymous:2022:PSa

[1647] Anonymous. Pages 1–210 (3 September 2022). Theoretical Computer Science, 928(??):1–210, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbf

[1648] Anonymous. Editorial Board. Theoretical Computer Science, 928(??): ii, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004728.

Ma:2022:OOA

[1649] Ran Ma, Sainan Guo, and Xiaoyan Zhang. An optimal online algorithm for single-processor scheduling problem with learning effect. Theoretical Computer Science, 928(??): 1–12, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003589.

Guo:2022:OFD

[1650] Xiangyu Guo, Kelin Luo, Zhihao Gavin Tang, and Yuhao Zhang. The online food delivery problem on stars. Theoretical Computer Science, 928(??): 13–26, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003607.

Yang:2022:OVC

[1651] Boting Yang. One-visibility cops and robber on trees: Optimal cop-win strategies. Theoretical Computer Science, 928(??):27–47, September 3, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003796.

Feng:2022:DWN

[1652] Wei Feng and Shiying Wang. The diagnosability of wheel networks with the condition: 2-extra. Theoretical Computer Science, 928(??):48– 60, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003784.

Li:2022:PPB

[1653] Xinghua Li, Suixiang Gao, Wenguo Yang, Yu Si, and Zhen Liu. Purchase preferences-based air passenger choice behavior analysis from sales transaction data. Theoretical Computer Science, 928(??):61– 70, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003802.

Dong:2022:PEB

[1654] Luobing Dong, Haobin Luo, Shan Zhang, Yanan Ren, Mingdong Duan, and Yifan Qin. Progressive-encoding-based transmission for DNN-enabled edge intelligence in unreliable network. Theoretical Computer Science, 928(??): 71–81, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003826.

Meng:2022:ESS

[1655] Xiangguang Meng, Jianxiong Guo, Xingjian Ding, and Xiujuan Zhang. Efficient scheduling strategy for mobile chargers in task-based wireless sensor networks. Theoretical Computer Science, 928(??):82– 93, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200384X.

Chen:2022:DIE

[1656] Sheng Chen and Yi Dai. Delta invariant for Eulerian digraphs. Theoretical Computer Science, 928(??):94– 103, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003851.

Zhang:2022:AIM

[1657] Yapu Zhang, Shengminjie Chen, Wenqing Xu, and Zhenning Zhang. Adaptive influence maximization under fixed observation time-step. Theoretical Computer Science, 928(??):104– 114, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003838.

Bridges:2022:MST

[1658] Douglas S. Bridges. Morse set theory as a foundation for constructive mathematics. Theoretical Computer Science, 928(??):115– 135, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003875.

Gaikwad:2022:DAG

[1659] Ajinkya Gaikwad and Soumen Maity.

Defensive alliances in graphs. Theoretical Computer Science, 928(??):136– 150, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003917.

Acan:2022:SNO

[1660] Hüseyin Acan, Sankardeep Chakraborty, Seungbum Jo, Kei Nakashima, Kunihiko Sadakane, and Srinivasa Rao Satti. Succinct navigational oracles for families of intersection graphs on a circle. Theoretical Computer Science, 928(??):151–166, September 3, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003930.

Bang-Jensen:2022:CSA

[1661] J. Bang-Jensen, S. Bessy, D. Gonçalves, and L. Picasarri-Arrieta. Complexity of some arc-partition problems for digraphs. Theoretical Computer Science, 928(??):167–182, September 3, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003899.

Bereg:2022:OPB

[1662] S. Bereg, J. M. Díaz-Báñez, M. Haghpanah, P. Horn, M. A. Lopez, N. Marín, A. Ramírez-Vigueras, F. Rodríguez, O. Solé-Pi, A. Stevens, and J. Urrutia. Optimal placement of base stations in border surveillance using limited capacity drones. Theoretical Computer Science, 928(??):183– 196, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522003929.

Wang:2022:TSS

[1663] Huizhen Wang and Hongwei Du. Time sensitive sweep coverage with minimum UAVs. Theoretical Computer Science, 928(??):197–209, September 3, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397522003905.

Anonymous:2022:PSb

[1664] Anonymous. Pages 1–206 (11 September 2022). Theoretical Computer Science, 929(??):1–206, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbg

[1665] Anonymous. Editorial Board. Theoretical Computer Science, 929(??): ii, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004996.

Liang:2022:AAP

[1666] Wei Liang and Zhao Zhang. Approximation algorithm for prize-collecting sweep cover with base stations. Theoretical Computer Science, 929(??):1– 10, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003942.

Cheng:2022:EME

[1667] Dongqin Cheng. Embedding mutually edge-disjoint cycles into locally twisted cubes. Theoretical Computer Science, 929(??):11– 17, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003966.

Dutta:2022:PIB

[1668] Priyanka Dutta, Willy Susilo, Dung Hoang Duong, and Partha Sarathi Roy. Puncturable identity-based and attributebased encryption from lattices. Theoretical Computer Science, 929(??):18– 38, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003954.

Liu:2022:FSB

[1669] Xujun Liu and Olgica Milenkovic. Finding the second-best candidate under the Mallows model. Theoretical Computer Science, 929(??):39– 68, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003978.

Bhore:2022:BCS

[1670] Sujoy Bhore, Satyabrata Jana, Supantha Pandit, and Sasanka Roy. The balanced connected subgraph problem for geometric intersection graphs. Theoretical Computer Science, 929(??):69– 80, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200398X.

Sekar:2022:IRM

[1671] K. R. Sekar, Naga Raju M, Mohd AnulHaq, Anil Kumar, R. Shalini, and S. Poojalaxmi. An improved ranking methodology for malignant carcinoma in multicriterian decision making using hesitant VIKOR fuzzy. Theoretical Computer Science, 929(??):81– 94, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004054.

Chandnani:2022:AAF

[1672] Neeraj Chandnani and Chandrakant N.

Khairnar. An analysis of architecture, framework, security and challenging aspects for data aggregation and routing techniques in IoT WSNs. Theoretical Computer Science, 929(??):95– 113, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004066.

Metilda:2022:PGU

[1673] M. I. Mary Metilda, D. Lalitha, and S. Vaithyasubramanian. Password generation using array generating interval timed colored Petri net (AG-ITCPN) for effective security. Theoretical Computer Science, 929(??):114– 123, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004078.

Roy:2022:LBP

[1674] Partha Sarathi Roy, Dung Hoang Duong, Willy Susilo, Arnaud Sipasseuth, Kazuhide Fukushima, and Shinsaku Kiyomoto. Lattice-based public-key encryption with equality test supporting flexible authorization in standard model. Theoretical Computer Science, 929(??):124–139, September 11, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004091.

Li:2022:EAR

[1675] Songhua Li, Minming Li, Lingjie Duan, and Victor C. S. Lee. Efficient algorithms for ride-hitching in UAV travelling. Theoretical Computer Science, 929(??):140–156, September 11, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200408X.

Gul:2022:OEO

[1676] Shai Gul, Eitan Tiktinsky, Slava Shamshanov, and Reuven Cohen. Online exploration outside a convex obstacle. Theoretical Computer Science, 929(??):157–163, September 11, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200411X.

Romashchenko:2022:CRI

[1677] Andrei Romashchenko. Clustering with respect to the information distance. Theoretical Computer Science, 929(??): 164–171, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004133.

Levin:2022:GKT

[1678] Leonid A. Levin. Gacs–Kucera theorem. Theoretical Computer Science, 929(??):172–173, September 11, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004145.

Liao:2022:FHM

[1679] Chao Liao, Jiabao Lin, Pinyan Lu, and Zhenyu Mao. An FPTAS for the hardcore model on random regular bipartite graphs. Theoretical Computer Science, 929(??):174– 190, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004169.

Das:2022:EDR

[1680] Archak Das, Kaustav Bose, and Buddhadeb Sau. Exploring a dynamic ring without landmark. Theoretical Computer Science, 929(??):191– 205, September 11, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004285.

Anonymous:2022:PSc

[1681] Anonymous. Pages 1–228 (21 September 2022). Theoretical Computer Science, 930(??):1–228, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbh

[1682] Anonymous. Editorial Board. Theoretical Computer Science, 930(??): ii, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005059.

Ikeda:2022:FCC

[1683] Masahiro Ikeda, Atsushi Miyauchi, Yuuki Takai, and Yuichi Yoshida.

Finding Cheeger cuts in hypergraphs via heat equation. Theoretical Computer Science, 930(??):1– 23, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004352.

Ban:2022:TES

[1684] Jung-Chao Ban, Chih-Hung Chang, Wen-Guei Hu, and Yu-Liang Wu. Topological entropy for shifts of finite type over Z and trees. Theoretical Computer Science, 930(??):24– 32, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200439X.

Kondor:2022:NHD

[1685] Gábor Kondor. NP-hardness of mdimensional weighted matching problems. Theoretical Computer Science, 930(??):33–36, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397522004315.

Brause:2022:PFG

[1686] Christoph Brause, Petr Golovach, Barnaby Martin, Daniël Paulusma, and Siani Smith. Partitioning H-free graphs of bounded diameter. Theoretical Computer Science, 930(??):37– 52, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004364.

Tripathi:2022:CPD

[1687] Vikash Tripathi, Ton Kloks, Arti Pandey, Kaustav Paul, and Hung-Lung Wang. Complexity of paired domination in AT-free and planar graphs. Theoretical Computer Science, 930(??): 53–62, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004297.

Gurram:2022:SEA

[1688] Girija Vani Gurram, Noorullah C.

Shariff, and Rajkumar L. Biradar. A Secure Energy Aware Meta-Heuristic Routing Protocol (SEAMHR) for sustainable IoT-Wireless Sensor Network (WSN). Theoretical Computer Science, 930(??):63–76, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004303.

Li:2022:SLT

[1689] Peng Li, Jianhui Shang, and Yi Shi. A simple linear time algorithm to solve the MIST problem on interval graphs. Theoretical Computer Science, 930(??): 77–85, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004340.

Fruchtman:2022:WFL

[1690] Aharon Fruchtman, Yoav Gross, Shmuel T. Klein, and Dana Shapira. Weighted forward looking adaptive coding. Theoretical Computer Science, 930(??):86–99, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004388.

Zhang:2022:WFA

[1691] Wenming Zhang, Huan Hu, Yongxi Cheng, Hongmei Li, and Haizhen Wang. The work function algorithm for the paging problem. Theoretical Computer Science, 930(??):100– 105, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004327.

Rukavicka:2022:PLR

[1692] Josef Rukavicka. Palindromic length and reduction of powers. Theoretical Computer Science, 930(??):106– 115, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004376.

Acharyya:2022:MCS

[1693] Ankush Acharyya, Ramesh K. Jallu, Vahideh Keikha, Maarten Löffler, and Maria Saumell. Minimum color spanning circle of imprecise points. Theoretical Computer Science, 930(??):116– 127, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004339.

Ducoffe:2022:EQB

[1694] Guillaume Ducoffe. Eccentricity queries and beyond using hub labels. Theoretical Computer Science, 930(??): 128–141, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004443.

Mira:2022:PTA

[1695] Frank Ángel Hernández Mira, Ernesto Parra Inza, José María Sigarreta Almira, and Nodari Vakhania. A polynomialtime approximation to a minimum dominating set in a graph. Theoretical Computer Science, 930(??):142– 156, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200442X.

Bhuiyan:2022:CLC

[1696] Mohammad Tawhidul Hasan Bhuiyan, Muhammad Rashed Alam, and M. Sohel Rahman. Computing the longest common almost-increasing subsequence. Theoretical Computer Science, 930(??):157–178, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004431.

Manurangsi:2022:AEF

[1697] Pasin Manurangsi and Warut Suksompong. Almost envy-freeness for groups: Improved bounds via discrepancy theory. Theoretical Computer Science, 930(??):179–195, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004546.

Ivan:2022:CGP

[1698] Maria-Romina Ivan, Imre Leader, and Mark Walters. Constructible graphs and pursuit. Theoretical Computer Science, 930(??):196–208, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004492.

Bampis:2022:OLM

[1699] Evripidis Bampis, Dimitris Christou, Bruno Escoffier, and Kim Thang Nguyen. Online learning for minmax discrete problems. Theoretical Computer Science, 930(??):209– 217, September 21, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004558.

Kortsarz:2022:AAE

[1700] Guy Kortsarz, Zeev Nutov, and Eli Shalom. Approximating activation edge-cover and facility location problems. Theoretical Computer Science, 930(??):218–228, September 21, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200456X.

Anonymous:2022:PSd

[1701] Anonymous. Pages 1–156 (29 September 2022). Theoretical Computer Science, 931(??):1–156, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbi

[1702] Anonymous. Editorial Board. Theoretical Computer Science, 931(??): ii, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005242.

Csaji:2022:CSH

[1703] Gergely Csáji. On the complexity of stable hypergraph matching, stable multicommodity flow and related problems. Theoretical Computer Science, 931(??):1– 16, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004534.

Melissinos:2022:ASS

[1704] Nikolaos Melissinos, Aris Pagourtzis, and Theofilos Triommatis. Approximation schemes for subsetsums ratio problems. Theoretical Computer Science, 931(??):17– 30, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004522.

Bilo:2022:AOA

[1705] Davide Bilò. Almost optimal algorithms for diameter-optimally augmenting trees. Theoretical Computer Science, 931(??):31–48, September 29, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004509.

Liu:2022:IAN

[1706] Zhicheng Liu, Jing Jin, Hong Chang, Donglei Du, and Xiaoyan Zhang. Improved algorithms for non-submodular function maximization problem. Theoretical Computer Science, 931(??):49– 55, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004510.

Kannari:2022:ABI

[1707] Phanindra Reddy Kannari, Noorullah Shariff Chowdary, and Rajkumar Laxmikanth Biradar. An anomalybased intrusion detection system using recursive feature elimination technique for improved attack detection. Theoretical Computer Science, 931(??): 56–64, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004479.

Li:2022:WTB

[1708] Tingyu Li, Jing Yu, and Haiping Zhang. Web of things based social media fake news classification with feature extraction using pretrained convoluted recurrent network with deep fuzzy learning. Theoretical Computer Science, 931(??):65– 77, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004467.

Liu:2022:GDU

[1709] Xiang Liu, Jinyu Zou, Pengpeng Chen, and Peng-Jun Wan. Gaussian downlink user selection subject to access limit, power budget, and rate demands. Theoretical Computer Science, 931(??): 78–92, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004455.

Weslin:2022:EWP

[1710] Devasirvatham Weslin and Thiyagarajan Joshva Devadas. Extricating web pages from deep web using deaima architecture. Theoretical Computer Science, 931(??):93– 103, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004480.

Martin:2022:CGC

[1711] Barnaby Martin, Daniël Paulusma, and Siani Smith. Colouring generalized claw-free graphs and graphs of large girth: Bounding the diameter. Theoretical Computer Science, 931(??):104– 116, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004571.

Le:2022:PMC

[1712] Van Bang Le and Jan Arne Telle. The perfect matching cut problem revisited. Theoretical Computer Science, 931(??): 117–130, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004595.

Alam:2022:MPN

[1713] Jawaherul Md. Alam, Michael A. Bekos, Martin Gronemann, Michael Kaufmann, and Sergey Pupyrev. The mixed page number of graphs. Theoretical Computer Science, 931(??):131– 141, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004583.

Narayanan:2022:EPT

[1714] S. Krishna Narayanan, S. Dhanasekaran, and V. Vasudevan. An effective parameter tuned deep belief network for detecting anomalous behavior in sensorbased cyber-physical systems. Theoretical Computer Science, 931(??):142– 151, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004613.

Calamoneri:2022:E

[1715] Tiziana Calamoneri. Editorial. Theoretical Computer Science, 931(??):152– 153, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522002912.

Anonymous:2022:ESIb

[1716] Anonymous. Editorial: Special issue on Algorithms and Computation. Theoretical Computer Science, 931(??): 154, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004108.

Gadducci:2022:SIT

[1717] Fabio Gadducci and Timo Kehrer.

Special issue on Theoretical Topics in Graph Transformation. Theoretical Computer Science, 931(??):155– 156, September 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004893.

Anonymous:2022:POa

[1718] Anonymous. Pages 1–116 (6 October 2022). Theoretical Computer Science, 932(??):1–116, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbj

[1719] Anonymous. Editorial Board. Theoretical Computer Science, 932(??): ii, October 6, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005370.

Gao:2022:STR

[1720] Xiangyu Gao, Jianzhong Li, and Dongjing Miao. Sublinear-time reductions for big data computing. Theoretical Computer Science, 932(??):

1–12, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004625.

Hong:2022:PAM

[1721] Weizhi Hong, Yingli Ran, and Zhao Zhang. Parallel algorithms for minimum general partial dominating set and maximum budgeted dominating set in unit disk graph. Theoretical Computer Science, 932(??): 13–20, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004601.

Mondal:2022:AHA

[1722] Debajyoti Mondal, Angelin Jemima Rajasingh, N. Parthiban, and Indra Rajasingh. APX-hardness and approximation for the k-burning number problem. Theoretical Computer Science, 932(??):21–30, October 6, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004819.

Miao:2022:NLB

[1723] Runjie Miao and Jinjiang Yuan. A note on LP-based approximation algorithms for capacitated facility location problem. Theoretical Computer Science, 932(??):31– 40, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004790.

Silva:2022:CBC

[1724] Janio Carlos Nascimento Silva and Uéverton S. Souza. Computing the best-case energy complexity of satisfying assignments in monotone circuits. Theoretical Computer Science, 932(??): 41–55, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004777.

Faria:2022:PPW

[1725] Luerbio Faria and Uéverton S. Souza.

On the probe problem for (r, ℓ)-well-coveredness: Algorithms and complexity. Theoretical Computer Science, 932 (??):56–68, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004789.

Chen:2022:MDR

[1726] Xujin Chen, Minming Li, Changjun Wang, Chenhao Wang, Mengqi Zhang, and Yingchao Zhao. Mechanisms for dual-role-facility location games: Truthfulness and approximability. Theoretical Computer Science, 932(??): 69–83, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004765.

Priyamvada:2022:ESC

[1727] Priyamvada and B. S. Panda. Exact square coloring of certain classes of graphs: Complexity and algorithms. Theoretical Computer Science, 932(??): 84–101, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004856.

Bilo:2022:NAA

[1728] Davide Bilò, Luciano Gualà, Stefano Leucci, Guido Proietti, and Mirko Rossi. New approximation algorithms for the heterogeneous weighted delivery problem. Theoretical Computer Science, 932(??): 102–115, October 6, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004844.

Anonymous:2022:POb

[1729] Anonymous. Pages 1–176 (14 October 2022). Theoretical Computer Science, 933(??):1–176, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbk

[1730] Anonymous. Editorial Board. Theoretical Computer Science, 933(??): ii, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200559X.

Reghizzi:2022:RLA

[1731] Stefano Crespi Reghizzi, Antonio Restivo, and Pierluigi San Pietro. Reducing the local alphabet size in tiling systems by means of 2D comma-free codes. Theoretical Computer Science, 933(??):1–20, October 14, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005424.

Nakashima:2022:PDE

[1732] Katsuhito Nakashima, Noriki Fujisato, Diptarama Hendrian, Yuto Nakashima, Ryo Yoshinaka, Shunsuke Inenaga, Hideo Bannai, Ayumi Shinohara, and Masayuki Takeda. Parameterized DAWGs: Efficient constructions and bidirectional pattern searches. Theoretical Computer Science, 933(??): 21–42, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005436.

Coronel:2022:RSS

[1733] Daniel Coronel, Alexander Frank, Mathieu Hoyrup, and Cristóbal Rojas. Realizing semicomputable simplices by computable dynamical systems. Theoretical Computer Science, 933(??): 43–54, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200528X.

Beal:2022:CWW

[1734] Marie-Pierre Béal and Maxime Crochemore.

Checking whether a word is Hammingisometric in linear time. Theoretical Computer Science, 933(??):55– 59, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005163.

Bao:2022:AAM

[1735] Xiaoguang Bao, Lei Xu, Wei Yu, and Wei Song. Approximation algorithms for the min-max clustered ktraveling salesmen problems. Theoretical Computer Science, 933(??): 60–66, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005126.

Sardar:2022:GES

[1736] Madhumita Sardar, Deepanwita Das, and Srabani Mukhopadhyaya. Grid exploration by a swarm of autonomous robots with minimum repetitions. Theoretical Computer Science, 933(??): 67–87, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005151.

Faro:2022:WAS

[1737] Simone Faro and Stefano Scafiti. A weak approach to suffix automata simulation for exact and approximate string matching. Theoretical Computer Science, 933(??):88– 103, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005138.

Cheng:2022:MOM

[1738] Siu-Wing Cheng. Multistage online maxmin allocation of indivisible entities. Theoretical Computer Science, 933(??):104–113, October 14, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005187.

Tian:2022:CDH

[1739] Yongcui Tian and Qiang Zhu. rcomponent diagnosability of hypercubes under the PMC model. Theoretical Computer Science, 933(??): 114–124, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005175.

Gaikwad:2022:FPA

[1740] Ajinkya Gaikwad and Soumen Maity.

Further parameterized algorithms for the F-free edge deletion problem. Theoretical Computer Science, 933(??): 125–137, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005205.

Sawada:2022:FSL

[1741] Joe Sawada, Aaron Williams, and Dennis Wong. Flip-swap languages in binary reflected Gray code order. Theoretical Computer Science, 933(??): 138–148, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200514X.

Bentert:2022:SCL

[1742] Matthias Bentert, Leon Kellerhals, and Rolf Niedermeier. The structural complexity landscape of finding balance-fair shortest paths. Theoretical Computer Science, 933(??):

149–162, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005102.

Chen:2022:NIB

[1743] Xinjian Chen, Qiong Huang, Hongbo Li, Zhijian Liao, and Willy Susilo. A novel identity-based multi-signature scheme over NTRU lattices. Theoretical Computer Science, 933(??): 163–176, October 14, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005114.

Anonymous:2022:POc

[1744] Anonymous. Pages 1–102 (23 October 2022). Theoretical Computer Science, 934(??):1–102, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbl

[1745] Anonymous. Editorial Board. Theoretical Computer Science, 934(??): ii, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005795.

Zhang:2022:TSW

[1746] Jinshan Zhang. Tight social welfare approximation of probabilistic serial. Theoretical Computer Science, 934 (??):1–6, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004807.

Hodges:2022:ACS

[1747] Reuven Hodges and Gidon Orelowitz.

Approximate counting of standard set-valued tableaux. Theoretical Computer Science, 934(??):7–20, October 23, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004832.

Asahiro:2022:MSS

[1748] Yuichi Asahiro, Ichiro Suzuki, and Masafumi Yamashita. Monotonic selfstabilization and its application to robust and adaptive pattern formation. Theoretical Computer Science, 934(??): 21–46, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200487X.

Das:2022:CSC

[1749] Avinandan Das, Lawqueen Kanesh, Jayakrishnan Madathil, Komal Muluk, Nidhi Purohit, and Saket Saurabh. On the complexity of singly connected vertex deletion. Theoretical Computer Science, 934(??):47– 64, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004820.

Aedo:2022:MAP

[1750] Ibai Aedo, Uwe Grimm, Yasushi Nagai, and Petra Staynova. Monochromatic arithmetic progressions in binary Thue–Morse-like words. Theoretical Computer Science, 934(??):

65–80, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004868.

Chen:2022:NSV

[1751] Meirun Chen, D. Frank Hsu, and Cheng-Kuan Lin. A new structure for a vertex to be locally t-diagnosable in large multiprocessor systems. Theoretical Computer Science, 934(??): 81–90, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522004959.

Choudhary:2022:SPT

[1752] Pratibha Choudhary and Venkatesh Raman. Structural parameterizations of Tracking Paths problem. Theoretical Computer Science, 934(??): 91–102, October 23, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005448.

Anonymous:2022:POd

[1753] Anonymous. Pages 1–200 (31 October 2022). Theoretical Computer Science, 935(??):1–200, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbm

[1754] Anonymous. Editorial Board. Theoretical Computer Science, 935(??): ii, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006041.

Li:2022:SRF

[1755] Yi Li, Xie Li, Yong Li, Xuechao Sun, Andrea Turrini, and Lijun Zhang. Synthesizing ranking functions for loop programs via SVM. Theoretical Computer Science, 935(??): 1–20, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200425X.

Hetzl:2022:URC

[1756] Stefan Hetzl and Jannik Vierling. Unprovability results for clause set cycles. Theoretical Computer Science, 935(??): 21–46, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004273.

Kahl:2022:SHD

[1757] Thomas Kahl. On symmetric higherdimensional automata and bisimilarity. Theoretical Computer Science, 935 (??):47–60, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004261.

Xu:2022:AMF

[1758] Ming Xu, Jianling Fu, Jingyi Mei, and Yuxin Deng. An algebraic method to fidelity-based model checking over quantum Markov chains. Theoretical Computer Science, 935(??):61– 81, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200490X.

Venhoek:2022:FCO

[1759] David Venhoek, Joshua Moerman, and Jurriaan Rot. Fast computations on ordered nominal sets. Theoretical Computer Science, 935(??): 82–104, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005291.

Barcelo:2022:ELP

[1760] Pablo Barceló, Nelson Higuera, Jorge Pérez, and Bernardo Subercaseaux. On the expressiveness of Lara: a proposal for unifying linear and relational algebra. Theoretical Computer Science, 935(??):105–127, October 31, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005308.

Petrakis:2022:CSB

[1761] Iosif Petrakis. Closed subsets in Bishop topological groups. Theoretical Computer Science, 935(??):128– 143, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005321.

Calautti:2022:QAI

[1762] Marco Calautti, Sergio Greco, Cristian Molinaro, and Irina Trubitsyna. Query answering over inconsistent knowledge bases: a probabilistic approach. Theoretical Computer Science, 935(??): 144–173, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200531X.

Baldan:2022:MES

[1763] Paolo Baldan and Alessandra Raffaetà.

Minimisation of event structures. Theoretical Computer Science, 935(??): 174–199, October 31, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005333.

Anonymous:2022:PNa

[1764] Anonymous. Pages 1–184 (10 November 2022). Theoretical Computer Science, 936(??):1–184, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBbn

[1765] Anonymous. Editorial Board. Theoretical Computer Science, 936(??): ii, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006144.

Sharma:2022:DDU

[1766] R. K. Sharma, P. R. Mishra, Yogesh Kumar, and Nupur Gupta. Differential δ-uniformity and non-linearity of permutations over Z_{n}. Theoretical Computer Science, 936(??):1– 12, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005199.

Morawietz:2022:CCG

[1767] Nils Morawietz, Niels Grüttemeier, Christian Komusiewicz, and Frank Sommer. Colored cut games. Theoretical Computer Science, 936(??): 13–32, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005539.

Lucke:2022:CMC

[1768] Felicia Lucke, Daniël Paulusma, and Bernard Ries. On the complexity of matching cut for graphs of bounded radius and H -free graphs. Theoretical Computer Science, 936(??):33– 42, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005497.

Jin:2022:RAB

[1769] Zhao Jin, Bowen Zhang, Tianyue Cao, Yongzhi Cao, and Hanpin Wang. Reasoning about block-based cloud storage systems via separation logic. Theoretical Computer Science, 936(??): 43–76, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005540.

Connor:2022:CCP

[1770] Matthew Connor, Othon Michail, and Igor Potapov. Centralised connectivity-preserving transformations for programmable matter: a minimal seed approach. Theoretical Computer Science, 936(??):77– 91, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005527.

Cai:2022:CCP

[1771] Justin Cai and Rafael Frongillo.

Computational complexity of problems for deterministic presentations of sofic shifts. Theoretical Computer Science, 936(??):92–115, November 10, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005515.

Scheffler:2022:RST

[1772] Robert Scheffler. On the recognition of search trees generated by BFS and DFS. Theoretical Computer Science, 936(??):116–128, November 10, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005503.

Angelini:2022:MLL

[1773] Patrizio Angelini, Michael A. Bekos, Philipp Kindermann, and Tamara Mchedlidze. On mixed linear layouts of series-parallel graphs. Theoretical Computer Science, 936(??):129– 138, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005485.

Tang:2022:BPS

[1774] Shaojie Tang. Beyond pointwise submodularity: Non-monotone adaptive submodular maximization subject to knapsack and k-system constraints. Theoretical Computer Science, 936(??): 139–147, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005643.

Gu:2022:SNV

[1775] Ran Gu, Gregory Gutin, Shasha Li, Yongtang Shi, and Zhenyu Taoqiu. Smallest number of vertices in a 2arc-strong digraph without good pairs. Theoretical Computer Science, 936(??): 148–171, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005667.

Lu:2022:ACG

[1776] Xiao-Tao Lü, Jin Chen, and Zhi-Xiong Wen. On the 2-abelian complexity of generalized Cantor sequences. Theoretical Computer Science, 936(??):172– 183, November 10, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005722.

Anonymous:2022:PNb

[1777] Anonymous. Pages 1–108 (18 November 2022). Theoretical Computer Science, 937(??):1–108, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBno

[1778] Anonymous. Editorial Board. Theoretical Computer Science, 937(??): ii, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006247.

Turkensteen:2022:ECT

[1779] Marcel Turkensteen and Gerold Jäger.

Efficient computation of tolerances in the sensitivity analysis of combinatorial bottleneck problems. Theoretical Computer Science, 937(??):1– 21, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005734.

Gascon:2022:MTP

[1780] Mathieu Gascon, Riccardo Dondi, and Nadia El-Mabrouk. MUL-tree pruning for consistency and optimal reconciliation — complexity and algorithms. Theoretical Computer Science, 937(??): 22–38, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005710.

Tan:2022:SAM

[1781] Jingjing Tan, Fengmin Wang, Weina Ye, Xiaoqing Zhang, and Yang Zhou. Streaming algorithms for monotone non-submodular function maximization under a knapsack constraint on the integer lattice. Theoretical Computer Science, 937(??):39– 49, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005709.

Li:2022:STS

[1782] Min Li, Hao Xiao, Qian Liu, and Yang Zhou. The submodularity of two-stage stochastic maximum-weight independent set problems. Theoretical Computer Science, 937(??):50– 62, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005692.

Tang:2022:SSP

[1783] Shaojie Tang. Stochastic submodular probing with state-dependent costs. Theoretical Computer Science, 937(??): 63–68, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005680.

Dong:2022:LSM

[1784] Rongcheng Dong, Yuichi Sudo, Taisuke Izumi, and Toshimitsu Masuzawa. Loosely-stabilizing maximal independent set algorithms with unreliable communications. Theoretical Computer Science, 937(??):69– 84, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005679.

Ning:2022:PSM

[1785] Xinya Ning, Nan Zhang, Zhenhua Duan, and Cong Tian. PPTL specification mining based on LNFG. Theoretical Computer Science, 937(??): 85–95, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005631.

Bensmail:2022:HDI

[1786] Julien Bensmail. On the hardness of determining the irregularity strength of graphs. Theoretical Computer Science, 937(??):96– 107, November 18, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005746.

Anonymous:2022:PNc

[1787] Anonymous. Pages 1–120 (26 November 2022). Theoretical Computer Science, 938(??):1–120, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2022:EBnp

[1788] Anonymous. Editorial Board. Theoretical Computer Science, 938(??): ii, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006302.

Bohm:2022:DCL

[1789] Martin Böhm and Bertrand Simon. Discovering and certifying lower bounds for the online bin stretching problem. Theoretical Computer Science, 938(??):1– 15, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200593X.

Zhang:2022:CCE

[1790] Jing Zhang and Yuan Li. Certificate complexity of elementary symmetric Boolean functions. Theoretical Computer Science, 938(??):16– 23, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005758.

Moselely:2022:OSP

[1791] Benjamin Moselely, Ruilong Zhang, and Shanjiawen Zhao. Online scheduling of parallelizable jobs in the directed acyclic graphs and speed-up curves models. Theoretical Computer Science, 938(??):24– 38, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005898.

Bhattacharya:2022:MLA

[1792] Arghya Bhattacharya and Rathish Das. Machine learning advised algorithms for the ski rental problem with a discount. Theoretical Computer Science, 938(??):39– 49, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005862.

Anselmo:2022:ACI

[1793] Marcella Anselmo, Manuela Flores, and Maria Madonia. On k-ary ncubes and isometric words. Theoretical Computer Science, 938(??):50– 64, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005874.

Kamiya:2022:DCW

[1794] Yuichi Kamiya. On dragon curves which have two corners just meeting. Theoretical Computer Science, 938(??): 65–80, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005928.

Luca:2022:LCS

[1795] Alessandro De Luca and Gabriele Fici. On the Lie complexity of Sturmian words. Theoretical Computer Science, 938(??):81–85, November 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005850.

Liang:2022:TCA

[1796] Ya-Chun Liang, Kuan-Yun Lai, Ho-Lin Chen, Kazuo Iwama, and Chung-Shou Liao. Tight competitive analyses of online car-sharing problems. Theoretical Computer Science, 938(??): 86–96, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005904.

He:2022:DSC

[1797] Meng He and Serikzhan Kazi. Data structures for categorical path counting queries. Theoretical Computer Science, 938(??):97–111, November 26, 2022. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005916.

Barcucci:2022:SLR

[1798] Elena Barcucci, Antonio Bernini, and Renzo Pinzani. Strings from linear recurrences and permutations: a Gray code. Theoretical Computer Science, 938(??):112– 120, November 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005886.

Anonymous:2023:PJa

[1799] Anonymous. Pages 1–302 (4 January 2023). Theoretical Computer Science, 939(??):1–302, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBa

[1800] Anonymous. Editorial Board. Theoretical Computer Science, 939(??): ii, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006697.

Georgiou:2023:WGS

[1801] Konstantinos Georgiou and Jesse Lucier. Weighted group search on a line & implications to the priority evacuation problem. Theoretical Computer Science, 939(??): 1–17, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005953.

Dolce:2023:BST

[1802] Francesco Dolce, L’ubomíra Dvořáková, and Edita Pelantová. On balanced sequences and their critical exponent. Theoretical Computer Science, 939(??): 18–47, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005941.

Ren:2023:GTP

[1803] Chunying Ren, Zijun Wu, Dachuan Xu, and Wenqing Xu. A game-theoretic perspective of deep neural networks. Theoretical Computer Science, 939(??): 48–62, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005965.

Bramas:2023:SIG

[1804] Quentin Bramas, Anissa Lamani, and Sébastien Tixeuil. Stand up indulgent gathering. Theoretical Computer Science, 939(??):63–77, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005990.

Castagnos:2023:BET

[1805] Guilhem Castagnos, Dario Catalano, Fabien Laguillaumie, Federico Savasta, and Ida Tucker. Bandwidth-efficient threshold EC-DSA revisited: Online/ offline extensions, identifiable aborts proactive and adaptive security. Theoretical Computer Science, 939(??): 78–104, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006004.

Bensmail:2023:ACD

[1806] Julien Bensmail, Hervé Hocquard, and Dimitri Lajou. On the algorithmic complexity of determining the AVD and NSD chromatic indices of graphs. Theoretical Computer Science, 939(??): 105–118, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005977.

Aloui:2023:BD

[1807] Karam Aloui, Damien Jamet, Hajime Kaneko, Steffen Kopecki, Pierre Popoli, and Thomas Stoll. On the binary digits of n and n^{2}. Theoretical Computer Science, 939(??): 119–139, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522005989.

Galby:2023:UEC

[1808] Esther Galby, Paloma T. Lima, Felix Mann, and Bernard Ries. Using edge contractions to reduce the semitotal domination number. Theoretical Computer Science, 939(??): 140–160, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006090.

Yang:2023:FFH

[1809] Yuxing Yang and Ningning Song.

Fault-free Hamiltonian paths passing through prescribed linear forests in balanced hypercubes with faulty links. Theoretical Computer Science, 939(??): 161–169, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006107.

Han:2023:DPS

[1810] Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. Deciding path size of nondeterministic (and input-driven) pushdown automata. Theoretical Computer Science, 939(??):170–181, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006193.

Martin:2023:FID

[1811] Barnaby Martin, Daniël Paulusma, Siani Smith, and Erik Jan van Leeuwen. Few induced disjoint paths for H-free graphs. Theoretical Computer Science, 939(??):182– 193, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200620X.

Galanis:2023:IIS

[1812] Andreas Galanis, Leslie Ann Goldberg, and Daniel Štefankovič. Implementations and the independent set polynomial below the Shearer threshold. Theoretical Computer Science, 939(??): 194–215, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006351.

Minamisawa:2023:MCC

[1813] Ko Minamisawa, Ryuhei Uehara, and Masao Hara. Mathematical characterizations and computational complexity of anti-slide puzzles. Theoretical Computer Science, 939(??): 216–226, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006363.

Bulgakova:2023:BAP

[1814] D. V. Bulgakova, N. Buzhinsky, and Y. O. Goncharov. On balanced and abelian properties of circular words over a ternary alphabet. Theoretical Computer Science, 939(??):227– 236, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006417.

Xu:2023:FDM

[1815] Xiaojia Xu, Yongcai Wang, Yu Zhang, and Deying Li. A fault diagnosis method to defend scapegoating attack in network tomography. Theoretical Computer Science, 939(??):

237–249, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006399.

Guo:2023:TDD

[1816] Jianxiong Guo, Xingjian Ding, Tian Wang, and Weijia Jia. Theoretical design of decentralized auction framework under mobile crowdsourcing environment. Theoretical Computer Science, 939(??):250– 260, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006375.

Castenow:2023:GEC

[1817] Jannik Castenow, Jonas Harbig, Daniel Jung, Till Knollmann, and Friedhelm Meyer auf der Heide. Gathering a Euclidean closed chain of robots in linear time and improved algorithms for chain-formation. Theoretical Computer Science, 939(??): 261–291, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006429.

Xiao:2023:ULB

[1818] Mingyu Xiao. Upper and lower bounds on approximating weighted mixed domination. Theoretical Computer Science, 939(??):292–302, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006430.

Anonymous:2023:PJb

[1819] Anonymous. Pages 1–296 (9 January 2023). Theoretical Computer Science, 940 (part A)(??):1–296, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBb

[1820] Anonymous. Editorial Board. Theoretical Computer Science, 940 (part A)(??):ii, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006892.

Lima:2023:CGC

[1821] Carlos V. G. C. Lima, Thiago Marcilon, Nicolas Martins, and Rudini Sampaio. The connected greedy coloring game. Theoretical Computer Science, 940 (part A)(??):1– 13, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0304397522006454.

An:2023:SHF

[1822] Zhiyuan An, Jing Pan, Yamin Wen, and Fangguo Zhang. Secret handshakes: Full dynamicity, deniability and lattice-based design. Theoretical Computer Science, 940 (part A) (??):14–35, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006466.

Giroire:2023:RGM

[1823] Frédéric Giroire, Stéphane Pérennes, and Thibaud Trolliet. A random growth model with any real or theoretical degree distribution. Theoretical Computer Science, 940 (part A) (??):36–51, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006478.

Wang:2023:CCB

[1824] Zichen Wang and Hongwei Du. Collaborative coalitions-based joint service caching and task offloading for edge networks. Theoretical Computer Science, 940 (part A)(??):52– 65, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397522006442.

Berglund:2023:RER

[1825] Martin Berglund and Brink van der Merwe. Re-examining regular expressions with backreferences. Theoretical Computer Science, 940 (part A) (??):66–80, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006570.

Kashaev:2023:SOC

[1826] Danish Kashaev and Richard Santiago.

A simple optimal contention resolution scheme for uniform matroids. Theoretical Computer Science, 940 (part A) (??):81–96, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006545.

Bilo:2023:PG

[1827] Vittorio Bilò, Laurent Gourvès, and Jérôme Monnot. Project games. Theoretical Computer Science, 940 (part A) (??):97–111, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006557.

Tsur:2023:FPAa

[1828] Dekel Tsur. Faster parameterized algorithms for two vertex deletion problems. Theoretical Computer Science, 940 (part A)(??):112– 123, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006569.

Buriola:2023:DTI

[1829] Gabriele Buriola, Domenico Cantone, Gianluca Cincotti, Eugenio G. Omodeo, and Gaetano T. Spartà. A decidable theory involving addition of differentiable real functions. Theoretical Computer Science, 940 (part A)(??): 124–148, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006533.

Philip:2023:CHI

[1830] Geevarghese Philip, Rani M. R., and Subashini R. On computing the Hamiltonian index of graphs. Theoretical Computer Science, 940 (part A)(??): 149–179, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006521.

Zhang:2023:ECC

[1831] Hong Zhang, Shuming Zhou, Xiaoqing Liu, and Zhenqin Yu. Extra (component) connectivity and diagnosability of bubble sort networks. Theoretical Computer Science, 940 (part A)(??): 180–189, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006582.

Goyal:2023:TFA

[1832] Dishant Goyal and Ragesh Jaiswal.

Tight FPT approximation for constrained k-center and k-supplier. Theoretical Computer Science, 940 (part A)(??):190–208, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0304397522006594.

Bandopadhyay:2023:SPB

[1833] Susobhan Bandopadhyay, Suman Banerjee, Aritra Banik, and Venkatesh Raman. Structural parameterizations of budgeted graph coloring. Theoretical Computer Science, 940 (part A)(??): 209–221, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397522006600.

Eiben:2023:PSS

[1834] Eduard Eiben, Gregory Gutin, Philip R.

Neary, Clément Rambaud, Magnus Wahlström, and Anders Yeo. Preference swaps for the stable matching problem. Theoretical Computer Science, 940 (part A)(??):222–230, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0304397522006612.

Lopaciuk:2023:BWT

[1835] Szymon Łopaciuk and Daniel Reidenbach. On Billaud words and their companions. Theoretical Computer Science, 940 (part A)(??):231– 253, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006636.

Faro:2023:CSA

[1836] Simone Faro and Stefano Scafiti. Compact suffix automata representations for searching long patterns. Theoretical Computer Science, 940 (part A)(??): 254–268, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006648.

Raskin:2023:PCL

[1837] Mikhail Raskin. Protocols with constant local storage and unreliable communication. Theoretical Computer Science, 940 (part A)(??):269– 282, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200665X.

Li:2023:NAA

[1838] Jiangkun Li and Peng Zhang. New approximation algorithms for the rooted Budgeted Cycle Cover problem. Theoretical Computer Science, 940 (part A)(??):283–295, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397522006752.

Anonymous:2023:PJc

[1839] Anonymous. Pages 1–224 (9 January 2023). Theoretical Computer Science, 940 (part B)(??):1–224, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBc

[1840] Anonymous. Editorial Board. Theoretical Computer Science, 940 (part B)(??):ii, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006958.

Kari:2023:E

[1841] Lila Kari.  Editorial.  Theoreti-

cal Computer Science, 940 (part B) (??):1, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006983.

Thompson:2023:ETS

[1842] Declan Thompson. Execution trace sets for real computation. Theoretical Computer Science, 940 (part B) (??):2–17, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004121.

Witt:2023:HMV

[1843] Carsten Witt. How majorityvote crossover and estimation-of-distribution algorithms cope with fitness valleys. Theoretical Computer Science, 940 (part B)(??):18– 42, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004881.

Phung:2023:ISR

[1844] Xuan Kien Phung. On invertible and stably reversible non-uniform cellular automata. Theoretical Computer Science, 940 (part B)(??):43– 59, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200545X.

Dinneen:2023:QFT

[1845] Michael J. Dinneen, Pankaj S. Ghodla, and Simone Linz. A QUBO formulation for the Tree Containment problem. Theoretical Computer Science, 940 (part B)(??):60– 79, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005461.

Jiang:2023:NSN

[1846] Suxia Jiang, Bowen Xu, Tao Liang, Xiaoliang Zhu, and Tingfang Wu. Numerical spiking neural P systems with production functions on synapses. Theoretical Computer Science, 940 (part B) (??):80–89, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005552.

Xu:2023:AME

[1847] Honghui Xu, Wei Li, and Zhipeng Cai. Analysis on methods to effectively improve transfer learning performance. Theoretical Computer Science, 940 (part B)(??):90–107, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005655.

Vereshchagin:2023:IDF

[1848] Nikolay Vereshchagin. Information disclosure in the framework of Kolmogorov complexity. Theoretical Computer Science, 940 (part B)(??):108– 122, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005825.

Heins:2023:SEN

[1849] Jonathan Heins, Jakob Bossek, Janina Pohl, Moritz Seiler, Heike Trautmann, and Pascal Kerschke. A study on the effects of normalized TSP features for automated algorithm selection. Theoretical Computer Science, 940 (part B)(??):123– 145, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006089.

Bozga:2023:VCB

[1850] Marius Bozga, Radu Iosif, and Joseph Sifakis. Verification of componentbased systems with recursive architectures. Theoretical Computer Science, 940 (part B)(??):146–175, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006181.

Hertling:2023:REJ

[1851] Peter Hertling. Remarks on the effective Jordan decomposition. Theoretical Computer Science, 940 (part B)(??): 176–188, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006491.

Istrate:2023:PWC

[1852] Gabriel Istrate and Cosmin Bonchiş. A parametric worst-case approach to fairness in cooperative games with transferable utility. Theoretical Computer Science, 940 (part B)(??):189– 205, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200648X.

Allender:2023:CHU

[1853] Eric Allender, John Gouwar, Shuichi Hirahara, and Caleb Robelle. Cryptographic hardness under projections for time-bounded Kolmogorov complexity. Theoretical Computer Science, 940 (part B)(??):206–224, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006508.

Anonymous:2023:PJd

[1854] Anonymous. Pages 1–300 (4 January 2023). Theoretical Computer Science, 941(??):1–300, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBd

[1855] Anonymous. Editorial Board. Theoretical Computer Science, 941(??): ii, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007290.

Srividya:2023:TEA

[1856] Putty Srividya, Lavadya Nirmala Devi, and A. Nageswar Rao. A trusted effective approach for forecasting the failure of data link and intrusion in wireless sensor networks. Theoretical Computer Science, 941(??): 1–13, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004753.

Kantamaneni:2023:SEN

[1857] Sravanthi Kantamaneni, A. Charles, and T. Ranga Babu. Speech enhancement with noise estimation and filtration using deep learning models. Theoretical Computer Science, 941(??): 14–28, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004935.

Ankam:2023:MDF

[1858] Sreejyothsna Ankam and Dr. N. Sudhakar Reddy. A mechanism to detecting flooding attacks in quantum enabled cloud-based lowpower and lossy networks. Theoretical Computer Science, 941(??):29– 38, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004923.

Perumalla:2023:MOA

[1859] Subhadra Perumalla, Santanu Chatterjee, and A. P. Siva Kumar. Modelling of oppositional Aquila Optimizer with machine learning enabled secure access control in Internet of Drones environment. Theoretical Computer Science, 941(??): 39–54, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004911.

Karthick:2023:COP

[1860] G. S. Karthick and P. B. Pankajavalli. Chronic obstructive pulmonary disease prediction using Internet of Thingsspiro system and fuzzy-based quantum neural network classifier. Theoretical Computer Science, 941(??): 55–76, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004947.

Savitha:2023:USM

[1861] M. Savitha and M. Senthilkumar. A unique secure multimodal biometricsbased user anonymous authenticated key management protocol (SMUAAKAP) based on block chain mechanism for generic HIoTNs. Theoretical Computer Science, 941(??): 77–90, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005473.

Bulteau:2023:SSF

[1862] Laurent Bulteau, Vincent Jugé, and Stéphane Vialette. On shuffledsquare-free words. Theoretical Computer Science, 941(??):91–103, January 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006405.

Li:2023:RMM

[1863] Wanting Li, Yongcai Wang, Deying Li, and Xiaojia Xu. A robust map matching method by considering memorized multiple matching candidates. Theoretical Computer Science, 941(??): 104–120, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006387.

Zou:2023:AAM

[1864] Wenjie Zou, Longkun Guo, Chunlin Hao, and Lei Liu. Approximation algorithm for MinSum linear barrier coverage with sink-based mobile sensors on the plane. Theoretical Computer Science, 941(??): 121–130, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200651X.

Nakprasit:2023:VAP

[1865] Kittikorn Nakprasit, Watcharintorn Ruksasakchai, and Pongpat Sittitrai. Vertex 2-arboricity of planar graphs without 4-cycles adjacent to 6-cycles. Theoretical Computer Science, 941(??): 131–139, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006624.

Zhang:2023:OBA

[1866] Feiteng Zhang, Bin Liu, Zhenming Liu, and Qizhi Fang. Order based algorithms for the core maintenance problem on edge-weighted graphs. Theoretical Computer Science, 941(??): 140–155, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006727.

Chakraborty:2023:CRI

[1867] Sankardeep Chakraborty and Seungbum Jo. Compact representation of interval graphs and circular-arc graphs of bounded degree and chromatic number. Theoretical Computer Science, 941(??): 156–166, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006788.

Zhang:2023:CAA

[1868] An Zhang, Liang Zhang, Yong Chen, Guangting Chen, and Xing Wang. Complexity and approximation algorithms for two parallel dedicated machine scheduling with conflict constraints. Theoretical Computer Science, 941(??):167– 179, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006764.

Sun:2023:LBA

[1869] Jian Sun, Gregory Gutin, Ping Li, Peihao Shi, and Xiaoyan Zhang. An LP-based approximation algorithm for the generalized traveling salesman path problem. Theoretical Computer Science, 941(??):180– 190, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006776.

Lichen:2023:DCM

[1870] Junran Lichen, Lijian Cai, Jianping Li, Suding Liu, Pengxiang Pan, and Wencheng Wang. Delayconstrained minimum shortest path trees and related problems. Theoretical Computer Science, 941(??): 191–201, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006739.

Queffelec:2023:CPC

[1871] Arthur Queffelec, Ocan Sankur, and François Schwarzentruber. Complexity of planning for connected agents in a partially known environment. Theoretical Computer Science, 941(??): 202–220, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200679X.

Berge:2023:IMC

[1872] Pierre Bergé and Lou Salaün. The influence of maximum (s, t)-cuts on the competitiveness of deterministic strategies for the Canadian Traveller Problem. Theoretical Computer Science, 941(??):221– 240, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006831.

Terai:2023:GPA

[1873] Satoshi Terai, Koichi Wada, and Yoshiaki Katayama. Gathering problems for autonomous mobile robots with lights. Theoretical Computer Science, 941(??): 241–261, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006855.

Altisen:2023:CEW

[1874] Karine Altisen, Pierre Corbineau, and Stéphane Devismes. Certification of an exact worst-case self-stabilization time. Theoretical Computer Science, 941(??): 262–277, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200682X.

Goedgebeur:2023:CGN

[1875] Jan Goedgebeur, Shenwei Huang, Yiao Ju, and Owen Merkel. Colouring graphs with no induced sixvertex path or diamond. Theoretical Computer Science, 941(??):278– 299, January 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006806.

Anonymous:2023:PJe

[1876] Anonymous. Pages 1–312 (9 January 2023). Theoretical Computer Science, 942(??):1–312, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBe

[1877] Anonymous. Editorial Board. Theoretical Computer Science, 942(??): ii, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007411.

Cicerone:2023:APF

[1878] Serafino Cicerone, Alessia Di Fonso, Gabriele Di Stefano, and Alfredo Navarra. Arbitrary pattern formation on infinite regular tessellation graphs. Theoretical Computer Science, 942(??): 1–20, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006843.

Ravelo:2023:CAM

[1879] Santiago Valdés Ravelo and Cristina G. Fernandes. Complexity and approximability of Minimum Path-Collection Exact Covers. Theoretical Computer Science, 942(??):21– 32, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007034.

Pan:2023:PAC

[1880] Zhiyong Pan, Baolei Cheng, Jianxi Fan, Yan Wang, and Xiajing Li. A parallel algorithm to construct edge independent spanning trees on the line graphs of conditional bijective connection networks. Theoretical Computer Science, 942(??): 33–46, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007009.

Abhinav:2023:PAF

[1881] Ankit Abhinav, Susobhan Bandopadhyay, Aritra Banik, and Saket Saurabh. Parameterized algorithms for finding highly connected solution. Theoretical Computer Science, 942(??): 47–56, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522006995.

Raveena:2023:TSF

[1882] Raveena and Krishnendra Shekhawat.

A theory of L-shaped floor-plans. Theoretical Computer Science, 942(??): 57–92, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007010.

Chen:2023:DBS

[1883] Jinmian Chen, Yukun Cheng, Zhiqi Xu, and Yan Cao. Decision on block size in blockchain systems by evolutionary equilibrium analysis. Theoretical Computer Science, 942(??): 93–106, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007022.

Zhang:2023:IAA

[1884] Zhen Zhang, Qilong Feng, Junyu Huang, and Jianxin Wang. Improved approximation algorithms for solving the squared metric k-facility location problem. Theoretical Computer Science, 942(??):107–122, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007046.

Bonomi:2023:OSS

[1885] Silvia Bonomi, Antonella Del Pozzo, Maria Potop-Butucaru, and Sébastien Tixeuil. Optimal self-stabilizing mobile byzantine-tolerant regular register with bounded timestamps. Theoretical Computer Science, 942(??):

123–141, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200706X.

Almethen:2023:DTH

[1886] Abdullah Almethen, Othon Michail, and Igor Potapov. Distributed transformations of Hamiltonian shapes based on line moves. Theoretical Computer Science, 942(??):142– 168, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007058.

Emiris:2023:NNP

[1887] Ioannis Z. Emiris, Vasilis Margonis, and Ioannis Psarros. Near-neighbor preserving dimension reduction via coverings for doubling subsets of ℓ_{1}. Theoretical Computer Science, 942(??): 169–179, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007150.

Karkkainen:2023:SIL

[1888] Juha Kärkkäinen, Marcin Pia̧tkowski, and Simon J. Puglisi. String inference from longest-common-prefix array. Theoretical Computer Science, 942(??): 180–199, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007162.

deMonterno:2023:SMD

[1889] Louis Penet de Monterno, Bernadette Charron-Bost, and Stephan Merz. Synchronization modulo P in dynamic networks. Theoretical Computer Science, 942(??):200–212, January 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007125.

Epstein:2023:SMA

[1890] Leah Epstein. Several methods of analysis for cardinality constrained bin packing. Theoretical Computer Science, 942(??):213– 229, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007137.

Friedrich:2023:PDC

[1891] Tobias Friedrich, Andreas Göbel, Martin S. Krejca, and Marcus Pappik. Polymer dynamics via cliques: New conditions for approximations. Theoretical Computer Science, 942(??): 230–252, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007113.

Ohsaka:2023:RTS

[1892] Naoto Ohsaka. On reconfigurability of target sets. Theoretical Computer Science, 942(??):253– 275, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007149.

Lanzinger:2023:TBB

[1893] Matthias Lanzinger. Tractability beyond β-acyclicity for conjunctive queries with negation and SAT. Theoretical Computer Science, 942(??): 276–296, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007216.

Liu:2023:RAG

[1894] Xiaoqing Liu, Shuming Zhou, Eddie Cheng, and Hong Zhang. Reliability analysis of the generalized balanced hypercube. Theoretical Computer Science, 942(??):297– 311, January 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007198.

Anonymous:2023:PJf

[1895] Anonymous. Pages 1–266 (17 January 2023). Theoretical Computer Science, 943(??):1–266, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBf

[1896] Anonymous. Editorial Board. Theoretical Computer Science, 943(??): ii, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007733.

Zhao:2023:RAD

[1897] Shu-Li Zhao and Jou-Ming Chang.

Reliability assessment of the divideand-swap cube in terms of generalized connectivity. Theoretical Computer Science, 943(??):1–15, January 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007204.

Bohnlein:2023:SPG

[1898] Toni Böhnlein, Oliver Schaudt, and Joachim Schauer. Stackelberg packing games. Theoretical Computer Science, 943(??):16–35, January 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007228.

Panda:2023:AMS

[1899] B. S. Panda and Juhi Chaudhary. Acyclic matching in some subclasses of graphs. Theoretical Computer Science, 943(??):36– 49, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007253.

Berzish:2023:TME

[1900] Murphy Berzish, Joel D. Day, Vijay Ganesh, Mitja Kulczynski, Florin Manea, Federico Mora, and Dirk Nowotka. Towards more efficient methods for solving regular-expression heavy string constraints. Theoretical Computer Science, 943(??):50– 72, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200723X.

Aichholzer:2023:GLT

[1901] Oswin Aichholzer, Matias Korman, Yoshio Okamoto, Irene Parada, Daniel Perz, André van Renssen, and Birgit Vogtenhuber. Graphs with large total angular resolution. Theoretical Computer Science, 943(??):73– 88, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007320.

Roditty:2023:ADO

[1902] Liam Roditty and Roei Tov. Approximate distance oracles with improved stretch for sparse graphs. Theoretical Computer Science, 943(??): 89–101, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006818.

Bensmail:2023:MBL

[1903] Julien Bensmail, Foivos Fioravantes, Fionn Mc Inerney, Nicolas Nisse, and Nacim Oijid. The Maker–Breaker Largest Connected Subgraph game. Theoretical Computer Science, 943(??): 102–120, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007356.

Civril:2023:NAA

[1904] A. Çivril. A new approximation algorithm for the minimum 2-edge-connected spanning subgraph problem. Theoretical Computer Science, 943(??): 121–130, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007368. See

corrigendum [2094].

Ito:2023:RNS

[1905] Takehiro Ito, Yuni Iwamasa, Yasuaki Kobayashi, Yu Nakahata, Yota Otachi, and Kunihiro Wasa. Reconfiguring (non-spanning) arborescences. Theoretical Computer Science, 943(??): 131–141, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007241.

Nemmaoui:2023:GTA

[1906] Saadia Nemmaoui, Mohamed Baslam, and Belaid Bouikhalene. Game theory analysis of biometric systems market: Users choice VS service providers revenue. Theoretical Computer Science, 943(??):142– 152, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007186.

Liu:2023:MCD

[1907] Depeng Liu, Bow-Yaw Wang, Chen Fu, and Lijun Zhang. Model checking differentially private properties. Theoretical Computer Science, 943(??):

153–170, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005837.

Kopp:2023:LAE

[1908] Nils Köpp and Helmut Schwichtenberg. Lookahead analysis in exact real arithmetic with logical methods. Theoretical Computer Science, 943(??): 171–186, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522005849.

Bezhanishvili:2023:RHP

[1909] G. Bezhanishvili, J. Harding, and P. J. Morandi. Remarks on hyperspaces for Priestley spaces. Theoretical Computer Science, 943(??):187– 202, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007174.

Anand:2023:DLN

[1910] M. Anand, Kishan Bhushan Sahay, Mohammed Altaf Ahmed, Daniyar Sultan, Radha Raman Chandan, and Bharat Singh. Deep learning and natural language processing in computation for offensive language detection in online social networks by feature selection and ensemble classification techniques. Theoretical Computer Science, 943(??): 203–218, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003887.

Lakshmi:2023:MVR

[1911] Takkolu Sree Lakshmi, Muthukumarasamy Govindarajan, and Asadi Sreenivasulu. Malware visual resemblance analysis with minimum losses using Siamese neural networks. Theoretical Computer Science, 943(??): 219–229, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004406.

Latha:2023:OSA

[1912] V. L. Padma Latha, N. Sudhakar Reddy, and A. Suresh Babu. Optimizing scalability and availability of cloud based software services using modified scale rate limiting algorithm. Theoretical Computer Science, 943(??): 230–240, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522004418.

Qian:2023:MOE

[1913] Chao Qian, Dan-Xuan Liu, Chao Feng, and Ke Tang. Multi-objective evolutionary algorithms are generally good: Maximizing monotone submodular functions over sequences. Theoretical Computer Science, 943(??): 241–266, January 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007332.

Anonymous:2023:Ja

[1914] Anonymous. 25 January 2023. Theoretical Computer Science, 944(??):

??, January 25, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBg

[1915] Anonymous. Editorial Board. Theoretical Computer Science, 944(??): ??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000282.

Mihelcic:2023:CRM

[1916] Matej Mihelčić and Adrian Satja Kurdija. On the complexity of redescription mining. Theoretical Computer Science, 944(??):??, January 25, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200754X.

Guo:2023:SMD

[1917] Longkun Guo, Kewen Liao, Di Xiao, and Pei Yao. Submodular maximization over data streams with differential privacy noise. Theoretical Computer Science, 944(??): ??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522006740.

Tang:2023:SAS

[1918] Shaojie Tang and Jing Yuan. Streaming adaptive submodular maximization. Theoretical Computer Science, 944(??):??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007101.

Kipouridis:2023:TBN

[1919] Evangelos Kipouridis, Paul G. Spirakis, and Kostas Tsichlas. Thresholdbased network structural dynamics. Theoretical Computer Science, 944 (??):??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007514.

Esperet:2023:DCL

[1920] Louis Esperet, Sébastien Julliot, and Arnaud de Mesmay. Distributed coloring and the local structure of unit-disk graphs. Theoretical Computer Science, 944(??):??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007563.

Georgiou:2023:EUD

[1921] Konstantinos Georgiou, Sean Leizerovich, Jesse Lucier, and Somnath Kundu. Evacuating from ℓ_{p} unit disks in the wireless model. Theoretical Computer Science, 944(??): ??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007575.

Ding:2023:RBD

[1922] Tongtong Ding and Min Xu. Relationship between diagnosability and non-inclusive diagnosability of trianglefree connected graphs under the PMC model. Theoretical Computer Science, 944(??):??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007587.

Li:2023:FGV

[1923] Weifeng Li, Longxin Lin, Zhen Zhang, and Shuqiang Huang. The family of generalized variational network of cube-connected cycles. Theoretical Computer Science, 944(??): ??, January 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007551.

Anonymous:2023:Fa

[1924] Anonymous. 4 February 2023. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBh

[1925] Anonymous. Editorial Board. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000385.

Nan:2023:ACC

[1926] Jiehui Nan, Haiming Zhu, Shigang Liu, and Honggang Hu. Adaptively codecorrelation robustness functions and its applications to private set intersection. Theoretical Computer Science, 945(??):??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752200768X.

Zhou:2023:CFE

[1927] Yu Zhou, Shengli Liu, and Nan Cui.

Computational fuzzy extractor from LWE. Theoretical Computer Science, 945(??):??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007691.

Lyons:2023:FON

[1928] Steven Lyons, Raju Rangaswami, and Ning Xie. Finding optimal non-datapath caching strategies via network flow. Theoretical Computer Science, 945(??):??, February 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007678.

Luo:2023:TOL

[1929] Chuanwen Luo, Ning Liu, Yunan Hou, Yi Hong, Zhibo Chen, and Deying Li. Trajectory optimization of lasercharged UAV to minimize the average age of information for wireless rechargeable sensor network. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007617.

Silva:2023:DSQ

[1930] Alexandre Silva, Carlos Pereira dos Santos, João Pedro Neto, and Richard J. Nowakowski. Disjunctive sums of quasi-nimbers. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752200737X.

Sun:2023:DSA

[1931] Xueli Sun, Jianxi Fan, Baolei Cheng, and Yan Wang. The t/m-diagnosis strategy of augmented k-ary n-cubes. Theoretical Computer Science, 945 (??):??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007538.

Misselbeck-Wessel:2023:RTS

[1932] Daniel Misselbeck-Wessel and Peter Schuster. Radical theory of Scott-open filters. Theoretical Computer Science, 945(??):??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007599.

Selivanov:2023:DCS

[1933] Victor Selivanov. Descriptive complexity of qcb_{0}-spaces. Theoretical Computer Science, 945(??):??, February 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007459.

Eberl:2023:HOC

[1934] Matthias Eberl. Higher-order concepts for the potential infinite. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007460.

Chen:2023:PSI

[1935] Yijia Chen, Pierre-Louis Curien, Yuxin Deng, and Min Zhang. Preface for the special issue of Theoretical Computer Science in honor of the 60 ^{th} birthday of Yuxi Fu. Theoretical Computer Science, 945(??): ??, February 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000336.

Anonymous:2023:Fb

[1936] Anonymous. 10 February 2023. Theoretical Computer Science, 946(??):??, February 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBi

[1937] Anonymous. Editorial Board. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000518.

Faro:2023:ICD

[1938] Simone Faro, Francesco Pio Marino, and Arianna Pavone. Improved characters distance sampling for online and offline text searching. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007642.

Caucchiolo:2023:HAM

[1939] Andrea Caucchiolo and Ferdinando Cicalese. Hardness and approximation of multiple sequence alignment with column score. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007654.

Galeana-Sanchez:2023:GPC

[1940] Hortensia Galeana-Sánchez, Felipe Hernández-Lorenzana, and Rocío Sánchez-López. A generalization of properly colored paths and cycles in edge-colored graphs. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007629.

Ducoffe:2023:DPW

[1941] Guillaume Ducoffe. Distance problems within Helly graphs and k-Helly graphs. Theoretical Computer Science, 946(??):??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000038.

Bekos:2023:RDP

[1942] Michael A. Bekos, Giordano Da Lozzo, Fabrizio Frati, Martin Gronemann, Tamara Mchedlidze, and Chrysanthi N. Raftopoulou. Recognizing DAGs with page-number 2 is NP-complete. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000026.

Chaubal:2023:TBS

[1943] Siddhesh Chaubal and Anna Gál. Tight bounds on sensitivity and block sensitivity of some classes of transitive functions. Theoretical Computer Science, 946(??):??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007782.

Kopra:2023:RLE

[1944] Johan Kopra. Rapid left expansivity, a commonality between Wolfram’s Rule 30 and powers of p/q. Theoretical Computer Science, 946(??): ??, February 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007502.

Doerr:2023:SDM

[1945] Benjamin Doerr and Amirhossein Rajabi. Stagnation detection meets fast mutation. Theoretical Computer Science, 946(??):??, February 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007526.

Anonymous:2023:Fc

[1946] Anonymous. 20 February 2023. Theoretical Computer Science, 947(??):??, February 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBj

[1947] Anonymous. Editorial Board. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000622.

Gomez:2023:IHR

[1948] Renzo Gómez, Flávio Keidi Miyazawa, and Yoshiko Wakabayashi. Improved -hardness results for the minimum tspanner problem on bounded-degree graphs. Theoretical Computer Science, 947(??):??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300004X.

Borowiecki:2023:CBT

[1949] Piotr Borowiecki, Dariusz Dereniowski, and Dorota Osula. The complexity of bicriteria tree-depth. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397522007666.

Berdinsky:2023:SCF

[1950] Dmitry Berdinsky, Sanjay Jain, Bakhadyr Khoussainov, and Frank Stephan. String compression in FA–presentable structures. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300018X.

Niu:2023:TDC

[1951] Ruichao Niu, Shujie Zhou, and Min Xu. Two-disjoint-cycle-cover vertex bipancyclicity of bipartite hypercube-like networks. Theoretical Computer Science, 947(??):??, February 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000166.

Wang:2023:SFT

[1952] Yu Wang and Jixiang Meng. Structure fault-tolerance of divide-and-swap k-ary n-cube. Theoretical Computer Science, 947(??):??, February 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000178.

Lan:2023:HCB

[1953] Ting Lan and Huazhong Lü. Hamiltonian cycles of balanced hypercube with more faulty edges. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300021X.

Richard:2023:ASS

[1954] Adrien Richard and Elisa Tonello. Attractor separation and signed cycles in asynchronous Boolean networks. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000191.

Ganesan:2023:PAD

[1955] Ashwin Ganesan. Performance analysis of distance-1 distributed algorithms for admission control under the 2-hop interference model. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000312.

Trotta:2023:DPG

[1956] Davide Trotta, Matteo Spadetto, and Valeria de Paiva. Dialectica principles via Gödel doctrines. Theoretical Computer Science, 947(??): ??, February 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000051.

Barmpalias:2023:KGT

[1957] George Barmpalias and Alexander Shen. The Kučera–Gács theorem revisited by Levin. Theoretical Computer Science, 947(??):??, February 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000063.

Anonymous:2023:Fd

[1958] Anonymous. 28 February 2023. Theoretical Computer Science, 948(??):??, February 28, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBk

[1959] Anonymous. Editorial Board. Theoretical Computer Science, 948(??): ??, February 28, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000798.

Rabinowitz:2023:OMS

[1960] Assaf Rabinowitz and Dror Rawitz.

Overflow management with selfeliminations. Theoretical Computer Science, 948(??):??, February 28, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000452.

Halldorsson:2023:SCC

[1961] Magnús M. Halldórsson and Alexandre Nolin. Superfast coloring in CON-GEST via efficient color sampling. Theoretical Computer Science, 948(??): ??, February 28, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000245.

Kamiyama:2023:PEM

[1962] Naoyuki Kamiyama. Pareto efficient matchings with pairwise preferences. Theoretical Computer Science, 948 (??):??, February 28, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000208.

Pitt:2023:TTT

[1963] Lenny Pitt. Turing Tumble is Turing-Complete. Theoretical Computer Science, 948(??):??, February 28, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000476.

Alistarh:2023:WFA

[1964] Dan Alistarh, Faith Ellen, and Joel Rybicki. Wait-free approximate agreement on graphs. Theoretical Computer Science, 948(??): ??, February 28, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000464.

Anonymous:2023:Ma

[1965] Anonymous. 9 March 2023. Theoretical Computer Science, 949(??):??, March 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBl

[1966] Anonymous. Editorial Board. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000932.

DAgostino:2023:ORL

[1967] Giovanna D’Agostino, Davide Martincigh, and Alberto Policriti. Ordering regular languages and automata: Complexity. Theoretical Computer Science, 949(??):??, March 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000221.

Fan:2023:DRR

[1968] Austen Z. Fan and Jin-Yi Cai. Dichotomy result on 3-regular bipartite non-negative functions. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000580.

Laber:2023:NTB

[1969] Eduardo Sany Laber and Lucas Saadi Murtinho. Nearly tight bounds on the price of explainability for the k-center and the maximumspacing clustering problems. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000579.

Zhang:2023:PSU

[1970] Nan Zhang, Chaofeng Yu, Zhenhua Duan, and Cong Tian. A proof system for unified temporal logic.

Theoretical Computer Science, 949 (??):??, March 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000154.

Bang-Jensen:2023:CFL

[1971] J. Bang-Jensen and A. Yeo. The complexity of finding low chromatic spanning sub(di)graphs with prescribed connectivity properties. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000713.

Zhang:2023:EIN

[1972] Long Zhang, Jiguo Yu, Yuzhong Zhang, Donglei Du, and Min Guo. Efficiency and inefficiency of Nash equilibrium for scheduling games on batchingmachines with activation cost. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000439.

Capuni:2023:TRC

[1973] Ilir Capuni, Jarkko Kari, and Alexander Shen. Taming randomness and complexity — Essays in honour of Professor Péter Gács. Theoretical Computer Science, 949(??): ??, March 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000890.

Bienvenu:2023:RD

[1974] Laurent Bienvenu, Valentino Delle Rose, and Wolfgang Merkle. Relativized depth. Theoretical Computer Science, 949(??):??, March 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523000075.

Anonymous:2023:Mb

[1975] Anonymous. 16 March 2023. Theoretical Computer Science, 950(??): ??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBm

[1976] Anonymous. Editorial Board. Theoretical Computer Science, 950(??): ??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001032.

Galby:2023:CBS

[1977] Esther Galby. The complexity of blocking (semi)total dominating sets with edge contractions. Theoretical Computer Science, 950(??):??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522007630.

Gotte:2023:BSM

[1978] Thorsten Götte, Christina Kolb, Christian Scheideler, and Julian Werthmann. Beep-and-sleep: Message and Energy Efficient Set Cover. Theoretical Computer Science, 950(??):

??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000695.

Bshouty:2023:OTL

[1979] Nader H. Bshouty. An optimal tester for k-linear. Theoretical Computer Science, 950(??):??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000725.

Lehre:2023:SIT

[1980] Per Kristian Lehre, Aneta Neumann, and Chao Qian. Special issue on theoretical foundations of evolutionary computation. Theoretical Computer Science, 950(??): ??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000981.

Bossek:2023:DAT

[1981] Jakob Bossek and Dirk Sudholt. Do additional target points speed up evolutionary algorithms? Theoretical Computer Science, 950(??): ??, March 16, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000701.

Anonymous:2023:Mc

[1982] Anonymous. 24 March 2023. Theoretical Computer Science, 951(??):

??, March 24, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBn

[1983] Anonymous. Editorial Board. Theoretical Computer Science, 951(??): ??, March 24, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001214.

Bazhenov:2023:LAS

[1984] Nikolay Bazhenov, Vittorio Cipriani, and Luca San Mauro. Learning algebraic structures with the help of Borel equivalence relations. Theoretical Computer Science, 951(??): ??, March 24, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000750.

Shang:2023:CRC

[1985] Yilun Shang. Concentration of rainbow k-connectivity of a multiplex random graph. Theoretical Computer Science, 951(??):??, March 24, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000841.

Yang:2023:HSF

[1986] Lulu Yang, Xiaohui Hua, and Yuxing Yang. Hyper star fault tolerance of bubble sort networks. Theoretical Computer Science, 951(??): ??, March 24, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300083X.

Chang:2023:DGP

[1987] Yi-Jun Chang, Jan Studený, and Jukka Suomela. Distributed graph problems through an automata-theoretic lens. Theoretical Computer Science, 951(??):??, March 24, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000233.

Tsin:2023:SCA

[1988] Yung H. Tsin. A simple certifying algorithm for 3-edge-connectivity. Theoretical Computer Science, 951 (??):??, March 24, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000737.

Peyriere:2023:MMD

[1989] Jacques Peyrière. Moore machines duality. Theoretical Computer Science, 951(??):??, March 24, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000877.

Branson:2023:FJR

[1990] Luke Branson and Andrew M. Sutton. Focused jump-and-repair constraint handling for fixed-parameter tractable graph problems closed under induced subgraphs. Theoretical Computer Science, 951(??):??, March 24, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000324.

Anonymous:2023:Md

[1991] Anonymous. 31 March 2023. Theoretical Computer Science, 952(??): ??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBo

[1992] Anonymous. Editorial Board. Theoretical Computer Science, 952(??): ??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001299.

Rusu:2023:CRS

[1993] Irena Rusu. On the complexity of recognizing Stick, BipHook and Max Point-Tolerance graphs. Theoretical Computer Science, 952(??): ??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000865.

Rajsbaum:2023:DCP

[1994] Sergio Rajsbaum. A distributed computing perspective of unconditionally secure information transmission in Russian cards problems. Theoretical Computer Science, 952(??): ??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000749.

Bu:2023:LUB

[1995] Gewu Bu, Zvi Lotker, Maria Potop-Butucaru, and Mikaël Rabie. Lower and upper bounds for deterministic convergecast with labeling schemes. Theoretical Computer Science, 952 (??):??, March 31, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000889.

Zhang:2023:CCA

[1996] Qifan Zhang, Shuming Zhou, and Eddie Cheng. Component connectivity of augmented cubes. Theoretical Computer Science, 952(??): ??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300097X.

Song:2023:SZK

[1997] Yongcheng Song, Jiang Zhang, Xinyi Huang, Wei Wu, and Haining Yang. Statistical zero-knowledge and analysis of rank-metric zero-knowledge proofs of knowledge. Theoretical Computer Science, 952(??):??, March 31, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000440.

Lecroq:2023:TSI

[1998] Thierry Lecroq and Svetlana Puzynina. TCS special issue: Combinatorics on Words — WORDS 2021. Theoretical Computer Science, 952 (??):??, March 31, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000014.

Abbott:2023:FCR

[1999] Alastair A. Abbott, Cezar Câmpeanu, Ludwig Staiger, Marius Zimand, and Arto Salomaa. Frontiers of computability, randomness, and complexity (dedicated to the 70th birthday of Professor Cristian Calude). Theoretical Computer Science, 952 (??):??, March 31, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001329.

Anonymous:2023:Aa

[2000] Anonymous. 10 April 2023. Theoretical Computer Science, 953(??): ??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBp

[2001] Anonymous. Editorial Board. Theoretical Computer Science, 953(??): ??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001494.

Cirpons:2023:PTM

[2002] R. Cirpons and J. D. Mitchell. Polynomial time multiplication and normal forms in free bands. Theoretical Computer Science, 953(??): ??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000968.

Wolnik:2023:DTN

[2003] Barbara Wolnik, Anna Nenca, and Bernard De Baets. A decomposition theorem for number-conserving multistate cellular automata on triangular grids. Theoretical Computer Science, 953(??):??, April 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001081.

Tsur:2023:FPAb

[2004] Dekel Tsur. Faster parameterized algorithms for Bicluster Editing and Flip Consensus Tree. Theoretical Computer Science, 953(??): ??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001093.

Sheng:2023:FPA

[2005] Bin Sheng and Gregory Gutin. Fixed parameterized algorithms for generalized feedback vertex set problems. Theoretical Computer Science, 953 (??):??, April 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001111.

Natsui:2023:FPB

[2006] Kei Natsui and Kenjiro Takazawa. Finding popular branchings in vertexweighted directed graphs. Theoretical Computer Science, 953(??):

??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001123.

Gluck:2023:RCP

[2007] Robert Glück and Tetsuo Yokoyama.

Reversible computing from a programming language perspective. Theoretical Computer Science, 953(??): ??, April 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397522003619.

Anonymous:2023:Ab

[2008] Anonymous. 18 April 2023. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBq

[2009] Anonymous. Editorial Board. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001652.

Bramas:2023:APD

[2010] Quentin Bramas, Anissa Lamani, and Sébastien Tixeuil. The agreement power of disagreement. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000853.

Ashkenazi:2023:LFS

[2011] Yotam Ashkenazi, Shlomi Dolev, Sayaka Kamei, Yoshiaki Katayama, Fukuhito Ooshita, and Koichi Wada. Location functions for self-stabilizing Byzantine tolerant swarms. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523000683.

Krithika:2023:SET

[2012] R. Krithika, Pranabendu Misra, and Prafullkumar Tale. A single exponential-time FPT algorithm for cactus contraction. Theoretical Computer Science, 954(??):??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001160.

Chen:2023:OTC

[2013] Hong-Bin Chen and Geng-Yu Chen. Optimal 2-traceability codes of length 4. Theoretical Computer Science, 954 (??):??, April 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001135.

Aravind:2023:PMS

[2014] N. R. Aravind and Roopam Saxena. Perfectly matched sets in graphs: Parameterized and exact computation. Theoretical Computer Science, 954 (??):??, April 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300110X.

Wang:2023:NDP

[2015] Biing-Feng Wang. A new dynamic programming algorithm for the simplified partial digest problem. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001147.

Cantone:2023:CAD

[2016] Domenico Cantone, Pietro Maugeri, and Eugenio G. Omodeo. Complexity assessments for decidable fragments of Set Theory. III: Testers for crucial, polynomial-maximal decidable Boolean languages. Theoretical Computer Science, 954(??): ??, April 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523000993.

Anonymous:2023:Ac

[2017] Anonymous. 26 April 2023. Theoretical Computer Science, 955(??): ??, April 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBr

[2018] Anonymous. Editorial Board. Theoretical Computer Science, 955(??): ??, April 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001810.

Xie:2023:GTR

[2019] Ning Xie, Shuai Xu, and Yekun Xu. A generalization of a theorem of Rothschild and van Lint. Theoretical Computer Science, 955 (??):??, April 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001159.

Munaro:2023:AAS

[2020] Andrea Munaro and Shizhou Yang. On algorithmic applications of simwidth and mim-width of (H_{1}, H_{2})-free graphs. Theoretical Computer Science, 955(??):??, April 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300138X.

Suzuki:2023:CCS

[2021] Yuma Suzuki and Ryuhei Miyashiro. Characterization of canonical systems with six types of coins for the change-making problem. Theoretical Computer Science, 955(??): ??, April 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001354.

Datta:2023:AME

[2022] Ajoy K. Datta, Stéphane Devismes, Colette Johnen, and Lawrence L. Larmore. Analysis of a memoryefficient self-stabilizing BFS spanning tree construction. Theoretical Computer Science, 955(??):??, April 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001172.

Baazizi:2023:NCJ

[2023] Mohamed-Amine Baazizi, Dario Colazzo, Giorgio Ghelli, Carlo Sartiani, and Stefanie Scherzinger. Negationclosure for JSON Schema. Theoretical Computer Science, 955(??): ??, April 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001366.

Anonymous:2023:Me

[2024] Anonymous. 4 May 2023. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBs

[2025] Anonymous. Editorial Board. Theoretical Computer Science, 956(??): ??, May 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001937.

Gomes:2023:DM

[2026] Guilherme C. M. Gomes, Bruno P.

Masquio, Paulo E. D. Pinto, Vinicius F. dos Santos, and Jayme L. Szwarcfiter. Disconnected matchings. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001342.

Konstantinidis:2023:ACP

[2027] Stavros Konstantinidis, António Machiavelo, Nelma Moreira, and Rogério Reis. On the average complexity of partial derivative transducers. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001433.

Nakanishi:2023:NVC

[2028] Misa Nakanishi. A note on vertices contained in the minimum dominating set of a graph with minimum degree three. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001445.

Leonardos:2023:BFT

[2029] Nikos Leonardos, Aris Pagourtzis, and Ioannis Papaioannou. Byzantine fault tolerant symmetric-persistent circle evacuation. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001524.

Hung:2023:RDV

[2030] Ruo-Wei Hung. Restrained domination and its variants in extended supergrid graphs. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001457.

Rizzi:2023:RWS

[2031] Romeo Rizzi and Stéphane Vialette.

On recognising words that are squares for the shuffle product. Theoretical Computer Science, 956(??): ??, May 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303134.

Orellana-Martin:2023:EFE

[2032] David Orellana-Martín, Luis Valencia-Cabrera, and Mario J. Pérez-Jiménez. The environment as a frontier of efficiency in tissue P systems with communication rules. Theoretical Computer Science, 956(??):??, May 4, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001251.

Anonymous:2023:Mf

[2033] Anonymous. 12 May 2023. Theoretical Computer Science, 957(??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBt

[2034] Anonymous. Editorial Board. Theoretical Computer Science, 957(??): ??, May 12, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002153.

Li:2023:SVR

[2035] Yiming Li, Shengli Liu, Shuai Han, Dawu Gu, and Jian Weng. Simulatable verifiable random function from the LWE assumption. Theoretical Computer Science, 957(??): ??, May 12, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001391.

Wang:2023:SCD

[2036] Cai-Xia Wang, Yu Yang, and Shou-Jun Xu. Secure connected domination and secure total domination in unit disk graphs and rectangle graphs. Theoretical Computer Science, 957(??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001378.

Gu:2023:PMS

[2037] Shuyang Gu, Chuangen Gao, Jun Huang, and Weili Wu. Profit maximization in social networks and nonmonotone DR-submodular maximization. Theoretical Computer Science, 957(??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001603.

Zhao:2023:NAS

[2038] Xueyang Zhao, Binghao Yan, and Peng Zhang. New algorithms for a simple measure of network partitioning. Theoretical Computer Science, 957(??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001597.

Liu:2023:MME

[2039] Hongxi Liu, Mingzu Zhang, and Xiaowei Xu. Many-to-many edge-disjoint paths in (n, k)-enhanced hypercube under three link-faulty hypotheses. Theoretical Computer Science, 957 (??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001688.

Diaz-Caro:2023:NCN

[2040] Alejandro Díaz-Caro and Gilles Dowek.

A new connective in natural deduction, and its application to quantum computing. Theoretical Computer Science, 957(??):??, May 12, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001536.

Kraus:2023:TTA

[2041] Nicolai Kraus, Fredrik Nordvall Forsberg, and Chuangjie Xu. Typetheoretic approaches to ordinals. Theoretical Computer Science, 957(??): ??, May 12, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001561.

Aishwaryaprajna:2023:VAR

[2042] Aishwaryaprajna and Jonathan E.

Rowe. The Voting algorithm is robust to various noise models. Theoretical Computer Science, 957(??): ??, May 12, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001573.

Anonymous:2023:Mg

[2043] Anonymous. 22 May 2023. Theoretical Computer Science, 958(??):??, May 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBu

[2044] Anonymous. Editorial Board. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002335.

Zhang:2023:FTC

[2045] Hong Zhang, Shuming Zhou, Qifan Zhang, and Tao Tian. Fault tolerance of composite graph based on disc-ring and folded hypercube. Theoretical Computer Science, 958 (??):??, May 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300169X.

Mrykhin:2023:HLG

[2046] Mikhail Mrykhin and Alexander Okhotin. The hardest language for grammars with context operators. Theoretical Computer Science, 958 (??):??, May 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001421.

Labarre:2023:SPB

[2047] Anthony Labarre. Sorting by prefix block-interchanges. Theoretical Computer Science, 958(??):??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001706.

Agrawal:2023:PCP

[2048] Akanksha Agrawal, Sutanay Bhattacharjee, Satyabrata Jana, and Abhishek Sahu. Parameterized complexity of perfectly matched sets. Theoretical Computer Science, 958 (??):??, May 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001743.

Qiao:2023:TDC

[2049] Hongwei Qiao and Jixiang Meng.

Two-disjoint-cycle-cover vertex pancyclicity of augmented cubes. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001731.

Abu-Khzam:2023:IFP

[2050] Faisal N. Abu-Khzam, Norma Makarem, and Maryam Shehab. An improved fixed-parameter algorithm for 2-Club Cluster Edge Deletion. Theoretical Computer Science, 958(??):

??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001779.

Yuan:2023:PDI

[2051] Jun Yuan, Shuyuan Ge, and Aixia Liu. The partial diagnosability of interconnection networks under the Hybrid PMC model. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300172X.

Hajebi:2023:PCA

[2052] Sahab Hajebi and Ramin Javadi. On the parameterized complexity of the acyclic matching problem. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001755.

Pasqua:2023:ACD

[2053] Michele Pasqua and Marino Miculan. AbU: a calculus for distributed event-driven programming with attribute-based interaction. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001548.

Alhazov:2023:NNC

[2054] Artiom Alhazov, Rudolf Freund, Sergiu Ivanov, and Sergey Verlan. Numerical networks of cells. Theoretical Computer Science, 958(??): ??, May 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300186X.

Dinneen:2023:SSS

[2055] Michael J. Dinneen, Alec Henderson, and Radu Nicolescu. Sublinear P system solutions to NP-complete problems. Theoretical Computer Science, 958(??):??, May 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001615.

Anonymous:2023:Mh

[2056] Anonymous. 30 May 2023. Theoretical Computer Science, 959(??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBv

[2057] Anonymous. Editorial Board. Theoretical Computer Science, 959(??): ??, May 30, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002621.

Niu:2023:FTA

[2058] Baohua Niu, Shuming Zhou, Hong Zhang, and Qifan Zhang. Fault tolerability analysis of folded crossed cubes based on g-component and ggood neighbor fault pattern. Theoretical Computer Science, 959(??): ??, May 30, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001718.

Dong:2023:ADN

[2059] Luobing Dong, Zhilong Sun, and Yongsong Qin. An attention-driven nonlinear optimization method for CSbased hyperspectral image reconstruction. Theoretical Computer Science, 959(??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001585.

DiIanni:2023:OEA

[2060] Miriam Di Ianni. Opinion evolution among friends and foes: the deterministic majority rule. Theoretical Computer Science, 959(??): ??, May 30, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001883.

Xiao:2023:VKP

[2061] Mingyu Xiao and Shaowei Kou. A 5k-vertex kernel for 3-path vertex cover. Theoretical Computer Science, 959(??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001858.

Nakai:2023:CRL

[2062] Takeshi Nakai and Kazumasa Shinagawa. Constant-round linear-broadcast secure computation with penalties. Theoretical Computer Science, 959 (??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001871.

Hatanaka:2023:FPA

[2063] Tatsuhiko Hatanaka, Felix Hommelsheim, Takehiro Ito, Yusuke Kobayashi, Moritz Mühlenthaler, and Akira Suzuki. Fixed-parameter algorithms for graph constraint logic. Theoretical Computer Science, 959 (??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523001767.

Horne:2023:WPF

[2064] Ross Horne, Sjouke Mauw, and Semen Yurkov. When privacy fails, a formula describes an attack: a complete and compositional verification method for the applied π-calculus. Theoretical Computer Science, 959 (??):??, May 30, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300155X.

Anonymous:2023:Jb

[2065] Anonymous. 7 June 2023. Theoretical Computer Science, 960(??):??, June 7, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBw

[2066] Anonymous. Editorial Board. Theoretical Computer Science, 960(??): ??, June 7, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002700.

Zhao:2023:BCD

[2067] Jun Zhao, Minghao Yang, Junqing Gong, Kai Zhang, and Haifeng Qian. Bounded-collusion decentralized ABE with sublinear parameters. Theoretical Computer Science, 960(??): ??, June 7, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002104.

Brandstadt:2023:CDA

[2068] Andreas Brandstädt and Raffaele Mosca. Combining decomposition approaches for the Maximum Weight Stable Set problem. Theoretical Computer Science, 960(??): ??, June 7, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300227X.

Chlebikova:2023:ISR

[2069] Janka Chlebíková, Clément Dallard, and Niklas Paulsen. Impact of soft ride time constraints on the complexity of scheduling in Dial-A-Ride Problems. Theoretical Computer Science, 960(??):??, June 7, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002360.

Nitaj:2023:NAS

[2070] Abderrahmane Nitaj, Willy Susilo, and Joseph Tonien. A new attack on some RSA variants. Theoretical Computer Science, 960(??): ??, June 7, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002116.

Bice:2023:LFP

[2071] Tristan Bice and Wiesław Kubiś.

Lattice-free and point-free: Vickers duality for subbases of stably locally compact spaces. Theoretical Computer Science, 960(??):??, June 7, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002086.

Yilmaz:2023:NFD

[2072] Damla Yılmaz. A note on fuzzy derivations of some ideals on d-algebras. Theoretical Computer Science, 960 (??):??, June 7, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002372.

Kumar:2023:SPD

[2073] Narendra Kumar, Siddhartha Siddhiprada Bhoi, and Abhay Kumar Singh. A study of primer design with w-constacyclic shift over F_{4}. Theoretical Computer Science, 960 (??):??, June 7, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002384. See

corrigendum [2379].

Anonymous:2023:Jc

[2074] Anonymous. 15 June 2023. Theoretical Computer Science, 961(??):??, June 15, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBx

[2075] Anonymous. Editorial Board. Theoretical Computer Science, 961(??): ??, June 15, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002839.

Li:2023:OFL

[2076] Fu Li, C. Gregory Plaxton, and Vaibhav B. Sinha. The obnoxious facility location game with dichotomous preferences. Theoretical Computer Science, 961(??):??, June 15, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002438.

Fu:2023:SAS

[2077] Bin Fu, Yumei Huo, and Hairong Zhao. Streaming approximation scheme for minimizing total completion time on parallel machines subject to varying processing capacity. Theoretical Computer Science, 961 (??):??, June 15, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002505.

Cameron:2023:RSV

[2078] Ben Cameron and Chính T. Hoàng.

A refinement on the structure of vertex-critical (P_{5}, gem)-free graphs. Theoretical Computer Science, 961 (??):??, June 15, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002499.

Du:2023:CBG

[2079] Liman Du, Wenguo Yang, and Suixiang Gao. Competition-based generalized self-profit maximization in dual-attribute network. Theoretical Computer Science, 961(??):??, June 15, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002530.

Bhore:2023:SEA

[2080] Sujoy Bhore and Rahul Jain. Spaceefficient algorithms for reachability in directed geometric graphs. Theoretical Computer Science, 961(??): ??, June 15, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002517.

Hermelin:2023:TIC

[2081] Danny Hermelin, Yuval Itzhaki, Hendrik Molter, and Rolf Niedermeier. Temporal interval cliques and independent sets. Theoretical Computer Science, 961(??):??, June 15, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523001986.

Fritz:2023:RMC

[2082] Tobias Fritz, Tomáš Gonda, Paolo Perrone, and Eigil Fjeldgren Rischel. Representable Markov categories and comparison of statistical experiments in categorical probability. Theoretical Computer Science, 961(??): ??, June 15, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002098.

Anonymous:2023:Jd

[2083] Anonymous. 22 June 2023. Theoretical Computer Science, 962(??):??, June 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBy

[2084] Anonymous. Editorial Board. Theoretical Computer Science, 962(??): ??, June 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002992.

Cheng:2023:OIS

[2085] Yunlong Cheng, Hao Zhou, Xiaofeng Gao, Jiaqi Zheng, and Guihai Chen. Optimizing incremental SDN upgrades for load balancing in ISP networks. Theoretical Computer Science, 962 (??):??, June 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002402.

Lang:2023:SSF

[2086] Rongling Lang, Hui Lei, and Junxue Zhang. On the saturation spectrum of families of cycle subdivisions. Theoretical Computer Science, 962(??):??, June 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002542.

DellaMonica:2023:ITL

[2087] Dario Della Monica, Angelo Montanari, and Pietro Sala. An interval temporal logic characterization of extended ω-regular languages. Theoretical Computer Science, 962 (??):??, June 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002426.

Zhang:2023:NLB

[2088] Pingyuan Zhang, Haining Yang, Liang Zhu, Yanhua Zhang, Hao Wang, and Qiuliang Xu. A new latticebased online/offline signatures framework for low-power devices. Theoretical Computer Science, 962(??): ??, June 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002554.

Garg:2023:CFE

[2089] Jugal Garg and Aniket Murhekar.

Computing fair and efficient allocations with few utility values. Theoretical Computer Science, 962(??): ??, June 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002451.

Goldberg:2023:LBQ

[2090] Paul W. Goldberg and Matthew J. Katzman. Lower bounds for the query complexity of equilibria in Lipschitz games. Theoretical Computer Science, 962(??):??, June 22, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300244X.

Zhang:2023:ASP

[2091] Meng Zhang and Yi Zhang. Arbitrarysize permutation networks using arbitrary-radix switches. Theoretical Computer Science, 962(??):??, June 22, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002529.

Anonymous:2023:Je

[2092] Anonymous. 26 June 2023. Theoretical Computer Science, 963(??):??, June 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBz

[2093] Anonymous. Editorial Board. Theoretical Computer Science, 963(??): ??, June 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003092.

Civril:2023:CNA

[2094] A. Çivril. Corrigendum to “A new approximation algorithm for the minimum 2-edge-connected spanning subgraph problem” [Theor. Comput. Sci. 943 (2023) 121–130]. Theoretical Computer Science, 963(??): ??, June 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002396. See

[1904].

Nickerl:2023:PNE

[2095] Julian Nickerl and Jacobo Torán. Pure Nash equilibria in a generalization of congestion games allowing resource failures. Theoretical Computer Science, 963(??):??, June 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002463.

Haraguchi:2023:PDE

[2096] Kazuya Haraguchi and Hiroshi Nagamochi. Polynomial-delay enumeration algorithms in set systems. Theoretical Computer Science, 963(??): ??, June 26, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002761.

Hashimoto:2023:CAT

[2097] Kenji Hashimoto and Sebastian Maneth. Characterizing attributed tree translations in terms of macro tree transducers. Theoretical Computer Science, 963(??):??, June 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397523002566.

Epstein:2023:KBP

[2098] Samuel Epstein. The Kolmogorov birthday paradox. Theoretical Computer Science, 963(??):??, June 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397523002773.

Ioannidis:2023:FNS

[2099] Stavros D. Ioannidis, Bart de Keijzer, and Carmine Ventre. Financial networks with singleton liability priorities. Theoretical Computer Science, 963(??):??, June 26, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002785.

Anonymous:2023:Jf

[2100] Anonymous. 10 July 2023. Theoretical Computer Science, 964(??):??, July 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBba

[2101] Anonymous. Editorial Board. Theoretical Computer Science, 964(??): ??, July 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003183.

Altisen:2023:SSS

[2102] Karine Altisen, Stéphane Devismes, Anaïs Durand, Colette Johnen, and Franck Petit. Self-stabilizing systems in spite of high dynamics. Theoretical Computer Science, 964(??): ??, July 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002797.

Lin:2023:LST

[2103] Bi-Shiang Lin, Kuo-Si Huang, and Chang-Biau Yang. Linear-space Stable algorithms for the longest common subsequence problem. Theoretical Computer Science, 964(??): ??, July 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002578.

Buno:2023:SSD

[2104] Kelvin Buño and Henry Adorna.

Solving 3-SAT in distributed P systems with string objects. Theoretical Computer Science, 964(??): ??, July 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300289X.

Anonymous:2023:Jg

[2105] Anonymous. 18 July 2023. Theoretical Computer Science, 965(??):??, July 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbb

[2106] Anonymous. Editorial Board. Theoretical Computer Science, 965(??): ??, July 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003341.

Huang:2023:MCV

[2107] Chien-Chung Huang and François Sellier. Matroid-constrained vertex cover. Theoretical Computer Science, 965 (??):??, July 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002906.

Feng:2023:SBP

[2108] Raymond Feng, Jesse Geneson, Andrew Lee, and Espen Slettnes. Sharp bounds on the price of bandit feedback for several models of mistakebounded online learning. Theoretical Computer Science, 965(??): ??, July 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002931.

Yang:2023:NSN

[2109] Qian Yang, Xin Xiong, Hong Peng, Jun Wang, and Xiaoxiao Song. Nonlinear spiking neural P systems with multiple channels. Theoretical Computer Science, 965(??): ??, July 18, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300292X.

Ipate:2023:MLB

[2110] Florentin Ipate, Ionuţ Niculescu, Raluca Lefticaru, Savas Konur, and Marian Gheorghe. A model learning based testing approach for kernel P systems. Theoretical Computer Science, 965(??):??, July 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002888.

Gazdag:2023:CPD

[2111] Zsolt Gazdag and Károly Hajagos. A characterisation of P by DLOGTIMEuniform families of polarizationless P systems using only dissolution rules. Theoretical Computer Science, 965 (??):??, July 18, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002876.

Anonymous:2023:Jh

[2112] Anonymous. 26 July 2023. Theoretical Computer Science, 966–967 (??):??, ???? 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbc

[2113] Anonymous. Editorial Board. Theoretical Computer Science, 966–967 (??):??, ???? 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003584.

Durand:2023:RAP

[2114] Anaïs Durand, Michel Raynal, and Gadi Taubenfeld. Reaching agreement in the presence of contentionrelated crash failures. Theoretical Computer Science, 966–967(??):

??, ????  2023.  CODEN TC-

SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002955.

Henning:2023:AAP

[2115] Michael A. Henning, Arti Pandey, and Vikash Tripathi. Algorithmic aspects of paired disjunctive domination in graphs. Theoretical Computer Science, 966–967(??):??, ???? 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003031.

Antonelli:2023:CPL

[2116] Melissa Antonelli, Ugo Dal Lago, and Paolo Pistone. On counting propositional logic and Wagner’s hierarchy. Theoretical Computer Science, 966–967(??):??, ???? 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002414.

Kutrib:2023:SCE

[2117] Martin Kutrib and Uwe Meyer. Syntax checking either way. Theoretical Computer Science, 966–967(??): ??, ???? 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003134.

Lidbetter:2023:SRG

[2118] Thomas Lidbetter and Yifan Xie. The search and rescue game on a cycle. Theoretical Computer Science, 966– 967(??):??, ???? 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003298.

Wang:2023:REC

[2119] Yihong Wang, Jianxi Fan, Yuejuan Han, Yan Wang, and Baolei Cheng. Reliability evaluation of complete graph-based recursive networks. Theoretical Computer Science, 966– 967(??):??, ???? 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003286.

Kutrib:2023:SCF

[2120] Martin Kutrib and Matthias Wendlandt. State complexity of finite partial languages. Theoretical Computer Science, 966–967(??): ??, ???? 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003146.

Anonymous:2023:Ad

[2121] Anonymous. 11 August 2023. Theoretical Computer Science, 968(??): ??, August 11, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbd

[2122] Anonymous. Editorial Board. Theoretical Computer Science, 968(??): ??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004000.

Chang:2023:TSN

[2123] Hong Chang, Jing Jin, Zhicheng Liu, Ping Li, and Xiaoyan Zhang. Twostage non-submodular maximization. Theoretical Computer Science, 968 (??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003304.

Lyon:2023:NSC

[2124] Oliver A. S. Lyon and Kai Salomaa. The nondeterministic state complexity of the site-directed deletion language operation. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003122.

Xu:2023:MKF

[2125] Kexin Xu, Benjamin Hong Meng Tan, Li-Ping Wang, Khin Mi Mi Aung, and Huaxiong Wang. Multi-key fully homomorphic encryption from NTRU and (R)LWE with faster bootstrapping. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003390.

Fuchs:2023:ANS

[2126] Michael Fuchs and Tsan-Cheng Yu. Asymptotic normality for the size of graph tries built from M-ary tree labelings. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003249.

Komander:2023:TDH

[2127] Dominique Komander and Carolin Rehs. Twin-distance-hereditary digraphs. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003225.

Botta:2023:DAZ

[2128] Vincenzo Botta and Ivan Visconti.

Doubly adaptive zero-knowledge proofs. Theoretical Computer Science, 968 (??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003274.

MacDonald:2023:RPH

[2129] Sophie MacDonald. The road problem and homomorphisms of directed graphs. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002943.

Priyamvada:2023:CAI

[2130] Priyamvada and B. S. Panda. Complexity and algorithms for injective edge coloring of graphs. Theoretical Computer Science, 968(??):??, August 11, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003237.

Bera:2023:APS

[2131] Somnath Bera, Atulya K. Nagar, Sastha Sriram, and K. G. Subramanian. An array P system based on a new variant of pure 2D context-free grammars. Theoretical Computer Science, 968(??): ??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003407.

Li:2023:SNP

[2132] Yanyan Li, Bosheng Song, and Xiangxiang Zeng. Spiking neural P systems with weights and delays on synapses. Theoretical Computer Science, 968 (??):??, August 11, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003419.

Anonymous:2023:Ae

[2133] Anonymous. 21 August 2023. Theoretical Computer Science, 969(??): ??, August 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbe

[2134] Anonymous. Editorial Board. Theoretical Computer Science, 969(??): ??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004152.

Montecchiani:2023:PCC

[2135] Fabrizio Montecchiani, Giacomo Ortali, Tommaso Piselli, and Alessandra Tappini. On the parameterized complexity of s-club cluster deletion problems. Theoretical Computer Science, 969(??):??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003547.

Ani:2023:TGC

[2136] Joshua Ani, Erik D. Demaine, Dylan H. Hendrickson, and Jayson Lynch. Trains, games, and complexity: 0/1/2-player motion planning through input/output gadgets. Theoretical Computer Science, 969 (??):??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300258X.

Dondi:2023:UTG

[2137] Riccardo Dondi. Untangling temporal graphs of bounded degree. Theoretical Computer Science, 969(??): ??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003535.

Elkind:2023:JGM

[2138] Edith Elkind, Piotr Faliszewski, Ayumi Igarashi, Pasin Manurangsi, Ulrike Schmidt-Kraepelin, and Warut Suksompong. Justifying groups in multiwinner approval voting. Theoretical Computer Science, 969(??):

??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003523.

Hospodar:2023:COS

[2139] Michal Hospodár and Viktor Olejár. The cut operation in subclasses of convex languages. Theoretical Computer Science, 969(??):??, August 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003638.

Lechenet:2023:ECA

[2140] Jean-Christophe Léchenet, Nikolai Kosmatov, and Pascale Le Gall. Efficient computation of arbitrary control dependencies. Theoretical Computer Science, 969(??):??, August 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003420.

Nourine:2023:HDI

[2141] Lhouari Nourine and Simon Vilmin. Hierarchical decompositions of implicational bases for the enumeration of meet-irreducible elements. Theoretical Computer Science, 969(??): ??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003432.

Kutrib:2023:IUF

[2142] Martin Kutrib, Andreas Malcher, Carlo Mereghetti, and Beatrice Palano. Iterated uniform finite-state transducers on unary languages. Theoretical Computer Science, 969(??): ??, August 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003626.

Anonymous:2023:Af

[2143] Anonymous. 29 August 2023. Theoretical Computer Science, 970(??): ??, August 29, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbf

[2144] Anonymous. Editorial Board. Theoretical Computer Science, 970(??): ??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004322.

Goldwurm:2023:LLL

[2145] Massimiliano Goldwurm, Jianyi Lin, and Marco Vignati. Local limit laws for symbol statistics in bicomponent rational models. Theoretical Computer Science, 970(??): ??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300364X.

Zhang:2023:EIS

[2146] Huanwen Zhang, Yan Wang, Jianxi Fan, and Chang Shu. Edgeindependent spanning trees in folded crossed cubes. Theoretical Computer Science, 970(??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523003663.

Reem:2023:PAS

[2147] Daniel Reem. The projector algorithm: a simple parallel algorithm for computing Voronoi diagrams and Delaunay graphs. Theoretical Computer Science, 970(??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003675.

Dev:2023:RBS

[2148] Subhadeep Ranjan Dev, Sanjana Dey, Florent Foucaud, Ralf Klasing, and Tuomo Lehtilä. The RED–BLUE SEPARATION problem on graphs. Theoretical Computer Science, 970 (??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003742.

Cheng:2023:ECA

[2149] Eddie Cheng, László Lipták, Ke Qiu, Zhizhang Shen, and Abhishek Vangipuram. On the g-extra connectivity of augmented cubes. Theoretical Computer Science, 970(??): ??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003729.

Liu:2023:BFS

[2150] Xiang Liu, Weiwei Wu, Wanyuan Wang, Yuhang Xu, Xiumin Wang, and Helei Cui. Budget-feasible Sybilproof mechanisms for crowdsensing. Theoretical Computer Science, 970 (??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523002918.

Liu:2023:NNL

[2151] Xiyu Liu, Yuzhen Zhao, and Liping Wang. Nonlinear neural-like P model for time series classification. Theoretical Computer Science, 970 (??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003687.

Alhazov:2023:SVR

[2152] Artiom Alhazov, Vincent Ferrari-Dominguez, Rudolf Freund, Nicolas Glade, and Sergiu Ivanov. A P systems variant for reasoning about sequential controllability of Boolean networks. Theoretical Computer Science, 970(??):??, August 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003699.

Anonymous:2023:Sa

[2153] Anonymous. 6 September 2023. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbg

[2154] Anonymous. Editorial Board. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004474.

Nan:2023:FTS

[2155] Lizhen Nan, Shiying Wang, and Lina Zhao. Fault-tolerant strong Menger connectivity of modified bubble-sort graphs. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397523003730.

Bianchini:2023:JLC

[2156] Riccardo Bianchini, Francesco Dagnino, Paola Giannini, and Elena Zucca. A Java-like calculus with heterogeneous coeffects. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397523003766.

Chen:2023:PMP

[2157] Lv Chen, Lingli Chen, and Qin Li. Practical multi-party quantum homomorphic encryption. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003808.

Martinez:2023:LBC

[2158] Virginia Ardévol Martínez, Marco Caoduro, Laurent Feuilloley, Jonathan Narboni, Pegah Pournajafi, and Jean-Florent Raymond. A lower bound for constant-size local certification. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300381X.

Chakraborty:2023:FGR

[2159] Dibyayan Chakraborty and Kshitij Gajjar. Finding geometric representations of apex graphs is NP-hard. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003778.

Kanellopoulos:2023:ASS

[2160] Panagiotis Kanellopoulos, Maria Kyropoulou, and Alexandros A. Voudouris. Not all strangers are the same: the impact of tolerance in Schelling games. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300378X.

Foerster:2023:GES

[2161] Klaus-Tycho Foerster, Tomasz Jurdziński, and Stefan Schmid. Guest editorial: Structural Information and Communication Complexity 2021. Theoretical Computer Science, 971(??):

??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523002268.

Kupferman:2023:LRS

[2162] Orna Kupferman and Asaf Petruschka. Lazy regular sensing. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003705.

Aman:2023:RRR

[2163] Bogdan Aman and Gabriel Ciobanu. Relating randomized right-hand sides to communicating rewriting rules. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003754.

Janett:2023:TDD

[2164] Duri Janett and Johannes Lengler. Two-dimensional drift analysis: Optimizing two functions simultaneously can be hard. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003857.

Doerr:2023:BED

[2165] Benjamin Doerr and Martin S. Krejca. Bivariate estimation-of-distribution algorithms can find an exponential number of optima. Theoretical Computer Science, 971(??):??, September 6, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003870.

Trinh:2023:TSB

[2166] Van-Giang Trinh, Belaid Benhamou, and Sylvain Soliman. Trap spaces of Boolean networks are conflict-free siphons of their Petri net encoding. Theoretical Computer Science, 971(??): ??, September 6, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003869.

Anonymous:2023:Sb

[2167] Anonymous. 13 September 2023. Theoretical Computer Science, 972(??):??, September 13, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbh

[2168] Anonymous. Editorial Board. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004619.

Pan:2023:SFT

[2169] Kaige Pan. Star fault tolerance of hypercube. Theoretical Computer Science, 972(??):??, September 13, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003651.

Duvignau:2023:SSB

[2170] Romaric Duvignau, Michel Raynal, and Elad Michael Schiller. Selfstabilizing Byzantine fault-tolerant repeated reliable broadcast. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003833.

Robert:2023:PZP

[2171] Léo Robert, Daiki Miyahara, Pascal Lafourcade, and Takaaki Mizuki. Physical ZKP protocols for Nurimisaki and Kurodoko. Theoretical Computer Science, 972(??):??, September 13, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003845.

Guo:2023:OSD

[2172] Sainan Guo, Ran Ma, Yuefang Sun, Xiaoyan Zhang, and Yong Zhang. Online scheduling with deterioration and unexpected processor breakdown. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003717.

Ghosal:2023:DRM

[2173] Pratik Ghosal, Adam Kunysz, and Katarzyna Paluch. The dynamics of rank-maximal and popular matchings. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003961.

Dey:2023:PG

[2174] Palash Dey. Priced gerrymandering.

Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003936.

Kim:2023:SCC

[2175] Sungmin Kim, Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. On Simon’s congruence closure of a string. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003912.

Hospodar:2023:NOC

[2176] Michal Hospodár and Viktor Olejár.

Nondeterministic operational complexity in subregular languages. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003882.

Ramirez-de-Arellano:2023:GCR

[2177] Antonio Ramírez de Arellano, David Orellana-Martín, and Mario J. Pérez-Jiménez. Generating, computing and recognizing with virus machines. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003900.

Konstantinidis:2023:ANU

[2178] Stavros Konstantinidis, Mitja Mastnak, Nelma Moreira, and Rogério Reis. Approximate NFA universality and related problems motivated by information theory. Theoretical Computer Science, 972(??): ??, September 13, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003894.

Anonymous:2023:Sc

[2179] Anonymous. 21 September 2023. Theoretical Computer Science, 973(??):??, September 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbi

[2180] Anonymous. Editorial Board. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004796.

deLiguoro:2023:STC

[2181] Ugo de’Liguoro and Riccardo Treglia. From semantics to types: the case of the imperative λ-calculus. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300395X.

Zhong:2023:UGA

[2182] Hao Zhong, Yong Tang, Qi Zhang, Ronghua Lin, and Weisheng Li. A unified greedy approximation for several dominating set problems. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003821.

Fujishige:2023:LTC

[2183] Yuta Fujishige, Yuki Tsujimaru, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Linear-time computation of DAWGs, symmetric indexing structures, and MAWs for integer alphabets. Theoretical Computer Science, 973(??):??, September 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004061.

Seiffarth:2023:MCS

[2184] Florian Seiffarth, Tamás Horváth, and Stefan Wrobel. Maximal closed set and half-space separations in finite closure systems. Theoretical Computer Science, 973(??):??, September 21, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004188.

Bang-Jensen:2023:CAC

[2185] Jørgen Bang-Jensen, Florian Hörsch, and Matthias Kriesell. Complexity of (arc)-connectivity problems involving arc-reversals or deorientations. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004103.

Rahman:2023:SIS

[2186] Md. Saidur Rahman, Petra Mutzel, and Slamin. Special issue on selected papers from the 16th International Conference and Workshops on Algorithms and Computation (WALCOM 2022). Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004085.

Zhu:2023:MSN

[2187] Mingzhe Zhu, Haoxin Sun, Wei Li, and Zhongzhi Zhang. Modeling spatial networks by contact graphs of disk packings. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523003791.

Delporte-Gallet:2023:OAS

[2188] Carole Delporte-Gallet, Hugues Fauconnier, Michel Raynal, and Mouna Safir. Optimal algorithms for synchronous Byzantine k-set agreement.

Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004115.

Ngondi:2023:SDH

[2189] Gerard Ekembe Ngondi. Semantics of dynamic hiding in mobile UTP-CSP. Theoretical Computer Science, 973(??): ??, September 21, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003924.

Anonymous:2023:Sd

[2190] Anonymous. 29 September 2023. Theoretical Computer Science, 974(??):??, September 29, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbj

[2191] Anonymous. Editorial Board. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005054.

Escardo:2023:HOG

[2192] Martín Escardó and Paulo Oliva.

Higher-order games with dependent types. Theoretical Computer Science, 974(??):??, September 29, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004243.

Inoue:2023:AHW

[2193] Yusuke Inoue, Kenji Hashimoto, and Hiroyuki Seki. An ambiguity hierarchy of weighted context-free grammars. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004255.

Cicerone:2023:VMV

[2194] Serafino Cicerone, Gabriele Di Stefano, Lara Drožek, Jaka Hedžet, Sandi Klavžar, and Ismael G. Yero. Variety of mutual-visibility problems in graphs. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004097.

Bilo:2023:CGP

[2195] Vittorio Bilò and Cosimo Vinci. Congestion games with prioritybased scheduling. Theoretical Computer Science, 974(??):??, September 29, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004073.

Lenhart:2023:MWG

[2196] William J. Lenhart and Giuseppe Liotta. Mutual witness Gabriel drawings of complete bipartite graphs. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004280.

Ancona:2023:CCS

[2197] Davide Ancona, Pietro Barbieri, and Elena Zucca. Checked corecursive streams: Expressivity and completeness. Theoretical Computer Science, 974(??):??, September 29, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523003948.

Broy:2023:SVC

[2198] Manfred Broy. Specification and verification of concurrent systems by causality and realizability. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300419X.

Fleischmann:2023:NUW

[2199] Pamela Fleischmann, Lukas Haschke, Jonas Höfer, Annika Huch, Annika Mayrock, and Dirk Nowotka. Nearly k-universal words — Investigating a part of Simon’s congruence. Theoretical Computer Science, 974(??): ??, September 29, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004267.

Anonymous:2023:Oa

[2200] Anonymous. 9 October 2023. Theoretical Computer Science, 975(??):

??, October 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbk

[2201] Anonymous. Editorial Board. Theoretical Computer Science, 975(??): ??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300525X.

Equi:2023:GCI

[2202] Massimo Equi, Veli Mäkinen, and Alexandru I. Tomescu. Graphs cannot be indexed in polynomial time for sub-quadratic time string matching, unless SETH fails. Theoretical Computer Science, 975(??):??, October 9, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004413.

Alpern:2023:FSG

[2203] Steve Alpern. The Faulty Satnav (GPS) Problem: Search for home in networks with unreliable directions. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300422X.

Bujtas:2023:CCA

[2204] Csilla Bujtás, Nima Ghanbari, and Sandi Klavžar. Computational complexity aspects of super domination. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004504.

DasGupta:2023:CDR

[2205] Bhaskar DasGupta, Elena Grigorescu, and Tamalika Mukherjee. On computing discretized Ricci curvatures of graphs: Local algorithms and (localized) fine-grained reductions. Theoretical Computer Science, 975(??): ??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004401.

He:2023:ACS

[2206] Yuqiao He, Guoliang Qiu, and Chihao Zhang. Approximability of the complementarily symmetric Holant problems on cubic graphs. Theoretical Computer Science, 975(??): ??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300453X.

Burkardt:2023:ILP

[2207] John Burkardt and Marcus R. Garvie.

An integer linear programming approach to solving the Eternity Puzzle. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004516.

Dong:2023:REH

[2208] Hui Dong, Mengjie Lv, Weibei Fan, and Guijuan Wang. Reliability evaluation of half hypercube networks. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004553.

Bai:2023:PAS

[2209] Tian Bai and Mingyu Xiao. A parameterized algorithm for subset feedback vertex set in tournaments. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004528.

Drabent:2023:RCA

[2210] Włodzimierz Drabent. A relaxed condition for avoiding the occur-check. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004206.

Brausse:2023:KCD

[2211] Franz Brauße, Konstantin Korovin, Margarita V. Korovina, and Norbert Th. Müller. The ksmt calculus is a δ-complete decision procedure for non-linear constraints. Theoretical Computer Science, 975(??): ??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004383.

Asahiro:2023:CCA

[2212] Yuichi Asahiro, Hiroshi Eto, Tesshu Hanaka, Guohui Lin, Eiji Miyano, and Ippei Terabaru. Corrigendum to “Complexity and approximability of the happy set problem” [Theor. Comput. Sci. 866 (2021) 123–144]. Theoretical Computer Science, 975 (??):??, October 9, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004279. See

[931].

Anonymous:2023:Ob

[2213] Anonymous. 17 October 2023. Theoretical Computer Science, 976(??): ??, October 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbl

[2214] Anonymous. Editorial Board. Theoretical Computer Science, 976(??): ??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300539X.

Faqeh:2023:CPD

[2215] Rasha Faqeh, André Martin, Valerio Schiavoni, Pramod Bhatotia, Pascal Felber, and Christof Fetzer. Capacity planning for dependable services. Theoretical Computer Science, 976(??):??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004395.

Darwich:2023:PTE

[2216] Omar Darwich, Ahmet-Sefa Ulucan, Quentin Bramas, Anissa Lamani, Anaïs Durand, and Pascal Lafourcade. Perpetual torus exploration by myopic luminous robots. Theoretical Computer Science, 976(??): ??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004565.

Guo:2023:CSC

[2217] Litao Guo and Gülnaz Boruzanlı Ekinci. Connectivity and super connectivity of folded hypercube-like networks. Theoretical Computer Science, 976(??):??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004644.

Li:2023:OIA

[2218] Weian Li, Qi Qi, Changjun Wang, and Changyuan Yu. Optimally integrating ad auction into e-commerce platforms. Theoretical Computer Science, 976(??):??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004541.

Abidha:2023:CDS

[2219] V. P. Abidha, Pradeesha Ashok, Avi Tomar, and Dolly Yadav. Colouring a dominating set without conflicts: q-Subset Square Colouring. Theoretical Computer Science, 976 (??):??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004735.

Magen:2023:MGA

[2220] Roey Magen and Moni Naor. Mirror games against an open book player. Theoretical Computer Science, 976 (??):??, October 17, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004723.

Gezer:2023:RTC

[2221] M. Utkan Gezer, Özdeniz Dolu, Nevzat Ersoy, and A. C. Cem Say. Real-time, constant-space, constantrandomness verifiers. Theoretical Computer Science, 976(??):??, October 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004681.

DArco:2023:MSP

[2222] Paolo D’Arco, Zahra Ebadi Ansaroudi, and Francesco Mogavero. Multistage proof-of-works: Properties and vulnerabilities. Theoretical Computer Science, 976(??):??, October 17, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004218.

Anonymous:2023:Oc

[2223] Anonymous. 25 October 2023. Theoretical Computer Science, 977(??): ??, October 25, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbm

[2224] Anonymous. Editorial Board. Theoretical Computer Science, 977(??): ??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005480.

Bein:2023:BCB

[2225] Wolfgang Bein and Lawrence L. Larmore. Breaking the 2-competitiveness barrier for two servers in a tree. Theoretical Computer Science, 977 (??):??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004656.

Shanjani:2023:CCG

[2226] Sima Hajiaghaei Shanjani and Valerie King. Communication costs in a geometric communication network. Theoretical Computer Science, 977 (??):??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004747.

Zhuang:2023:MCC

[2227] Hongbin Zhuang, Wanling Lin, Xiao-Yan Li, and Jou-Ming Chang. Matroidal connectivity and conditional matroidal connectivity of star graphs. Theoretical Computer Science, 977 (??):??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004863.

Bramas:2023:OEP

[2228] Quentin Bramas, Pascal Lafourcade, and Stéphane Devismes. Optimal exclusive perpetual grid exploration by luminous myopic opaque robots with common chirality. Theoretical Computer Science, 977(??): ??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004759.

Albayrak:2023:QES

[2229] Seda Albayrak and Jason P. Bell.

Quantitative estimates for the size of an intersection of sparse automatic sets. Theoretical Computer Science, 977(??):??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004577.

Neraud:2023:LAG

[2230] Jean Néraud. Loopless algorithms to generate maximum length gray cycles wrt. k-character substitutions. Theoretical Computer Science, 977 (??):??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004887.

Castellano:2023:WWG

[2231] Ilaria Castellano, Anna Giordano Bruno, and Nicolò Zava. Weakly weighted generalised quasi-metric spaces and semilattices. Theoretical Computer Science, 977(??):??, October 25, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004425.

Diaz-Caro:2023:EPP

[2232] Alejandro Díaz-Caro and Gilles Dowek. Extensional proofs in a propositional logic modulo isomorphisms. Theoretical Computer Science, 977(??): ??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004851.

Cairoli:2023:GAM

[2233] Francesca Cairoli, Fabio Anselmi, Alberto d’Onofrio, and Luca Bortolussi. Generative abstraction of Markov population processes. Theoretical Computer Science, 977(??): ??, October 25, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004826.

Anonymous:2023:Na

[2234] Anonymous. 2 November 2023. Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbn

[2235] Anonymous. Editorial Board. Theoretical Computer Science, 978(??): ??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300556X.

Albouy:2023:ABR

[2236] Timothé Albouy, Davide Frey, Michel Raynal, and François Taïani. Asynchronous Byzantine reliable broadcast with a message adversary. Theoretical Computer Science, 978(??): ??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004231.

Horsch:2023:OMT

[2237] Florian Hörsch. On orientations maximizing total arc-connectivity. Theoretical Computer Science, 978(??): ??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004899.

Dahiya:2023:SDT

[2238] Yogesh Dahiya and Meena Mahajan.

On (simple) decision tree rank. Theoretical Computer Science, 978(??): ??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004905.

Galby:2023:PCM

[2239] Esther Galby, Dániel Marx, Philipp Schepper, Roohani Sharma, and Prafullkumar Tale. Parameterized complexity of multicut in weighted trees. Theoretical Computer Science, 978 (??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004875.

Blumrosen:2023:HBM

[2240] Liad Blumrosen and Yehonatan Mizrahi. How bad is the merger paradox? Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300470X.

Ito:2023:SBW

[2241] Takehiro Ito, Jun Kawahara, Shin ichi Minato, Yota Otachi, Toshiki Saitoh, Akira Suzuki, Ryuhei Uehara, Takeaki Uno, Katsuhisa Yamanaka, and Ryo Yoshinaka. Sorting balls and water: Equivalence and computational complexity. Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004711.

Busch:2023:FST

[2242] Costas Busch, Bogdan S. Chlebus, Maurice Herlihy, Miroslav Popovic, Pavan Poudel, and Gokarna Sharma. Flexible scheduling of transactional memory on trees. Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523004978.

Bonato:2023:OVL

[2243] Anthony Bonato, Trent G. Marbach, Michael Molnar, and JD Nir. The one-visibility localization game. Theoretical Computer Science, 978(??): ??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004991.

Pramanick:2023:DAF

[2244] Subhajit Pramanick, Sai Vamshi Samala, Debasish Pattanayak, and Partha Sarathi Mandal. Distributed algorithms for filling MIS vertices of an arbitrary graph by myopic luminous robots. Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005005.

Alzamel:2023:MDP

[2245] Mai Alzamel, Christopher Hampson, Costas S. Iliopoulos, Zara Lim, Solon Pissis, Dimitrios Vlachakis, and Steven Watts. Maximal degenerate palindromes with gaps and mismatches. Theoretical Computer Science, 978 (??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004954.

Omeradzic:2023:PAM

[2246] Amir Omeradzic and Hans-Georg Beyer. Progress analysis of a multirecombinative evolution strategy on the highly multimodal Rastrigin function. Theoretical Computer Science, 978(??):??, November 2, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004929.

Anonymous:2023:Nb

[2247] Anonymous. 10 November 2023. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbo

[2248] Anonymous. Editorial Board. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005765.

Oostveen:2023:SDP

[2249] Jelle J. Oostveen and Erik Jan van Leeuwen. Streaming deletion problems parameterized by vertex cover. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523004917.

Karia:2023:HHS

[2250] Neel Karia, Faraaz Mallick, and Palash Dey. How hard is safe bribery? Theoretical Computer Science, 979(??):

??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004693.

Bousquet:2023:FVS

[2251] Nicolas Bousquet, Felix Hommelsheim, Yusuke Kobayashi, Moritz Mühlenthaler, and Akira Suzuki. Feedback vertex set reconfiguration in planar graphs. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005017.

Martynova:2023:HIH

[2252] Olga Martynova and Alexander Okhotin.

Homomorphisms and inverse homomorphisms on graph-walking automata. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523005108.

Mitani:2023:FTL

[2253] Kazuki Mitani, Takuya Mieno, Kazuhisa Seto, and Takashi Horiyama. Finding top-k longest palindromes in substrings. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523004966.

Cai:2023:GME

[2254] Jianping Cai, Ximeng Liu, Jiayin Li, and Shuangyue Zhang. Generation matrix: an embeddable matrix representation for hierarchical trees. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004930.

Ito:2023:RAU

[2255] Takehiro Ito, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, Yuta Nozaki, Yoshio Okamoto, and Kenta Ozeki. On reachable assignments under dichotomous preferences. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005091.

Geneson:2023:OLS

[2256] Jesse Geneson and Ethan Zhou. Online learning of smooth functions. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005169.

Jaykrishnan:2023:EDS

[2257] G. Jaykrishnan and Asaf Levin. EP-TAS for the dual of splittable bin packing with cardinality constraint. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005157.

Gottlob:2023:FCH

[2258] Georg Gottlob, Matthias Lanzinger, Reinhard Pichler, and Igor Razgon. Fractional covers of hypergraphs with bounded multi-intersection. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005170.

Liu:2023:QBS

[2259] Xujun Liu, Olgica Milenkovic, and George V. Moustakides. Querybased selection of optimal candidates under the Mallows model. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005194.

Chan:2023:FLG

[2260] Hau Chan, Zifan Gong, Minming Li, Chenhao Wang, and Yingchao Zhao. Facility location games with ordinal preferences. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005212.

Feng:2023:BCA

[2261] Junkai Feng, Ruiqi Yang, Haibin Zhang, and Zhenning Zhang. A bi-criteria algorithm for online nonmonotone maximization problems: DR-submodular+concave. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005200.

Chakraborty:2023:ACG

[2262] Dibyayan Chakraborty, Harmender Gahlawat, and Bodhayan Roy. Algorithms and complexity for geodetic sets on partial grids. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005303.

Drewes:2023:HTA

[2263] Frank Drewes, Richard Mörbitz, and Heiko Vogler. Hybrid tree automata and the yield theorem for constituent tree automata. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300498X.

Okumura:2023:OAR

[2264] Takashi Okumura, Koichi Wada, and Xavier Défago. Optimal L-algorithms for rendezvous of asynchronous mobile robots with external-lights. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300511X.

He:2023:IAB

[2265] Yuchen He and Chihao Zhang. Improved algorithms for bandit with graph feedback via regret decomposition. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523005133.

Fouche:2023:LTM

[2266] Willem L. Fouché and Safari Mukeru.

On local times of Martin–Löf random Brownian motion. Theoretical Computer Science, 979(??):??, November 10, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005121.

Kaufmann:2023:SAP

[2267] Marc Kaufmann, Maxime Larcher, Johannes Lengler, and Xun Zou. Selfadjusting population sizes for the (1, λ)-EA on monotone functions. Theoretical Computer Science, 979(??): ??, November 10, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523004942.

Anonymous:2023:Nc

[2268] Anonymous. 20 November 2023. Theoretical Computer Science, 980(??):??, November 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2023:EBbp

[2269] Anonymous. Editorial Board. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005984.

Cohen:2023:LAO

[2270] Alon Cohen, Argyrios Deligkas, and Moran Koren. Learning approximately optimal contracts. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005327.

Mandal:2023:ELE

[2271] Subhrangsu Mandal, Anisur Rahaman Molla, and William K. Moses. Efficient live exploration of a dynamic ring with mobile robots. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005145.

Zhou:2023:OSE

[2272] Xiaotian Zhou, Haoxin Sun, Wei Li, and Zhongzhi Zhang. Optimization on the smallest eigenvalue of grounded Laplacian matrix via edge addition. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005339.

Carvalho:2023:ACF

[2273] André Carvalho. Algebraic and context-free subsets of subgroups. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300542X.

Goldberg:2023:PCA

[2274] Paul W. Goldberg and Matthew Katzman. PPAD-complete approximate pure Nash equilibria in Lipschitz games. Theoretical Computer Science, 980(??):??, November 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005315.

Li:2023:SDH

[2275] Ping Li, Shurong Zhang, Xiaomin Hu, and Weihua Yang. Structural diagnosability of hypercubes under the PMC and MM* models. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005443.

Edalat:2023:RSI

[2276] Abbas Edalat, Amin Farjudian, and Yiran Li. Recursive solution of initial value problems with temporal discretization. Theoretical Computer Science, 980(??):??, November 20, 2023. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005340.

Hunter:2023:FAD

[2277] Ethan Hunter, Jessica Enright, and Alice Miller. Feasibility assessments of a dynamical approach to compartmental modelling on graphs: Scaling limits and performance analysis. Theoretical Computer Science, 980(??): ??, November 20, 2023. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005601.

Anonymous:2024:Ja

[2278] Anonymous. 4 January 2024. Theoretical Computer Science, 981(??): ??, January 4, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBa

[2279] Anonymous. Editorial Board. Theoretical Computer Science, 981(??): ??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006138.

Bullinger:2024:TDG

[2280] Martin Bullinger and Warut Suksompong. Topological distance games. Theoretical Computer Science, 981 (??):??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005510.

Caragiannis:2024:RMI

[2281] Ioannis Caragiannis and Shivika Narang. Repeatedly matching items to agents fairly and efficiently. Theoretical Computer Science, 981(??): ??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005595.

Chen:2024:SFA

[2282] Shengminjie Chen, Donglei Du, Ruiqi Yang, Wenguo Yang, and Yapu Zhang. A single factor approximation ratio algorithm for DR-submodular maximization on integer lattice beyond nonnegativity and monotonicity. Theoretical Computer Science, 981(??): ??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005674.

Huang:2024:QAL

[2283] Xiaowei Huang, Shihao Zhang, and Lvzhou Li. Quantum algorithms for learning hidden strings with applications to matroid problems. Theoretical Computer Science, 981(??): ??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005686.

Zhou:2024:REC

[2284] Qianru Zhou, Baolei Cheng, Jingya Zhou, Jia Yu, Yan Wang, and Jianxi Fan. Reliability evaluation for a class of recursive match networks.

Theoretical Computer Science, 981 (??):??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005662.

Wang:2024:OSE

[2285] Liguan Wang, Yuan Li, and Haibin Kan. Obtaining simulation extractable NIZKs in the updatable CRS model generically. Theoretical Computer Science, 981(??):??, January 4, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005650.

Amir:2024:RPS

[2286] Amihood Amir, Eitan Kondratovsky, Gad M. Landau, Shoshana Marcus, and Dina Sokol. Reconstructing parameterized strings from parameterized suffix and LCP arrays. Theoretical Computer Science, 981 (??):??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005431.

delaCruz:2024:STH

[2287] Ren Tristan A. de la Cruz, Francis George C. Cabarle, and Henry N. Adorna. Steps toward a homogenization procedure for spiking neural P systems. Theoretical Computer Science, 981(??):??, January 4, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005637.

Anonymous:2024:Jb

[2288] Anonymous. 8 January 2024. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBb

[2289] Anonymous. Editorial Board. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006230.

Tanaka:2024:DAH

[2290] Yuzuru Tanaka. Designing aperture 2^{2k} hexagonal grids and their indexing as factor rings of Eisenstein integers. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005704.

Rizzo:2024:EFG

[2291] Nicola Rizzo, Massimo Equi, Tuukka Norri, and Veli Mäkinen. Elastic founder graphs improved and enhanced. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005820.

Zhao:2024:RGN

[2292] Xiaojun Zhao, Qingying Deng, and Xiaowang Li. Relating g-good-neighbor connectivity and g-good-neighbor diagnosability of strong digraph network. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005716.

Wang:2024:GCN

[2293] Di Wang and Jinhui Xu. Gradient complexity and non-stationary views of differentially private empirical risk minimization. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005728.

Huang:2024:OBD

[2294] Sumin Huang and Sergey Kitaev. On ordering of β-description trees. Theoretical Computer Science, 982 (??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005868.

Gourves:2024:FCV

[2295] Laurent Gourvès, Ararat Harutyunyan, Michael Lampis, and Nikolaos Melissinos. Filling crosswords is very hard. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005881.

Duchene:2024:BII

[2296] Eric Duchêne, Nacim Oijid, and Aline Parreau. Bipartite instances of INFLU-ENCE. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300587X.

Henning:2024:AHR

[2297] Michael A. Henning, Arti Pandey, Gopika Sharma, and Vikash Tripathi. Algorithms and hardness results for edge total domination problem in graphs. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005832.

Vu:2024:RMS

[2298] Hoa T. Vu. Revisiting maximum satisfiability and related problems in data streams. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005844.

Yang:2024:CCH

[2299] Peng Yang, Yuan Huang, and Zhiguo Fu. The computational complexity of Holant problems on 3-regular graphs. Theoretical Computer Science, 982 (??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005698.

Bachtler:2024:ADP

[2300] Oliver Bachtler, Tim Bergner, and Sven O. Krumke. Almost disjoint paths and separating by forbidden pairs. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005856.

Gai:2024:SPP

[2301] Ling Gai, Weiwei Zhang, and Zhao Zhang. Selfish bin packing with punishment. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005893.

Jain:2024:SEA

[2302] Rahul Jain and Raghunath Tewari. Space efficient algorithm for solving reachability using tree decomposition and separators. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005649.

Chang:2024:ECD

[2303] Yi-Jun Chang. The energy complexity of diameter and minimum cut computation in bounded-genus networks. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005923.

Dahal:2024:DHI

[2304] Sameep Dahal and Jukka Suomela.

Distributed half-integral matching and beyond. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523005911.

Kiyak:2024:ECR

[2305] Fırat Kıyak and A. C. Cem Say. Energy complexity of regular languages. Theoretical Computer Science, 982 (??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006047.

Bekos:2024:CGD

[2306] Michael A. Bekos, Martin Gronemann, Fabrizio Montecchiani, and Antonios Symvonis. Convex grid drawings of planar graphs with constant edge-vertex resolution. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006035.

Nasre:2024:PCM

[2307] Meghana Nasre, Prajakta Nimbhorkar, Keshav Ranjan, and Ankita Sarkar. Popular critical matchings in the many-to-many setting. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005947.

Castelnovo:2024:SCA

[2308] Davide Castelnovo, Fabio Gadducci, and Marino Miculan. A simple criterion for M, N -adhesivity. Theoretical Computer Science, 982(??): ??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005935.

Remy:2024:MBF

[2309] Élisabeth Remy and Paul Ruet. From multivalued to Boolean functions: Preservation of soft nested canalization. Theoretical Computer Science, 982(??):??, January 8, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300590X.

Anonymous:2024:Fa

[2310] Anonymous. 1 February 2024. Theoretical Computer Science, 983(??): ??, February 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBc

[2311] Anonymous. Editorial Board. Theoretical Computer Science, 983(??): ??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006497.

Berend:2024:TGH

[2312] Daniel Berend, Liat Cohen, and Omrit Filtser. A tour of general Hanoi graphs. Theoretical Computer Science, 983(??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006023.

Bruse:2024:WFL

[2313] Florian Bruse, Maurice Herwig, and Martin Lange. Weights of formal languages based on geometric series with an application to automatic grading. Theoretical Computer Science, 983 (??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006084.

Kanzaki:2024:CCJ

[2314] Masaaki Kanzaki, Yota Otachi, Giovanni Viglietta, and Ryuhei Uehara. Computational complexity of jumping block puzzles. Theoretical Computer Science, 983(??): ??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006059.

Jana:2024:PSC

[2315] Satyabrata Jana, Souvik Saha, Abhishek Sahu, Saket Saurabh, and Shaily Verma. Partitioning subclasses of chordal graphs with few deletions. Theoretical Computer Science, 983 (??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006011.

Chen:2024:AAD

[2316] Xianrun Chen, Sai Ji, Chenchen Wu, Yicheng Xu, and Yang Yang. An approximation algorithm for diversityaware fair k-supplier problem. Theoretical Computer Science, 983(??): ??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006187.

Berend:2024:PAV

[2317] D. Berend and S. Mamana. A probabilistic algorithm for vertex cover. Theoretical Computer Science, 983 (??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006199.

Wang:2024:USS

[2318] Xiwen Wang, Kai Zhang, Junqing Gong, Shi-Feng Sun, and Jianting Ning. Updatable searchable symmetric encryption: Definitions and constructions. Theoretical Computer Science, 983(??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006175.

Mirjalali:2024:MPS

[2319] Kian Mirjalali and Hamid Zarrabi-Zadeh. Massively parallel and streaming algorithms for balanced clustering. Theoretical Computer Science, 983 (??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006096.

Maruyama:2024:SSM

[2320] Syohei Maruyama, Yuichi Sudo, Sayaka Kamei, and Hirotsugu Kakugawa. Selfstabilizing 2-minimal dominating set algorithms based on loop composition. Theoretical Computer Science, 983(??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006278.

Greugny:2024:SMM

[2321] Eléa Thibault Greugny, François Fages, Ovidiu Radulescu, Peter Szmolyan, and Georgios N. Stamatas. A skin microbiome model with AMP interactions and analysis of quasi-stability vs stability in population dynamics. Theoretical Computer Science, 983 (??):??, February 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006072.

Anonymous:2024:Fb

[2322] Anonymous. 12 February 2024. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBd

[2323] Anonymous. Editorial Board. Theoretical Computer Science, 984(??): ??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006722.

Boyar:2024:ACA

[2324] Joan Boyar, Kim S. Larsen, and Denis Pankratov. Advice complexity of adaptive priority algorithms. Theoretical Computer Science, 984(??): ??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300631X.

Kakimura:2024:RCB

[2325] Naonori Kakimura and Riku Nitta. Randomized counter-based algorithms for frequency estimation over data streams in O(log log N) space. Theoretical Computer Science, 984(??): ??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006308.

Parshina:2024:FIC

[2326] Olga Parshina and Svetlana Puzynina. Finite and infinite closed-rich words. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752300628X.

Tsur:2024:ACC

[2327] Dekel Tsur. Algorithms for 2-club cluster deletion problems using automated generation of branching rules. Theoretical Computer Science, 984 (??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006345.

Tang:2024:GEA

[2328] Zhongzheng Tang, Jingwen Chen, and Chenhao Wang. Greedy+Singleton: an efficient approximation algorithm for k-submodular knapsack maximization. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006333.

Bhagat:2024:DRI

[2329] Subhash Bhagat and Andrzej Pelc.

Deterministic rendezvous in infinite trees. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006266.

Sun:2024:IDA

[2330] Xiaoming Sun, Jialin Zhang, Shuo Zhang, and Zhijie Zhang. Improved deterministic algorithms for non-monotone submodular maximization. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006060.

Bai:2024:EAR

[2331] Tian Bai and Mingyu Xiao. Exact algorithms for restricted subset feedback vertex set in chordal and split graphs. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006394.

Almalki:2024:GSC

[2332] Nada Almalki and Othon Michail. On geometric shape construction via growth operations. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006370.

Hermelin:2024:ASQ

[2333] Danny Hermelin, Leon Kellerhals, Rolf Niedermeier, and Rami Pugatch. Approximating sparse quadratic programs. Theoretical Computer Science, 984(??):??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006321.

Ascone:2024:FPA

[2334] Rocco Ascone, Giulia Bernardini, and Luca Manzoni. Fixed points and attractors of reactantless and inhibitorless reaction systems. Theoretical Computer Science, 984(??): ??, February 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006357.

Anonymous:2024:Fc

[2335] Anonymous. 21 February 2024. Theoretical Computer Science, 985(??):??, February 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBe

[2336] Anonymous. Editorial Board. Theoretical Computer Science, 985(??): ??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300693X.

Poovanandran:2024:CSC

[2337] Ghajendran Poovanandran, Jamie Simpson, and Wen Chean Teh. Counting subwords in circular words and their Parikh matrices. Theoretical Computer Science, 985(??):??, February 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006576.

Rabinovich:2024:ORM

[2338] Dmitry Rabinovich, Michael Amir, and Alfred M. Bruckstein. Optimally reordering mobile agents on parallel rows. Theoretical Computer Science, 985(??):??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006436.

Wu:2024:AAF

[2339] Di Wu, Qilong Feng, and Jianxin Wang. Approximation algorithms for fair k-median problem without fairness violation. Theoretical Computer Science, 985(??):??, February 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300645X.

Ma:2024:TMT

[2340] Hengzhao Ma, Jianzhong Li, Tianpeng Gao, and Xiangyu Gao. Turing machines with two-level memory: New computational models for analyzing the input/output complexity. Theoretical Computer Science, 985 (??):??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006606.

Lan:2024:FSS

[2341] Yan Lan, Yuan Yuan, Yinling Wang, Xin Han, and Yong Zhou. Flow shop scheduling problems with transportation constraints revisited. Theoretical Computer Science, 985(??): ??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300662X.

Coy:2024:RSH

[2342] Sam Coy, Artur Czumaj, Christian Scheideler, Philipp Schneider, and Julian Werthmann. Routing schemes for hybrid communication networks. Theoretical Computer Science, 985 (??):??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006655.

Zhang:2024:LPS

[2343] Shenxing Zhang. On the linearity of the periods of subtraction games. Theoretical Computer Science, 985 (??):??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006631.

Lei:2024:LUL

[2344] Qin Lei, Hong Jin, Jia Lee, and Jiang Zhong. Land use and land cover change simulation enhanced by asynchronous communicating cellular automata. Theoretical Computer Science, 985(??):??, February 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006448.

Anonymous:2024:F

[2345] Anonymous. 27 February 2024. Theoretical Computer Science, 986(??):??, February 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBf

[2346] Anonymous. Editorial Board. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000094.

Tu:2024:FTP

[2347] Binbin Tu, Yu Chen, Hongrui Cui, and Xianfang Wang. Fast twoparty signature for upgrading ECDSA to two-party scenario easily. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006382.

Zhang:2024:IAP

[2348] Xiaojin Zhang. Improved algorithm for permutation testing. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006291. See

corrigendum [2488].

Wang:2024:FEO

[2349] Yong Wang. Finding the edges in optimal Hamiltonian cycles based on frequency quadrilaterals. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006369.

Chan:2024:MMG

[2350] T. H. Hubert Chan, Zhihao Gavin Tang, and Quan Xue. Max-min greedy matching problem: Hardness for the adversary and fractional variant. Theoretical Computer Science, 986 (??):??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006424.

Dvorak:2024:PDP

[2351] Tomáš Dvořák and Mei-Mei Gu. Paired 2-disjoint path covers of burnt pancake graphs with faulty elements. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006400.

Yin:2024:DAC

[2352] Wen Yin, Jiarong Liang, Yuhao Xie, and Changzhen Li. The t/kdiagnosability of m-ary n-cube networks. Theoretical Computer Science, 986(??):??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006588.

Lu:2024:BCG

[2353] Xiao-Tao Lü, Jin Chen, Zhi-Xiong Wen, and Wen Wu. On the 2binomial complexity of the generalized Thue–Morse words. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006552.

Dvorakova:2024:SAE

[2354] L’ubomíra Dvořáková. String attractors of episturmian sequences. Theoretical Computer Science, 986(??): ??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006540.

Li:2024:NLC

[2355] Shuo Li. A note on the Lie complexity and beyond. Theoretical Computer Science, 986(??):??, February 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006564.

Berthomieu:2024:SNT

[2356] Bernard Berthomieu and Dmitry A. Zaitsev. Sleptsov nets are Turingcomplete. Theoretical Computer Science, 986(??):??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752300659X.

Liu:2024:EFT

[2357] Xiaoqing Liu, Baolei Cheng, Yan Wang, Jia Yu, and Jianxi Fan. Enhancing fault tolerance of balanced hypercube networks by the edge partition method. Theoretical Computer Science, 986(??):??, February 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006539.

Gerontitis:2024:IZN

[2358] Dimitrios Gerontitis, Changxin Mo, Predrag S. Stanimirović, and Vasilios N. Katsikis. Improved zeroing neural models based on two novel activation functions with exponential behavior. Theoretical Computer Science, 986 (??):??, February 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006412.

Anonymous:2024:Ma

[2359] Anonymous. 1 March 2024. Theoretical Computer Science, 987(??):??, March 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBg

[2360] Anonymous. Editorial Board. Theoretical Computer Science, 987(??): ??, March 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000288.

Nutov:2024:NCN

[2361] Zeev Nutov. 2-node-connectivity network design. Theoretical Computer Science, 987(??):??, March 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397523006801.

Luo:2024:FCR

[2362] Fucai Luo, Saif Al-Kuwari, Haiyan Wang, and Xingfu Yan. Fully collusion resistant trace-and-revoke functional encryption for arbitrary identities. Theoretical Computer Science, 987(??):??, March 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006813.

Gao:2024:ASA

[2363] Tianpeng Gao and Jianzhong Li. Approximate sorting and its applications in I/O model. Theoretical Computer Science, 987(??): ??, March 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006618.

Zhang:2024:SBO

[2364] Kuize Zhang. State-based opacity of labeled real-time automata. Theoretical Computer Science, 987 (??):??, March 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006862.

Lv:2024:ACF

[2365] Yali Lv, Cheng-Kuan Lin, D. Frank Hsu, and Jianxi Fan. An algorithm for conditional-fault local diagnosis of multiprocessor systems under the MM^{⋆} model. Theoretical Computer Science, 987(??):??, March 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006850.

Grandjean:2024:IDL

[2366] Étienne Grandjean, Théo Grente, and Véronique Terrier. Inductive definitions in logic versus programs of real-time cellular automata. Theoretical Computer Science, 987(??): ??, March 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006680.

Lakshmi:2024:PCA

[2367] M. Gayathri Lakshmi, G. Arul Freeda Vinodhini, T. Kamaraj, and D. G. Thomas. Parallel contextual array insertion deletion grammars, pure 2D context-free grammars and associated P systems. Theoretical Computer Science, 987(??):??, March 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006643.

Anonymous:2024:Mb

[2368] Anonymous. 12 March 2024. Theoretical Computer Science, 988(??): ??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBh

[2369] Anonymous. Editorial Board. Theoretical Computer Science, 988(??): ??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000410.

Wu:2024:ETR

[2370] Yutong Wu, Ali Khodabakhsh, Bo Li, Evdokia Nikolova, and Emmanouil Pountourakis. Eliciting truthful reports with partial signals in repeated games. Theoretical Computer Science, 988(??):??, March 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006849.

V:2024:GPA

[2371] Ullas Chandran S. V., Sandi Klavžar, Neethu P. K., and Rudini Sampaio. The general position avoidance game and hardness of general position games. Theoretical Computer Science, 988 (??):??, March 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006837.

Azais:2024:DCS

[2372] Romain Azaïs and Florian Ingels. Detection of common subtrees with identical label distribution. Theoretical Computer Science, 988(??): ??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006795.

Qin:2024:DMC

[2373] Xiao-Wen Qin, Rong-Xia Hao, and Sheng-Lung Peng. Diagnosability of multigraph composition networks. Theoretical Computer Science, 988 (??):??, March 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006886.

Skoviera:2024:DWF

[2374] Martin Škoviera and Peter Varša. Deciding whether four perfect matchings can cover the edges of a snark is NP-complete. Theoretical Computer Science, 988(??):??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006874.

Rass:2024:TMW

[2375] Stefan Rass. Tell me who you are friends with and I will tell you who you are: Unique neighborhoods in random graphs. Theoretical Computer Science, 988(??):??, March 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400001X.

Zhou:2024:NAA

[2376] Aizhong Zhou, Haodi Feng, Jiong Guo, Haitao Jiang, Nan Liu, Binhai Zhu, and Daming Zhu. New approximation algorithms for RNA secondary structures prediction problems by local search. Theoretical Computer Science, 988(??):??, March 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006679.

Shallit:2024:PPS

[2377] Jeffrey Shallit. Proving properties of some greedily-defined integer recurrences via automata theory. Theoretical Computer Science, 988(??): ??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752300676X.

Zhang:2024:LGN

[2378] Luping Zhang and Fei Xu. Languages generated by numerical P systems with thresholds. Theoretical Computer Science, 988(??):??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006898.

Kumar:2024:CSP

[2379] Narendra Kumar, Siddhartha Siddhiprada Bhoi, and Abhay Kumar Singh. Corrigendum to “A study of primer design with wconstacyclic shift over F_{4} ” [Theor. Comp. Sci. 960 (2023) 113925]. Theoretical Computer Science, 988(??): ??, March 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523005182. See

[2073].

Anonymous:2024:Mc

[2380] Anonymous. 21 March 2024. Theoretical Computer Science, 989(??): ??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBi

[2381] Anonymous. Editorial Board. Theoretical Computer Science, 989(??): ??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000537.

Schmid:2024:WPS

[2382] Stefan Schmid, Jakub Svoboda, and Michelle Yeo. Weighted packet selection for rechargeable links in cryptocurrency networks: Complexity and approximation. Theoretical Computer Science, 989(??):??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397523006667.

Lee:2024:GSL

[2383] Keenan Lee and André van Renssen.

Generalized sweeping line spanners. Theoretical Computer Science, 989 (??):??, March 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000057.

Lundstrom:2024:SSI

[2384] Oskar Lundström, Michel Raynal, and Elad M. Schiller. Self-stabilizing indulgent zero-degrading binary consensus. Theoretical Computer Science, 989(??):??, March 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000021.

Dai:2024:EEP

[2385] Sijia Dai, Xinru Guo, Huahua Miao, Guichen Gao, Yicheng Xu, and Yong Zhang. The existence and efficiency of PMMS allocations. Theoretical Computer Science, 989(??): ??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000033.

Gaikwad:2024:OAG

[2386] Ajinkya Gaikwad and Soumen Maity. Offensive alliances in graphs. Theoretical Computer Science, 989(??): ??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000161.

Hamada:2024:RCC

[2387] Koki Hamada and Shuichi Miyazaki. Refined computational complexities of Hospitals/Residents problem with regional caps. Theoretical Computer Science, 989(??):??, March 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397524000045.

Lobe:2024:MEB

[2388] Elisabeth Lobe and Annette Lutz. Minor embedding in broken chimera and derived graphs is NP-complete. Theoretical Computer Science, 989 (??):??, March 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397523006825.

Ding:2024:NSD

[2389] Cunsheng Ding. New support 5designs from lifted linear codes. Theoretical Computer Science, 989(??): ??, March 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400015X.

Anonymous:2024:Aa

[2390] Anonymous. 1 April 2024. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBj

[2391] Anonymous. Editorial Board. Theoretical Computer Science, 990(??): ??, April 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000719.

Brandstadt:2024:FDI

[2392] Andreas Brandstädt and Raffaele Mosca. Finding dominating induced matchings in P_{10}-free graphs in polynomial time. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000197.

Capra:2024:MRP

[2393] Lorenzo Capra and Michael Köhler-Bußmeier. Modular rewritable Petri nets: an efficient model for dynamic distributed systems. Theoretical Computer Science, 990(??): ??, April 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000124.

Feghali:2024:TRW

[2394] Carl Feghali and Malory Marin. Three remarks on W_{2} graphs. Theoretical Computer Science, 990(??): ??, April 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000185.

Baste:2024:FAN

[2395] Julien Baste and Dimitri Watel. An FPT algorithm for node-disjoint subtrees problems parameterized by treewidth. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397524000215.

Shi:2024:AAM

[2396] Yishuo Shi and Xiaoyan Lai. Approximation algorithm of maximizing non-monotone non-submodular functions under knapsack constraint. Theoretical Computer Science, 990 (??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000240.

Acharyya:2024:CHS

[2397] Ankush Acharyya, Vahideh Keikha, Diptapriyo Majumdar, and Supantha Pandit. Constrained hitting set problem with intervals: Hardness, FPT and approximation algorithms. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000173.

Duchene:2024:SGC

[2398] Éric Duchêne, Valentin Gledel, Sylvain Gravier, Fionn Mc Inerney, Mehdi Mhalla, and Aline Parreau. Smash and grab: the 0 · 6 scoring game on graphs. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400032X.

Wang:2024:ECB

[2399] Luping Wang, Jie Chen, Huan Dai, and Chongben Tao. Efficient codebased fully dynamic group signature scheme. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000227.

Anonymous:2024:PSI

[2400] Anonymous. Preface: Special issue on theoretical aspects of computing. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000239.

Geffert:2024:SCB

[2401] Viliam Geffert, Dominika Pališínová, and Alexander Szabari. State complexity of binary coded regular languages. Theoretical Computer Science, 990(??):??, April 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000148.

Anonymous:2024:Ab

[2402] Anonymous. 12 April 2024. Theoretical Computer Science, 991(??): ??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBk

[2403] Anonymous. Editorial Board. Theoretical Computer Science, 991(??): ??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000884.

Maiti:2024:QCT

[2404] Arnab Maiti and Palash Dey. Query complexity of tournament solutions. Theoretical Computer Science, 991 (??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000379.

Wang:2024:ESF

[2405] Lusheng Wang and Boting Yang. Edge searching and fast searching with constraints. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000318.

Du:2024:CRR

[2406] Jiao Du, Lin Li, Shaojing Fu, Longjiang Qu, and Chao Li. Constructions of 2-resilient rotation symmetric Boolean functions with odd number of variables. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000446.

Casteigts:2024:SSP

[2407] Arnaud Casteigts, Timothée Corsini, and Writika Sarkar. Simple, strict, proper, happy: a study of reachability in temporal graphs. Theoretical Computer Science, 991(??): ??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000495.

Lu:2024:ACA

[2408] Tianhang Lu, Han Xiao, and Qizhi Fang. Approximate core allocations for edge cover games. Theoretical Computer Science, 991(??): ??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000483.

Wu:2024:THF

[2409] Xiaoyu Wu, Lili Mei, and Guochuan Zhang. Two homogeneous facility location games with a minimum distance requirement on a circle. Theoretical Computer Science, 991(??):

??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000136.

Jana:2024:AAD

[2410] Saswata Jana and Partha Sarathi Mandal. Approximation algorithms for drone delivery scheduling with a fixed number of drones. Theoretical Computer Science, 991(??): ??, April 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000574.

Bliznets:2024:FDM

[2411] Ivan Bliznets, Anton Bukov, and Danil Sagunov. Fair division with minimal withheld information in social networks. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000616.

Zhu:2024:PMP

[2412] Yizhao Zhu, Lanxiang Chen, and Yi Mu. Practical and malicious private set intersection with improved efficiency. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000586.

Contente:2024:CMF

[2413] Michele Contente and Maria Emilia Maietti. The compatibility of the minimalist foundation with homotopy type theory. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000367.

Hemery:2024:MOA

[2414] Mathieu Hemery and François Fages.

On a model of online analog computation in the cell with absolute functional robustness: Algebraic characterization, function compiler and error control. Theoretical Computer Science, 991(??):??, April 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000471.

Anonymous:2024:Ac

[2415] Anonymous. 21 April 2024. Theoretical Computer Science, 992(??): ??, April 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBl

[2416] Anonymous. Editorial Board. Theoretical Computer Science, 992(??): ??, April 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001002.

Dallant:2024:HFC

[2417] Justin Dallant and John Iacono.

How fast can we play Tetris greedily with rectangular pieces? Theoretical Computer Science, 992(??):

??, April 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000203.

Chen:2024:TSL

[2418] Meirun Chen, Cheng-Kuan Lin, and Kung-Jui Pai. A tree structure for local diagnosis in multiprocessor systems under the comparison model. Theoretical Computer Science, 992 (??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000598.

Kusum:2024:SNA

[2419] Kusum and Arti Pandey. Some new algorithmic results on co-secure domination in graphs. Theoretical Computer Science, 992(??):??, April 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000665.

De:2024:OHU

[2420] Minati De and Satyam Singh. Online hitting of unit balls and hypercubes in R^{d} using points from Z^{d}. Theoretical Computer Science, 992 (??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000677.

Poggiali:2024:QCK

[2421] Alessandro Poggiali, Alessandro Berti, Anna Bernasconi, Gianna M. Del Corso, and Riccardo Guidotti. Quantum clustering with k-means: a hybrid approach. Theoretical Computer Science, 992(??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397524000811.

Wu:2024:IMP

[2422] Hailun Wu, Lina Ba, and Heping Zhang. Integer k-matching preclusion of some interconnection networks. Theoretical Computer Science, 992 (??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000641.

Liu:2024:FAI

[2423] Zhaoman Liu, Yanbo Chen, Jianting Ning, and Yunlei Zhao. Fully anonymous identity-based broadcast signcryption with public verification. Theoretical Computer Science, 992 (??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000793.

Mitrana:2024:JCF

[2424] Victor Mitrana, Andrei Păun, Mihaela Păun, and José Ramón Sánchez Couso. Jump complexity of finite automata with translucent letters. Theoretical Computer Science, 992 (??):??, April 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000653.

Anonymous:2024:Ad

[2425] Anonymous. 27 April 2024. Theoretical Computer Science, 993(??): ??, April 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBm

[2426] Anonymous. Editorial Board. Theoretical Computer Science, 993(??): ??, April 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001129.

Poudel:2024:OSC

[2427] Pavan Poudel, Shishir Rai, and Swapnil Guragain. Ordered scheduling in control-flow distributed transactional memory. Theoretical Computer Science, 993(??):??, April 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397524000781.

Ju:2024:NOC

[2428] Yiao Ju and Shenwei Huang. Near optimal colourability on hereditary graph families. Theoretical Computer Science, 993(??):??, April 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400080X.

Zhang:2024:CTO

[2429] Xiaoqi Zhang, Tengxiang Lin, Cheng-Kuan Lin, Zhen Chen, and Hongju Cheng. Computational task offloading algorithm based on deep reinforcement learning and multi-task dependency. Theoretical Computer Science, 993(??):??, April 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400077X.

Chen:2024:SLL

[2430] Jianer Chen, Zirui Chu, Ying Guo, and Wei Yang. Space limited lineartime graph algorithms on big data. Theoretical Computer Science, 993 (??):??, April 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000835.

Fritsch:2024:ICC

[2431] Robin Fritsch, Younjoo Lee, Adrian Meier, Kanye Ye Wang, and Roger Wattenhofer. The impact of core constraints on truthful bidding in combinatorial auctions. Theoretical Computer Science, 993(??):??, April 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000823.

Chalopin:2024:FOL

[2432] Jérémie Chalopin, Manoj Changat, Victor Chepoi, and Jeny Jacob. First-order logic axiomatization of metric graph theory. Theoretical Computer Science, 993(??):??, April 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000756.

Qiu:2024:DGA

[2433] Daowen Qiu, Le Luo, and Ligang Xiao. Distributed Grover’s algorithm. Theoretical Computer Science, 993 (??):??, April 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000768.

Anonymous:2024:Md

[2434] Anonymous. 1 May 2024. Theoretical Computer Science, 994(??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBn

[2435] Anonymous. Editorial Board. Theoretical Computer Science, 994(??): ??, May 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400135X.

Schlotter:2024:RWP

[2436] Ildikó Schlotter. Recognizing when a preference system is close to admitting a master list. Theoretical Computer Science, 994(??): ??, May 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000604.

Sun:2024:NMT

[2437] Zhe Sun, Jincheng Zhuang, Zimeng Zhou, and Fang-Wei Fu. A new McEliece-type cryptosystem using Gabidulin–Kronecker product codes. Theoretical Computer Science, 994 (??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000951.

Guo:2024:PDH

[2438] JunYi Guo and Po-Chun Yu. Puzzle and dragons is hard. Theoretical Computer Science, 994(??): ??, May 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000926.

Jia:2024:HML

[2439] Wenjuan Jia, Jiang Zhang, Binwu Xiang, and Baocang Wang. Hardness of (M)LWE with semi-uniform seeds. Theoretical Computer Science, 994(??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000963.

Chang:2024:TSB

[2440] Hong Chang, Jing Jin, Zhicheng Liu, Donglei Du, and Xiaoyan Zhang. Two-stage BP maximization under p-matroid constraint. Theoretical Computer Science, 994(??): ??, May 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000847.

Kato:2024:SAC

[2441] Toru Kato and Yoichi Hirashima. The synchronized ambient calculus. Theoretical Computer Science, 994 (??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400094X.

Kim:2024:SCP

[2442] Sungmin Kim, Sang-Ki Ko, and Yo-Sub Han. Simon’s congruence pattern matching. Theoretical Computer Science, 994(??):??, May 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524000938.

Anonymous:2024:Me

[2443] Anonymous. 12 May 2024. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBo

[2444] Anonymous. Editorial Board. Theoretical Computer Science, 995(??): ??, May 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001415.

Dvorakova:2024:UBA

[2445] L’ubomíra Dvořáková and Edita Pelantová. An upper bound on asymptotic repetition threshold of balanced sequences via colouring of the Fibonacci sequence. Theoretical Computer Science, 995(??):

??, May 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001051.

Choudhury:2024:AST

[2446] Ashish Choudhury. Almost-surely terminating asynchronous Byzantine agreement against general adversaries with optimal resilience. Theoretical Computer Science, 995(??): ??, May 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001075.

Chakraborty:2024:RGI

[2447] Dibyayan Chakraborty, Kshitij Gajjar, and Irena Rusu. Recognizing geometric intersection graphs stabbed by a line. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001038.

Liu:2024:BWA

[2448] Tianyu Liu and Xiongxin Yang. Beyond windability: Approximability of the four-vertex model. Theoretical Computer Science, 995(??): ??, May 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001063.

Dutta:2024:RPC

[2449] Subhranil Dutta, Tapas Pal, and Ratna Dutta. Reinforcing privacy in cloud computing via adaptively secure non-zero inner product encryption and anonymous identitybased revocation in unbounded setting. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001178.

Okhotin:2024:GOB

[2450] Alexander Okhotin, Maria Radionova, and Elizaveta Sazhneva. GF(2)- operations on basic families of formal languages. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439752400104X.

Csuhaj-Varju:2024:NWC

[2451] Erzsébet Csuhaj-Varjú and György Vaszil. Networks of Watson–Crick D0L systems with communication by substrings. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001087.

Anonymous:2024:Mf

[2452] Anonymous. 21 May 2024. Theoretical Computer Science, 996(??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBp

[2453] Anonymous. Editorial Board. Theoretical Computer Science, 996(??):

??, May 21, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001609.

Dourado:2024:CHI

[2454] Mitre C. Dourado, Marisa Gutierrez, Fábio Protti, and Silvia Tondato. Computing the hull and interval numbers in the weakly toll convexity. Theoretical Computer Science, 996 (??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001166.

Kutrib:2024:PPS

[2455] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. On the power of pushing or stationary moves for input-driven pushdown automata. Theoretical Computer Science, 996 (??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400118X.

Fraigniaud:2024:PTB

[2456] Pierre Fraigniaud, Maël Luce, and Ioan Todinca. On the power of threshold-based algorithms for detecting cycles in the CONGEST model. Theoretical Computer Science, 996 (??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001154.

Hanaka:2024:GDP

[2457] Tesshu Hanaka, Hirotaka Ono, Yota Otachi, and Saeki Uda. Grouped domination parameterized by vertex cover, twin cover, and beyond. Theoretical Computer Science, 996 (??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001221.

Cicerone:2024:MPF

[2458] Serafino Cicerone, Alessia Di Fonso, Gabriele Di Stefano, and Alfredo Navarra. Molecular pattern formation on grids in the Moblot model. Theoretical Computer Science, 996 (??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001257.

Zakzok:2024:CFW

[2459] Mohammad Zakzok and Kai Salomaa. Converting finite width AFAs to nondeterministic and universal finite automata. Theoretical Computer Science, 996(??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439752400121X.

Comet:2024:CKS

[2460] Jean-Paul Comet, Hélène Collavizza, and Laetitia Gibart. Constrained Kripke structure for identifying parameters of biological models. Theoretical Computer Science, 996(??):??, May 21, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001208.

Anonymous:2024:Mg

[2461] Anonymous. 27 May 2024. Theoretical Computer Science, 997(??):??, May 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBq

[2462] Anonymous. Editorial Board. Theoretical Computer Science, 997(??): ??, May 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001737.

Fomin:2024:PCB

[2463] Fedor V. Fomin, Pierre Fraigniaud, and Petr A. Golovach. Parameterized complexity of broadcasting in graphs. Theoretical Computer Science, 997(??):??, May 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001233.

Matsuoka:2024:CCN

[2464] Tatsuya Matsuoka and Naoto Ohsaka.

Computational complexity of normalizing constants for the product of determinantal point processes. Theoretical Computer Science, 997(??): ??, May 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001282.

Harris:2024:TCT

[2465] Pamela E. Harris, Erik Insko, and Florian Lehner. Tipsy cop and tipsy robber: Collisions of biased random walks on graphs. Theoretical Computer Science, 997(??): ??, May 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001270.

Hara:2024:LSA

[2466] Keisuke Hara. A logarithmic-sized accountable ring signature scheme in the standard model. Theoretical Computer Science, 997(??): ??, May 27, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001312.

Wu:2024:NAF

[2467] Xiaoliang Wu, Qilong Feng, Jinhui Xu, and Jianxin Wang. New algorithms for fair k-center problem with outliers and capacity constraints. Theoretical Computer Science, 997 (??):??, May 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001300.

Rios-Wilson:2024:IUA

[2468] Martín Ríos-Wilson and Guillaume Theyssier. Intrinsic universality in automata networks I: Families and simulations. Theoretical Computer Science, 997(??):??, May 27, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001269.

Anonymous:2024:Jc

[2469] Anonymous. 1 June 2024. Theoretical Computer Science, 998(??):??, June 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBr

[2470] Anonymous. Editorial Board. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001841.

Araujo:2024:GCI

[2471] Samuel N. Araújo, João Marcos Brito, Raquel Folz, Rosiane de Freitas, and Rudini M. Sampaio. Graph convexity impartial games: Complexity and winning strategies. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400149X.

Poirier:2024:FBA

[2472] Léo Poirier and Wolfgang Steiner.

Factor-balanced S-adic languages. Theoretical Computer Science, 998 (??):??, June 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001506.

Konemann:2024:CNC

[2473] Jochen Könemann, Justin Toth, and Felix Zhou. On the complexity of nucleolus computation for bipartite b-matching games. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524000914.

Pasqua:2024:BEA

[2474] Michele Pasqua and Marino Miculan. Behavioral equivalences for AbU: Verifying security and safety in distributed IoT systems. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400152X.

Frati:2024:UPT

[2475] Fabrizio Frati. Upward planarity testing of biconnected outerplanar DAGs solves partition. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001567.

Buning:2024:CPU

[2476] Hans Kleine Büning. Classes of propositional UMU formulas and their extensions to minimal unsatisfiable formulas. Theoretical Computer Science, 998(??):??, June 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001531.

Naquin:2024:FSF

[2477] Émile Naquin and Maximilien Gadouleau.

Factorisation in the semiring of finite dynamical systems. Theoretical Computer Science, 998(??): ??, June 1, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001245.

Honkala:2024:RPS

[2478] Juha Honkala. Rational power series in several noncommuting variables and the Skolem–Mahler–Lech theorem. Theoretical Computer Science, 998(??):??, June 1, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001555.

Anonymous:2024:Jd

[2479] Anonymous. 12 June 2024. Theoretical Computer Science, 999(??):??, June 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBs

[2480] Anonymous. Editorial Board. Theoretical Computer Science, 999(??): ??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524002020.

Feng:2024:PPF

[2481] Yansong Feng, Abderrahmane Nitaj, and Yanbin Pan. Partial prime factor exposure attacks on some RSA variants. Theoretical Computer Science, 999(??):??, June 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001646.

Batra:2024:CRC

[2482] Prashant Batra and Vikram Sharma. Complexity of a root clustering algorithm for holomorphic functions. Theoretical Computer Science, 999 (??):??, June 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001191.

Lin:2024:HEM

[2483] Hao Lin, Mingqiang Wang, Jincheng Zhuang, and Yang Wang. Hardness of Entropic Module-LWE. Theoretical Computer Science, 999(??): ??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001683.

Fernau:2024:CCG

[2484] Henning Fernau, Lakshmanan Kuppusamy, and Indhumathi Raman. On the computational completeness of generalized forbidding matrix grammars. Theoretical Computer Science, 999(??):??, June 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001543.

Dore:2024:DFT

[2485] François Doré, Enrico Formenti, Antonio E. Porreca, and Sara Riva. Decomposition and factorisation of transients in functional graphs. Theoretical Computer Science, 999(??): ??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001294.

Li:2024:ODQ

[2486] Guanzhong Li and Lvzhou Li. Optimal deterministic quantum algorithm for the promised element distinctness problem. Theoretical Computer Science, 999(??):??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439752400166X.

Fazekas:2024:TCC

[2487] Szilárd Zsolt Fazekas, Naoya Iwano, Yu Kihara, Ryuichi Matsuoka, Shinnosuke Seki, and Hinano Takeuchi. Towards composable computations by RNA co-transcriptional folding: a proof-of-concept demonstration of nested loops in oritatami. Theoretical Computer Science, 999(??): ??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001658.

Zhang:2024:CIA

[2488] Xiaojin Zhang. Corrigendum to “Improved Algorithm for Permutation Testing” [Theoretical Computer Science 986 (2024) 114316]. Theoretical Computer Science, 999(??): ??, June 12, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001518. See

[2348].

Anonymous:2024:Je

[2489] Anonymous. 24 June 2024. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2024:EBt

[2490] Anonymous. Editorial Board. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524002184.

Behera:2024:UNZ

[2491] Bishnu Charan Behera and Somindu C. Ramanna. Unbounded non-zero inner product encryption. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001634.

Ichida:2024:SSD

[2492] Hiroyuki Ichida and Keiichi Kaneko. Set-to-set disjoint paths in a folded hypercube. Theoretical Computer Science, 1000(??):??, June 24, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397524001774.

Jiang:2024:BBT

[2493] Mingming Jiang, Shengli Liu, Shuai Han, and Dawu Gu. Biometricbased two-factor authentication scheme under database leakage. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001671.

Taghikhani:2024:HMP

[2494] Sepideh Taghikhani, Fahimeh Baroughi, and Behrooz Alizadeh. A hybrid modified PSO algorithm for the inverse p-median location problem in fuzzy random environment. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001890.

Brodal:2024:PQD

[2495] Gerth Stølting Brodal. Priority queues with decreasing keys. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397524001786.

Bonfante:2024:GRL

[2496] Guillaume Bonfante and Florian Deloup. The genus of regular languages and directed graph emulators. Theoretical Computer Science, 1000 (??):??, June 24, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001944.

Wu:2024:SNP

[2497] Ruina Wu and Yuzhen Zhao. Spiking neural P systems with structural plasticity and mute rules. Theoretical Computer Science, 1000(??): ??, June 24, 2024. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397524001695.